A Bibliography of Publications about the Python Scripting and Programming Language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

27 April 2024 Version 2.201

Abstract

This bibliography records books about the Python scripting and programming Language and related software.

Title word cross-reference

#105 [Bri12a]. #106 [Bri12b].

1 [BRF^{+}23, DB17, Hug18, RKVL14, Sch21]. 1.5 [SAA18]. $105 [Liu22]. $136.46 [Sau23]. 2 [NT24, SLDF^{+}21]. $24.95 [Lee17]. 3 [Bra13, CC20, ML16a, NS22, Pet02, PBN^{+}09, Ras20, SDS00, ZL20]. $34.95 [Ano00a]. $44.95 [Ano97a]. $63.96 [Lip22]. $99.95 [Hor22, Lip21]. ^{1}Σ [BGHC23]. β [WBS21]. D [HH17]. Γ [DEMM19]. μ [Bie21]. N [HH17]. ν [Bie21]. Q [JWHS16, MSP^{+}22]. UnBlocks^{gen} [Ras20].

-amino [WBS21]. -d [CC20, Pet02, RKVL14]. -from- [Bie21]. -LSKUM [MSP^{+}22]. -method [DEMM19]. -peptide [WBS21]. -Statistics [JWHS16].

. [Lip22]. .NET [SKS08, HF06, SM04, Stu07].

/Python [LWH12, SV14].

1 [ddSNX22]. 1.0 [Bal18, BS19, DCRF23, KPK^{+}17]. 1.1.0-pyfao56 [BTDT23]. 1.x [Sal18]. 10 [SV14]. 1004 [RG20]. 1016 [HFB21]. 1025 [PLRG22]. 10th [ACM97]. 11 [Ano18]. 11th [IEE97a]. 16 [HHJCRB21]. 1997 [Ano97d].

2 [DPSD08, FNX22, GNPP23, JNN13, LS17, MP19a]. 2-D [GNPP23]. 2.0 [Ano00c, MBW07, Seg07, SJL18, Yes15]. 2.0-pyFIRI [ZRK22]. 2.2 [Chu02b]. 2.6 [Pay11]. 2000 [Ano00c]. 2001 [Ano01a]. 2009 [Jan10]. 2020 [CLT20]. 23 [ESM98]. 251pp. [Gve09]. 27th [CLT20]. 2D [ddSNX22].

3 [Bea09a, Bea12, BJ14a, Fas18, HB16, Kiu13, LD19, Lee17, MP19a, MCGK19, MSV23, MH15, MOM21, Per14b, Phi10, Phi15, Pil09, Sa’23, SW15, SS21, Sum09, Sum10]. 3.0 [MMEH08, Per14b, RSPJ21, Wei09]. 3.1 [Pay11]. 3D [FNX22, PAB^{+}97, SML06].

4 [CSZ^{+}19, Cla15]. 4th [Ass96, USE00].

5 [Dan18a, Liu22]. 56 [BTDT23, Tho22, TBD23].

6 [Ano13, Hor22]. 6th [Ano97c].

7 [Ano15a]. 77 [GH18, GHN19].

8 [Lip21]. 839 [Kir04].

9 [Ano15b]. 910 [Kor11]. 923 [Wim12]. ’96 [HK95, Ass96]. ’97 [ACM97]. 978 [Ano13, Ano15a, Ano15b, Cla15, Hor22, Lee17, Lip21, Liu22, Sa’23]. 978-0-12-418676-7 [Ano14]. 978-0-12-418683-5 [Ano14]. 978-0-367-74845-6 [Hor22]. 978-0-521-72596-5 [Gve09]. 978-0-691-17050-3 [Lee17]. 978-1-032-04174-2 [Lip22]. 978-1-108-49433-5 [Liu22]. 978-1-138-48395-8 [Lip21]. 978-1-4665-0455-4 [Cla15].

978-1-59327-590-7 [Ano15a]. 978-1-59327-640-9 [Ano15b]. 978-1-59749-957-6 [Ano13]. 978-981-16-8615-3 [Sa’23]. 997 [Spe19].

ab-initio [PP23]. Abaqus [GTC21, SMD23]. ABCal [Le 24]. absolute [Daw03, Daw06, Daw10, MTS^{+}18]. Abstract [CBW^{+}21]. Abstraction [Hen08, LP19]. Abstractions [RHM^{+}17]. accelerate [LHM14]. accelerated [CWLG^{+}21, ECS23, MSP^{+}22, NT24]. Accelerating [ZLBF14, YPB16]. Acceleration [BKH^{+}22, JS24, Wu13]. Accelerator [RSRT19]. accelerators [Mar18]. acceptance [Sal14]. access [BB17]. Accessing [Buc14]. Accord [Sta19]. Account [Luk23]. Accuracy [LV20]. accurate [LJ23, LN23]. acid [WBS21]. Acquisition [CWM^{+}21, Myr15]. activation [SH19b]. Active [BBB02, CHG^{+}16, RKR21, Mit00]. Active-subspaces [CHG^{+}16]. ActiveState [Ano00c]. ACTS [DGMP09]. Ada [Bri12a, Bri12b, Mar21, Men09, Mos20b, Och09, RAH^{+}01, Ver22b]. Adaptive [CHM24, CV22, SVY09]. Adding [IOC^{+}12, Tab10]. Additivity [WV22]. Adjoint [HA20]. Administration [Bru09, GJ08, Sil14]. adsorbed [SSH16]. Advanced [Chu07a, LT03, Sha24, DDK19, Lan13]. Advances [Poo16, Spo12]. Adventures [Ric15]. affine [RCAE^{+}20]. After [DF21a]. again [Cas20]. against [Pre03]. age [LL08]. Agent [EVV12, KFV20, NZPWR22, GG21]. Agent-Based [KFV20, NZPWR22, GG21]. agents [HMD^{+}96, Lot14b]. Agglomerative [Mül13]. aggregate [VFMM08]. Agile [Suc13, VB08, You08]. Agoren [Ano01a]. Agreement [EVV12]. Agresti [Hor22, Sau23]. agricultural [ZTC^{+}21]. AI [Sav23, ZGL20]. AIL [vd91]. Ajax [RDS07]. Ajax-powered [RDS07]. aktuell [EK08]. aktuelles [Lin06b]. Alan [Hor22, Sau23]. Albany [LPH23]. Alberta [ACM97]. Alexander [Ish17]. Algebra [PBB22, BCG05, JČMG11, LKJC21, RV20, Sah15]. AlgoPy [WL13]. Algorithm [BTR14, HWJ^{+}21, Luk23, Wit15, AWO^{+}21, HFB21, Kir04, Kor11, PLRG22, RG20, Spe19, Tau16, Wim12]. Algorithm-Centric [HWJ^{+}21]. Algorithmic [WL13]. Algorithms [BH12, DD15, GMP^{+}20, Het10, Het14, Jan10, KBLJ18, OOM^{+}23, CFPS23, Di 13, DV21, KG23, RZ09, SUM21, VBM^{+}18, Wep15]. Alias [GLS^{+}10]. Alice [PAB^{+}97, UVa95]. alike [Thu22]. all-atom [HQF^{+}20]. alle [Lin06b, The09]. Allen [Gve09]. aller [DF00]. allow [BM18]. alternative [MSS^{+}19]. always [Buc15]. amino [WBS21]. Amit [Ano15b]. AMLCS [NRC23]. AMLCS-DA [NRC23]. amplification [SAD22]. amplitudes [HJK^{+}24]. AmPyfier [SAD22]. ana cont [KH23]. analyse [VM22]. analyses [Le 24, MFL^{+}22, Spi18]. Analysing [ZTC^{+}21]. Analysis [BHS18, CRL21, CZZ19, CH17, GdGB^{+}18, GDP18, HU17, IEE20, ISMA18, KSB12, LV20, LS16, MAFM21, Mar21, McK12, Mil18, MB07, OZW18, PBK22, Sha24, Sta21, TGEA09, TLR21, VPO19, WBR^{+}22, XL18, AM19, ACS23, BB13, Bow15, BYL^{+}21, CC20, Cue13, DLC23, FPSZ21, Gal22, GFB^{+}14, GLS^{+}10, GHKW22, GEH19, HS22, HOB23, Hug18, Idr14, Ish17, IOC^{+}12, IHT23, ICVG14, JKKN23, KSS20, KS21, KXK^{+}23, LRL23, LYZ^{+}22, MP19a, MSP^{+}22, MPMC21, Mül23, NMGB17, Nel15, OLRLB21, PM23, PFLG21, Ras20, RJAL09, RGP^{+}22, RRRGVD19, Sal18, SHFJ18, SLDF^{+}21, SL20, Sto13, SH19b, Tom15, VCLS21, WPM^{+}20, Wes15, Wet20, WKS22, Yes15, ZWCQ22, LFT21, PARS22, Vel23]. analysts [O’C13]. analytic [KH23, MD22a, TF23]. Analytical [NCS17, MS15, She15b]. Analytics [EPF^{+}23, Lip22, Piñ24, ÇA23, CFÁA^{+}20, DDK19, Mil14, Mil15, Nel15, RO15a, SK19]. analyze [Lay15, Lot14b, MB17, OBWM23, Wes15, ZAPS20]. analyzer [MTPHH18]. Analyzing [GRG21, GPEM23, HFF^{+}17, MK16, PMBF17, YHX22]. Andrew [Ish17]. Android [Bar11]. Andy [Ano00a]. angewendet [Wei06a]. angle [MTPHH18, OLRLB21, RGP^{+}22, SL20]. angle-resolved [SL20]. animation [vdH18]. anisotropic [BVV22]. Annotated [Di 13, Wep15, Bro00]. Annotator [HYG16]. Annual [ACM97, Ass96, USE00]. anomalies [MTS^{+}18]. anomaly [BJM23]. Ansible [BB22a, DDT20]. AnsibleMetrics [DDT20]. antenna [SDP^{+}20]. Anthology [Anoxxb]. Antlab [GMS17]. Any [Pre03]. Apache [TFAL21, TSC02]. APEX [Wai21]. API [Ano98b, AJH^{+}12, BG17, FSZD20, HZ23, Kuc98c, Men09, MBA19, SGPHD^{+}17, Smi17, van98c]. APIs [BLS^{+}23, Meh15]. App [Pip15, San15]. Application [Ano01b, BSSz^{+}20, EML21, Gut16, Orb18, Bah15, BL97a, Buc15, CS21, HQF^{+}20, HCPF95, KXK^{+}23, OE21, Ree04, Rem01b, SV14, CMS22]. application-programming [SV14]. Applications [AJH^{+}12, Ble20, CBST24, DGMP09, Has16, ML16a, MSR03, Nil07a, PBK22, Pul00, RM19, SRS^{+}23, Shi07, AFL23, BL97b, BCG05, CV15, CB96, DAJ^{+}15, Di 13, GMP^{+}20, Gup15, HF06, Har12, Hig03, Jos15, Kar14, Lot14a, MLB22, Pip15, RDS07, RG10, RCRS06, SW15, ST^{+}07, SML06, Seg07, SLA^{+}23, Ull15, Wei15, Wep15, Yan14, YHA^{+}16, Zac15]. Applied [Sha24]. apply [Idr14]. applying [Sal14]. Apprendre [Swi09]. apprenticeships [Lon08]. Approach [CHM24, EVV12, Joh15, Liu22, Mat16, Mis19, Osa17, SMF^{+}23, Arb14, BBHB22, Bro06, FMPS17, HHK^{+}09, PP23, SML06, SWD15, Sha03]. Approaches [KLM14, KLM15, Bra13, MD15, RMZG06]. Approximate [FV18]. approximation [AYI^{+}24]. approxposterior [FV18]. apps [Sal18]. April [IEE97a]. arbitrary [JT14]. arbitrary-precision [JT14]. ARC [RSPJ21, Ano00c]. ArcGIS [Tom15]. ArcGIS-geospatial [Tom15]. Architecture [HM18, LLL^{+}20]. Architectures [SBC^{+}17]. archival [Ver22a]. Archive [Wal97]. archives [Bro00]. ArcPy [Tom15]. Arduino [DAJ^{+}15, KSH14]. Area [SYGY21]. ARITH [CLT20]. Arithmetic [CLT20, KJ14, Bri06, JT14]. Arizona [Ano97d]. ARL’s [War18]. ARM [HC16]. ARPIST [LJ23]. array [Fas18]. Arrays [HH17, JM20]. art [MSV23, Wei15, FLS^{+}20, HC16]. Article [Wat95]. Artificial [Tan06, DSB^{+}23, Sa’23]. ASCII [IG19]. Asia [IEE97b]. aspiring [Lot14b]. assembly [TBJ18]. Assessing [LJ19]. assessment [SUM21]. assets [TTHZ23]. Assimilation [JCL^{+}20, NRC23]. assisted [WRBT21]. Associates [Jan10]. Assurance [WL20]. AST [HWW^{+}15, ZLBF14]. Astroinformatics [MAC14]. astronomical [HJPB17]. astronomy [Ish17, ICVG14]. astrophysics [MAC^{+}21]. asymmetric [CFSK14]. Asynchronous [PLRG22]. Atlanta [IEE20, USE00]. Atmospheric [vdOJP^{+}20, NRC23]. ATNF [Pit18]. atom [HQF^{+}20]. Atomic [OPA^{+}14, Ren23, RSPJ21]. atomistic [HHP^{+}16, WBK^{+}24]. Attacks [OOM^{+}23]. Attribute [TLR21]. Auction [EVV12]. Auction-Based [EVV12]. AUGMECON [FPN^{+}22]. AUGMECON-Py [FPN^{+}22]. auroral [ZRK21, ZRK22]. Ausdrücke [SKS08]. Authentication [SST23]. author [Le 24]. AutoCNet [LRPD18]. autoDIAS [SH19b]. automan [Ram18a]. Automate [Swe14, Tom15]. Automated [ABCC22, Bea98, HAB^{+}20, LMW12, LLL^{+}20, MCF^{+}22, TS14, TLR21, AAYK23, Law15, LKF23, MBA^{+}22, SH19b, LW10].

Automatic [LV20, MAC14, MO14, RCAE^{+}20, RJAL09, AG23, SL21]. Automatically [CYKG23]. Automating [Ble20, RHM^{+}17, JAGP14]. Automation [OZW18, Ram18a, Ste06]. autonomous [FMPS17, Len15]. AutoParallel [RCAE^{+}20]. avancées [Chu07a]. awarded [Bal12]. aware [ABF13, VC18]. Awk [Bla02, Bla04]. AWS [Ano20]. axisymmetric [MFQ^{+}21]. Azure [TCE21].

B [Gve09, Lip21]. b/w [Lip21]. backed [YHA^{+}16]. Bad [Bea09a]. badcrossbar [JM20]. Banded [Wim12]. Banff [ACM97]. BANSHEE [KMLP^{+}23]. Based [ABCC22, Ang00a, Aya14, BBB02, CHM24, DGMP09, EVV12, FLS^{+}20, HKGvS21, HWJ^{+}21, KFV20, Nor17, PBK22, Ram18a, RSRT19, SN12, SST23, Sta21, VBI^{+}23, ZGL20, AZH^{+}24, ABGD^{+}20, AZLY18, BBHB22, BKMY03, BKMY04, BWMS22, BB17, CS21, CFMR19, CM20, CJYH23, CB96, CKC23, Dav21, Dec04, FNX22, Gal22, GP22, GCBDDBF23, GPP^{+}21, GG21, GYWG23, Gup15, JAGP14, JLP23, KG23, KSH14, KXK^{+}23, Lac06, LHH^{+}21, LYX^{+}18, LYZ^{+}22, LDAL20, NL23, NZPWR22, OL17, Ott18, OBWM23, PFH^{+}16, PBN^{+}09, RS17, Rad06, Rad08, RBP^{+}21, SUM21, SLB^{+}23, SABEh20, SFC23, SLA^{+}23, TEG18, TOPC23, Ver22a, VCLS21, WBS22, WPH^{+}22, YKKD19, YMCF23, ZXX23, ddSNX22, dPPD^{+}23, FLR22]. bases [Swi09]. BASEveg [CVV23]. Bash [Bla02, Bla04, Ano12]. Basic [Het14, Tro96a, Het10]. basierenden [Lac06]. basierte [Ric14]. basis [Kir04]. Bastard [RAH^{+}01]. Batteries [Dub07]. Battery [GPKL^{+}20, HBA^{+}20, KBC^{+}20, SMT^{+}21]. Battery-free [KBC^{+}20]. Battery-Testers [GPKL^{+}20]. Bayes [Sto13]. Bayesian [KMLP^{+}23, DF21b, JG17, PHF10, PLRG22, SAB^{+}20, Sto13, SBMD14, Laz22]. bayesint [DF21b]. BDD [Sal14]. BeagleBone [Hia15, Ric13, Ric14, Buc14]. beam [THG^{+}23]. Beautiful [Nai14]. become [BM15]. Bed [Ngu08a]. Beedle [Jan10]. BEEP [HBA^{+}20]. beginner [Daw03, Daw06, Daw10, Swe13, Sha24]. beginners [Can14, Cha15, Fla08, She15a, Swe14]. Beginning [Het05, Het08, KM15, McG07a, Mue14, Nag17, Pay11, Vai09, Vai14]. behandelt [LAG07]. behavior [WMM18]. Behavioral [Yaa23]. Behaviors [GRG21]. Beispiele [The09]. Beispielen [The09]. believe [KI19]. BEM [ŚBA^{+}15]. benchmarking [AES^{+}22, JLP23]. Benchmarks [DLC23]. Benefits [Kra22, CEI^{+}12]. Best [Hug97, Sco17, LS98, ZKB^{+}08]. best-bet [LS98]. bet [LS98]. Beta [DF21b]. Better [CCM^{+}23, Sit02, ZXX23, SW14, Sla15]. between [BLN^{+}21, DOS17, EAMSR^{+}23, LFT09, MRT^{+}22, Sta17]. beyond [BMK03, GL08b, GL09, SB15]. BFree [KBC^{+}20]. bias [Le 24]. bibliographic [Vel23]. bibliometric [Mül23]. bibliometrics [RK19]. bifold [Kar23, Kar23]. Big [AAB22, SBC^{+}17, VSS17]. BigMPI4py [AAB22]. BIL [BM18]. Bildbearbeitung [DF00]. Bilder [DF00]. BiLSTM [PH23]. Binary [AVS20]. Bind [Cot03]. binding [JPJ^{+}23, KVSC21]. Bindings [LFN^{+}11, RJAL09]. Biocomputing [HK95]. Bioinformatics [AVS20, Ant15, Fla08, Bas08, Bas10, Kin09, SNB^{+}19, SB15]. Biological [AJH^{+}12, VBB18, XL18, ZV19]. Biological-Modeling [AJH^{+}12]. biologists [LHB14]. Biology [MGS07, Di 13, LHM14, SB15, Wep15]. biomolecular [CBB14, CCFB16]. bionic [KI19]. Bivariate [Sha24]. Black [BLN^{+}21, Sei14, Buc14, Ano15a]. Black-box [BLN^{+}21]. Blackboard [UVa95]. BLAKE [AMPH14]. blame [VSS17]. BLAS [FSZD20]. BLASFEO [FSZD20]. Blaxxun [Ano00c]. Bleeding [Wil97a]. blender [TTHZ23]. Blindspots [BLS^{+}23]. Blitz [AJJF14]. Blockchain [PBK22, SST23]. Blockchain-Based [PBK22]. Blockchain-Enabled [SST23]. Blood [DB17]. blueprints [DDT20]. Boca [Hor22, Lip21, Lip22, Sau23]. body [Gra18]. Book [Ano97b, Ano97a, Ano00b, Ano00a, Ano13, Ano14, Ano15a, Ano15b, Bax01, Gve09, Hor22, Hug01, Ish17, Ish19, Jan10, Joh97, Lak17, Laz22, Lee17, Lip21, Lip22, Liu22, LMW12, Mat16, Orb18, Roo98, Sa’23, Sau23, Cha15, GHT06, GHT08]. Bookshelf [McG98c, Wil97b, Wil00, Wil01, Wil02]. Boost [Moo21, Lan13]. boring [Swe14]. Borland [Ano00c]. Bose [Nil07a]. Both [Hug97, Sma12, HB16]. Bottleneck [PFLG21]. Boundary [CFCB17, CCFB16, ŚBA^{+}15, CBB14, KRH18a, KRH18b]. Boundary-integral [CFCB17]. Box [RAH^{+}01, BLN^{+}21]. Boy [RAH^{+}01]. BPMN [Fun23]. BPpy [Yaa23]. brain [KG23]. breadth [Sha03]. breadth-first [Sha03]. Bridging [ACZ^{+}22]. Bringing [Mos20b]. broken [MWM20]. Browser [BP17, CB96]. browser-based [CB96]. BSD [DF00]. BSP [Hin03, HLSØ06]. Buch [vF97]. Buchs [The09]. BuckinghamPy [KS21]. bug [HZ23, KCS11, LWH^{+}10]. bugs [LFYX23, Ros08]. Build [RM19, Ada14, Bah15, Jos15, Lot14a, Nai14, Pip15, SML06, Tos09, Yan14, Zac15]. Builder [Dav21]. Building [BL97b, CR15, Gup15, Kar14, KDC^{+}18, Rem01b, Ste06, DPH16, HF06, Hig03, HWW^{+}15, RG10, Sal18, Seg07, VBM^{+}18, WBS21]. Buildings [ZGL20]. built [DTM^{+}18, Hos12]. built-in [DTM^{+}18]. bulk [SLDF^{+}21]. Business [Ano01b, Fun23].

C [DF00, Jan10, Ano18, Bai08, Bea96, Bea98, BFM18, Bri06, BM18, CSZ^{+}19, Cot03, Dal01, EPM09, GH18, GHN19, HZ23, HAGH18, Kor11, LPH23, LWH12, MSP^{+}22, Mar21, MBA19, Nag06, Ngu08c, NZPWR22, OE21, Pre03, Pul00, RZ09, Sah12, SV14, SML06, SGPHD^{+}17, SMD23, Smi18, SM04, Stu07, WX18, WMA12, Wil00, Yes15, Zho97, van98c]. C# [SKS08]. C/C [Bea98, Mar21, Pul00, WMA12]. CA [Ano97a, Ano00a, Ass96]. CAD [Ano00c, CC20, Ngu08d]. Cadabra [CFPS23]. Cait [WBR^{+}22]. Calculate [HKGvS21, MWM20]. Calculating [DF21b, BGHC23, Gre16, SP23]. Calculation [RG20, TLR21, VBI^{+}23, CFSK14, DCRF23, JKST22, Kar23, MR18, QZWU19]. Calculations [Kor11, OZW18, AM21, BLN^{+}21, CMS22, Cri18, Gra18, JPJ^{+}23, KPK^{+}17, RSPJ21, VCF22, WZZ^{+}23]. Calculus [BH12, Sah15]. calibration [CWC^{+}23]. California [ACM92, Ano97c]. caloric [SAA18]. CALPHAD [OL17]. CALPHAD-based [OL17]. Cambridge [Gve09, Liu22]. cameras [CWC^{+}23]. CameraTransform [GRW^{+}19]. Can [Dör08, CBB14]. Canada [ACM97, USE96]. canonical [PHH^{+}12]. can’t [KI19]. capabilities [She97]. capacity [Zie19b]. Capsule [PH23]. capture [SGPHD^{+}17]. Careers [Ano20, Cas20]. Carlo [ABGD^{+}20, DMD^{+}21, DEMM19, HQF^{+}20, JKKN23, KXK^{+}23, Mor17, Nil07b, PHH^{+}12, WAN^{+}22]. carrier [TQGE23]. Case [JWHS16, Kra22, AJJF14, HPH12, LJ19, LHM14, OMGDG14, RCRS06, SSS22, TFAL21, ÜK12, BB13]. Case-Control [JWHS16]. Cassandra [DMD^{+}21]. catalogue [Pit18, Var16]. Causal [ZL23, SMM^{+}22]. cavities [CBB14]. CD [DF00, Lin06b, The09]. CDFTPY [VCF22]. cell [RKVL14]. cellML [CM07]. census [Smi17]. Centralities [PKDM21]. Centric [HWJ^{+}21]. CESSIPy [CR22]. CFD [MBLA19, TGEA09]. CG [AJH^{+}12]. CGI [Wei06a, Kuc98b, RG00, Van97b].

CGI-Scripting [Wei06a]. cgimodel [RG00]. chain [HQF^{+}20, Mau02, You08]. Chains [GHXZ24]. Challenges [Sie17]. Chameleon [SVY09]. Change [KFV20, Cas17]. Changes [RAH^{+}01, YHX22]. Changing [SYGY21]. channel [SGPHD^{+}17]. Chaospy [FL15]. Chaotic [VBI^{+}23]. Chapman [Ano01a, Lip21, Sau23]. characteristic [LYZ^{+}22]. characteristics [CSRV13]. Characterizing [GHXZ24]. charge [ABS20]. charged [SLP^{+}22]. charges [FH22]. Charts [SJL18, Ada14]. ChebTools [Ano18]. Chebyshev [Ano18, SST23]. checking [Hen08, RF16]. chemical [HPT^{+}16, NCS17]. chemically [Gal22]. chemistry [BLN^{+}21, Dah18a, KRSD^{+}23, PSGL21, VAP^{+}21]. ChemPy [Dah18a]. Chet [Ano14]. CHICOM [GHN19]. Chimera [HCPF95]. Chityala [Myr15]. CHIWEI [GH18]. Chlorophyll [RKR21]. Chlorophyll-a [RKR21]. choice [AJJF14]. Choose [CLMM20]. Choosing [LS97]. Christian [Lak17]. Chromatin [CSZ^{+}19]. ChromStruct [CSZ^{+}19]. cij [LDW^{+}21]. Ciphers [Swe13]. Circuit [MAFM21]. circular [CJYH23]. Circulation [NRC23]. city [SUM21]. civil [CR22]. Class [Ngu08a, Hig03, LYZ^{+}22, RS17, TSD^{+}12]. classes [How98, TSD^{+}12, Ngu08c]. classical [NM22, VCF22]. classifiability [ZWCQ22]. Classification [AVS20, BHF16, PH23, VBI^{+}23, CFSK14, Sta17]. classifier [CKC23]. clear [Ram15]. Client [Lac06]. Climate [HKGvS21, GWM18, Ros18, WHG17]. CLIMLAB [Ros18]. Clinamen2 [WBK^{+}24]. closed [Ver22a, XGW23]. closed-loop [Ver22a, XGW23]. Cloud [Kra22, Suc13, BB22a, Kla11, NEGZG18]. Cloud-Native [Kra22]. Cluster [BUS21, BLE21]. Clustering [JWHS16, Mül13, TOPC23]. Clusters [GHXZ24]. CMInject [WAK22]. CMIstark [CFSK14]. CMU [Mac91, Mac92a, Mac92b].

Cnerator [OE21]. CNNs [SH19a]. co [Wu13]. co-designed [Wu13]. coastal [MAS^{+}22]. CoastalImageLib [MAS^{+}22]. coaxial [YMCF23]. Cobol [Ano01b]. coco [Sta17]. Code [Ano00c, Ano01b, BLS^{+}23, CSZ^{+}19, Fun23, GHN19, Har15, HAB^{+}20, Mit00, The09, VMFG17, XMW10, ABGD^{+}20, AM10, ABCV21, Ber13, BC09, Buc15, CZA^{+}23, Cas17, CBLI22, Day07b, DBdFdSR21, ESM24, GH18, JKST22, Kar23, Lot15b, LDW^{+}21, MCF^{+}22, MPMC21, MO14, OE21, PP23, RBV16, Ros08, SML06, SP23, SMM^{+}22, SZNW23, SGZ23, WMM18, WZ18, YHX22, DDT20]. Code-Beispiele [The09]. Codes [Anoxxd, Hig22]. codice [Day07a]. Coding [MAFM21, Cha15, DV21, ZKB^{+}08]. Codon [SRS^{+}23]. coefficient [LN23]. coeur [Chu07b, Chu07a]. COFFEE [DFSW19]. Cognitive [KXK^{+}23]. coherent [MR22]. coils [MFQ^{+}21]. cold [MTPHH18]. cold-neutron [MTPHH18]. Collaboration [Ano01b]. Collaborative [SYGY21, VSS17]. Collation [Tau16]. Collected [Ngu08a]. collecting [Mit15, PMBF17]. Collection [GMN21, Ngu08d, Ngu08e]. collections [BDT13, DTM^{+}18, SVY09]. collective [Seg07]. Collector [Roo97]. color [Lip22]. Combinatorial [ZTT22]. combined [AM10]. Combining [Zie19a, MS15]. Comfort [ZGL20, BDQ^{+}22, TS20]. Command [CC20, IG19, CWM^{+}21]. commandi [Day07a]. commands [Day07b]. Comments [Roo97, VMFG17]. commercial [Tab10]. Common [Mac91, Mac92a, Mac92b, MBA19, RO15b, RAH^{+}01, MAS^{+}22]. Communication [SST23]. Communications [Ano00c]. Communities [GPEM23]. community [Gar09]. Comparative [VPO19, Kak08, Mau02, ZXX24]. comparing [GHN19, MD15]. Comparison [Pre00, Zho97]. Competitive [DV21]. compilable [LKJC21]. Compilation [NM22, MD15, WMA12]. Compiler [Ano01b, SRS^{+}23, CEI^{+}12, HS12, IOC^{+}12, Mac92b, Ott18, TTS^{+}10]. compilers [ZXX24]. Compiling [CG23, Fee16, CGK11]. complete [BS19, DMD^{+}21, Meh15, Sum09, Sum10]. completely [Orl97]. completion [SZW^{+}22]. complex [LR14, She15b, WM21, YMCF23]. Complexity [DD15, ABCV21, KCS11]. Compliant [Ano01b]. components [HCPF95]. Composable [MLGW18]. Composing [RHM^{+}17]. composition [BWMS22]. composition-structure [BWMS22]. COMPOSTI [NL23]. compound [AAYK23]. compounds [DMC^{+}15]. comprehend [Cox14]. Comprehensive [RO15b, RG10, Tel06, You08]. compressible [GP22, Par22]. Compressive [OPA^{+}14]. Computation [AMGM20, BKH^{+}22, Coe17, Gut13, Gut16, HA20, Laz22, LLL^{+}20, Orb18, SL21, Wim12, CM20, Gir21, Guz03, Le 24, OC20, SHTE23]. Computational [Ano21, Aya14, Bäc07, Bar21, Ble20, HM18, LT03, Lan08, MS07, OL17, VMFG17, Vir16, Bor07, HHK^{+}09, Lan06, Lan07, MDRN18, PSGL21, PGM19, RGS^{+}21, SMM^{+}22, TBA^{+}17, VAP^{+}21, Gal22]. Computations [KJ14, AM10, CLM05, RHLTG^{+}22, YFD98]. compute [EAMSR^{+}23]. Computer [BCE^{+}22, CLT20, Dow15, IEE97b, MR07, Osa17, Can14, Dow09, Jos15, JČMG11, MTS^{+}18, MH15, Rad06, RMZG06, San13, Swe13]. Computers [BT06, JLP23]. Computing [AJYH18, BKRT21, BSSz^{+}20, DD15, HTA^{+}97, HM18, IEE20, LHB14, LD07, NM22, Oli07, PG07, PGH11, Ram18a, TT21, VB08, BCRS15, BS19, BL97b, CSRV13, Cla15, Dan18a, DMC^{+}15, DDMS14, DCOC^{+}19, EBNS22, FKA^{+}17, HHVB21, Her14, JM20, Kir04, LW10, Lub14, LS17, Mar18, Meh15, MSP^{+}17, MBK09, Nil07a, Ras18, Ros13, RFG13, Ros14, SLB^{+}23, SAA18, SFC23, YPB16]. concentrations [DMC^{+}15, TQGE23]. concentré [Mar07]. concepts [BM15, ÜK12]. concise [Joh12, Ram15]. concurrency [Tab10]. concurrent [WO14]. condensates [Nil07a]. condensed [Sch21]. Conditional [Sco17]. Conditioned [CJ22]. conditions [CMS22]. Conference [ACM92, Ano97c, Eur91, HTA^{+}97, IEE97b, IEE20, USE96, USE00]. Confidence [PRH17]. configuration [Hos12, LJ19, QZWU19, VC18]. confusing [AAYK23]. confusion [HJHZ18]. ConIII [LD19]. Connolly [Ish17]. consistent [FH22, SL21]. console [Rem01b]. constants [DCOC^{+}19]. constrained [MWK^{+}20]. constructed [HCPF95]. Constructing [ZV19, CB96, Ver22a]. contact [OLRLB21]. content [GFB^{+}14, May21]. context [MR09, MBK09]. contiguous [NO23]. continuation [KH23, VY15]. contract [Plo97]. contraction [SG18]. Control [BBB02, JWHS16, KB07, OOM^{+}23, AG23, BB22a, BST^{+}17, BE20, CJGA23, GCBDDBF23]. control-oriented [CJGA23]. Controlled [WX18, OE21]. Controlling [ZGL20, KSH14]. Convenient [LD19]. conversations [FSMCFP23]. conversion [McF16]. convert [BRF^{+}23, GZT^{+}18]. converter [SMD23, Sta17]. converting [Sta17, VAP^{+}21]. Convex [Ble20]. Cookbook [Mar02, Roh16, XDR21, Ant15, BJ14a, Buc15, Cox14, Law15, MMA05, O’C13, Per14b, Ros14, Sar14, Zac15]. cool [Jos15]. cooling [EBNS22]. COOTS [USE96]. Copatrec [KF23]. Copperhead [CGK11]. Copula [YH21]. CORBA [MSR03]. Core [Bax01, Chu01, Chu07c, Chu07b, Chu07a]. Corner [Ano01a]. corner.py [FM16]. corpus [Joh08]. Correction [Ano02]. Corrections [Spe19, GRW^{+}19]. correctness [Hen08]. correlated [SP23]. Correlation [FNX22, KF23, MR18, RCT23, CWLG^{+}21, ddSNX22]. correspondence [LRPD18, RS17]. cosmology [CFPS23]. cottoncandy [NEGZG18]. count [VM22]. Counterexample [LP19]. counting [GdGB^{+}18]. country [Sta17]. Coupled [HM18, CVV23, Roa23]. Coupling [HM18]. Course [WX18, Fla08, GL07a, Guz03, MS15, Mau02, MSW08]. courses [RMZG06]. Cover [KFV20]. cp [Zie19b]. cp-tools [Zie19b]. CppyABM [NZPWR22]. Cracking [GAS^{+}16]. crafting [Vai09, Vai14]. CRAPPY [CWM^{+}21]. CRC [Cla15, Hor22, Lip21, Lip22, Sau23]. Create [Ano01a, Cha01, Kno08, Lay15, LPG^{+}23, RDS07, Ull15]. create-modify-reuse [Kno08]. Creating [Fun23, MD17, CR15, KL97, MAS^{+}22, Mur18]. creation [Mül23, WAN^{+}22]. creativity [Gal14]. Creator [Chu02b]. Credible [DF21b]. Crispyn [BW22, BWS23]. criteria [BW22, BWS23, SKS23]. cross [CFW17, HPT^{+}16, Ull15]. cross-machine [CFW17]. cross-platform [HPT^{+}16, Ull15]. crossbar [JM20]. Crossing [Gue18]. crunching [Wil05]. Cryogenic [WBR^{+}22]. cryptography [Swe13]. crystal [FPSZ21, CZA^{+}23]. crystalline [Zie19b]. crystallography [TV13]. CRYSTALpytools [CZA^{+}23]. CS [Sha03]. CS1 [EPM09, GL07a, GL08c, Rad08]. CS2 [EPM09]. Cube [ÇA23]. CUDA [JS24]. Current [PBB22]. currents [JM20]. curriculum [HRS06]. customizable [Bah15]. Customized [TGEA09]. Cutting [TTHZ23, RO15a]. cutting-edge [RO15a]. CV [Len15]. Cyber [Ano21, OOM^{+}23]. Cyber-Attacks [OOM^{+}23]. Cyber-Physical [Ano21]. CycFlowDec [BS21]. cycles [BS21, Hug18]. Cython [Smi15].

d [CC20, Bra13, BRF^{+}23, DB17, GNPP23, Hug18, ML16a, NS22, NT24, Pet02, PBN^{+}09, RKVL14, Ras20, SDS00, Sch21, SAA18, SLDF^{+}21, ZL20]. D3GB [BP17]. DAEs [SL21]. DAG [IG19]. dark [May21]. DarkNews [AZH^{+}24]. Data [Ano97d, AAB22, AMB19, Bad20, CBST24, CSZ^{+}19, CZZ19, EPF^{+}23, FLS^{+}20, GPKL^{+}20, GDP18, GRKN^{+}19, Gut16, HW19, HC16, HHJCRB21, Hor22, HMB23, JCL^{+}20, KJ15, KHD^{+}16, MAFM21, McK12, MT18, MT19, ML16a, Mil18, Ngu08a, PL20, RZ09, RKR21, Sau23, SBC^{+}17, Sha24, TAD23, Van16, WL20, Wil05, ZGV^{+}23, Ada14, ASAA20, BB13, BST^{+}17, BE20, BB17, BM15, BYL^{+}21, BRF^{+}23, CGK11, CFÁA^{+}20, CR15, Cue13, DB22, DRFB^{+}23, DDK19, FKA^{+}17, FRdN21, GdGB^{+}18, GZT^{+}18, GEH19, Hac23a, HHB^{+}22, HPT^{+}16, HFF^{+}17, Idr14, Ish17, ICVG14, JAGP14, Jos16, JKKN23, LRPD18, Lay15, Lob19, Lot14b, Mad15, MB17, MFL^{+}22, MTPHH18, MMG19, Mil14, Mil15, Mit15, NMGB17, Nel15, NRC23, NO23, OLRLB21, PFLG21, Ras18, Ren23, RGP^{+}22, Ros13, Ros14, RC18, Sal18, SML06]. data [She15b, Smi17, SH19a, SLA^{+}23, Tom15, Var16, VM22, Ver22a, VCLS21, Wes15, Wet20, ASAA20, Orb18, Ish17, Jan10, Liu22]. data-driven [BE20]. data-intensive [FKA^{+}17]. Database [Ano00c, DDK19, SCAK^{+}19, SMD23, YHA^{+}16]. database-backed [YHA^{+}16]. Databases [GRG21]. dataset [RS17]. Datasets [HH17, HJPB17, Vel23]. Dateien [DF00]. Datenbanken [The09, Wei06b]. David [Jan10]. day [Cha15, GL08b, GL09]. DB [Ano98b, Kuc98c]. DB-API [Ano98b, Kuc98c]. dc [CFSK14]. dcor [RCT23]. Debian [DF00]. debt [TFAL21]. Debugging [Par11, BL97a]. December [IEE97b]. Decision [HGHR20, RBV16, SSS22, SKS23, WKS22, WKS23]. decision-making [SSS22, SKS23, WKS23]. declarative [FMPS17]. Decoding [Hig22]. decomposing [BS21]. decomposition [TEG18, DTR18]. Deep [GHXZ24, JCY^{+}19, LV20, RM19, BCM21, CJGA23, dPPD^{+}23, FLR22]. deeper [RO15a]. defect [DMC^{+}15, TQGE23]. Defects [KCVM22, AM21, BMZ^{+}18, TR22]. Deferred [Spe19]. defined [Dah18b]. defining [RS17]. Definitive [HKM08, Gar09, Lot15b, Sum08]. Demethylation [AAB22]. demonstration [SCAK^{+}19]. demos [Jos15]. Dense [Wim12]. density [Kar23, MWK^{+}20, VCF22]. density-in [Kar23]. density-in/dependent [Kar23]. Dependency [CCM^{+}23]. Dependent [KSB12, Hor23, Kar23]. Derivative [CFMR19]. Derivative-free [CFMR19]. Derivatives [AMGM20, HM18]. derived [ZAPS20]. description [Dec04]. Design [Ano01b, BBB02, Bro06, KL97, KB07, MMT09, MH18, Plo97, VEV^{+}19, VKSB15, VCR17, AGMFGE23, BSS16, BCC^{+}18, Dow09, Kas15, KXK^{+}23, Len15, LHM14, MRG18, OA17, Wei15, Gve09]. designed [Wu13]. Designing [KWBB22, Mit00, CG17, FL15, ZKB^{+}08]. desktop [Bah15, Law15]. desukutoppu [SM04]. detailed [Sar14]. Detecting [GAS^{+}16]. Detection [OOM^{+}23, PH23, SJK^{+}21, TLR21, AAYK23, BJM23, ÇA23, CJYH23, ESM24, JKST22, RMM23, SAB^{+}20]. Detectors [WBR^{+}22, ABGD^{+}20, LWH^{+}10]. determination [LN23, SL21]. determine [SGZ23]. determining [BW22, BWS23]. detrended [GHKW22]. Develop [CBST24, BSS16, DAJ^{+}15, DV21, Gup15, Jos15, Pal14, SW14, Sar14, She15b]. Developed [MOM21]. Developer [Hug01, RAH^{+}01, dos01, Ano19]. developers [Tos09]. developing [CJGA23, Har12, RCRS06, RH15]. Development [Ano00c, Ano01b, DF21a, GdGB^{+}18, Hin03, HKM08, KM15, LD07, Bah15, BL97a, CFÁA^{+}20, Con95, FBC09, Gar09, Gov15, Mau02, McG07a, OGA^{+}20b, OGA^{+}20a, Per14a, Pip15, Ree04, SUM21, You08, Lac06]. Devices [ABCC22, Por03, KPK^{+}17]. Devito [LLL^{+}20]. DEVSimPy [CS21]. DEVSimPy-mob [CS21]. DFT [VMRFC23]. DGB [Par11]. Diagnosis [SJK^{+}21, SL21]. dialect [Men09]. Diatomic [BGHC23]. Diatomic-py [BGHC23]. DIETERpy [GMKRS21]. Different [Bra13, KLM14, KLM15, Mar21, MD22b, Orl97, PBN^{+}09, Sta17, Wep15]. Differential [BT06, GWW09, LT03, LMW12, MSL^{+}07, DFSW19]. differentiation [FSMCFP23, SL21, WL13]. Diffpack [LT03, LM03]. Diffusion [CZZ19, LN23]. Diffusive [BFM18]. Digital [Rad08, Hos14, LL08, Rad06, CWLG^{+}21, FNX22, MBA^{+}22, ddSNX22]. Dimensional [Pat16, KS21, NL23, ZAPS20]. dimensionality [MSV23]. Direct [JCL^{+}20, BE20, JKST22, ML16b]. direpack [MSV23]. Discloses [AAB22]. discontinuous [CBLI22]. discover [BS19]. discrete [CS21, DPH16, JEC18, vdH18]. discrete-event [CS21]. Discretization [BH12]. Disease [JWHS16]. Diseases [SJK^{+}21]. Disengagement [GHXZ24]. Dispatch [GMKRS21]. dispel4py [FKA^{+}17]. display [Wes15]. Distance [BSSz^{+}20, RCT23]. distortion [SH19b]. distortion/interaction [SH19b]. Distributed [Ano98c, ABC97, Eur91, SZW^{+}22, BCM21, GNPP23, KI19, RCAE^{+}20, RFG13, SSH08, SCAK^{+}19]. Distributed-memory [SZW^{+}22, SSH08]. distributing [ZKB^{+}08]. Distribution [Ano00c]. Distributions [DF21b, FV18]. districts [SUM21]. dit [JEC18]. Dive [Orr05, Pil04, Pil09]. Django [EK08, Alc09, Ben08, FBC09, Gup15, HKM08, MBW07]. DJMol [PSGL21]. DMN [Fun23]. DNA [AAB22, HM22]. DNest4 [BFM18]. DNN [FLR22]. DNN-Tuner [FLR22]. Docker [ZMD21]. Document [Ano00c]. documents [KL97]. Does [She97]. Doing [Sah15, Ano15b]. DOLFIN [LW10, LWH12]. domain [HWW^{+}15]. domain-specific [HWW^{+}15]. Domains [GHXZ24]. Done [HKM08, LRvE17]. données [Swi09]. Doping [Ano21]. dot [KPK^{+}17]. double [Kar23]. double-folded [Kar23]. Downey [Gve09]. download [HM22]. Downscaling [LVH^{+}18]. dozen [Ros08]. draw [RC18]. Dreaming [Ros08]. driven [BE20, Gov15, Kin05, LKK23, LGS10, Per14a]. DropPy [OLRLB21]. DSLs [SRS^{+}23]. DSN [Ver22a]. DssPyLib [GNPP23]. dtwParallel [EAMSR^{+}23]. dummies [MM06, Mue14]. duo [Els23]. dust [Gre18]. dustmaps [Gre18]. Dutch [vdOJP^{+}20]. Dyer [Ano00b]. Dynamic [DTR18, Ada14, AES^{+}22, BC09, BII^{+}20, CEI^{+}12, DTM^{+}18, EAMSR^{+}23, FhDAF09, GLS^{+}10, LWH^{+}10, MRG18, OMGDG14, RF16, SAA18, Wu13, YHA^{+}16, YPB16]. dynamical [Mar17, Wie18]. Dynamically [Kla99, BDT13, IOC^{+}12]. dynamically-typed [IOC^{+}12]. Dynamics [AMB19, CVV23, CMM14, GP22, IHT23, JNN12, JNN13, KMK^{+}21, LHH^{+}21, LN23, PW17, SV14, Sch21, SSD^{+}22, THG^{+}23]. dynamism [BCC^{+}18].

e-book [Ano14]. E-Business [Ano01b]. eadf [SDP^{+}20]. Early [HBA^{+}20, Sev15a]. Earth [GKK^{+}23, ZRK21, ZRK22]. easily [SW15]. EAST [BYWW23]. Eastman [RE22]. Easy [CX23, RG00, SJL18, Arb14, Bea96, Cox14, Kla99, NEGZG18, PAB^{+}97, Tos09]. easy-to-comprehend [Cox14]. EASYMORE [GKK^{+}23]. eBook [Haj08]. EC [SHTE23]. EC-KitY [SHTE23]. EcoDynElec [LPG^{+}23]. Econometrics [CS09]. Economics [CLMM20]. Ecosystem [Bar21, PGH11, PM23, TFAL21]. Eddy [vdOJP^{+}20]. Eddylicious [ML18]. Edge [Wil97a, RO15a, SFC23]. EdgeSimPy [SFC23]. Edit [BSSz^{+}20]. Editing [Ano00c]. Edition [Ano00b, Orb18]. Editor [RAH^{+}01, CB96, Dub07, Gar98]. Education [Bäc07, MS07, VMFG17, AG23, DDMS14]. educational [GP22, MDR^{+}22]. EEG [CGHGRB21, VBI^{+}23]. eeglib [CGHGRB21]. effect [CFSK14]. Effective [Sla15, Cri18, Ram15, Wha18]. effects [SAA18]. efficiency [AM10]. Efficient [CSRV13, CFÁA^{+}20, DD15, GHKW22, KRH18a, KRH18b, RG20, SN12, TEG18, Wim12, BM15, DBdFdSR21, JPJ^{+}23, LYZ^{+}22, McF16, Pal14, SW14, Sie17, SK19, Zac15, WSK22]. Efficiently [PL20, EAMSR^{+}23, Lot14a]. effmass [Wha18]. effort [DLC23]. EFIT [BYWW23]. egtplot [MWS18]. Eighth [HTA^{+}97]. Einführung [EK08, LAG07]. Einsteiger [Ric14]. Einstein [Nil07a]. Einstieg [The09]. einsum [SG18]. einsum-like [SG18]. elastic [SLDF^{+}21]. electric [CFSK14]. electrical [XGW23, LYX^{+}18]. electricity [LPG^{+}23]. electro [Wie18]. electro-dynamical [Wie18]. electrochemical [VCLS21]. electrodynamics [FH22]. Electromagnetic [LFN^{+}11, BVV22]. Electron [ABGD^{+}20, GDP18, SP23]. Electronic [BLN^{+}21, GRG21, HTH^{+}20, JPJ^{+}23]. electrostatic [ABS20]. electrostatics [CFCB17, CBB14, CCFB16, KRH18a, KRH18b]. Elektronik [Ric14]. Elektronik-Projekte [Ric14]. Element [LMW12, RHM^{+}17, TT21, AM10, KRH18a, KRH18b, Kir04, LW10, LWH12, ZMS18]. Elementary [MS15]. Elements [Ano02, Ple02, CBB14, CCFB16, GTC21]. ellipsometry [RGP^{+}22]. embedded [Cas17, CGK11, LHM14, Par11]. Embedding [MOM21, Nag06, Pul00, van98b, GF18, Rie09, van95a]. Embo [PFLG21]. emgr [Him23]. Emipy [OBWM23]. emissions [OBWM23]. emphasizing [Gar09]. Empirical [ACS23, CASA22, KCVM22, Pre00, SS13, HZ23, KVSC21, LKF23, MP19a, PFLG21, Him23]. employer [Ano20]. employing [VMRFC23]. ENA [BG17]. Enabled [SST23]. Enabling [GDP18, KBC^{+}20, PHH^{+}12]. ENASearch [BG17]. encoding [SGPHD^{+}17]. encrypt [Lot14b]. End [LD07, GHT06]. ended [YMCF23]. Endogenous [GMKRS21]. Endoscopy [PH23]. ends [MP19b]. energetics [TQGE23]. Energy [BB13, LFT21, GDP18, HHM18, MWM20, MTPHH18, QZWU19, RCT23, SUM21, WZZ^{+}23, WM21, GPEM23]. Engine [Ano00c, Ang99, SCAK^{+}19, Pip15, San15]. Engineering [IEE97b, STS^{+}18, VMFG17, CR22, Kiu10, Kiu13, MP19b, SML06, Yan14]. Engineers [MA11, Nag17, O’C13, Sei09]. engines [BKC14]. English [VPO19, ZV19]. enhanced [BBC^{+}24, WO14]. enhancement [SSS22]. Ensemble [AVS20]. ensure [Buc15]. ensuring [RS17]. entangled [ECS23]. enter [BCRS15]. Enterprise [Hig03]. Entity [KDC^{+}18, ESM24]. Entropy [LD19, VBI^{+}23, SR19]. Entropy-Based [VBI^{+}23]. Entwicklung [Lac06]. Enumeration [SN12]. ENVI [BM18]. ENVI-BIL [BM18]. Environment [Ano01b, Bar21, BKMY03, BMK03, BKMY04, OGA^{+}20b, OGA^{+}20a, Pal14, VFMM08, WKS23]. environmental [LPG^{+}23]. Environments [CHM24, McG98c]. EpyNN [MDR^{+}22]. eqtools [CFW17]. equation [Bra13, GNPP23, KSS20, MD21]. Equations [BT06, GWW09, LT03, LMW12, Luk23, MSL^{+}07, TT21, CKK^{+}13, Dah18b, DFSW19, WZ18]. equilibria [CM20, CFW17]. Equilibrium [San13, BYWW23, DMC^{+}15]. Equivalent [MH18]. Era [Sev15b, BB13]. ergodicity [MWM20]. Errata [Ano01a]. error [JKKN23]. Errors [Luk23]. ErsatzPasswords [GAS^{+}16]. ESB [Suc13]. Essential [Cop08, Joh08, Bea00, Bea01, Bea06, Bea09b, Bow15, Day07b, Lot14b]. Essentials [PR02, Bah15, BM15, CV15, FL06, Lot15b, Moh15, Wes15]. Estimate [CSZ^{+}19]. estimating [WV22, YMCF23]. Estimation [Nor17, FRdN21, LWV20, ZTC^{+}21]. esy [PL20]. esy-osmfilter [PL20]. èUbungsaufgaben [The09]. EUR [Sa’23]. Europe [OBWM23]. EurOpen [Eur91]. Evaluating [Dör08, GVLD^{+}23, TAD23, AES^{+}22, WPH^{+}22]. Evaluation [ABCV21, GMKRS21, HBA^{+}20, KLM14, KLM15, RO15b, TTS^{+}10, BHJ^{+}18, CSRV13, IOC^{+}12, MD15, TR22, VKSB15, ZXX24]. evaluator [Huc18]. evapotranspiration [BTDT23, Tho22, TBD23]. even [LR14]. événements [Swi09]. Event [Kin05, TLR21, AZH^{+}24, CS21, DPH16, HQF^{+}20, vdH18]. event-chain [HQF^{+}20]. Event-driven [Kin05]. events [LRvE17, WRBT21]. everyday [Wil05]. evidence [LFT09]. Evolution [TFAL21, DFSW19, HZ23, Yes15]. Evolutionary [MWS18, WBK^{+}24, Wie18, SHTE23]. evolving [BJM23]. exact [Bri06, TF23]. ExactPack [TF23]. Examining [Ang99, CK00]. example [Jos15, RH15]. Examples [Lip22, Piñ24]. exascale [GVLD^{+}23]. excited [LHH^{+}21]. exciting [BCRS15, Gup15]. Execution [Fun23, JCY^{+}19, AFL23, BB22a, BKC14, DM20, RCAE^{+}20, Spi18]. exemple [MRA06]. Exercise [KWBB22]. exercises [TTHZ23]. existing [Lot15b]. ExoData [Var16]. exoplanet [Var16]. expanded [RSPJ21]. Expanding [LTT^{+}24]. Expansion [HJJ^{+}22]. expansions [Ano18]. Experience [CM07, Pop10]. Experiences [SBC^{+}17, ZGV^{+}23]. Experiment [WX18, BST^{+}17]. experimental [CWM^{+}21, KSH14]. Experiments [AMB19, MD17, MAFM21, CG17, dPPD^{+}23]. Expert [Sma12, ZKB^{+}08]. Explain [VMFG17]. explainable [DSB^{+}23, PDS^{+}22]. explaining [LFT09]. Explicit [HPH12]. Explorative [HLR15]. Exploratory [HGHR20]. explore [Dan18a, Jos16, Mad15, Sah15]. Explorer.py [WM21]. exploring [BMK03, Sev13]. Exporting [Ver22b]. Exposure [JWHS16]. express [Har07]. expression [CKC23, Stu03, Stu07]. expressions [LR14, Rom14, SG18, SM04, Stu07]. extendable [KVSC21]. Extended [CBW^{+}21]. Extending [Dar12, DY96a, DY96b, DY99, NM22, van95a, van98b, CEI^{+}12, PHH^{+}12]. extensibility [Kla99]. Extensible [BL97a, BCM21, BJM23, CFW17, HCPF95]. Extension [Ngu08c, BSG^{+}16, WBS21]. Extensions [Ano98c, Bea98, Dal01, DPSD08, Ras18]. extensive [YMA23]. Exterior [BH12]. Extract [PL20]. Extracting [HGMC^{+}97]. extraction [CGHGRB21]. Extractive [VPO19]. extremely [CKC23]. extrinsic [DMC^{+}15]. Eyringpy [DCOC^{+}19]. EZFF [KMK^{+}21].

F [Ano00b]. Faceted [Kra22]. facilitating [Spi18]. factor [SZNW23]. factorisation [LB22]. Factorization [CJ22]. factorizations [Roa23]. FAD.js [BB17]. fake [RMM23]. Fall [MH18]. Fan [RDBC23]. Fan-Slicer [RDBC23]. Fanpy [KRSD^{+}23]. FAO [BTDT23, Tho22, TBD23]. FAO-56 [BTDT23, Tho22, TBD23]. FAQ [Ano98e]. far [SDP^{+}20]. far-field [SDP^{+}20]. Faraway [Lip21]. Fast [BKH^{+}22, LV20, Mül13, RDBC23, SMF^{+}23, BB17, SSD^{+}22, Yes15, MBA19]. fastcluster [Mül13]. Faster [SN12]. fastmat [WSK22].

FastSPM [BRF^{+}23]. father [LS97]. fault [WPH^{+}22]. faults [MLB22, WPH^{+}22]. FDT [RMM23]. FDTD [CKK^{+}13]. feature [CGHGRB21]. features [KMLP^{+}23, LTT^{+}24, LR14, Pip15]. Featuring [SJL18]. Federated [DSB^{+}23]. Feedback [BE20]. FEM [GNPP23]. FEniCS [LMW12, TT21]. fenicsR13 [TT21]. fgivenx [Han18]. FIAT [Kir04]. Field [Pat16, Cri18, MFQ^{+}21, OC20, SDP^{+}20, Wie18, WMA^{+}22]. fields [CMS22, CFSK14, Mur18]. Fifth [CI96]. file [VAP^{+}21]. files [KI19, Ver22a, VAP^{+}21, VC18]. filled [CBB14]. Filter [BHF16]. filtering [CV22]. Finally [Aya14]. Finance [Vir16, Di 13, Hil15b, Wei15, Wep15, Yan14]. financial [Yan14]. find [MRT^{+}22]. fine [YHX22]. fine-grained [YHX22]. Finite [Ano02, LMW12, Ple02, RHM^{+}17, TT21, AM10, KRH18a, KRH18b, Kir04, LW10, LWH12, ZMS18]. FiPy [GWW09]. Firedrake [RHM^{+}17]. First [CCM^{+}23, AM21, GL08b, GL09, HRS06, MSW08, Rad06, SL08, Sha03, TSD^{+}12]. first-class [TSD^{+}12]. first-principles [AM21]. fit [GH18]. fitting [ENCS20]. fix [YHX22]. FL [Hor22, Lip21, Lip22, Sau23]. Flask [Gup15]. Flexibility [CFMR19, Gar09, PHH^{+}12]. Flexible [KB07, BL97b, CSRV13, SAA18, YMCF23]. floating [BII^{+}20, JT14]. floating-point [BII^{+}20, JT14]. Flow [DB17, HWJ^{+}21, MOM21, RDB21, TGEA09, BS21, Par22, YHA^{+}16]. flows [ML16b]. Flowtracks [ML16a]. fluctuation [GHKW22]. Fluent [Ram15]. Fluid [AMB19, CM20, GP22]. FluidDyn [AMB19]. FluidFFT [MBA19]. FluidSim [MBLA19]. Fluorescence [RKR21]. fly [Cas17]. fmas [MD22a]. fMRI [WL20]. foci [CMS22]. Fock [AYI^{+}24]. focused [CMS22]. focusing [DLC23]. folded [Kar23]. Follow [STS^{+}18]. fondamentales [Chu07b]. Force [OPA^{+}14]. forcefields [KMK^{+}21]. Forecasting [HKGvS21]. foreign [LWH^{+}10]. forensic [Hos14, O’C13]. Forensics [Hos14, Ano14]. Forests [PRH17]. Form [ZTT22, GS12]. formalism [SP23]. format [Fas18]. formats [VAP^{+}21]. Formula [AJJF14]. Formulation [Ble20]. Fortran [AJJF14, DY99, GH18, GHN19, LC11, MSP^{+}22, MMEH08]. Fortran-77 [GH18, GHN19]. forward [MD22a]. Fotos [DF00]. Foundation [Ngu08c]. Foundational [Osa17]. Foundations [RG10, RG14, You08, Hor22, Sau23]. Fourier [MBA19, BKH^{+}22]. FPGA [HWJ^{+}21]. FPGAs [AJYH18]. fracture [ZL20]. fragments [OA17]. Frames [HHJCRB21]. Framework [ABCC22, AMB19, FLS^{+}20, HA20, Him23, JCL^{+}20, JPOB20, LV20, MMT09, Nor17, Ram18a, RBP^{+}21, Zuk97, AES^{+}22, BCM21, BSG^{+}16, BB13, BK19, CM20, DMC^{+}15, FMPS17, FKA^{+}17, FPN^{+}22, GMKRS21, Gar09, GQCP^{+}18, IHT23, JNN12, JNN13, JKKN23, KVSC21, KSH14, Kuc98b, Law15, LC11, LGS10, LYX^{+}18, Mar17, NT24, OMGDG14, PGM19, RGS^{+}21, SV14, SUM21, SAA18, SL20, Tay18, VCLS21, WAK22, ZMS18, dPPD^{+}23]. frameworks [Dal07, Gup15, MBW07, RCRS06]. Francis [Lip21, Lip22]. Francisco [ACM92]. Franklin [Jan10]. Free [Ano00c, Tro96c, Tro96a, Tro96b, Tro97, CFMR19, HFF^{+}17, KBC^{+}20, OC20, QZWU19, WZZ^{+}23, ZRK21, ZRK22]. Freeway [SYGY21]. French [Chu07b, Chu07a]. friendly [BCM21, DDK19, GPP^{+}21, LYZ^{+}22]. front [MP19b]. front-ends [MP19b]. Frontier [Gue18]. FSDP [ZGV^{+}23]. full [MGFO23, Pip15, PMM^{+}13, Wie18, WMA^{+}22]. full-field [Wie18, WMA^{+}22]. full-waveform [MGFO23]. Fully [ZGV^{+}23, AG23]. Function [AMPH14, Kor11]. Functional [ACM92, Lot15a, MB07, WBK^{+}24, BBH^{+}15, Cla15, Han18, Her14, MWK^{+}20, MCF^{+}22, VCF22]. Functional-style [WBK^{+}24]. functionalities [WMA^{+}22]. functionality [BBC^{+}24]. functionals [BT21]. functions [BT21, CSRV13, Kir04, May17, MR18, SZW^{+}22]. Fundamental [Chu07b]. fundamentals [Chu08, Lee11]. Funktionen [Wei06b]. Further [LFT09, She15b]. fuzzy [CKC23, KXK^{+}23, WKS23]. fvextra [Poo16].

Gala [PW17]. galactic [PW17]. Galerkin [CBLI22]. Game [Gut04, Har07, McG07a, RH15]. Games [KM15, MWS18, RH15, Swe09, Ull15]. Garbage [GMN21]. Gas [Mor17, OOM^{+}23, DCOC^{+}19, MFQ^{+}21, WZZ^{+}23]. gaseous [ABGD^{+}20]. gate [JLP23]. gate-based [JLP23]. Gaussian [SBMD14]. Ge [Wei06b]. Ge-Packt [Wei06b]. Gem [Bri12a, Bri12b, GF18]. gem5 [RSRT19]. Gemma [Cas17]. gene [CKC23]. General [AJYH18, LS16, NRC23, AZLY18, CFPS23, DCRF23, Myr15, OGA^{+}20b, OGA^{+}20a, DBdFdSR21]. General-Purpose [AJYH18, LS16, DCRF23]. generalised [LB22]. generalized [MD21, MD22b, SZW^{+}22]. Generate [Bie97, GTC21, TTHZ23, ZAPS20]. Generating [May21, MFQ^{+}21, VAP^{+}21]. Generation [Ano01b, HAB^{+}20, JPOB20, AM10, Dal07, DRFB^{+}23, FPSZ21, HOB23, LKF23, ML18, OE21, Ras20, RJAL09]. Generator [Bri02, AZH^{+}24, HR20, MO14, NCS17, WZ18, vd91, Hag17]. genesis [TV13]. Geneva [IEE97a]. Genome [BP17]. genomic [GQCP^{+}18]. gentle [MS13]. Geodesy [Ngu08c]. geometric [PBN^{+}09]. geometries [WAN^{+}22]. Geometry [LDCM23, TEG18]. Geomstats [LDCM23]. Georges [THG^{+}23]. Georgia [IEE20, USE00]. Geospatial [KHD^{+}16, Bah15, Law15, Tom15, Wes15, Tom15].

Geostatistical [SMF^{+}23]. Germ [AAB22]. German [Lac06]. gestalten [DF00]. Get [Wil02, CR15, Cox14, SW15]. Getting [BCRS15, Hos12, MAFM21, McG07b, Nai14, Nix15, Ric13]. GIFT [SGPHD^{+}17]. GIFT-Grab [SGPHD^{+}17]. GIMP [DF00, DF00]. GIMP- [DF00]. Ginzburg [Hor23]. GIS [Bah15, CC20, Law15, Shi07]. Glade [Ano01a, Cha01, Ree04]. glass [WMM18]. glasses [BWMS22]. Glider [GRKN^{+}19]. GliderTools [GRKN^{+}19]. Global [LVH^{+}18, SGZ23]. Globus [Jac02]. GLR [Dej22]. GlusterFS [Dar12]. GMES [CKK^{+}13]. GMPY [Hor15]. gmx2qmmm [GPP^{+}21]. GNLStools.py [MD22b]. GNU [DF00]. Go [Ano98d, Ram18b, Dan18b, VCR17]. Golang [BB22a]. goniometry [OLRLB21]. Good [Bea09a, Pre03, HB16]. goodness [GH18]. Google [Pip15, San15]. GPR [CKC23]. GPU [BKH^{+}22, CFCB17, CWLG^{+}21, ECS23, JS24, MSP^{+}22, NT24, TMH^{+}23]. GPU-accelerated [ECS23, NT24]. GPU-acceleration [JS24]. GPUCorrel [CWLG^{+}21]. GPUs [AJYH18, CBB14, CCFB16, DLC23, VY15]. Grab [SGPHD^{+}17]. Gradual [TSD^{+}12, Sie17, VKSB15, VSS17]. Grafik [Lin06b]. Grafik-Modul [Lin06b]. Grafiken [DF00]. grained [YHX22]. GRamian [Him23]. Grammar [MP19b]. grand [PHH^{+}12]. grand-canonical [PHH^{+}12]. grants [Bal12]. Graph [BSSz^{+}20, Hag17, JCY^{+}19, LS16, Boe17, GF18, KPKS22]. Graph-Mining [LS16]. graph-theoretic [Boe17]. GraphBLAS [PKDM21]. Graphical [SJL18, Kar14, MTS^{+}18]. Graphics [SDS00, GL08b, GL09, PAB^{+}97, SML06]. graphs [BJM23]. grasp [Lot14a]. Gravitational [CFPS23, MAC^{+}21, Wet20]. gravitational-wave [MAC^{+}21, Wet20]. gravity [MTS^{+}18]. Gray [Ish17, Sei09].

Great [RAH^{+}01]. green [ADP21]. Grids [MSS^{+}19]. groff [RAH^{+}01]. GROMACS [PHH^{+}12, GPP^{+}21]. GromPy [PHH^{+}12]. große [Haj08]. Grounding [MSS^{+}19]. Group [Lip21, Lip22, SST23, WV22, FSMCFP23, ZC20]. Grundlagen [DF00]. GSAS [TV13]. GSAS-II [TV13]. GSimPy [ZC20]. Gsolve [MTS^{+}18]. GTK [DF00]. GTK-Versionen [DF00]. Guest [Dub07]. GUI [Wei06a, Bri02, Con95, GCBDDBF23, Har12, LS98, Sum08, Tal00]. GUI-Programmierung [Wei06a]. Guide [GRG21, HAB^{+}20, HKM08, Lee17, Mit00, Sha24, BS19, Feh02, Gar09, Ish17, Ish19, ICVG14, KN15, Lob19, Lot15b, Meh15, Mil14, RO15a, RG10, RH15, Sum08, Swe13, Tel06]. Guided [LP19, FhDAF09, MCF^{+}22, Ott18]. Guido [Chu02b, Dou98, Kuc98a, Sev15a, Sev15b]. GUIs [EK08]. Gumbie [Bri02]. gut [SKS08]. Guttag [Orb18]. GWpy [MAC^{+}21].

H [AYI^{+}24, HA20]. H-Revolve [HA20]. H-wave [AYI^{+}24]. Hack [Ott18]. hackers [O’C13, Sei09, Sei14]. Hacking [Swe13]. Haddad [Ano00b]. half [TMH^{+}23]. Hall [Lip21, Sau23]. Hall/CRC [Lip21, Sau23]. Halloween [Cas17]. Hammond [Ano00a]. hand [Buc15]. Handbook [Piñ24, Van16, dos01, Hug01, Lip22]. Handbuch [EK08, EK09]. Handle [GZT^{+}18, CBB14, Var16, VC18]. Handling [RDB21, WKS23]. Handook [RAH^{+}01]. Hands [Mil18, Bro06, Cha15, Ros14]. Hands-On [Mil18, Bro06, Cha15, Ros14]. HappyJIT [HS12]. Hardback [Lip21, Liu22]. Hardware [GMN21, JPOB20, Dec04, Tab10]. harmonic [GEH19]. harness [Lay15]. Harnessing [VEV^{+}19]. Hartree [AYI^{+}24]. Hash [AMPH14]. Haskell [CM07, Pop10]. Hawaii [HK95]. hbk [Lip22]. hcb [Sau23]. HEART [LKJC21, CRL21]. Heat [VEV^{+}19, SAA18, Zie19b]. Heatrapy [SAA18]. heavy [AZH^{+}24]. Hector [ABC97, WHG17]. Height [MH18]. Held [HTA^{+}97]. HELM [MOM21]. HELMpy [MOM21]. Help [Dör08, HR00, HS13, Law15, Len15]. Herman [Cla15]. Heterogeneous [BKRT21, HM18, SBC^{+}17]. HexagDLy [SH19a]. hexagonally [SH19a]. HHVM [Ott18]. Hi [CSZ^{+}19]. Hi-C [CSZ^{+}19]. Hierarchical [HA20, Mül13, DB22]. HIFIR [CJ22]. High [AMGM20, BKRT21, BUS21, CBST24, Dog15, EPF^{+}23, GPKL^{+}20, GO14, Hin02, Hin03, IEE20, KWBB22, Lew09, LD07, MBLA19, ML16b, NT24, OLRLB21, SRS^{+}23, BCRS15, BSS16, BS19, CZ22, CFÁA^{+}20, GQCP^{+}18, GVLD^{+}23, HF06, Lan13, Mar18, Par22, Ros13, Ros14, SNB^{+}19, BB13, SSH08]. High-Level [Hin02, Hin03, GVLD^{+}23]. High-order [AMGM20]. High-Performance [BKRT21, CBST24, KWBB22, MBLA19, SRS^{+}23, BCRS15, BS19, CZ22, HF06, Mar18, Par22, Ros13, Ros14, SNB^{+}19, SSH08]. High-Precision [GPKL^{+}20]. High-Productivity [BUS21]. High-throughput [OLRLB21, GQCP^{+}18]. higher [MO14]. highlighter [BC09]. highly [AAYK23]. Hill [Lak17]. Hindi [VPO19]. histograms [GH18, GHN19]. histolab [MBA^{+}22]. Historic [TS14]. historical [LPG^{+}23]. Hits [Sco17]. hmis [MB17]. Holistic [VFMM08]. Holographic [BDF^{+}20]. Holomorphic [MOM21]. HoloPy [BDF^{+}20]. Home [ABCC22, Ste06]. Homogeneity [HMB23]. Homogeneous [SBC^{+}17]. homotopy [VY15]. Hong [IEE97b]. Hosmer [Ano14]. hot [ÇA23, OMGDG14]. hot-reprogramming [OMGDG14]. Houches [Sta19]. Hour [Ngu08b]. HPC [MBA19, SLA^{+}23]. HSImage [BM18]. HTML [Bie97, KL97]. HTML-supporting [KL97]. HTML5 [Ada14]. HTML5-ready [Ada14]. HTMLgen [Ham98]. Hub [ZMD21]. huge [Cox14]. Hughes [Ano00b]. humanleague [Smi18]. HW [Wu13]. HW/SW [Wu13]. Hybrid [CJ22, LP19]. hydraulic [ZL20]. Hydrodynamics [RBP^{+}21]. hyogen [SM04]. Hyperbolic [LeV09]. hyperbox [KG23, KG23]. hyperbox-based [KG23]. hyperbox-brain [KG23]. hyperfine [BGHC23]. hyperparameters [FLR22]. hyperspectral [BM18]. HyPhy [Spi18]. HYSPLIT [War18].

i-PI [CMM14]. IAST [SSH16]. Ibrahim [Ano00b]. iCorrVision [FNX22, ddSNX22]. iCorrVision-2D [ddSNX22]. iCorrVision-3D [FNX22]. IDE [Ano01b, SW14]. Ideal [SSH16, GP22]. ideas [Chu07b, Chu07a]. identification [CR22, LRPD18]. identifiers [SGZ23]. IEEE [CLT20]. II [Ang00c, CFPS23, TV13]. Iisignature [RG20]. Ill [CJ22]. Ill-Conditioned [CJ22]. illustrations [Lip21, Lip22]. Image [Ble20, CWLG^{+}21, FNX22, SLGB14, ddSNX22, GRW^{+}19, LRPD18, MAS^{+}22, RMM23, Myr15]. Images [ISMA18, OPA^{+}14, PH23, BM18, TEG18, ZMD21]. Imaging [MB07, SBC^{+}17, CG17]. immediately [Cha15]. immunoassays [VCLS21]. impact [LJ19]. impacts [LPG^{+}23]. Impedance [LYX^{+}18]. Imperative [JCY^{+}19, MCF^{+}22]. imple [VMRFC23]. Implement [Mil18, Wei15]. Implementation [BBB02, DD15, HAC^{+}23b, JWHS16, Ngu08a, AWO^{+}21, ABS20, BISC14, CKC23, DEMM19, KL97, KMLP^{+}23, LB22, MR22, OA17, Sco17, Tau16, ZXX24]. Implementations [RO15b]. Implemented [BHF16, Bra13, Kar14]. Implementing [Bri06, EVV12, Nil07b, Zuk97, MD22b, RNR17]. improve [ZAPS20]. Improved [BBC^{+}24, Sco17]. improvements [DPSD08, WMA^{+}22]. Improving [Ago01, CFMR19]. impulse [FRdN21]. impulseest [FRdN21]. In-Code [VMFG17]. In-Cognitive [KXK^{+}23]. In-database [SCAK^{+}19, DDK19]. in-plane [FNX22, ddSNX22]. In-situ [MFL^{+}22]. in/dependent [Kar23]. Included [Dub07]. includes [Gar09, SML06]. Including [MOM21]. Incomplete [CJ22]. independent [PFH^{+}16, She97]. Index [HKGvS21, SL21]. Indicator [BLE21]. indices [BDQ^{+}22]. indispensable [Buc15]. individual [NS22]. industrial [OBWM23]. Industry [Joh15, Mat16]. Infected [PH23]. Infer [VBB18]. Inference [CHM24, SN12, SBMD14, DB22, SAB^{+}20, SMM^{+}22, ZL23, AWO^{+}21]. inferential [Hac23a]. inflationary [MR18]. inflow [ML18]. informatics [Sev13]. Information [LDCM23, PFLG21, BII^{+}20, Gra18, HGMC^{+}97, JEC18, Sev13, YHA^{+}16]. Informix [Ano00c]. infrastructure [CZA^{+}23, DDT20]. Infrastructure-as-Code [DDT20]. Inhibiting [GAS^{+}16]. InitDAE [SL21]. Initiation [Bai08]. initio [CMM14, JPJ^{+}23, KRSD^{+}23, PP23]. Injecting [MLB22]. injection [WAK22]. inkl [The09]. Input [Sta21, FRdN21, MD22b, VAP^{+}21]. Input-Output [Sta21]. inputs [CSRV13]. ins [DF00, Est01]. Insert [HW19]. insertions [WBS21]. insightful [Lay15]. insights [RO15a]. inspired [VBM^{+}18]. Instabilities [CYKG23]. Instant [WMA12]. Instantiation [Ano01b]. Instrument [BCE^{+}22]. Instrumentino [KSH14]. instruments [KSH14]. insulators [BMZ^{+}18]. Integral [RG20, ŚBA^{+}15, CMM14, CFCB17]. integrals [BHJ^{+}18]. integrate [CV15, NZPWR22]. Integrated [GMN21, FNX22, LFT21, ZWCQ22, ddSNX22]. integrating [Bea96, TSC02]. Integration [HHJCRB21, Bar22, Dah18c, Els23, LJ23, SHTE23]. integrations [Suc13]. integrators [MO14]. integrity [RS17]. Intelligence [Sa’23, Tan06, DSB^{+}23, Lot14b, Seg07]. Intelligent [RM19, ZGL20]. intensive [FKA^{+}17]. Inter [AAC^{+}98]. Inter-Language [AAC^{+}98]. interacting [BG17]. interaction [BM18, HHVB21, SH19b]. interactions [Kar23, MMEH08]. Interactive [BP17, PG07, RAH^{+}01, SDS00, Dan18a, DDK19, GP22, Len15, PAB^{+}97, RE22, Ros18, Ros13, Ros14, Ull15, Ano00c]. interatomic [DBdFdSR21, KMK^{+}21]. interdisciplinary [SWD15]. Interface [ACM97, DGMP09, LD19, SJL18, BLN^{+}21, CMM14, DOS17, DMD^{+}21, GPP^{+}21, Gre18, GTC21, HJPB17, HHP^{+}16, Jac02, LPH23, MTS^{+}18, Pit18, PHH^{+}12, PWFM17, RK19, Spo12, VMRFC23, WZ18, WHG17, vdOJP^{+}20]. Interfaces [Ano01a, Cha01, LWH^{+}10, Smi17, Smi18, Wet20]. interfacial [CM20, SHFJ18]. interferometeric [HJPB17]. interferometers [BJR^{+}20]. intermetallic [DMC^{+}15]. International [Ano97c, CI96, IEE97b, IEE97a, IEE20]. Internet [DF00, Joh97, The09, Wei06b, Ano97b, DAJ^{+}15, McG98b, Shi07, SST23, WvA96]. Internet-Programmierung [Wei06b]. Interpolation [BKH^{+}22, CZ22]. interpretation [ZXZX22]. interpreted [BKC14]. Interpreter [ZXX23, van98b, CM07, PSGL21, van95a]. interpreters [HWW^{+}15, LFYX23, MD15, ZLBF14]. interstellar [Gre18]. interval [SABEh20]. interval-valued [SABEh20]. Intervals [DF21b, PRH17]. Interviews [Kuc98a]. intricacies [Lot14a]. intrinsic [DMC^{+}15]. Introducing [BC09, GL07a, HRS06, Lub14]. Introduction [Bau96, Dub07, Ful96, Gut13, Gut16, Har12, Has16, Ngu08b, SWD15, SMM^{+}22, Tan06, ÜK12, War18, vRD05b, Can14, Joh12, MS13, Poo16, Rad06, Sto13, Sum09, Sum10, vRD03, Orb18]. Introductory [KLM14, KLM15, MMP^{+}22, WX18, LFT09, MBK09, RMZG06]. Intuitionistic [WKS23]. Intuitive [Hug14, KRH18a, KRH18b]. invariant [LGS10]. invariant-driven [LGS10]. inventing [Hos14]. Inverse [LD19]. inversion [MGFO23, NL23]. Investigates [Ano95]. Investigation [SS13]. Investment [GMKRS21]. invited [Sie17]. involving [SAA18]. IO [Buc14]. IOData [VAP^{+}21]. ion [BFT20]. ionization [SZNW23]. ionosphere [ZRK21, ZRK22]. IoT [ABCC22, TAD23, ZGL20]. IPython [Ber13, PG07, Ros13, Ros14]. irt [LR23]. ISBN [Ano13, Ano14, Ano15a, Ano15b, Cla15, Hor22, Lee17, Lip21, Lip22, Liu22, Sa’23]. Ising [LD19, LD19]. Ising-Type [LD19]. isoconversional [RHLTG^{+}22]. isosurface [MZL^{+}23]. isothermal [RHLTG^{+}22]. isotropic [Mur18]. issues [GMP^{+}20]. Item [LR23]. Iterated [RG20]. Iterated-Integral [RG20]. Iterative [CJ22, BHA18]. iterators [ZLBF14]. Ivezić [Ish17]. Izzet [Ano01a].

J [Ano00b, Ano13, Ish17]. J. [Lip21]. Jacob [Ish17]. Jacobian [NCS17]. Jacquez [JWHS16]. January [HK95]. Janus [KFV20]. Java [SML06, SKS08, AAYK23, Ang99, Ano00c, Ano01b, Ano20, ADP21, BLS^{+}23, Dan18a, HF06, Hig03, Hug97, MSR03, Pre03, Rem01b, Roo97, SM04, Stu07, Tro97, Wil05, Zho97]. Java-Compliant [Ano01b]. JavaScript [Ric14, CASA22, Dan18a, Gar09, ZMD21]. JCanvas [Ano01b]. JDBC [Tan01]. jedem [Wei06a]. JeLLyFysh [HQF^{+}20]. Jesse [Lee17]. Jinn [LWH^{+}10]. JIT [ABF13, BBH^{+}15, BB17, CEI^{+}12, HS12, Ott18, PFH^{+}16, ZXX24]. JIT-based [BB17, PFH^{+}16]. John [Jan10, Orb18]. Jose [Ano97c, Liu22]. JPython [Ang99]. JSON [BB17]. Jug [Coe17]. Julia [BCRS15, BSS16, Bal18, BS19, BCC^{+}18, CLMM20, Dan18a, DD15, GMP^{+}20, GVLD^{+}23, Jos16, KRH18a, KRH18b, Lob19, MSP^{+}22, MT18, MT19, MSS^{+}19, Nag17, PKDM21, Roh16, Sal18, She15b, SNNB17, Ver22b, ZV19]. Julian [HKGvS21, Lip21]. Julian-Oscillation [HKGvS21]. July [Ano01a, Ass96]. Jump [MH18]. Junction [SN12]. June [ACM92, CLT20, USE96]. Jupyter [Bar21, Dan18a, Moo21]. JupyterLab [Dan18a]. Just [NM22, RF16, Wil97a, LJ19, May21, TTS^{+}10, WMA12]. Just-in-time [NM22, RF16, LJ19, TTS^{+}10, WMA12]. Justin [Ano15a]. Jython [PR02, Bri02, Hig03].

Kapitel [Wei06a]. Kateri [Hor22, Sau23]. KDE [Ano00c]. Keeping [Chu02a]. Keith [Lip22]. Kernel [Ago01, Ano01a]. kernels [SZW^{+}22]. key [BM15, RS17]. key-based [RS17]. keynote [McK16]. Kids [Lin06b]. Kinder [Lee17]. Kinect [TCE21]. kinetic [LWV20]. KineticPy [MWM20]. kinetics [MWM20, NCS17, RHLTG^{+}22]. KinZ [TCE21]. kisses [Bri12a, Bri12b]. kit [SHTE23]. KitY [SHTE23]. Kivy [Ull15]. Klassen [Wei06a, Wei06b]. kMap.py [BYL^{+}21]. Know [MAFM21]. Knowbot [HMD^{+}96]. Knowledge [CHM24]. Knowledge-Based [CHM24]. Kokkos [GVLD^{+}23]. Kong [IEE97b]. Korner [Ago01]. Kubo [SP23]. kurz [SKS08].

L [Har07]. lab [Rad06, AG23]. lab-based [Rad06]. labeled [HH17]. labeling [DRFB^{+}23]. laboratories [MCGK19]. laboratory [AG23]. Lady [Bri12a, Bri12b]. lagrangians [May21, May21]. LALSuite [Wet20]. LAMMPS [DCRF23]. Land [KFV20]. Landau [Hor23]. Landscape [CBST24]. landscapes [MWM20, WM21]. Lane [SYGY21]. Langages [Bla02]. Language [AAC^{+}98, Ano95, BKL09, CL06, Ful96, HC16, Het10, Het14, LD07, Lut98a, PFH^{+}16, RM19, Rie09, Roo97, Sco09, SS13, Tro96c, ZV19, van98a, vRD05a, AJJF14, BCRS15, BS19, BBH^{+}15, Bor07, CLM05, Can14, CEI^{+}12, CGK11, Dec04, DY96a, Fee16, GMP^{+}20, HWW^{+}15, IOC^{+}12, KRH18a, KRH18b, LS97, LWH^{+}10, LGS10, MR07, Nel15, Ngu08d, Och09, Orl97, Per14b, SL08, SNB^{+}19, She97, Sum09, Sum10, vd91]. Language-independent [PFH^{+}16]. Languages [Ano98d, CG23, Dör08, ESM98, Mar21, Pre00, Pre03, PBB22, Bea96, BL97b, BDT13, BKC14, Cas20, CEI^{+}12, DTM^{+}18, FhDAF09, GLS^{+}10, HPH12, LS97, Mau02, MRG18, Ngu08e, OMGDG14, RF16, Wu13, YPB16]. Large [vdOJP^{+}20, BL97b, BCG05, Var16]. Large-Eddy [vdOJP^{+}20]. large-scale [BL97b]. laser [EBNS22]. Lasso [HHJCRB21]. LAST [Sco17]. Latency [PBK22]. latest [Cas20]. Lattice [DBdFdSR21, Mor17]. Lattice-Gas [Mor17]. lattice mc [Mor17]. Laughs [Wil97a]. LAVA [DCRF23]. Layer [AAB22]. layers [CBB14]. Lazily [Har15]. LazySorted [Har15]. Learn [Cha15, Don14, Ros13, CV15, HS13, Idr14, Mad15, Nai14, Sal18, Swe09, ZKB^{+}08]. Learned [Wit21]. Learning [Ada14, AVS20, Arb14, BW14, CBST24, Dan18a, Duf15, GHXZ24, Gun14, Hia15, Hil15a, Hil20, Hug99, JCY^{+}19, LV20, Lay15, Len15, LA99, LA04, Lut08, Lut09a, Lut09b, Lut13, MH15, OOM^{+}23, RM19, Ros13, SW15, TAD23, BSS16, BCM21, Bow15, CR15, DSB^{+}23, GMP^{+}20, Har07, Ish17, ICVG14, Kar14, KG23, PDS^{+}22, RS17, RO15a, Ras18, Sal18, SCAK^{+}19, SHTE23, WRBT21, ZL23, dPPD^{+}23, Roa23, Lak17, Laz22]. learning-assisted [WRBT21]. LEED [GDP18]. LEEM [GDP18]. Lefever [MD21]. LegacyJ [Ano01b]. Legate [BLP^{+}21]. LEGOs [BMK03]. Leitfaden [Haj08]. lepton [AZH^{+}24]. leptons [SLP^{+}22]. Lessons [CG23]. lets [Cas17]. Letter [Gar98]. Letters [RAH^{+}01]. Level [Hin02, Hin03, GVLD^{+}23, LFT21, NT24, Myr15]. leverage [BSS16, LR14, PM23, SW14]. LHC [BB13]. Libraries [MLGW18, MBA19, PBB22, Bah15, Hig03, She97, Wes15, Wet20]. Library [AMGM20, BTR14, BKH^{+}22, CZZ19, CHG^{+}16, CR12, HGHR20, HC16, HAGH18, ISMA18, LR23, LS16, Lun01, Moo21, PL20, RG20, TLR21, ZTT22, AM19, Ada14, BW22, BWS23, BG17, BJM23, BB22b, BDQ^{+}22, BM18, Cri18, DM20, EPF^{+}23, FPSZ21, Gre16, HJHZ18, HTH^{+}20, HU17, HOB23, HBA^{+}20, HB16, JAGP14, JLP23, JT14, KL97, KRSD^{+}23, KMK^{+}21, LRPD18, LB22, LRvE17, LWH12, MBA^{+}22, May17, MRT^{+}22, MMP11, MK16, NZPWR22, DDT20, PBN^{+}09, Ras20, RV20, RJAL09, RFG13, RC18, SKS23, SLDF^{+}21, SH17, THG^{+}23, TOPC23, TF23, Ull15, VAP^{+}21, WAN^{+}22, WKS22, Yes15, YMCF23, YMA23, ZMS18, Zie19b, van95b, BLE21]. Life [Has16, Yan14]. light [MR22]. lightweight [Bah15, CBLI22, Gar09, RGS^{+}21, RFG13]. Like [Dow15, Dow09, SG18]. Line [CC20, IG19, Har07, Har07]. Linear [Luk23, PBB22, SBMD14, TT21, AGMFGE23, BCG05, CFSK14, Dah18b, FPN^{+}22, LKJC21, RV20, SLB^{+}23, WSK22, Lip21]. linear-operator [RV20]. Lines [HAB^{+}20]. linguistics [Joh08]. Linking [vd91]. Linux [Ano00b, DF00, Ric14, USE00, Ago01, Ano00b, Ano00c, Bla02, Bla04, GJ08, Lac06, MS^{+}00, Par11, RAH^{+}01, TSC02]. Linux-basierte [Ric14]. LIon [BFT20]. lipid [WRBT21]. liquids [VCF22]. Lisp [ACM92, Mac91, Mac92a, Mac92b]. List [Har15]. lists [Ano20]. literature [HSD^{+}22]. litstudy [HSD^{+}22]. little [VSS17]. LJ [Kuc98a, RAH^{+}01]. LLNL [Ano98c]. LLNL-Distributed [Ano98c]. LLVM [RSRT19]. LLVM-Based [RSRT19]. Load [MOM21]. Local [LFT21]. locality [BII^{+}20]. localization [IHT23, WPH^{+}22]. Log [RG20]. Logging [Kra22]. logic [Bro06, CKC23]. long [MWM20]. long-time [MWM20]. Look [CCM^{+}23]. Loop [ABF13, LS17, RCAE^{+}20, Ver22a, XGW23]. Loop-aware [ABF13]. Loosa [RAH^{+}01]. Lorentz [FH22]. loss [SZW^{+}22]. Love [Wit21]. Low [Fun23, GDP18, ZAPS20]. Low-Code [Fun23]. low-dimensional [ZAPS20]. Low-energy [GDP18]. lower [ZRK21, ZRK22]. LR [Dej22]. LR/GLR [Dej22]. LSKUM [MSP^{+}22]. LSPR [CFCB17]. Lua [Gut04]. Lugiato [MD21]. luminescence [TR22]. Lunch [Ang00a]. Lutz [Ano97a].

M [Jan10, Lee17, Wep15]. Mac [SML06, Wal07]. Machine [Bow15, CBST24, OOM^{+}23, TAD23, CFW17, CR15, GMP^{+}20, Ish17, ICVG14, Kar14, KG23, MRG18, PDS^{+}22, RS17, RO15a, Ras18, Sal18, SCAK^{+}19, SHTE23, WRBT21, ZL23]. machines [MRG17, PFH^{+}16]. Madden [HKGvS21]. Made [CX23, RG00, AW06, Cox14]. Magallanes [Liu22]. magcoilcalc [MFQ^{+}21]. MAGICC [GWM18]. magnet [MFQ^{+}21]. Magnetic [MB07, CFW17, HHVB21, MFQ^{+}21, OC20]. Magni [OPA^{+}14]. Magpylib [OC20]. maintainability [LTT^{+}24]. maintaining [RS17]. major [WMA^{+}22]. majors [Guz03, HHK^{+}09]. make [Mad15]. makes [Kla99]. Making [Dal01, DDK19, HGHR20, SSS22, SKS23, Swe09, WKS23]. Man [OOM^{+}23]. Man-in-the-Middle [OOM^{+}23]. manage [RC18]. Management [Ano97d, Ano00c, BBB02, CCM^{+}23, IG19, KDC^{+}18, ML16a, AW06, Hos12, JAGP14, SFC23]. Managing [Por03, PMBF17]. Manifolds [HAGH18, ZAPS20]. manipulate [GZT^{+}18]. Manipulation [Mil18, RDB21]. ManPy [DPH16]. Manual [DF00, van98c, vRD05a, Mac91, Mac92a, van95c, HYG16]. Manuel [Liu22]. manufacturing [DPH16]. Many [HGHR20, Kra22, Gra18]. many-body [Gra18]. Many-Faceted [Kra22]. Many-Objective [HGHR20]. map [KXK^{+}23]. Maple [GS12]. mapped [Fas18]. Mapping [PBB22, ZTT22, GRW^{+}19, HK15, MZL^{+}23, Tom15, WM21]. maps [Gre18]. March [HTA^{+}97]. Marginalization [SN12]. Maria [Hor22, Sau23]. Mark [Ano97a, Ano00a]. Markdown [XDR21, LKJC21]. Marketing [Mil15]. Mass [SN12, Ras20, Wha18]. Massive [DRFB^{+}23]. massively [BSG^{+}16, Mos20a]. master [Ada14, Meh15, Pip15, Zac15]. Mastering [Dog15, Het14, Kas15, LR14, Lot14a, Mad15, Meh15, Rom14, She15b, Wei15, BJ14a, Het10, Tid07]. Matching [Hig22, KDC^{+}18, KB18]. MatchingTools [Cri18]. MatchPy [KB18]. MatCoupLy [Roa23]. material [DCRF23, PP23, SNNB17]. Materials [VEV^{+}19, PSGL21, SLDF^{+}21, TQGE23, VMRFC23, WM21]. Math [Ano15b, Sah15]. Mathematica [Sta19]. Mathematical [SdSS16, SMT^{+}21, Wei15]. Mathematics [LL08]. MATLAB [KMLP^{+}23, MSS^{+}19, CLMM20, TCE21]. Matplotlib [RC18, Tos09, Nel15]. Matrices [FSZD20, Wim12, FM16, MRT^{+}22]. matrix [HJHZ18, LB22, Roa23, SH17]. matter [May21, Sch21]. maximize [MRT^{+}22]. Maximum [LD19, SR19]. Maxwell [CKK^{+}13]. May [Ano97d, Ano01a, Eur91]. Maya [Gal14]. Maybe [How06]. mcda [WBS22]. McNulty [Lip22]. MCycle [Hug18]. MD2D [LN23]. Mead} [PP23]. means [SLA^{+}23]. measurements [FNX22, MTS^{+}18, ddSNX22]. Measuring [ÅW16, KCS11, DDT20, ZC20]. mechanical [SLDF^{+}21]. Mechanics [Ble20, BWMS22, MS15, GPP^{+}21]. mechanics/Gromacs [GPP^{+}21]. mechanochemical [DyH21]. MechElastic [SLDF^{+}21]. media [BVV22, Guz03]. Mediator [Far02]. Mediator/Python [Far02]. Medical [ISMA18, SBC^{+}17, SML06]. medium [LRvE17]. medium-sized [LRvE17]. Meep [LFN^{+}11]. meets [ZXX24]. membranes [WRBT21]. memory [Fas18, SSH08, SZW^{+}22, Tab10]. memory-mapped [Fas18]. mental [SG07, TTHZ23]. Merging [SYGY21]. mesh [GS12, HR20]. meshfree [MSP^{+}22]. message [BL97a]. Meta [Tro96b, Le 24, MD15]. meta-analyses [Le 24]. meta-compilation [MD15]. MetaCluster [TOPC23]. metagraphs [RNR17]. metaheuristic [TOPC23]. metaheuristic-based [TOPC23]. Method [HFB21, LMW12, MOM21, RHM^{+}17, BE20, CWC^{+}23, CKK^{+}13, DEMM19, KXK^{+}23, LKK23, LB22, SNNB17]. Methodologists [Dör08]. Methods [Gez20, Gez23, LT03, MS07, Nor17, AES^{+}22, AM10, BW22, BWS23, FL15, GF18, KRH18a, KRH18b, KRSD^{+}23, Kiu10, Kiu13, MSV23, VCLS21, WKS22, You08]. metrics [Gre16]. MFDFA [GHKW22]. MGRIT [HFB21]. MGtoolkit [RNR17]. microcontrollers [Ano19]. microframework [VBM^{+}18]. Microscopy [BDF^{+}20, OPA^{+}14]. MicroStructPy [HR20]. microstructure [HR20]. microsynthesis [Smi18]. mics [BLE21]. mics-library [BLE21]. Middle [OOM^{+}23]. migration [Fee16]. migrations [Lot15b]. Milestone [SV14]. minimal [May21]. minimal-lagrangians [May21]. Minimum [Hig22]. Minimum-Weight [Hig22]. Mining [LS16, sR11, Rus14, YHX22, Ish17, ICVG14, Lay15]. Minkowski [BT21]. Minneapolis [HTA^{+}97]. Minolta [Ano00c]. Misconception [RAH^{+}01]. Mitch [Ano01a]. Mitchell [Mit00]. MitM [OOM^{+}23]. Mix [Est01]. Mix-ins [Est01]. Mixed [TT21]. mixes [LPG^{+}23]. ML [ZL23]. MLLPA [WRBT21]. MLxtend [Ras18]. mmappickle [Fas18]. MN [HTA^{+}97]. MO [Cas17]. mob [CS21]. Mobile [ST^{+}07, CS21, HMD^{+}96, Len15]. MobileNetV2 [PH23]. MobileNetV2-BiLSTM [PH23]. Modal [PARS22]. Mode [CR12, DTR18]. Model [BTR14, CFMR19, EML21, PH23, War18, ABS20, CVV23, GWM18, Hen08, JPJ^{+}23, May21, RBV16, WZZ^{+}23, WHG17, Bie21]. Model-based [CFMR19]. Modelica [DRFB^{+}23, Ern99]. ModelicaGridData [DRFB^{+}23]. Modeling [Ano01b, AJH^{+}12, DDMS14, HGHR20, ISMA18, JPOB20, KFV20, KB07, Laz22, Lip22, Mil14, Piñ24, RSRT19, Tie16, YH21, ZV19, BK19, BMZ^{+}18, GG21, HHM18, HCPF95, KN15, Mar17, MFQ^{+}21, Mil15, NZPWR22, PSGL21, PGM19, Ros18, SFC23, WMM18, YMCF23, YMA23, Lee17, Ish19]. Modelling [PHF10, SMT^{+}21, BVV22, BJR^{+}20, LFT21, PBN^{+}09, Tay18]. Models [HM18, LD19, Lip21, SBMD14, ZV19, DPH16, DSB^{+}23, DB22, GKK^{+}23, GVLD^{+}23, Kar14, KVSC21, Lay15, MD22b, MD22a, SG07, NRC23]. Modern [Sev15b, AJJF14, Lub14, MS15, Mit15, TV13]. moderne [LAG07]. modernize [Lot15b]. modified [KSS20]. modify [Kno08]. modred [BTR14]. Modul [Lin06b]. Modular [CFW17, HYG16, MBLA19, BST^{+}17, CJGA23, HPT^{+}16, KSH14, THG^{+}23]. Modularisierung [EK08]. Module [AAB22, Ham98, KB07, MD17, Mor17, SBMD14, TGEA09, Wei06b, BS21, BGHC23, CGHGRB21, CR22, CWC^{+}23, CWM^{+}21, DyH21, Fas18, LN23, MCGK19, MD22b, PARS22, PHH^{+}12, Tom15, WRBT21]. modules [Idr14]. moisture [MPMC21]. molecular [CMM14, DFC18, DPF^{+}21, FC17, HCPF95, KMK^{+}21, LHH^{+}21, LN23, SHFJ18, SSD^{+}22, SZNW23, VCF22, Yes15]. molecule [IHT23]. molecules [BGHC23]. moment [SR19]. MongoDB [HW19]. monitor [GCBDDBF23]. monitoring [MGFO23]. Monograph [Anoxxc]. Monte [ABGD^{+}20, DMD^{+}21, DEMM19, HQF^{+}20, JKKN23, KXK^{+}23, Mor17, Nil07b, PHH^{+}12, WAN^{+}22]. MontePython [Nil07b]. Monterey [Ass96]. monty [PMM^{+}13]. Moodle [Rad08]. morphodynamics [CVV23]. MoSDeF [DMD^{+}21]. most [Cox14, LR14]. Motifs [AAB22]. movies [BRF^{+}23]. moving [FH22]. MPI [DPS05, DPSD08, DFSW19, NO23]. MPI-2 [DPSD08]. MPI-parallelized [DFSW19]. mpi4py [DF21a]. mpmath [JT14]. mpnum [SH17]. Multi [EVV12, GMS17, Kra22, MLGW18, Och09, Pul00, Sta21, BHJ^{+}18, Fee16, FPN^{+}22, KMK^{+}21, LRPD18, LYZ^{+}22, MTPHH18, QZWU19, SUM21, SGPHD^{+}17, SKS23]. Multi-Agent [EVV12]. Multi-Case [Kra22]. multi-channel [SGPHD^{+}17]. multi-class [LYZ^{+}22]. multi-configuration [QZWU19]. multi-criteria [SKS23]. multi-energy [SUM21]. multi-image [LRPD18]. Multi-language [Och09, Fee16]. multi-objective [FPN^{+}22, KMK^{+}21]. Multi-Processing [MLGW18]. Multi-Regional [Sta21]. Multi-Robot [GMS17]. multi-scale [BHJ^{+}18]. Multi-Threaded [Pul00]. Multi-Threading [MLGW18]. Multiclass [HJHZ18]. Multicomplex [AMGM20]. multicoordinate [LKK23]. multideterminant [KRSD^{+}23]. multidimensional [Bar22]. Multidisciplinary [MMT09, HHK^{+}09]. Multidual [AMGM20]. MultiFLEXX [MTPHH18]. multiflexxlib [MTPHH18]. multifractal [GHKW22]. multigrid [BISC14]. Multilane [SYGY21]. multilayered [BVV22]. multimedia [GFB^{+}14]. multimodal [AES^{+}22, SL20]. multiphysics [BSG^{+}16, LPH23]. Multiple [Kor11, MP19b, BLE21]. Multiple-Precision [Kor11]. Multiprocessing [SMF^{+}23, AFL23]. Multiscale [LWV20]. Multithreaded [MAG21]. multithreading [Swi09]. Multivariate [Sha24, ASAA20, CSRV13]. MultiZ [AMGM20]. munuSSM [Bie21]. music [HOB23, MBK09]. Musicaiz [HOB23]. Musterlösungen [Wei06a]. MVTS [ASAA20]. MVTS-Data [ASAA20]. MWA [HYG16]. MyHDL [Dec04]. Myro [Men09]. MySQL [Bis01, HF06, TSC02]. Mython [Rie09].

N [VMRFC23]. named [ESM24]. Names [RAH^{+}01, AAYK23, Sta17]. nano [WMA^{+}22]. nano-optics [WMA^{+}22]. NanoNET [KVSC21]. nanoparticle [WAK22]. Nanophotonics [BVB^{+}07]. nanostructures [Wie18]. Nansat [KHD^{+}16]. NAS [DLC23]. Native [Kra22]. Natter [SLGB14]. Natural [BKL09, DD15, RM19, SLGB14, Per14b]. nature [VBM^{+}18]. nature-inspired [VBM^{+}18]. NCBI [HM22]. Ncpol2sdpa [Wit15]. need [Cha15]. Needs [Mar21]. negative [LB22]. Nelder [PP23]. Nelson [Lee17]. Neo4j [Gup15]. Neper2CAE [GTC21]. NESSie.jl [KRH18a, KRH18b]. Nest [Dub05]. Nested [BFM18]. nests [RCAE^{+}20]. Net [Bis01]. NetLogo [GG21]. Network [Ago01, Ano00c, BBB02, GRG21, HW19, LS16, Ngu08a, SYGY21, SST23, VBI^{+}23, XL18, AW06, CC20, FL06, GL07a, Gre16, HK15, RG10, RG14, Sar14, SW15, WZ18]. Networking [IEE20]. Networks [KSB12, OOM^{+}23, SJK^{+}21, VBB18, BS21, Boe17, BB22b, CJGA23, Mül23, FLR22, KMLP^{+}23, MDR^{+}22]. Netzwerkkommunikation [EK08]. Neural [FLR22, MDR^{+}22, SJK^{+}21, VBI^{+}23, BB22b, CJGA23, MCF^{+}22]. neural-guided [MCF^{+}22]. neurophysiology [NMGB17]. Neuropsydia.py [MD17]. NeuTomPy [MMG19]. neutral [AZH^{+}24, SLP^{+}22]. neutrino [AZH^{+}24]. neutrino-nucleus [AZH^{+}24]. neutron [MTPHH18]. neutrosophic [SABEh20]. NeuVis [Ano01b]. New-Style-Klassen [Wei06b]. News [HPT^{+}16, LRvE17, PHH^{+}12, Sav23, She15a, Yes15]. Next [Dal07]. Next-generation [Dal07]. NiaPy [VBM^{+}18]. NiceProp [GP22]. NIST [Ano95]. NJ [Lee17]. NL4Py [GG21]. NLKT [Per14b]. NLP [PF22]. NLTK [Per14b]. NNetEn [VBI^{+}23]. No [Ano15a, Ano15b, Wat95]. NOAA [War18]. noble [MFQ^{+}21]. node [KPKS22]. nodes [GVLD^{+}23, PP23, PP23]. noise [MD22b]. noisy [May17]. Noisyopt [May17]. nomisweb [Smi17]. non [Dah18b, FRdN21, GP22, GEH19, Guz03, LB22, NO23, RHLTG^{+}22, SLB^{+}23, ZRK21, ZRK22, KMLP^{+}23]. non-auroral [ZRK21, ZRK22]. non-contiguous [NO23]. non-ideal [GP22]. non-isothermal [RHLTG^{+}22]. non-linear [Dah18b, SLB^{+}23]. non-majors [Guz03]. non-negative [LB22]. non-parametric [FRdN21, KMLP^{+}23]. non-uniformly [GEH19]. Noncommuting [Wit15]. Nonlinear [MS07, RPL^{+}23, KF23, MD22b, MMP11]. nonlocal [KRH18a, KRH18b]. nonsense [Huc18]. normalization [BBHB22, SABEh20]. Norway [Eur91]. Nostril [Huc18]. Notation [TS14]. Notions [Chu07b, Chu07a]. November [IEE20]. Novice [Sma12, Het05, Het08, McG07a]. Novonix [GPKL^{+}20]. noWorkflow [PMBF17]. npm [CASA22]. NSDPY [HM22]. nuclear [WZ18]. nucleon [Kar23]. nucleus [AZH^{+}24]. Numba [GVLD^{+}23, KWBB22].

Numbers [AMGM20]. Numeric [MLGW18]. Numerical [Bar22, BT06, CYKG23, DHH96, Gez20, Gez23, HJK^{+}24, HM18, Joh15, Kiu10, Kiu13, LT03, Ram18a, Wim12, BHJ^{+}18, Bra13, Dah18c, DFSW19, LJ23, MS15, ML16b, Ros13, Ros14, WAK22, Mat16]. numerically [Dah18b]. NumPy [AJJF14, DOS17, Mil18, RC18, Fas18]. NURBS [BK19]. NURBS-Python [BK19]. nutshell [Mar03, Mar06].

Object [CI96, DY96a, ESM98, GL08a, HAGH18, Lut98a, Lut01, MMT09, MBLA19, USE96, AM19, BK19, GL07b, GL08c, HPH12, HHP^{+}16, JKST22, Kak08, KCS11, Lot14a, NS22, Phi10, Phi15, SML06, She97, YFD98]. Object-Orientation [CI96]. Object-Oriented [ESM98, HAGH18, Lut01, MMT09, MBLA19, USE96, DY96a, GL08a, AM19, BK19, GL07b, GL08c, HPH12, JKST22, Kak08, Lot14a, NS22, Phi15, SML06, YFD98]. Objective [HGHR20, BW22, BWS23, FPN^{+}22, KMK^{+}21]. Objects [ABC97, AAB22, She97, Kak08, Phi15, WBS22]. Objekte [Wei06a]. objektorientierte [Wei06b, Wei06a, Wei09]. objektorientierter [The09]. Objektorientierung [EK08]. objet [Swi09]. Observability [Kra22]. observables [Ver22a]. observations [Pop10]. ObsPy [TLR21]. obtain [VMRFC23]. obtaining [DBdFdSR21]. O’Connor [Ano13]. Octave [Wet20]. October [ACM97, Ano97c, CI96, USE00]. ODE [Dah18c]. OECP [YMCF23]. Off [Luk23]. Offended [RAH^{+}01]. Offline [GAS^{+}16]. oflibnumpy [RDB21]. oflibpytorch [RDB21]. Ogg [RAH^{+}01]. Oil [OOM^{+}23]. OLR [HKGvS21]. OLR-Based [HKGvS21]. OMI [HKGvS21]. On-Ramp [SYGY21]. One [GS12, Ngu08b, Pat16, Cha15, Ros08, TMH^{+}23]. One-Dimensional [Pat16]. One-Hour [Ngu08b]. One-peak [GS12]. online [JAGP14]. only [Cha15, KI19]. OO [LAG07]. OO-Programmierung [LAG07]. OOP [She97]. Open [AMB19, CX23, EML21, Eur91, HYG16, JPOB20, JČMG11, MMP^{+}22, MOM21, Ray98, AG23, Bah15, BK19, BLN^{+}21, CR22, CV22, CFW17, CG17, CBLI22, DPH16, FL15, FNX22, FH22, GMP^{+}20, GNPP23, HAC^{+}23b, HU17, Idr14, JNN12, JNN13, KPK^{+}17, KSH14, MAS^{+}22, MDRN18, MBP^{+}22, NZPWR22, OBWM23, PSGL21, SV14, SSS22, TOPC23, TV13, Ull15, VSS17, YMCF23, You08, ZRK21, ZRK22, ddSNX22, LFN^{+}11, LPG^{+}23, Len15]. Open-CV [Len15]. open-ended [YMCF23]. Open-Source [AMB19, CX23, HYG16, JPOB20, MMP^{+}22, AG23, BK19, BLN^{+}21, CR22, CV22, CFW17, CG17, CBLI22, DPH16, FNX22, FH22, GNPP23, HAC^{+}23b, HU17, JNN12, KPK^{+}17, KSH14, MAS^{+}22, MDRN18, MBP^{+}22, NZPWR22, OBWM23, PSGL21, TOPC23, TV13, YMCF23, ddSNX22]. open-world [VSS17]. OpenCL [Di 14]. OpenCV [Jos15, MH15]. OpenFL [DSB^{+}23]. OpenFL-XAI [DSB^{+}23]. OpenFOAM [MFL^{+}22]. OpenMechanochem [DyH21]. OpenMG [BISC14]. OpenMoist [MPMC21]. OpenSees [ZMS18]. OpenSeesPy [GYWG23, ZMS18]. OpenSeesPyView [GYWG23]. OpenStreetMap [Boe17, PL20]. Operating [CI96, Mos20b, LYZ^{+}22]. Operational [PARS22]. operations [BBHB22, SABEh20]. operator [RV20]. opt einsum [SG18]. Optical [RDB21, AGMFGE23, BJR^{+}20, CMS22, MD22a, SLB^{+}23]. optics [WMA^{+}22]. optimisation [FPN^{+}22, RFG13, SUM21]. optimisation-based [SUM21]. Optimization [CFMR19, CX23, EML21, HAGH18, JG17, MMT09, Mir18, PLRG22, RPL^{+}23, SdSS16, Wit15, ZTT22, AES^{+}22, GLS^{+}10, MFQ^{+}21, RV20, Rie09, SLA^{+}23, THG^{+}23, WBK^{+}24, Wie18]. optimizations [ABF13, BB17]. Optimized [LH20, TMH^{+}23]. Optimizer [EML21]. Optimizing [FSZD20, RKVL14, BB22b, FLR22, HWW^{+}15, LJ19, MD15, May17, SG18, ZLBF14]. options [LS98]. OPTN [HW19]. optoelectronic [VMRFC23]. OPUS [MSR03]. Oracle [Ano00c]. ORBKIT [HPT^{+}16]. orchestration [BB22a]. order [AMGM20, FSMCFP23, Lot14a, MO14, PGM19, SG18]. ordered [LYZ^{+}22]. Ordering [Ang00a]. ordinal [BBHB22]. Oregon [CLT20]. O’Reilly [Ano97a, Ano00a]. Organ [HW19]. Orientated [KHD^{+}16]. Orientation [CI96]. Oriented [ESM98, HAGH18, Lut98a, Lut01, MMT09, MBLA19, USE96, AM19, BK19, CJGA23, DY96a, GL07b, GL08a, GL08c, HPH12, HHP^{+}16, JKST22, Kak08, KCS11, Lot14a, NS22, Phi10, Phi15, Ros18, SML06, YFD98]. OS/2 [DF00]. Oscillation [HKGvS21]. oscillators [FH22]. osmfilter [PL20]. OSMnx [Boe17]. otebook [VMRFC23]. Other [Wil00, Hos12]. Our [Cas20]. out-of-plane [FNX22]. Outcomes [WX18]. outlier [SAB^{+}20]. Output [Sta21, FRdN21, SMD23]. OutPyR [SAB^{+}20]. Overdetermined [Luk23]. overhead [ZXZX22]. overview [LS98]. own [Nai14, RH15, SML06]. oxide [BWMS22].

P [VMRFC23, VMRFC23]. P2NMAP [HK15]. PaCAL [KJ14]. Pacific [HK95, IEE97b]. pack [Rad08]. Package [AVS20, CRL21, CX23, DB17, DF21b, GHXZ24, GPKL^{+}20, HFB21, Hig22, HKGvS21, HMB23, IG19, KFV20, KHD^{+}16, KJ14, LDCM23, LD19, LVH^{+}18, ML16a, MWS18, MBLA19, MOM21, OPA^{+}14, RDBC23, SdSS16, VEV^{+}19, VBI^{+}23, WL20, XMW10, XL18, YH21, AGMFGE23, ASAA20, AME^{+}21, AYI^{+}24, BYWW23, BFT20, Bie21, BLE21, Boe17, BE20, BHA18, BRF^{+}23, BJR^{+}20, CVV23, CMS22, CFSK14, CJGA23, CZ22, CG17, CKK^{+}13, Dah18a, DFC18, DPF^{+}21, DFSW19, EBNS22, EAMSR^{+}23, FSMCFP23, FH22, FRdN21, FC17, Gal22, GRW^{+}19, GKK^{+}23, GdGB^{+}18, Gir21, GL08b, GL09, GF18, Gra18, Hac23a, HHB^{+}22, HAC^{+}23b, Han18, HHVB21, HM22, HSD^{+}22, HHM18, Hug18, JEC18, JPJ^{+}23, KH23, KF23, KPK^{+}17, KPKS22, Le 24, LPG^{+}23, LHH^{+}21, LKK23, MWK^{+}20, MAC^{+}21, MGFO23, MAS^{+}22, MD22a, MTPHH18, MFQ^{+}21, MSV23, MMG19]. package [Mos20a, ML18, MR18, Mur18, NMGB17, NRC23, NEGZG18, OGA^{+}20b, OGA^{+}20a, OC20, PF22, PFLG21, PW17, QZWU19, RG18, RHLTG^{+}22, RNR17, RE22, RGP^{+}22, SHFJ18, SSH16, Smi18, SG18, Spi18, Sta17, TS20, TR22, TV13, TTHZ23, TMH^{+}23, VCF22, Var16, VM22, VC18, WZZ^{+}23, WBS22, Wha18, WKS23, WO14, ZTC^{+}21, ZC20, ZL23, ZRK21, ZRK22]. Packages [BUS21, CASA22, Mil18, ACS23, Lub14, ZMD21]. Packt [Wei06b]. Padasip [CV22]. PageRank [PKDM21]. Pages [Ang00b, Ang00c, Bie97, Ano97a, Ano00a, Liu22, Mit00]. pairs [AAYK23]. palette [TEG18]. palette-based [TEG18]. Palm [Ano00c, Ano00c]. Panda [Els23]. Pandas [Mil18, Nel15]. paper [Lee17, MDRN18]. Papers [Anoxxa, Anoxxb, Anoxxc, Anoxxd, ACM92]. paradigm [Kir04]. Parallel [BT06, CBW^{+}21, Coe17, DLC23, Di 14, DGMP09, HFB21, HTA^{+}97, Hin03, Hin07, IEE97a, Luk23, MAG21, MRG17, Pal14, PLRG22, RPL^{+}23, Sah12, TBA^{+}17, ZGV^{+}23, BSG^{+}16, CLM05, CGK11, GFB^{+}14, HHB^{+}22, HLSØ06, MRG18, PWFM17, Yes15, Zac15, MMP11, Par22]. Parallel-in-time [HFB21]. parallelisation [RCAE^{+}20]. Parallelism [BUS21, HBS^{+}20, SLA^{+}23]. parallelizable [GG21]. Parallelization [AAB22, DTM^{+}18, MAC14]. Parallelized [BTR14, CWM^{+}21, DFSW19]. Parallelizing [CL06]. Parameter [EML21, LWV20]. parameterization [KMK^{+}21]. parameters [DBdFdSR21, HHVB21, LJ19, RKVL14]. Parametric [LDCM23, FRdN21, KMLP^{+}23]. ParamGULP [DBdFdSR21]. Parapint [RPL^{+}23]. ParaView [SMD23]. Parglare [Dej22]. Parquet [KI19]. parser [Dej22]. ParShift [FSMCFP23]. parsing [Spi18]. Parsl [CBW^{+}21]. Part [Ang00b, Bri12a, Bri12b, Ang00c, FNX22, ddSNX22]. Partial [BT06, GWW09, LT03, MSL^{+}07, MD15]. Partially [Har15, MR22]. Particle [Hug14, ML16a, Mir18, Nor17, RBP^{+}21, WBR^{+}22, May21, RKVL14]. Particle-Based [Nor17]. particle-in-cell [RKVL14]. PARyOpt [PLRG22]. Passages [Gro20]. passing [BL97a]. Passive [HK15, JM20]. Password [GAS^{+}16]. PASTA [KBLJ18]. path [AZLY18, BSS16, CMM14]. path-based [AZLY18]. Pathology [MBA^{+}22]. Pattern [CJYH23, KB18, KF23]. Pattern-based [CJYH23]. Patterns [VCR17, Chr02, HZ23, Kas15, YHX22]. PBTK [EML21]. PC [Wal07, Ano00c]. PCA [ZAPS20]. PCA-derived [ZAPS20]. PCAfold [ZAPS20]. PDE [CL06]. PDF [DF00]. PDF-Dateien [DF00]. peak [GS12]. Penetration [ABCC22, Duf15, Buc15, Moh15, O’C13]. pentesters [Sei14]. Penultimate [RAH^{+}01]. People [Lip22, Piñ24]. peptide [WBS21]. perceptions [SL08]. Perfect [Hig22]. Perfectly [Sch17]. Performance [BKRT21, CBST24, DPSD08, Dog15, EPF^{+}23, FSZD20, IEE20, KWBB22, LH20, LLL^{+}20, LD07, MSP^{+}22, MBLA19, SRS^{+}23, SdSS16, TS14, ZXX23, BCRS15, BSS16, BS19, BCC^{+}18, BBC^{+}24, CLM05, CZ22, CFÁA^{+}20, DLC23, GVLD^{+}23, GO14, HF06, IOC^{+}12, Lan13, Lew09, LJ19, Mar18, ML16b, Par22, Ros13, Ros14, SNB^{+}19, SSH08]. performing [BSS16, VCF22]. periodic [HAC^{+}23b]. Perl [DF00, SKS08, Bla02, Bla04, HAE00, HF06, Kak08, LS97, MSR03, Ngu08d, Ngu08e, Pre03, Roo97, Sit02, SM04, Stu07]. permittivity [YMCF23]. Perprof [SdSS16]. Perprof-py [SdSS16]. Persistence [Ano01b]. Perspective [Eur91, KDC^{+}18, GRW^{+}19, Wep15]. perturbation [DB22, Gal22]. Perusing [Wil97b]. Petascale [Wit21]. Pfaffian [Wim12]. pgm [WZZ^{+}23]. pGrAdd [WV22]. phantoms [CG17]. Pharo [ADP21]. phase [AYI^{+}24, CM20, DCOC^{+}19, TQGE23, WRBT21]. phase-specific [WRBT21]. Phasepy [CM20]. phenology [Tay18]. Phil [Ano00b]. Philip [Lee17]. phonon [WZZ^{+}23]. photoemission [BYL^{+}21, SL20]. Photophysiology [RKR21]. PHP [SKS08, HF06, HS12, Ott18, SM04, Stu07, TTS^{+}10]. phrasebook [Day07b]. Phylogemetric [Gre16]. phylogenetic [Gre16]. phyphy [Spi18]. Physical [Ano21, Lee17, Tie16, Ish19, KN15, RKVL14, WAN^{+}22]. Physics [Aya14, Bäc07, Gez20, Gez23, MAFM21, MMP^{+}22, BLN^{+}21, Bor07, Di 13, EBNS22, MCGK19, RGS^{+}21, RSPJ21, San13, SSD^{+}22, Wep15, YMA23, BB13]. Phytoplankton [RKR21]. PI [CMM14, AG23, BW14, Cox14, Don14, MAFM21, Nix15]. Piano [TS14]. pickle [Fas18]. pICNIK [RHLTG^{+}22]. Pierro [Wep15]. PII [DPF^{+}21]. Pipeline [MSR03]. pipelines [WAK22]. pitfalls [CEI^{+}12]. PIV [NT24]. plan [Lot15b]. planar [ZL20]. plane [FNX22, ddSNX22]. Planes [RDBC23]. planetary [LRPD18]. planner [Boe23]. planning [LFT21]. planning-level [LFT21]. plant [PBN^{+}09, Tay18]. PlantGL [PBN^{+}09]. plasma [BYWW23, SSD^{+}22]. Platform [Ano00c, BBB02, Fun23, Lac06, SBC^{+}17, TT21, BLN^{+}21, HPT^{+}16, MDRN18, PSGL21, ST^{+}07, She97, Ull15, Lac06]. platform-independent [She97]. Platforms [HA20, GFB^{+}14, Par11]. Playdoh [RFG13]. Playing [Swe09]. PLEASE [GDP18]. plots [Tos09]. Plotting [OZW18, Han18, HFF^{+}17, JM20, Le 24]. Plug [DF00]. Plug-ins [DF00]. Plus [BUS21]. PLW [LD07]. PmlBeta [WBS21]. pMuTT [LWV20]. Pocket [Lut98b, Lut02, Lob19, Lut05, Lut14, Stu03, Stu07]. POGIL [HS13]. point [AM21, BMZ^{+}18, BII^{+}20, DMC^{+}15, FH22, JT14, SNNB17]. Poisson [Bra13, GNPP23]. polarizers [MFQ^{+}21]. policies [SFC23]. Policy [Liu22]. Political [Dör08]. pollution [RC18]. polycrystals [GTC21]. Polyhedral [LP19]. polymers [ECS23]. Polymorphie [Wei06a]. polymorphism [ÅW16]. Polynomial [SST23, VY15, Wit15]. Polynomial-Based [SST23]. pools [PWFM17]. popular [Idr14]. portability [GVLD^{+}23]. Portable [Di 14, Kor11, Bah15]. Portland [CLT20]. ports [Buc14]. posets [GS12]. positive [GS12]. Post [AM21, SJL18, GYWG23]. Post-processing [AM21, GYWG23, HTH^{+}20]. Post-treatment [SJL18]. Posterior [FV18, Han18]. postprocessing [HPT^{+}16]. posture [GCBDDBF23]. potential [ABS20, Cox14, DBdFdSR21, Hia15]. potentials [Kar23]. Power [BHS18, Day14, MOM21, BSS16, DRFB^{+}23, Gal14, Hug18, Lay15, MH15, Mur18, Phi15, SW14, SUM21, Tel06]. powerbox [Mur18]. powered [RDS07]. powerful [Els23, Idr14, Law15, Lot14a, Zac15]. pp [Hor22, Lee17, Lip21, Lip22, Sa’23, Sau23]. Practical [Ben08, Cue13, Gal14, Het02, Joh15, Mat16, CR15, DAJ^{+}15, Fla08, Ish17, ICVG14, Per14b, Swe14]. Practice [TS14]. practices [ZKB^{+}08]. practitioner [BM15]. pragmatic [RH15]. Pragmatics [Sco09]. praktisch [Wei06a]. pratical [Sar14]. Praxis [EK08]. Praxisbuch [Haj08]. pre [HTH^{+}20]. pre/post [HTH^{+}20]. pre/post-processing [HTH^{+}20]. Precise [YHA^{+}16]. Precision [GPKL^{+}20, Kor11, LV20, BJR^{+}20, BII^{+}20, JT14]. Preconditioning [CJ22]. predicting [AZLY18, Zie19b]. Prediction [HBA^{+}20, BWMS22]. predictions [ECS23]. predictive [Bow15, Lay15, Mil14, Mil15, RO15a]. preferences [DMC^{+}15]. PREMER [VBB18]. preparation [EPM09]. Preparing [TLR21]. Preprocess [GPKL^{+}20]. Preprocessing [CRL21, ASAA20, HHB^{+}22, LRL23, MBA^{+}22, PF22]. presentation [Kak08]. presented [ACM92]. Preserving [IG19]. Press [Ano15a, Ano15b, Cla15, Gve09, Hor22, Lip22, Liu22]. Presses [Ray98]. Price [Lee17]. Prime [Ano98d]. primer [Lan09, Lan11, Lan12]. Princeton [Lee17]. principles [AM21, LHB14]. print [Ano14]. priority [BBHB22]. prisms [LDAL20]. Pro [Alc09, BA14, Sil14]. probabalistic [Kar14]. Probabilistic [LP19, RBV16, DB22]. Probabilities [Pat16]. probe [YMCF23]. Problem [Bar21, Mis19, PBB22]. Problem-Solving [Bar21]. Problems [AVS20, Ble20, DB17, LeV09, ŚBA^{+}15, Wit15, ZTT22, Kar14, SKS23, She15b, TOPC23, Wil05]. ProbLog [FLR22]. ProbLog-based [FLR22]. procedure [HF06]. Proceedings [ACM92, ACM97, Ano97c, Ano97d, CI96, ESM98, Eur91, HTA^{+}97, IEE97b, IEE97a, IEE20, USE96, USE00, CLT20]. Process [Fun23, OOM^{+}23, RE22, Ros18, Wes15]. process-oriented [Ros18]. processes [SSS22, SAA18]. Processing [APS10, AMB19, BKL09, Ble20, FLS^{+}20, GRKN^{+}19, HTA^{+}97, Hug14, IEE97a, KHD^{+}16, MLGW18, Mil18, RM19, RKR21, SBC^{+}17, SH19a, AM21, BST^{+}17, GYWG23, HTH^{+}20, How98, McG00, Mer03, MMG19, Myr15, Per14b, PWFM17, SGPHD^{+}17, She15b]. processor [Tab10]. Procrustes [MRT^{+}22]. Procurement [HW19]. Produced [GPKL^{+}20]. product [SH17]. production [AZH^{+}24, ZTC^{+}21]. Productive [CBW^{+}21]. Productivity [Ano01b, BUS21, Moo21]. Products [Ano00c, Ano01b, MAS^{+}22]. Professional [MS^{+}00, MBW07, Het05, Het08, McG07a]. professionellen [DF00]. Profi [Haj08]. Profi-Leitfaden [Haj08]. Profile [Ano19, FhDAF09, SdSS16, Ott18]. Profile-guided [FhDAF09, Ott18]. profiles [LPG^{+}23]. Profiling [CYKG23, SLA^{+}23]. Program [DDLW02, AZLY18, BYL^{+}21, BII^{+}20, DEMM19, DCOC^{+}19, HS13, Len15, MCF^{+}22, MTS^{+}18, NL23, SLP^{+}22, Swe09, DBdFdSR21]. Programmable [OA17]. Programmanfänger [The09]. Programmation [Zia06, Bai08, Swi09]. Programmbeispiele [Lin06b]. Programmer [Ang99, Ano00b, McG98c, Wil97b, Wil00, Wil01, Wil02, Swi09]. Programmering [Lin06a]. Programmers [HR00, Ros08, Sma12, Cox14, Meh15]. Programmierer [Haj08]. Programmierung [EK08, LAG07, The09, Wei06a, Wei06b, Wei09]. Programming [ACM92, Ano95, Ano97b, Ano00b, Bal18, CL06, CG23, CBW^{+}21, DY96b, DY99, Ful96, Gar98, Gra00, Gut13, Gut16, HR00, HC16, Hin02, HWJ^{+}21, HMD^{+}96, Joh97, KLM14, KLM15, LT03, Lut96, Lut01, Lut06, MAG21, McG98b, McK16, MSS^{+}19, Nag17, Orb18, Pet02, Pre00, RG00, Roo98, San15, SDS00, Sco09, Seg07, SNNB17, SS13, Sum09, Sum10, Tro96a, WX18, Wit15, Ano00a, BCRS15, BSS16, BS19, BKMY03, BKMY04, Bro06, CLM05, Can14, Cas20, Chr02, Chu01, Chu07c, CKC23, Daw03, Daw06, Daw10, DAJ^{+}15, DLC23, Don14, DV21, Els23, FL06, FMPS17, Fla08, Gal14, GS12, GVLD^{+}23, GL07a, GL07b, GL08a, GYWG23, Gut04, Har07, HF06, Har12, Hia15, Hig03, Hil15a, Hil20, HLSØ06, HPH12, Joh12, Kin05, Lak17, Lam15, Lan13, Lan09, Lan11]. programming [Lan12, Law15, Lee11, LHB14, LFT09, LL08, Lob19, Lot14a, Lot15a, MS^{+}00, Mau02, McG98a, MRG18, MR07, MR09, MBW07, MSW08, Mue14, Nel15, Och09, Pal14, Phi10, Phi15, Ram15, RG10, RG14, RH15, Ros13, Sah12, Sah15, Sal18, SL08, SV14, Sar14, SW15, SWD15, Sei09, Sei14, She15b, SB15, Sum08, Swe13, Swe14, TSC02, ÜK12, WvA96, Wep15, WO14, Yaa23, Zac15, ZKB^{+}08, Bax01, Ano97a]. programming-based [GYWG23]. Programs [Di 14, JCY^{+}19, ÅW16, AAYK23, BSS16, Lan13, Nag06]. progress [Sie17]. project [BB13, Cha15, Pop10, You08]. Projects [CCM^{+}23, KCVM22, VCR17, Arb14, Bal12, Ben08, SW14, Sal18, WPH^{+}22]. Projekte [Ric14]. propagate [SLP^{+}22]. propagation [MR22, MD22a]. Properties [AGR24, AZLY18, CM20, DCRF23, SP23, SLDF^{+}21, TR22, VMRFC23, WV22]. proposals [SSS22]. Protein [SN12, KRH18a, KRH18b, MMEH08]. proteins [MMEH08]. prototype [Len15]. prototypes [DAJ^{+}15, YPB16]. Prototyping [BKC14, KBC^{+}20, Spe19, UVa95, CBLI22, KRSD^{+}23, RCRS06, ST^{+}07, vd91]. Provably [LJ23]. provenance [PMBF17, SGZ23]. provided [Pip15]. Providing [Ras18]. ProxySource [Ano01b]. pruned [BB22b]. psrqpy [Pit18]. Pteros [Yes15]. PTRAIL [HHB^{+}22]. public [SLP^{+}22]. publication [BRF^{+}23, JAGP14, Tos09]. publications [RRRGVD19]. Publishable [SJL18]. published [Gve09]. Pudipeddi [Myr15]. pulsar [Pit18]. pulse [MD22b, MD22a]. pure [SSD^{+}22]. pure-Python [SSD^{+}22]. Purpose [AJYH18, LS16, DCRF23, OGA^{+}20b, OGA^{+}20a, TV13]. put [Cas20]. puts [Ang99]. pvlib [HHM18]. py [BFT20, PP23, BYWW23, FLS^{+}20, FPN^{+}22, HC16, BGHC23, LR23, MD22a, SdSS16]. Py-ART [FLS^{+}20, HC16]. Py-EFIT [BYWW23]. py-fmas [MD22a]. py-irt [LR23]. PY-Nodes [PP23]. PY-SUMMA [AVS20]. Py2neo [Gup15]. PyACTS [DGMP09]. PyAlbany [LPH23]. PyARPES [SL20]. PYATB [JPJ^{+}23]. PyBaMM [SMT^{+}21]. PyBanshee [KMLP^{+}23]. pyBaram [Par22]. PyBEST [BLN^{+}21, BBC^{+}24]. pybliometrics [RK19]. PyblioNet [Mül23]. pycalphad [OL17]. PYCASO [CWC^{+}23]. PyCDFT [MWK^{+}20]. PyCDT [BMZ^{+}18]. PyCharge [FH22]. PyCiGen [GTC21]. pycity scheduling [SUM21]. Pycket [BBH^{+}15]. pyCLAMs [ZWCQ22]. PyCM [HJHZ18]. PyCOMPSs [BCE^{+}22, TBA^{+}17]. PyCSP [Gal22]. Pycuda [RDBC23]. PyDDA [JCL^{+}20]. PyDDRBG [AES^{+}22]. PyDEC [BH12]. PyDEF [SJL18]. pydiffusion [CZZ19]. PyDII [DMC^{+}15]. PyDMD [DTR18]. pyDSM [ECS23]. PyDTNN [BCM21]. PyECOM [GPEM23]. pyEIA [VCLS21]. pyEIT [LYX^{+}18]. pyerrors [JKKN23]. PyExaFMM [KWBB22]. PyF95 [LC11]. pyfao56 [BTDT23, Tho22, TBD23]. PyfastSPM [BRF^{+}23]. pyFDM [WKS22]. pyFFS [BKH^{+}22]. pyFIRI [ZRK21, ZRK22]. pyfMRIqc [WL20]. PyFocus [CMS22]. PyFR [Wit21]. PyFrac [ZL20]. PyFWI [MGFO23]. Pyg4ometry [WAN^{+}22]. pygal [Ada14]. Pygame [KM15, McG07a]. PyGBe [CFCB17, CCFB16]. PyGBe-LSPR [CFCB17]. pyGDM [WMA^{+}22, Wie18]. pyGFC [Ngu08c]. pyGLLE [MD21]. pyGlobus [Jac02]. pyGNMF [LB22]. Pygpc [WPM^{+}20]. pyGPGO [JG17].

PyGran [AM19]. pygrank [KPKS22]. PyGTED [BSSz^{+}20]. PyGtk [CK00]. pyhector [WHG17]. pyHomogeneity [HMB23]. pyIAST [SSH16]. PyIFDM [WKS23]. PyIVNS [SABEh20]. pyJac [NCS17]. pyJacqQ [JWHS16]. Pykat [BJR^{+}20]. PyLCP [EBNS22]. PyLESA [LFT21]. PyLighter [BC09]. PyLlama [BVV22]. PyLog [HWJ^{+}21]. Pylons [Gar09]. PyLops [RV20]. Pymagicc [GWM18]. PyMatching [Hig22]. PyMaxEnt [SR19]. PyMC [PHF10]. pyMCD [LKK23]. pymcdm [SKS23]. pyMDO [MMT09]. PyMGRIT [HFB21]. PyMikor [Bar22]. PyMIP [AW06]. pymodconn [CJGA23]. PyModPDE [KSS20]. PyMOL [Moo21, WBS21]. Pymrio [Sta21]. PyMTL3 [JPOB20]. pyneqsys [Dah18b]. pynucastro [WZ18]. PyNumero [RPL^{+}23]. pyodesys [Dah18c]. PyOECP [YMCF23]. PyOMA [PARS22]. PyOMA GUI [PARS22]. PyOMP [MAG21]. PyOPAsolver [BBHB22]. PyOpenCL [MR22]. PyOphidia [EPF^{+}23]. pyparsing [McG07b]. pyParticleEst [Nor17]. pyparty [Hug14]. PyPathway [XL18]. PypeR [XMW10]. pyPhenology [Tay18]. PyPhotonics [TR22]. PyPI [GHXZ24]. PyProcar [HTH^{+}20, LTT^{+}24]. PyPSA [BHS18]. PyPy [ABF13, LJ19]. PyQBench [JLP23]. PyQT [Har12, Rem01a, Sum08]. PyQUBO [ZTT22]. PyR [LS17, SS21]. PyRad [AME^{+}21, FLS^{+}20]. PyRDM [JAGP14]. pyrepo [WBS22]. pyrepo-mcda [WBS22]. Pyret [NMGB17]. PyRETIS [LRvE17]. Pyrgg [Hag17]. Pyro [BKMY03, BKMY04]. pyROM [PGM19]. PySCF [HAC^{+}23b]. pySDC [Spe19]. pySDC-Prototyping [Spe19]. pySecDec [BHJ^{+}18, HJJ^{+}22, HJK^{+}24]. pysimm [DFC18, DPF^{+}21, FC17]. PySPH [RBP^{+}21]. PySpike [MK16]. PySPLIT [War18]. pysrim [OZW18]. PySSM [SBMD14]. PySTPrism [LDAL20]. PyStructureFactor [SZNW23]. pySuStaIn [AWO^{+}21]. PySwarms [Mir18]. PySy [WO14]. pyTDGL [Hor23]. pyTEP [RE22]. PyTES [XGW23]. pythermalcomfort [TS20]. Python [Ano97c, Ano00b, Ano02, AJH^{+}12, BBB02, Bax01, BK19, CR15, DFC18, DPF^{+}21, DF00, EK08, Haj08, Hil20, Hug01, Ish17, Ish19, Jan10, Joh15, Lac06, Laz22, Lee17, Lin06b, LAG07, MDR^{+}22, Mat16, Mil18, Mue14, Orb18, Ram18b, RAH^{+}01, Ric14, Sar14, SML06, Ste14, SKS08, Swi09, The09, Wei06a, Wei09, Wep15, BKMY04, Swe14, ddSNX22, ÅW16, AZH^{+}24, Ada14, ACZ^{+}22, AGMFGE23, ASAA20, AES^{+}22, AVS20, AWO^{+}21, ABGD^{+}20, ABS20, ACS23, AAYK23, APS10, AW06, Ang99, Ang00b, Ang00c, Ano95, Ano97b, Ano98a, Ano98b, Ano98c, Ano98e, Ano00c, Ano01b, Ano12, Ano18, Ano19, Ano20, Anoxxa, Anoxxb, Anoxxc, Anoxxd, Ant15, AYI^{+}24, ADP21, Arb14, AG23, AFL23, ABC97, AM21, AAB22, AMB19, Aya14, Bäc07, BW22, BWS23, BB22a, Bah15]. Python [Bai08, BYWW23, Bar21, BKRT21, BDF^{+}20, BP17, Bar11, BUS21, Bas08, Bas10, BG17, Bau96, BSG^{+}16, BLP^{+}21, BVV22, BL97a, Bea00, Bea01, Bea06, Bea09b, Bea09a, Bea12, BJ14a, BJ14b, BHF16, BBHB22, BS21, BJM23, BISC14, BKH^{+}22, BB13, Bie21, Bie97, BVB^{+}07, BST^{+}17, BK19, BKL09, Bis01, BLE21, BGHC23, Bla02, Bla04, BKMY03, BMK03, BWMS22, Boe17, BE20, BT21, Boe23, BLN^{+}21, Bor07, BSSz^{+}20, BM15, BHA18, Bow15, BW14, BB22b, BYL^{+}21, Bra13, BTDT23, BFM18, BRF^{+}23, BDQ^{+}22, Bri12a, Bri12b, BMZ^{+}18, BCG05, Bro00, BM18, BHS18, BJR^{+}20, Bro06, BA14, BLS^{+}23, Bru09, Buc15, Buc14, CGHGRB21, CLM05, CL06, ÇA23, CZA^{+}23, Can14, CCM^{+}23, CVV23, CMS22, CR22, CWC^{+}23, Cas17, Cas20, CG23, CFPS23, CBST24, CSZ^{+}19, CV22, CMM14, CRL21, Cha15].

Python

[CV15, CFSK14, CM20, CYKG23, CJGA23, CZZ19, CZ22, CJYH23, CX23, CHM24, CFW17, CBLI22, CS09, Chr02, CSRV13, Chu01, Chu02a, Chu02b, Chu07b, Chu07a, Chu07c, Chu08, CKK^{+}13, CFÁA^{+}20, CFCB17, Coe17, CH17, CLMM20, CHG^{+}16, CR12, Con95, CBB14, CCFB16, CC20, Cot03, CWM^{+}21, CWLG^{+}21, Cox14, Cri18, CKC23, Dah18a, Dah18c, DPS05, DPSD08, Dal07, Dan18a, DCRF23, DSB^{+}23, Dar12, Dav21, Daw03, Daw06, Daw10, Day14, Day07b, Day07a, DyH21, DMD^{+}21, DEMM19, Dec04, DM20, DDLW02, Dej22, DTR18, DAJ^{+}15, DLC23, Di 13, Di 14, DB17, DMC^{+}15, DD15, Dog15, DDMS14, Don09, Don14, DRFB^{+}23, Dör08, Dou98, DFSW19, Dow09, Dow15, DGMP09, DY96a, DHH96, DY96b, DY99, Dub07, Duf15, DF21b, DV21, DBdFdSR21, EBNS22, EML21, ESM24, EPF^{+}23, Els23, EPM09, EK08, EK09, Ern99]. Python [EAMSR^{+}23, Est01, ECS23, EVV12, Far02, Fas18, Feh02, FSMCFP23, FMPS17, FKA^{+}17, FNX22, FH22, FRdN21, Fla08, FV18, FBC09, FM16, FPN^{+}22, FC17, FLR22, FPSZ21, Ful96, Fun23, Gad09, GMKRS21, GH18, GHN19, Gal14, Gal22, Gar98, Gar09, GS12, GRW^{+}19, Gez20, Gez23, GKK^{+}23, GWM18, GJ08, GdGB^{+}18, Gir21, GP22, GQCP^{+}18, GVLD^{+}23, GHT06, GHT08, GZT^{+}18, GL07b, GL08a, GL08b, GL08c, GPEM23, GFB^{+}14, GCBDDBF23, GPKL^{+}20, GNPP23, GO14, GHKW22, GPP^{+}21, Gov15, GF18, GDP18, GEH19, Gra18, Gra00, Gre18, Gre07, Gre16, GRKN^{+}19, Gue18, GG21, Gun14, GYWG23, Gup15, Gut04, Gut13, Gut16, GWW09, Hac23a, HBS^{+}20, HGHR20, HS22, Hag17, HJHZ18, HFB21, HHB^{+}22, Ham98, HR00, Han18, HAE00, Har15, HF06, HR20, HW19]. Python [Har12, Has16, HJPB17, HHVB21, HM22, HSD^{+}22, HC16, HTH^{+}20, Her14, HU17, HPT^{+}16, HOB23, HBA^{+}20, Het02, Het05, Het08, Het10, Het14, HFF^{+}17, Hia15, Hig22, Hig03, Hil15a, Hil15b, Hin02, Hin03, HLSØ06, Hin07, HHJCRB21, HKGvS21, HQF^{+}20, HAB^{+}20, HRS06, HB16, Hor23, Hos12, Hos14, HK15, How06, How98, HH17, HZ23, HWJ^{+}21, Huc18, Hug99, Hug14, HLR15, Hug18, Hug97, HMB23, HHP^{+}16, Idr14, IHT23, ICVG14, ISMA18, Jac02, JAGP14, JLP23, JEC18, JKST22, JPOB20, JG17, JPJ^{+}23, JWHS16, JNN12, JNN13, JT14, Joh08, Joh12, JM20, JD02, Jos15, JKKN23, KWBB22, KFV20, Kak08, Kar23, KSS20, KS21, Kar14, Kas15, KH23, KJ15, KCVM22, KF23, KG23, KRSD^{+}23, Kin05, KN15, Kin09, KM15, KPK^{+}17, Kiu10, Kiu13, Kla99, Kla11, KVSC21, Kno08, KSH14]. Python [KMLP^{+}23, KHD^{+}16, KBC^{+}20, KJ14, KXK^{+}23, KPKS22, KB07, KB18, KMK^{+}21, Kuc97, Kuc98b, Kuc98c, KD99, KBLJ18, Lac06, LS97, LR23, LFN^{+}11, Lam15, Lan13, LM03, Lan06, Lan07, Lan08, Lan09, Lan11, Lan12, LRPD18, Law15, Lay15, Le 24, LeV09, LPG^{+}23, Lee11, LD19, LHH^{+}21, LKK23, LB22, Len15, LRvE17, Lew09, LVH^{+}18, LN23, LHB14, LPH23, Lin06a, Lin06b, LL08, LYX^{+}18, LYZ^{+}22, LFYX23, LHM14, LWH12, LR14, Lot14a, Lot14b, Lot15a, Lot15b, Lub14, LKF23, Lun01, LH20, LDW^{+}21, LD07, Lut96, Lut98a, Lut98b, LA99, Lut99, Lut01, Lut02, LA04, Lut05, Lut06, Lut08, Lut09a, Lut09b, Lut13, Lut14, LFT21, LWV20, LS17, MWK^{+}20, MZL^{+}23, Mac92b, MAC^{+}21, Mad15, MB17, MD17, MP19a, MP19b, MS15, MSP^{+}22, MAFM21, MMP^{+}22, MBA^{+}22, MSL^{+}07, MGFO23, Mar17, Mar18]. Python [MLB22, Mar02, Mar03, MMA05, Mar06, MRA06, Mar07, Mar21, MCGK19, MM06, MAG21, MFL^{+}22, May17, MAS^{+}22, MTS^{+}18, McF16, McG98b, McG98c, McG98a, McG00, McG07a, McK12, Meh15, MMEH08, MD21, MD22b, MD22a, MPMC21, MTPHH18, MFQ^{+}21, MRT^{+}22, MSV23, Mer03, MSP^{+}17, MMG19, MMP11, MSR03, MR07, MR09, Mil14, Mil15, MB07, MA11, MH15, Mir18, MWS18, Mit15, MBA19, MBLA19, Moh15, MOM21, MBW07, Mor17, ML16b, Mos20a, MSW08, MSS^{+}19, ML18, MK16, MAC14, Mül13, MR18, Mur18, MS07, MGS07, NS22, NMGB17, Nag06, NT24, Nel15, Ngu08d, Ngu08a, Ngu08b, Ngu08c, Ngu08e, NM22, Nil07b, Nil07a, NRC23, NL23, Nix15, NO23, Nor17, NZPWR22, NEGZG18, O’C13, Oli07, OGA^{+}20b, OGA^{+}20a, Orl97, OMGDG14, OE21, OC20, Osa17, OL17, OBWM23, OPA^{+}14, Pal14]. Python [DDT20, PP23, PM23, Par22, Par11, PARS22, Pat16, PHF10, Pay11, PKDM21, Per14a, PGH11, Per14b, Pet02, PF22, Phi10, Phi15, PFLG21, Pil04, Pil09, PMBF17, Piñ24, PDS^{+}22, Pip15, Pit18, Ple02, Plo97, PL20, PRH17, PMM^{+}13, PHH^{+}12, Pop10, PBN^{+}09, PSGL21, Pre03, PW17, PWFM17, Pul00, QZWU19, RS17, Rad06, Rad08, RKVL14, RMM23, Ram18a, RBP^{+}21, Ram15, RG18, RHLTG^{+}22, RDS07, RCAE^{+}20, RCT23, RG00, RNR17, RMZG06, RO15a, Ras18, Ras20, RV20, RDB21, RO15b, Ree04, RZ09, RE22, Rem01a, Rem01b, Ren23, RG10, RG14, Ric15, RGS^{+}21, Roa23, RSPJ21, RGP^{+}22, RH15, Rom14, Roo97, Ros18, RK19, Ros13, RFG13, Ros14, RC18, RKR21, SR19, SW14, SLP^{+}22, Sah12, Sah15, SSH08, Sal14, San13, SL08, San15, SV14, SW15, Sav23]. Python [ST^{+}07, Sch21, Sch17, SMF^{+}23, SAD22, SL21, SUM21, Sco17, SWD15, SHFJ18, Sei09, Sei14, SDP^{+}20, SBC^{+}17, Sev13, Sev15a, Sev15b, SLB^{+}23, SGPHD^{+}17, Sha24, Sha03, SSS22, She15a, She97, SMD23, Shi07, SP23, Sil14, SAA18, SSD^{+}22, SSH16, SLDF^{+}21, SZW^{+}22, SLGB14, SHTE23, SdSS16, Sit02, Sla15, SABEh20, Sma12, Smi17, Smi18, SG18, SMM^{+}22, SZNW23, SFC23, SK19, Spi18, Spo12, Sta17, Sta21, Sta19, Ste06, SB15, SBMD14, SM04, Stu07, Suc13, SH17, SMT^{+}21, Sum08, Sum09, Sum10, SGZ23, Suz97, SH19b, Swe09, Swe13, SLA^{+}23, Tab10, Tan01, TFAL21, Tan06, TS20, Tau16, TR22, Tay18, TBA^{+}17, Tel06, TCE21, THG^{+}23, TOPC23, TSC02, Tho22, TBD23, TF23, Thu22, Tie16, TGEA09, Tom15, TQGE23, Tos09, TTHZ23, TMH^{+}23, Tro96c, Tro96a, Tro96b, Tro97, TSG^{+}23]. Python [TBJ18, ÜK12, Ull15, Vai09, Vai14, VCF22, VB08, Van97b, Van16, Var16, VEV^{+}19, Vel23, VBI^{+}23, VM22, Ver22a, Ver22b, VAP^{+}21, VC18, Vir16, VCLS21, VKSB15, VBM^{+}18, WBR^{+}22, WSK22, WX18, Wal07, WAN^{+}22, Wal97, WL13, WRBT21, WZZ^{+}23, WBK^{+}24, War18, Wat95, WvA96, Wei06b, Wei15, WPM^{+}20, WAK22, Wes15, Wet20, WPH^{+}22, Wie18, WMA^{+}22, WKS22, WMA12, WMM18, WO14, Wil97a, WHG17, Wil00, Wil05, Wit21, WV22, XMW10, XL18, XGW23, Yaa23, YKKD19, Yan14, YHX22, Yes15, You08, YMA23, YH21, Zac15, ZTT22, ZTC^{+}21, ZAPS20, ZC20, ZWCQ22, ZXZX22, ZXX23, ZXX24, ZL23, Zho97, ZMS18, Zia06, ZKB^{+}08, Zie19a, Zie19b, ZRK21, ZRK22, Zuk97, dPPD^{+}23, dos01, vd91, van95a, van95b, van95c, van95d, van97a, van98b, van98c, van98a, vRDK99]. Python [vRD03, vRD05a, vRD05b, vdOJP^{+}20, vdH18, vF97, Bri06, HHM18, Ano97a, Ano13, Ano15a, Ano15b, Hor22, Joh97, Lak17, Lip21, Lip22, Liu22, Myr15, Orr05, Roo98, Sa’23, Sau23, ZMD21, Ano00a, Ano14, Cla15, Gve09]. Python-3 [MCGK19]. Python-Based [BBB02, BKMY04, ddSNX22, Aya14, HWJ^{+}21, Ram18a, AZH^{+}24, ABGD^{+}20, BKMY03, Dav21, Dec04, FNX22, GP22, GCBDDBF23, GPP^{+}21, Gup15, JAGP14, Lac06, LHH^{+}21, NL23, OBWM23, PBN^{+}09, RBP^{+}21, SLB^{+}23, SFC23, SLA^{+}23, Ver22a, VCLS21]. Python-Buch [vF97]. Python/C [HZ23, van98c]. Python/Jupyter [Bar21]. Python/Numba [GVLD^{+}23]. PythonFOAM [MFL^{+}22]. Pythonic [BLN^{+}21, Dal01, ENCS20, JCL^{+}20, SRS^{+}23].

Pythons [Dub05]. PythonTEX

[MS13, Poo16]. Pythran [Gue18]. Pytim [SHFJ18]. PyTOPS [YKKD19]. PyTorch [JS24, Mis19, RM19, SH19a, ZGV^{+}23]. PyTracer [CYKG23]. PyTransport [MR18]. PyTrilinos [SSH08, Spo12]. PytuTester [MBP^{+}22]. pyuca [Tau16]. PyUnfold [BHA18]. PyUNIxMD [LHH^{+}21]. pyuvdata [HJPB17]. PyVecContour [MZL^{+}23]. pyvine [YH21]. pyvrft [BE20]. PyVT [LRL23]. pywidgets [VMRFC23]. PyWolf [MR22]. PyXtal [FPSZ21].

Q&A [Chu02b]. Qball [AME^{+}21]. QCOBJ [VC18]. QEDv2 [SV14]. QExpy [MCGK19]. QGIS [ÇA23, Law15]. qha [QZWU19]. QmeQ [KPK^{+}17]. QOptCraft [AGMFGE23]. Qt [Sum08]. quadratic [GS12]. Quail [CBLI22, HFF^{+}17]. Quality [WL20, BRF^{+}23, Tos09]. quantification [FL15, KMK^{+}21, OGA^{+}20b, OGA^{+}20a, TSG^{+}23]. Quantifying [ZXZX22]. QuantImPy [BT21]. Quantitative [KLM14, KLM15, Yan14]. quantity [VC18]. quantity-aware [VC18]. QuantLib [Vir16]. QuantLib-Python [Vir16]. Quantum [Hig22, NM22, Nil07b, AGMFGE23, GPP^{+}21, Gra18, HPT^{+}16, JLP23, JNN12, JNN13, KRSD^{+}23, KPK^{+}17, MD22b, SV14, TMH^{+}23]. Quantum-classical [NM22]. quasiharmonic [LDW^{+}21, QZWU19]. quaternions [RG18]. QUBO [ZTT22]. Qudi [BST^{+}17]. QuDPy [SLB^{+}23]. Quelltext [DF00]. Query [Bis01]. querying [Pit18]. quest [Ros08]. Questionnaires [MD17]. queueing [DDMS14]. quick [Lob19]. quickly [SW15]. quickstart [Feh02]. quimb [Gra18]. quivers [GS12]. QuTiP [JNN12, JNN13].

R [Hor22, Lip22, Liu22, Sau23, AVS20, BP17, Dan18a, GH18, GHN19, Mil14, Mil15, Mül13, Piñ24, Smi17, Smi18, SMM^{+}22, XMW10, XDR21]. R/PY [AVS20]. R/PY-SUMMA [AVS20]. R/Python [AVS20]. R13 [TT21]. Radar [FLS^{+}20, HC16]. radiation [WAN^{+}22]. radiolysis [AME^{+}21]. Ramp [SYGY21]. Random [Hag17, KJ14, Pat16, PRH17, AYI^{+}24, CSRV13]. range [Pip15, Sar14]. ranging [She97]. ranking [KPKS22]. rankings [Cas20]. Rapid [Ano01b, GDP18, RDS07, Ree04, RCRS06, Sum08, UVa95, Gar09, ST^{+}07]. rare [LRvE17]. Raspberry [AG23, BW14, Cox14, Don14, MAFM21, Nix15]. RaspberryPi [MBP^{+}22]. RaspyControl [AG23]. Rate [CRL21, DCOC^{+}19, McF16]. rates [SZNW23, WZ18]. Rational [Ano00c]. Ratios [DF21b]. Raton [Hor22, Lip21, Lip22, Sau23]. Ravishankar [Myr15]. Raw [WL20]. ray [CG17]. raytracing [Gir21]. RC [MAFM21]. RCLAda [Mos20b]. RDBMS [DDK19]. Reaching [Gre07]. reacting [Gal22]. reaction [WZ18]. reader [Sta19]. readership [Myr15]. reading [HB16, VAP^{+}21]. ready [Ada14]. reagent [Pop10]. Real [CWM^{+}21, FLS^{+}20, RAH^{+}01, SGPHD^{+}17, VCR17, WPH^{+}22, AG23, Bri06, GHT08, Gup15, HAC^{+}23b, Jos15, Kar14, Lot14a, Sar14, Yan14]. real-life [Yan14]. Real-Time [CWM^{+}21, FLS^{+}20, SGPHD^{+}17, AG23, HAC^{+}23b]. Real-World [VCR17, GHT08, Gup15, Jos15, Kar14, Lot14a, Sar14]. Realistic [CHM24, CMS22]. Reality [UVa95]. realtime [MDRN18]. recall [HFF^{+}17]. receiver [LYZ^{+}22]. Receptury [BJ14b]. Recipes [BJ14a, Mis19, Buc15, Cox14, Law15, Per14b, Ros14, Sar14]. Reciprocal [Sco17]. recognition [ESM24]. recognizer [KF23]. recoloring [TEG18]. Recommending [STS^{+}18]. reconciled [BCC^{+}18]. Reconfigurable [Kla99]. reconnecting [Lon08]. Reconstruction [OPA^{+}14, BYWW23, MMG19, SR19].

Recordings [TS14]. Records [GRG21]. recursion [SG07]. reduced [PGM19]. Reduction [BTR14, MTPHH18, MSV23]. Reed [Jan10]. refellips [RGP^{+}22]. Reference [Ano00b, BE20, Lut98b, Lut02, WBS22, van98c, vRD05a, BS19, Bea00, Bea01, Bea06, Bea09b, CJYH23, GHT08, Lob19, Lut05, Lut14, Stu03, Stu07, van95b, van95c]. Referenz [EK08]. Refinement [CJ22, LP19]. RegCPython [ZXX23]. region [Ott18]. region-based [Ott18]. Regional [Sta21, LPG^{+}23]. regions [HJJ^{+}22]. Register [ZXX23]. Register-based [ZXX23]. Regression [BHF16, Lip22, Piñ24]. Regular [Stu03, Stu07, YH21, LR14, Rom14, SM04]. Reguläre [SKS08]. reinforcement [dPPD^{+}23]. Related [DD15, KCVM22, ML16a, Wes15]. relations [KF23]. relationship [LFT09]. Relative [HAB^{+}20, MTS^{+}18]. relativity [CFPS23]. relaxation [WMM18]. Relaxations [Wit15]. RelaxPy [WMM18]. Release [Chu02a, vRD03]. Released [Ano00c]. Releases [Ano00c]. reliability [LTT^{+}24]. Religions [Wil00]. remapping [GKK^{+}23]. remarkable [Tos09]. remediation [TFAL21]. Remote [PLRG22, AG23]. Renewables [GMKRS21]. renormalization [JS24]. repair [BRM10]. replacement [Ano12]. replay [BRM10]. Replication [Ano21]. Report [Ano21, CM07, Pop10]. Representation [SDP^{+}20, AZLY18, HPH12, SH17]. Reproducible [Coe17, LeV09, BLE21, MBA^{+}22, SMM^{+}22]. reprogramming [OMGDG14]. requests [CV15]. resampy [McF16]. Research [AMB19, BVB^{+}07, LeV09, Liu22, Mir18, Par22, TS20]. Researcher [Moo21]. réseau [Swi09]. reservoir [MGFO23]. Residual [SJK^{+}21]. resilience [ZTC^{+}21]. ResiPy [ZTC^{+}21]. Reslicing [RDBC23]. RESNET [SJK^{+}21]. Resolution [Ble20]. resolved [SL20]. Resonance [MB07]. Resource [Bru09, SFC23, VFMM08]. Resources [Cas17]. Response [LR23, FRdN21]. responses [SDP^{+}20, SLB^{+}23]. responsive [Ull15]. REST [Suc13]. restrained [ABS20]. Result [BLS^{+}23]. Results [Ano21, Pop10]. RETCON [BRM10]. Retrievals [JCL^{+}20]. retrofitted [TTS^{+}10]. retuschieren [DF00]. Reusable [Mit00]. reuse [Kno08]. Revamping [Sav23]. reverse [Sei09]. Review [Ano97b, Ano97a, Ano00a, Ano13, Ano14, Ano15a, Ano15b, Cla15, Gve09, Hor22, Ish17, Ish19, Jan10, Joh97, Lak17, Laz22, Lee17, Lip21, Lip22, Liu22, Mat16, Orb18, Orr05, Roo98, Sa’23, Sau23, Wep15]. Reviews [Ano00b, Bax01, Hug01, HSD^{+}22, Le 24]. Revisiting [CHM24]. Revolve [HA20]. Rexx [Pre03]. Reyes [Liu22]. rft1d [Pat16]. RGBXY [TEG18]. RGBXY-space [TEG18]. RGEs [LS17]. rheology [ECS23]. Rhodium [HGHR20]. RhythmCount [VM22]. rhythmicity [VM22]. Rich [Lac06, Lac06]. richtig [DF00]. Riemannian [HAGH18]. rifarensu [SM04]. Right [HKM08, Buc15]. Rigidity [HS22]. rigidPy [HS22]. Rigorous [LV20]. riparian [CVV23]. river [CVV23]. RNA [SAB^{+}20]. RNA-Seq [SAB^{+}20]. Robinson [Ano00a]. Roborueda [GCBDDBF23]. Robot [GMS17, Len15, Mos20b]. robotics [BKMY03, BMK03, BKMY04, Els23, Len15, Men09, OMGDG14]. Robots [KB07, FMPS17, Men09]. Robust [CX23, HGHR20, Pal14, SK19]. Robustness [CFMR19]. rock [Ras20, YMA23]. rockphypy [YMA23]. Rong [Sa’23]. Rookies [MSW08]. ROOT [DOS17]. root numpy [DOS17]. roots [GS12]. ROPTLIB [HAGH18]. ROS [Len15]. Rossum [Chu02b, Dou98, Kuc98a, Sev15a, Sev15b]. rotational [BGHC23]. Rotten [ADP21].

Round [Luk23]. Round-Off [Luk23]. Routines [Mül13]. rowan [RG18]. RPC [Por03]. RPython [MRG17]. RSOME [CX23]. Ruby [SKS08, Bla02, Bla04, Gut04, Ngu08d, Ngu08e, Stu07, ZMD21]. rule [Sto13]. rules [SK19]. run [Pip15]. Running [Ber13, SW15]. Runtime [CHM24, VSS17]. Russell [Ano00b]. Rust [ABCV21, LH20].

S [VMRFC23]. S2352711016300395 [DPF^{+}21]. safe [Men09]. Saha [Ano15b]. salabim [vdH18]. SALib [HU17]. Salt [Hos12]. sample [McF16]. sampled [GEH19, SH19a]. Sampling [BFM18, OPA^{+}14, SN12, YH21]. San [ACM92, Ano97c]. Sarkas [SSD^{+}22]. SATLAS [GdGB^{+}18]. SC’20 [IEE20]. Scalable [ENCS20, GFB^{+}14, LR23, Lon08, RPL^{+}23, RSRT19, Mos20a, Pip15, RV20]. scale [BL97b, BHJ^{+}18, BCG05, EPF^{+}23]. scales [PBN^{+}09]. Scaling [MAC14, ZGV^{+}23]. scattering [AZH^{+}24, HJK^{+}24]. Scatterplot [FM16]. scheduling [SUM21, dPPD^{+}23]. Scheme [DF00, GNPP23]. Schemes [BHF16, Sta17]. schlably [dPPD^{+}23]. schneller [Wei06b]. Schrödinger [MD22b]. schwimmbad [PWFM17]. Science [Bad20, BCE^{+}22, Bar21, CBST24, HHJCRB21, HKGvS21, IEE97b, Lan08, MT18, Van16, VMFG17, BM15, HHK^{+}09, Jos16, Lan06, Lan07, Lob19, Mad15, MT19, MR07, Mil14, Mil15, Nel15, PSGL21, Rad06, RMZG06, Ras18, Ros14]. Sciences [Has16, SLA^{+}23]. Scientific [BKRT21, DGMP09, DY96b, DY99, HTA^{+}97, Hin02, Oli07, PG07, PGH11, SDS00, Sma12, VB08, BL97b, Ber13, CLM05, CSRV13, DDMS14, FKA^{+}17, Hil15a, Hil20, JAGP14, Lak17, Lan09, Lan11, Lan12, Meh15, Nil07a, NEGZG18, Ras18, RCRS06, RRRGVD19, SML06, Wep15, YFD98]. Scientist [Dow15, KHD^{+}16, Dow09, DDK19]. Scientist-Orientated [KHD^{+}16]. Scientists [Hor22, MA11, Nag17, Sau23, Ste14]. scientometric [Le 24, RRRGVD19]. ScientoPy [RRRGVD19]. Scikit [HLR15]. Scikit-spectra [HLR15]. SciPy [Ber13]. SCO [DF00]. Scope [Myr15]. Scopus [RK19]. Scott [Ano19, Mit00]. scraper [Nai14]. scraping [Mit15, Nai14]. scratch [Jos16, Len15]. screening [DB22]. Scriptable [RK19]. Scripting [Ano98d, Bea98, Dör08, Hin07, Kak08, Lan08, Lut98a, Lut01, McG98c, Ngu08e, Pre03, Rem01b, Wei06a, AW06, Bea96, BL97b, CEI^{+}12, FhDAF09, LS97, Lan06, Lan07, Ngu08d, She97]. Scripts [Bla04, GTC21, HW19, LM03, She97, Van97b, Ano12, Bla02, Ngu08d, PMBF17, Vai09, Vai14]. sDNA [CC20]. seamless [Els23, SHTE23, THG^{+}23]. seamlessly [CV15]. searches [WBK^{+}24]. Searching [KBLJ18, LKK23, PP23]. Seattle [CI96]. Sebastopol [Ano97a, Ano00a]. secml [PDS^{+}22]. Second [Ano00b, Orb18, USE96, Ano00c]. Secret [Swe13, Lot14b]. Secure [PDS^{+}22, SST23]. Security [Dan18b, ACS23, O’C13, Ste06]. SED [Dav21, Bla02, Bla04]. SEDBYS [Dav21]. Seiki [SM04]. Seismic [TLR21, NL23]. Seitz [Ano15a]. selection [SVY09]. Selective [Zuk97]. Selenium [Gun14]. Self [VMFG17, FH22, HWW^{+}15, MD15]. self-consistent [FH22]. Self-Explain [VMFG17]. self-optimizing [HWW^{+}15, MD15]. semantic [OA17]. Semi [Ano97d, LV20, KVSC21]. Semi-Automatic [LV20]. semi-empirical [KVSC21]. Semi-Structured [Ano97d]. semiconductors [AM21, BMZ^{+}18]. Semidefinite [Wit15]. semistructured [HGMC^{+}97]. SEML [ZV19]. sendmail [Ano00b]. sense [Mad15]. Sensitivity [HU17, WPM^{+}20]. Sensor [KBC^{+}20]. Seq [SAB^{+}20, SNB^{+}19]. sequence [CJGA23]. sequence-to-sequence [CJGA23]. sequences [HM22, WBS21]. serial [CLM05]. Series [BKH^{+}22, HMB23, KSB12, VBI^{+}23, ASAA20, EAMSR^{+}23]. Server [Ang00b, Ang00c, Ano00c, Ano01b, GMS17, Mit00]. serverless [AFL23]. Servers [Kla99]. services [CV15]. setups [CWM^{+}21]. Seven [Pre00]. Shaped [RDBC23]. Sharded [ZGV^{+}23]. shared [Ren23]. sharing [Hos14]. sharpen [Ros14]. Shawcroft [Ano19]. shell [AME^{+}21, Bla02, Bla04]. shmlast [Sco17]. short [MDRN18]. Should [How06]. Showcase [USE00]. SIAM [HTA^{+}97]. side [Jos16]. Signal [APS10, CRL21, VBI^{+}23, MD22a]. signals [JKST22]. Signatures [RG20]. significance [BW22, BWS23]. Silicon [ACZ^{+}22]. similar [AAYK23]. similarity [MRT^{+}22, ZC20]. Simple [MAFM21, MMP^{+}22, AW06, BS21, CKC23, GWM18, Lub14, She97, WHG17, ZXX24]. simplex [PP23]. simplification [Gal22]. Simplified [ZV19]. simplify [GQCP^{+}18, HLSØ06, BB22b]. Simplifying [NO23]. SimPrily [GQCP^{+}18]. simulate [Len15]. Simulating [MMEH08, AME^{+}21, BFT20, GPEM23, MR22, Mar17, SV14, Sch21, TMH^{+}23]. Simulation [CZZ19, JPOB20, Osa17, AM19, ABGD^{+}20, BSG^{+}16, BYL^{+}21, CS21, DPH16, DFC18, DPF^{+}21, DRFB^{+}23, FC17, KXK^{+}23, SFC23, WAK22, vdH18, vdOJP^{+}20]. Simulations [AMB19, MBLA19, MSS^{+}19, SMF^{+}23, CMM14, DyH21, DEMM19, FH22, GQCP^{+}18, HHP^{+}16, ML16b, RKVL14, RE22, San13, SHFJ18, THG^{+}23, Wie18, WMA^{+}22]. Simulator [LFN^{+}11, NS22, ZL20]. simultaneously [TQGE23]. SimuPy [Mar17]. single [IHT23]. single-molecule [IHT23].

Singular [CJ22, Gal22]. singularities [SL21]. Sismic [DM20]. site [DMC^{+}15]. sites [RDS07]. situ [MFL^{+}22]. sized [LRvE17]. sizing [Hug18]. Skew [Wim12]. Skew-Symmetric [Wim12]. Ski [MH18]. skijumpdesign [MH18]. Skill [Ngu08d]. Skills [Ano20, DV21, LFT09, Ros14, She15b]. Skin [SJK^{+}21]. Slicer [RDBC23]. Slitflow [IHT23]. Small [FSZD20]. Smart [ABCC22, LFT21, Seg07]. Smart-Home-Based [ABCC22]. Smells [CCM^{+}23]. SMMP [MMEH08]. Smooth [Pat16]. Smoothed [RBP^{+}21]. Snake [Wit21]. SNAP [LS16]. Snippet [Har15, Moo21, XMW10]. SOA [Suc13]. SoC [LHM14]. Social [Liu22, sR11, Rus14, SLA^{+}23]. SODA [ACZ^{+}22]. Soft [AJYH18]. Software [ACM97, Ano01b, AGR24, BH12, Coe17, Gve09, HPT^{+}16, Hin03, IEE97b, KWBB22, LRvE17, MMP^{+}22, ML16a, Mül23, PLRG22, PHH^{+}12, RRRGVD19, STS^{+}18, SdSS16, SJL18, TFAL21, Tro96c, Tro96a, Tro96b, Tro97, Ver22b, WL20, Yes15, AME^{+}21, BWMS22, CG17, CWLG^{+}21, DPH16, DMD^{+}21, Dow09, GNPP23, HAC^{+}23b, Hen08, JAGP14, KSS20, KS21, KCS11, KI19, LYZ^{+}22, MLB22, PARS22, Ros08, SR19, SGZ23, TV13, WBS22, WV22, YMCF23, ZAPS20, ZKB^{+}08, ZRK21, ZRK22, ddSNX22, Ano00c, Ano01b, FNX22]. solar [HHM18]. Solaris [DF00]. Sold [RAH^{+}01]. Soldier [RAH^{+}01]. Soloff’s [CWC^{+}23]. solute [DMC^{+}15]. Solution [BT06, LMW12, Bra13, DCOC^{+}19, Mis19, SSH16]. Solutions [Tro96c, Tro96a, Tro96b, Tro97, Sar14, TF23]. Solve [Dah18b, DB17, MSL^{+}07, GNPP23, Kar14, She15b, Wil05]. solvent [CBB14]. solvent-filled [CBB14]. Solver [TT21, CFCB17, CBB14, LPH23, Par22]. Solvers [CL06, CFMR19, MOM21, MSP^{+}22, SSH08].

Solving [Bar21, LD19, Luk23, ŚBA^{+}15, CKK^{+}13, MD21, SKS23]. something [Orl97]. Sometimes [Wil02]. sophisticated [She97]. Sorted [Har15]. soundness [VSS17]. Soup [Nai14]. Source [Anoxxd, AMB19, CX23, EML21, HYG16, JPOB20, LFN^{+}11, MMP^{+}22, MOM21, Ray98, AG23, Bah15, BK19, BLN^{+}21, CR22, CV22, CFW17, CG17, CBLI22, DPH16, ESM24, FL15, FNX22, FH22, GNPP23, HAC^{+}23b, HU17, Idr14, JNN12, JČMG11, KPK^{+}17, KSH14, MAS^{+}22, MDRN18, MBP^{+}22, NZPWR22, OE21, OBWM23, PSGL21, SML06, SSS22, SGZ23, TOPC23, TV13, Ull15, YHX22, YMCF23, You08, ZRK21, ZRK22, ddSNX22]. sous [Bla02, Bla04]. Space [JWHS16, STS^{+}18, SBMD14, LDAL20, TEG18, ÇA23]. Space-Time [JWHS16, LDAL20]. Sparse [Wit15, LRPD18, SZW^{+}22]. Spatial [LVH^{+}18, CZ22, CC20, GEH19]. spatio [LRL23]. spatio-temporal [LRL23]. Special [Bal12, Kor11]. Special-Function [Kor11]. species [WV22]. Specific [AAB22, HWW^{+}15, Sla15, WRBT21]. Specified [MH18]. spectra [Mur18, HLR15]. Spectral [CH17, Spe19]. Spectrometry [SN12]. Spectrometry-Based [SN12]. spectroscopic [RGP^{+}22]. spectroscopies [SL20]. Spectroscopy [HLR15]. spectrum [WPH^{+}22, CH17]. Speculative [JCY^{+}19, BB17]. SpharaPy [GEH19]. sphere [Hac23a]. spherical [LJ23]. sphstat [Hac23a]. spike [MK16]. spin [MFQ^{+}21, TMH^{+}23, VMRFC23]. spin-half [TMH^{+}23]. spin-one [TMH^{+}23]. spin-polarizers [MFQ^{+}21]. Spinney [AM21]. Spinsim [TMH^{+}23]. spot [ÇA23]. Sprachgrundlagen [EK08]. Spring [Eur91]. Springer [Sa’23]. SQL [Bad20, SCAK^{+}19, Ano20]. SQLAlchemy [Cop08, Gar09]. squeezing [Ano19]. Sridevi [Myr15]. SRIM [OZW18]. SSI [PBK22]. stability [TQGE23]. stable [BVV22, LJ23]. stack [Ras18, Hos12]. Stage [AWO^{+}21]. Standard [Bie21, Lun01, Gar09, HCPF95, OE21, Spi18]. Standardization [RKR21]. Starbase [Ano00c]. Starch [Ano15a, Ano15b]. Stark [CFSK14]. Stark-effect [CFSK14]. Stars [Gre07, Dav21]. start [Cha15, Cla15, Her14]. Started [Nai14, BCRS15, Hos12, McG07b, Nix15, Ric13]. Starting [Gad09]. Stata [HHJCRB21, SMM^{+}22]. State [PBB22, SBMD14, LHH^{+}21, MSV23, Wei15]. state-of-the-art [MSV23, Wei15]. statechart [DM20]. states [LKK23, KBLJ18]. Static [Mar21, AES^{+}22, FhDAF09, RJAL09, RF16]. statically [CEI^{+}12, IOC^{+}12]. statically-typed [IOC^{+}12]. statistical [BWMS22, CSRV13, HR20, MSV23, San13, Wei15]. Statistics [Has16, Hor22, ICVG14, JWHS16, Sau23, Sha24, SLGB14, Hac23a, RCT23, Sah15, Ish17]. StatMechGlass [BWMS22]. Status [DF21a]. Steering [YFD98]. Stencil [LLL^{+}20]. Step [Liu22]. Step-by-Step [Liu22]. Stern [CBB14]. stimulation [XGW23]. stk [TBJ18]. Stochastic [CX23, PHF10, CR22, OE21]. Stop [Ray98, Wit21]. Storage [BDT13, IEE20, KI19, NEGZG18, PFH^{+}16]. store [Fas18]. stored [HF06]. Stplanpy [Boe23]. Straightforward [Dah18c, Arb14]. strain [SH19b]. strategies [BDT13, PFH^{+}16]. Strategy [MWS18, SYGY21]. streamline [GKK^{+}23]. streams [BRF^{+}23]. street [Boe17]. string [Huc18]. Structure [BLN^{+}21, CSZ^{+}19, BGHC23, BWMS22, FPSZ21, HTH^{+}20, JPJ^{+}23, SZNW23, WBK^{+}24]. Structured [Ano97d, DB22, Kra22, Mur18]. Structures [ISMA18, Jan10, RZ09]. Structuring [Ngu08a]. stub [vd91]. Student [Lee17, SG07, SL08, WX18, Ish19, KN15]. Students [SDS00, HS13, Lon08]. Studies [JWHS16, BLE21, ÜK12]. Studio [SW14, SW14]. Study [CASA22, KCVM22, Kra22, TS14, AGMFGE23, AJJF14, FSMCFP23, HZ23, LJ19, LHM14, LKF23, OMGDG14, Rad08, RCRS06, SSS22, TFAL21, WRBT21]. Studying [MAFM21, May21]. stuff [Swe14]. Stunden [Wal07]. Style [Wei06b, WBK^{+}24]. Styles [AGR24]. Subglacial [SMF^{+}23]. subspaces [CHG^{+}16]. substances [Zie19b]. Subtype [AWO^{+}21]. Successful [RMZG06]. such [Gup15]. suitability [OMGDG14]. suite [BST^{+}17, SSD^{+}22, GDP18]. SUMMA [AVS20]. Summarization [VPO19]. Summit [Ray98]. Supercomputing [BLP^{+}21]. Supersymmetric [Bie21]. Supply [GHXZ24]. support [DDK19, HB16, HMD^{+}96, IOC^{+}12, MCGK19, Tab10]. Supporting [HBS^{+}20, KL97]. supports [SML06]. supramolecular [TBJ18]. survey [Ish17, ICVG14, MTS^{+}18, RRRGVD19, BLE21]. SuSe [RAH^{+}01]. sustainable [Boe23]. SVG [Ada14]. SW [Wu13]. Swallowed [CG23]. Swarm [Mir18]. SWIG [Bea96, Bea98, Cot03]. SWIGLAL [Wet20]. Switching [GRG21]. Switzerland [IEE97a]. Symbolic [FLR22, JCY^{+}19, AM10, BKC14, Cri18, HOB23, MSP^{+}17]. symbolically [Dah18b]. Symmetric [Wim12, CFSK14]. symmetry [CFSK14, FPSZ21]. Symposium [ACM97, CLT20, HK95, IEE97a]. SymPy [JČMG11, MSP^{+}17]. synchronizing [DTM^{+}18]. Synchronous [HA20]. synchrony [MK16]. Syngress [Ano13, Ano14]. Syntax [SS13, Lob19]. Synthesis [HWJ^{+}21, MCF^{+}22]. synthesizing [LWH^{+}10]. System [Ang00a, Ano00c, BHS18, Bru09, EVV12, IG19, Kor11, LLL^{+}20, PG07, ZGL20, CS21, CR22, DRFB^{+}23, FLR22, GJ08, Hos12, HMD^{+}96, KCS11, KL97, Sil14, Ste06, GKK^{+}23, Mos20b]. Systems [Ano21, BKRT21, CI96, CJ22, ESM98, Eur91, KDC^{+}18, LP19, Luk23, LFT21, MS07, MGS07, USE96, AGMFGE23, AG23, CR15, DPH16, Dah18c, Dah18b, DFC18, DPF^{+}21, DDMS14, FC17, Gal22, HAC^{+}23b, HHM18, JNN12, JNN13, JČMG11, LHM14, Mar17, MMP11, Pal14, QZWU19, SUM21, SP23, SAA18, SLA^{+}23, TMH^{+}23, VSS17, ZTC^{+}21].

T [Ano00b, Ano13, Ish17]. tailor [Cox14]. tailor-made [Cox14]. Taint [CR12]. Taking [Luk23]. tale [GHT06]. talk [Sie17]. Taming [Els23]. tasca [Day07a]. Task [GMS17, Fee16, HBS^{+}20, SLA^{+}23]. Tasks [MD17, PF22, Sar14]. TauRunner [SLP^{+}22]. Taylor [Lip21, Lip22]. TB2J [HHVB21]. Tcl [SML06, Ass96, DF00, Bla04, Ngu08d, Ngu08e, Bla02, LS97, McG98c, Pre03]. Tcl/Tk [Ass96, McG98c]. TDD [Sal14]. TDDFT [HAC^{+}23b]. TE [LS17, SS21]. teacher [Myr15]. Teaching [AMB19, GL07b, GL08c, KLM14, KLM15, BKMY03, BKMY04, CBLI22, MBK09, Par22, RMZG06]. tech [Cas17]. TechMiner [Vel23]. Technical [KDC^{+}18, PM23, BCRS15, BS19, TFAL21]. technique [YMCF23]. Techniques [Joh15, Mat16, Bow15, Idr14, Lan13, MS15, Mil14, Mil15, Per14b, TSC02, WPH^{+}22]. Technologies [USE96]. Technology [ACM97, AGR24, ESM98, Hos14, Lon08]. Ted [Cla15]. teens [Lam15]. tegdet [BJM23]. Teik [Sa’23]. Teleconferencing [Ago01]. TeMIP [AW06]. Template [How98]. templating [LC11]. Temporal [LVH^{+}18, BII^{+}20, LRL23]. Tennessee [RE22]. tensor [JS24, SZW^{+}22]. Tensorial [TT21]. Teoh [Sa’23]. Teradata [SCAK^{+}19]. terms [MD22a]. Test [Gov15, HAB^{+}20, HMB23, Ngu08a, Per14a, Ang99, LKF23, SAD22, TTHZ23]. Test-driven [Gov15, Per14a]. tester [MBP^{+}22]. Testers [GPKL^{+}20, O’C13]. Testing [ABCC22, Duf15, Sal14, YH21, Arb14, Buc15, DM20, Gun14, LC11, MBA^{+}22, Moh15]. Tests [Ano21, ADP21, GH18]. Tethys [LVH^{+}18]. TEX-alike [Thu22]. Text [Aya14, DF00, Mer03, VPO19, Per14b]. textbook [Hor22, Myr15]. TextCL [PF22]. TFInterpy [CZ22]. their [GS12]. theKompany.com [Ano01b]. theoretic [Boe17]. Theory [GRG21, LR23, Cri18, JEC18, MWK^{+}20, SSH16, VCF22]. Thermal [ZGL20, BDQ^{+}22, TS20]. thermochemical [LWV20, WV22]. Thermochemistry [LWV20]. thermodynamic [Hug18]. Thermodynamics [OL17]. thermoelasticity [LDW^{+}21]. Thermofeel [BDQ^{+}22]. they’re [LS97]. things [McK16, DAJ^{+}15, SST23]. Think [Dow15, Dow09]. thinking [HHK^{+}09]. Thomas [KSB12]. thoroughly [BM15]. Threaded [Pul00]. Threading [MLGW18]. Threads [Wei06a]. Three [Bea12, MWS18, Ros08]. Three-Strategy [MWS18]. throughput [GQCP^{+}18, OLRLB21]. tight [JPJ^{+}23, KVSC21]. tight-binding [JPJ^{+}23, KVSC21]. Time [Ano98d, ÇA23, CWM^{+}21, FLS^{+}20, Hor23, HMB23, JWHS16, KSB12, VBI^{+}23, Wil01, ASAA20, AG23, BJM23, EAMSR^{+}23, HFB21, HAC^{+}23b, LJ19, LDAL20, MWM20, NM22, RF16, SGPHD^{+}17, TTS^{+}10, WMA12]. Time-dependent [Hor23]. tip [GHT06]. Tits [GS12]. Tk [Ass96, Bla02, Bla04, McG98c]. Tkinter [Ano00b, Wei06b, Gra00]. TM [Jac02]. Today [Bar21, LS98]. Toe [Sa’23]. tokamak [BYWW23]. tomographic [MMG19]. tomography [BYL^{+}21, LYX^{+}18]. tongue [GHT06]. Too [RAH^{+}01]. Tool [BLN^{+}21, GMKRS21, HYG16, MH18, VBB18, Bea96, BBHB22, Con95, DPH16, DRFB^{+}23, FL15, GP22, GPEM23, GYWG23, JM20, Law15, LYZ^{+}22, Lot14b, LFT21, LS17, MB17, MWM20, NS22, OBWM23, PMBF17, RRRGVD19, SLB^{+}23, SHTE23, SABEh20, SH19b, Ver22a, YKKD19, You08]. Toolbox [GRKN^{+}19, LWV20, RKR21, SLGB14, Sta21, BHJ^{+}18, CV22, GEH19, HPT^{+}16, HFF^{+}17, KG23, KMLP^{+}23, MMG19, RSPJ21, Sch21, WPM^{+}20, XGW23]. Toolchain [ACZ^{+}22]. Toolchest [Ang99]. Toolkit [CK00, HC16, ISMA18, Jac02, Rem01a, Tal00, WBR^{+}22, War18, BVV22, BMZ^{+}18, DCRF23, LRL23, MZL^{+}23, MD21, Mir18, RMM23, Ros18, SML06, TCE21, TQGE23, TBJ18, Wie18, WMA^{+}22, ZWCQ22, ASAA20]. toolkits [LS98]. Tools [Ano00c, Ano01b, Cas17, DGMP09, LeV09, Ano18, CFW17, Gar09, Gun14, Meh15, SW14, Wes15, ESM98, LDAL20, Zie19b]. TOOLS-23 [ESM98]. toolset [LC11]. Top [CG23, Ano20, Cas20, CFSK14]. top-again [Cas20]. Topics [LT03, RRRGVD19]. Topography [SMF^{+}23]. TOPSIS [YKKD19]. torcpy [HBS^{+}20]. toroidal [CMS22]. Toronto [USE96]. total [Swe14]. Tour [Ngu08b, van98a]. Tracing [MD15, ABF13, BBH^{+}15, HS12, LFT09, PFH^{+}16]. TRACK [SP23]. Tracking [ML16a, Ver22a]. tradeoffs [AJJF14, IOC^{+}12]. train [MK16]. Training [Wal07]. trajectories [BFT20, LRL23]. trajectory [HHB^{+}22]. trans [NL23]. trans-dimensional [NL23]. Transactional [Ano01b, BRM10, Tab10]. transcendent [Ros08]. transcranial [XGW23]. transfer [MPMC21, NO23, SAA18]. transform [MTS^{+}18, MBA19]. transformation [Vai09, Vai14]. transformations [LGS10, MRT^{+}22, Tid07]. transforms [WSK22]. transient [MPMC21]. Transition [KBLJ18, LKK23, MP19a]. translating [Men09]. translation [AJJF14, GS12]. Translational [BCE^{+}22]. Transparent [AFL23]. transpilation [MCF^{+}22]. Transpiling [LH20].

Transplant2Mongo [HW19].

Transplantation [HW19]. transport [ABGD^{+}20, KPK^{+}17, SP23, WAN^{+}22]. transportation [Boe23]. trapped [BFT20]. TRAPping [Ern99]. traveltime [Gir21]. Traversal [BSSz^{+}20]. treatment [SJL18]. tree [Mos20a, NS22]. Trees [SN12, RBV16]. TreeSim [NS22]. TreeSwift [Mos20a]. trend [RRRGVD19]. Triangle [PKDM21]. triangles [LJ23]. Trilinos [Spo12]. Trivial [CASA22]. Tromsø [Eur91]. Trust [TAD23]. ttcrpy [Gir21]. Tubular [ISMA18]. Tucson [Ano97d]. Tuner [FLR22]. Tuning [BII^{+}20, BE20]. tunneling [SZNW23]. Tuplex [SK19]. TurboGears [MBW07, RDS07]. TurboPy [RGS^{+}21]. turbulent [ML16b, ML18]. turn [Law15]. Tutorial [Wat95, vRD05b, Hig03, KD99, SMM^{+}22, Sto13, van95d, vRDK99]. Tutorials [AGR24]. TV [Wal07]. Twisted [FL06, Kin05]. Twitter [STS^{+}18]. Two [Ros08, LS17]. two-loop [LS17]. Type [KCVM22, LD19, Men09, RF16, VSS17]. Type-Related [KCVM22]. type-safe [Men09]. typed [BDT13, CEI^{+}12, IOC^{+}12]. types [Ren23, VSS17]. typesetter [Thu22]. typing [FhDAF09, Sie17, TSD^{+}12, VKSB15]. typy [OA17].

U.P [Lee17]. u.v.m [EK08, The09].

Ubiquitous [AJH^{+}12]. Übungen [Wei06a]. udkm1Dsim [Sch21]. Ugly [Bea09a]. UI [Ull15]. UI/UX [Ull15]. UIST [ACM97]. UK [Smi17]. UKCensusAPI [Smi17]. ultrafast [Sch21, SLB^{+}23]. ultrashort [MD22a]. Ultrasound [RDBC23]. umfassende [EK08, EK09]. Umgang [DF00]. Umsteiger [The09]. uncertain [McK16]. Uncertainty [TSG^{+}23, FL15, FPN^{+}22, KXK^{+}23, KMK^{+}21, OGA^{+}20b, OGA^{+}20a, WPM^{+}20, WKS22]. uncover [Lot14b]. undergraduate [MCGK19, Myr15]. understand [Wei15]. understandability [ABCV21]. Understanding [Gut16, Orb18, BM15, LFYX23]. Underwater [GRKN^{+}19]. Undo [Zuk97]. Unearthing [SSS22]. unfolding [BHA18]. Unicode [Swi09, Tau16]. Unification [AAC^{+}98]. Unified [Kra22]. uniform [MFQ^{+}21, PWFM17]. uniformly [GEH19]. uniqueness [SGZ23]. unit [Hac23a, LC11, LKF23, Sal14]. unit-testing [LC11]. units [GZT^{+}18]. Univariate [Sha24]. universal [SKS23, SMD23]. University [Gve09, Liu22]. Unix [DF00, Eur91, SML06, GJ08, Ngu08e]. unleash [Cox14, Gal14, Hia15, MH15, Phi15]. Unleashed [MGS07]. unlock [Gal14, RO15a]. Unparametrized [KSB12]. Unsupervised [AVS20]. Untapped [Bru09]. Untriviality [CASA22]. unweighted [GH18, GHN19]. unyt [GZT^{+}18]. Upcrossing [Pat16]. Update [Ano00c, DF21a, DFC18, DPF^{+}21, Kuc97]. updated [KMLP^{+}23]. Updates [Ano00c, Ano01b, HPT^{+}16, LRvE17, PHH^{+}12, Yes15]. uPy [AJH^{+}12]. UQpy [OGA^{+}20b, OGA^{+}20a, TSG^{+}23]. USA [CLT20, HK95, USE00]. Usage [MSS^{+}19, ZMD21]. Use [How06, KFV20, SJL18, Bea96, HPH12, PAB^{+}97, Pop10, Sah15, Tom15]. use-case [HPH12]. used [Wei15]. useful [Dah18a]. User [ACM97, Ano01a, Cha01, DF00, GRG21, SJL18, BCM21, GCBDDBF23, GPP^{+}21, LYZ^{+}22, Mac91, Mac92a, MTS^{+}18]. user-friendly [BCM21, GPP^{+}21, LYZ^{+}22]. Using [AMGM20, APS10, Bie97, Bis01, BCG05, CL06, CSRV13, Cot03, DD15, Est01, GAS^{+}16, Gut13, Gut16, HLSØ06, HAB^{+}20, HS13, IG19, Jan10, KJ15, LM03, Liu22, Lut99, MSL^{+}07, OOM^{+}23, Orb18, Osa17, PH23, RM19, SBC^{+}17, Sha24, SJK^{+}21, SGZ23, TT21, Vir16, XMW10, Ada14, AG23, Ber13, BS21, BJM23, BWMS22, BB17, BII^{+}20, CV15, CKK^{+}13, CBB14, Cox14, Dan18a, DRFB^{+}23, FMPS17, Gar09, GFB^{+}14, GNPP23, Gup15, Hia15, HCPF95, JS24, JPJ^{+}23, Jos15, Kar14, Lan13, Len15, Lot14b, Lot15b, MS15, MCF^{+}22, Meh15, MH15, Myr15, Nel15, Nil07b, OLRLB21, Pal14, PP23, Pay11, Per14b, PFLG21, RDS07, RJAL09, RZ09, RK19, SCAK^{+}19, SL21, Sha03, SP23, SZW^{+}22, SMM^{+}22, Tab10, TTHZ23, Ull15, Vel23, Wes15, Wil05, Zac15]. Utilities [RKR21, Ras18]. Utility [CHG^{+}16, DBdFdSR21, Ano12]. utilize [Bah15, SW15]. utilizing [LC11]. UX [Ull15].

V [Orb18, MMEH08]. V2X [SYGY21]. v4.1 [TSG^{+}23]. Validation [Pre03]. valued [SABEh20]. values [MTS^{+}18, SL21]. VaMpy [DB17]. VanderPlas [Ish17]. Variability [CRL21]. variable [AAYK23, RGP^{+}22]. Variables [KJ14, Wit15, GKK^{+}23]. Variational [Ble20]. Vascular [ISMA18]. VASP [DCRF23]. VCF [HB16]. VCFPy [HB16]. vectorial [CMS22, Hac23a]. vectorized [MZL^{+}23]. vegetation [CVV23]. Velocimetry [ML16a]. ventilator [MBP^{+}22]. verbosity [ABCV21]. Vererbung [Wei06a]. Verification [JPOB20]. versatile [BVV22, BKMY03, BKMY04]. Version [BWS23, BTDT23, Him23, HQF^{+}20, SKS23, TBD23, ZRK22, KMLP^{+}23]. Versionen [DF00]. vessel [LRL23]. via [CR12, LKK23, NM22, TAD23, TEG18, YHX22]. Video [Wal07, RMM23, SGPHD^{+}17]. Video-Training [Wal07]. vielen [The09]. Viewer [Ano00c]. Vine [YH21]. Violent [O’C13, Ano13]. Virtual [MRG18, UVa95, MRG17, PFH^{+}16, BE20]. vision [Jos15, MH15]. viskillz [TTHZ23]. viskillz-blender [TTHZ23]. VisTrails [TGEA09]. Visual [Ano01b, Feh02, Vai09, Vai14, YPB16, SW14]. Visualization [APS10, IG19, Liu22, XL18, Ada14, DDK19, GYWG23, HOB23, MTPHH18, Mül23, NS22, Ros13, Ros14, Sal18, SML06, Vai09, Vai14]. visualize [MB17, SML06]. Visualizing [SMD23, TQGE23]. vital [RO15a]. VizGen [YPB16]. voltages [JM20]. vorbereiten [DF00]. voxel [LDAL20]. voxel-based [LDAL20]. VPython [SDS00]. VR [UVa95]. vs [MD15, Ram18b, WX18]. VTAnDeM [TQGE23]. VTK [SML06]. vulnerabilities [ACS23]. vulnerability [ESM24]. Vulnerable [BLS^{+}23].

w [Lip21]. Wacky [RAH^{+}01]. waLBerla [BSG^{+}16]. Want [Wil02]. wanted [Ano20]. warping [EAMSR^{+}23]. Warps [Wil01]. Washington [CI96]. WatchMan [BB13]. Water [LVH^{+}18]. wave [MAC^{+}21, Wet20, AYI^{+}24]. waveform [MGFO23]. Waveforms [TLR21]. wavefunction [HPT^{+}16]. wavefunctions [CFSK14]. waves [CFPS23]. way [Tos09]. ways [Sla15]. Wearable [Cas17]. Weather [FLS^{+}20, HC16, HKGvS21]. Web [EK08, Mit15, Seg07, Buc15, CS21, CV15, Dal07, Gar09, Gup15, HF06, KXK^{+}23, Nai14, Pip15, sR11, Rus14, Swi09, Ang00a, Ano00c, CB96, FBC09, HGMC^{+}97, Hig03, HKM08, KL97, MBW07, RDS07, Sal18, TSC02]. web-based [CS21, KXK^{+}23, Ang00a]. Web-Programmierung [EK08]. WebWriter [CB96]. Weight [Hig22]. weighted [GH18, GHN19]. weighting [BW22, BWS23]. WEKA [BHF16]. WekaPyScript [BHF16]. well [Cha15, LRvE17]. well-done [LRvE17]. wheelchair [GCBDDBF23]. Where [Wat95]. Which [Sit02]. Whisker [HYG16]. Who [STS^{+}18, Wat95]. whole [Ano98e]. wide [MTPHH18, Sar14]. Widgets [Tro96b]. WIMP [JKST22]. WimPyDD [JKST22]. Win32 [HR00, Ano00a]. Wind [JCL^{+}20, RC18]. window [Rem01b]. Windows [DF00, SML06, Ano00c, HR00]. windrose [RC18, RC18]. WIP [MDRN18]. Wireless [Ago01, PH23]. wish [Ano20]. Withdrawals [LVH^{+}18]. within [TGEA09, WZZ^{+}23]. Without [BUS21, BRM10]. Wits [HRS06, SG07, SL08]. WordPress [BP17]. work [Boe17]. workbench [Hos14]. Workbook [MMP^{+}22]. Workflow [IG19]. workflows [TBA^{+}17]. Working [HC16, Ano18, CFW17, RG18]. Workshop [Ano97d, CI96, Ass96]. workspaces [GG21]. World [CG23, Sav23, VCR17, BCRS15, GHT08, Gup15, Jos15, Jos16, Kar14, Lot14a, Mad15, McK16, Sar14, VSS17, WPH^{+}22]. Worlds [Hug97]. Worrying [Wit21]. WPTherml [VEV^{+}19]. Wrangling [KJ15]. Wrapper [LD07, Tan01, GWM18]. write [Sla15]. Writing [Mit00, Van97b, VMFG17, HB16, LFT09, VAP^{+}21]. written [CWLG^{+}21, Dah18a, Huc18]. WSG [Gar09]. wxPython [Tal00].

X [CG17, Tro96a]. X-ray [CG17]. XAI [DSB^{+}23]. xarray [HH17]. XDesign [CG17]. XDK [Ano00c]. XERIS [Wai21]. XERIS/APEX [Wai21]. XFree86 [DF00]. XFree86/23.3.6 [DF00]. XML [Wei06a, Wei06b, JD02, McG98b, McG98a, McG00, Por03, Tid07, VB08]. XML-RPC [Por03]. xSLHA [Sta19]. XSLT [Tid07]. xturtle [Lin06b]. xv [Sau23]. xvi [Lip22].

year [HRS06]. Years

[Bea12, DF21a, Sev15a, Ros08].

YesSoftware [Ano01b]. yex [Thu22]. Young [Dav21]. ython [VMRFC23].

Zato [Suc13]. Zeljko [Ish17]. Zelle [Jan10]. zfit [ENCS20]. Zheng [Sa’23]. Zope [HAE00, Pel01]. Zugriff [Wei06b].

References

Ascension:2022:BPM

[AAB22] Alex M. Ascensión and Marcos J. Araúzo-Bravo. BigMPI4py: Python module for parallelization of big data objects discloses germ layer specific DNA demethylation motifs. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 19 (3):1507–1522, May 2022. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic). URL https:/

/dl.acm.org/doi/10.1109/
TCBB.2020.3043979.

Achtzehnter:1998:ILU

[AAC^{+}98] Joachim Achtzehnter, Judy Anderson, Antony Courtney, Doug Cutting, Mark Davidson, Jeanette Figueroa, Ken Fishkin, Scott Hassan, Rob Head, Chris Jacobi, Bill Janssen, Swen Johnson, Dan Larner, Bill Nell, Denis Severson, Bridget Spitznagel, Mike Spreitzer, Mark Stefik, Martin von Löwis, Farrell Wymore, and Rick Yardumian. Inter-language unification. Technical Report ??, Xerox Palo Alto Research Center, Palo Alto, CA, USA, May 12, 1998. URL ftp:/

/ftp.parc.xerox.com/pub/
ilu/ilu.html. The ILU sys-

tem can be used from C++, Common Lisp, Guile, Java, Modula-3, Perl, Python, Scheme, and Standard C.

Aman:2023:ADC

[AAYK23] Hirohisa Aman, Sousuke Amasaki, Tomoyuki Yokogawa, and Minoru Kawahara. An automated detection of confusing variable pairs with highly similar compound names in Java and Python programs. Empirical Software Engineering, 28 (5):??, September 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10664-023-10339-2.

Arnold:1997:HDO

[ABC97] David Arnold, Andy Bond, and Martin Chilvers. Hector: Distributed objects in Python. Dr. Dobb’s Sourcebook, ??(??):??, January/ February 1997. ISSN 1077-

                                        9019. URL http://www.
                                        ddj.com/ddsbk/1997/1997_
                                        01/arno.htm.

Akhilesh:2022:APT

[ABCC22] Rohit Akhilesh, Oliver Bills, Naveen Chilamkurti, and Mohammad Jabed Morshed Chowdhury. Automated penetration testing framework for smart-home-based IoT devices. Future Internet, 14(10):276, September 27, 2022. CODEN ???? ISSN 1999-5903. URL https:/

                                        /www.mdpi.com/1999-5903/
                                        14/10/276.

Ardito:2021:ERC

[ABCV21] Luca Ardito, Luca Barbato, Riccardo Coppola, and Michele Valsesia. Evaluation of Rust code verbosity, understandability and complexity. PeerJ Computer Science, 7:e406:1–e406:33, February 2021. ISSN 2167-8359.

Ardo:2013:LAO

[ABF13] Håkan Ardö, Carl Friedrich Bolz, and Maciej Fijałkowski. Loop-aware optimizations in PyPy’s tracing JIT. ACM SIGPLAN Notices, 48(2):63– 72, February 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

AlAtoum:2020:ETG

[ABGD^{+}20] B. Al Atoum, S. F. Biagi, D. González-Díaz, B. J. P. Jones, and A. D. McDonald. Electron transport in gaseous detectors with a Pythonbased Monte Carlo simulation code. Computer Physics Communications, 254(??): Article 107357, September 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520301533.

Alenaizan:2020:PIR

[ABS20] Asem Alenaizan, Lori A.

Burns, and C. David Sherrill. Python implementation of the restrained electrostatic potential charge model. International Journal of Quantum Chemistry, 120(2):e26035:1– e26035:??, January 15, 2020. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

ACM:1992:PAC

[ACM92] ACM, editor. Proceedings of the 1992 ACM Conference on Lisp and Functional Programming: papers presented at the conference, San Francisco, California, June 22–24, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-483-X, 0-89791-481-3. LCCN QA76.73.L23A26 1992. Also published as LISP Pointers, vol. V, no. 1, January-March, 1992. ACM order no. 552920.

ACM:1997:PAS

[ACM97] ACM, editor. Proceedings of the ACM Symposium on User Interface Software and Technology. 10th Annual Symposium. UIST ’97: Banff, Alberta, Canada, 14–17 October 1997. ACM Press, New York, NY 10036, USA, 1997. ISBN 0-89791-881-9. LCCN ????

Alfadel:2023:EAS

[ACS23] Mahmoud Alfadel, Diego Elias Costa, and Emad Shihab. Empirical analysis of security vulnerabilities in Python packages. Empirical Software Engineering, 28(3):??, May 2023. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10278-4.

Agostini:2022:BPS

[ACZ^{+}22] Nicolas Bohm Agostini, Serena Curzel, Jeff Jun Zhang, Ankur Limaye, Cheng Tan, Vinay Amatya, Marco Minutoli, Vito Giovanni Castellana, Joseph Manzano, David Brooks, Gu-Yeon Wei, and Antonino Tumeo. Bridging Python to silicon: The SODA toolchain. IEEE Micro, 42(5):78–88, September/October 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Adams:2014:LPD

[Ada14]  Chad Adams.  Learning

Python data visualization: master how to build dynamic HTML5-ready SVG charts using Python and the pygal library. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78355-333-2, 1-78355-334-0 (e-book). iii + 195 pp. LCCN QA76.73.P98. URL http://

proquest.tech.safaribooksonline.
de/9781783553334.

Aranega:2021:RGT

[ADP21] Vincent Aranega, Julien Delplanque, and Guillermo Polito. Rotten green tests in Java, Pharo and Python.

Empirical Software Engineering, 26(6):??, November 2021. CODEN ES-ENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10664-021-10016-2.

Ahrari:2022:PPF

[AES^{+}22] Ali Ahrari, Saber Elsayed, Ruhul Sarker, Daryl Essam, and Carlos A. Coello Coello. PyDDRBG: a Python framework for benchmarking and evaluating static and dynamic multimodal optimization methods. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021001850.

Arjona:2023:TSE

[AFL23] Aitor Arjona, Gerard Finol, and Pedro García López. Transparent serverless execution of Python multiprocessing applications. Future Generation Computer Systems, 140(??):436–449, March 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X22003612.

Ariza:2023:RLF

[AG23] Jonathan Álvarez Ariza and Christian Nomesqui Galvis. RaspyControl Lab: a fully open-source and real-time remote laboratory for education in automatic control systems using Raspberry Pi and Python. HardwareX, 13:??, March 2023. CODEN ???? ISSN 2468-0672. URL http:

//www.sciencedirect.com/
science/article/pii/S2468067223000032.

Aguado:2023:QPP

[AGMFGE23] Daniel Gómez Aguado, Vicent Gimeno, Julio José Moyano-Fernández, and Juan Carlos Garcia-Escartin. QOptCraft: a Python package for the design and study of linear optical quantum systems. Computer Physics Communications, 282(??): Article 108511, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522002302.

Agoren:2001:KKL

[Ago01] Izzet Agoren. Kernel korner: Linux teleconferencing: Improving the wireless network. Linux journal, 85: 24, 26, 28, 30, May 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). See erratum [Ano01a].

Arya:2024:PSS

[AGR24] Deeksha M. Arya, Jin L. C.

Guo, and Martin P. Robillard. Properties and styles of software technology tutorials. IEEE Transactions on Software Engineering, 50(2): 159–172, February 2024. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Autin:2012:UUC

[AJH^{+}12] Ludovic Autin, Graham Johnson, Johan Hake, Arthur Olson, and Michel Sanner. uPy: a ubiquitous CG Python API with biological-modeling applications. IEEE Computer Graphics and Applications, 32 (5):50–61, September/October 2012. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Arabas:2014:FTB

[AJJF14] Sylwester Arabas, Dorota Jarecka, Anna Jaruga, and Maciej Fijalkowski. Formula translation in Blitz++, NumPy and modern Fortran: A case study of the language choice tradeoffs. Scientific Programming, 22(3):201–222, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

AlKadi:2018:GPC

[AJYH18] Muhammed Al Kadi, Benedikt Janssen, Jones Yudi, and Michael Huebner. Generalpurpose computing with soft GPUs on FPGAs. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 11(1):5:1– 5:??, March 2018. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic).

Alchin:2009:PD

[Alc09] Marty Alchin. Pro Django.

Springer eBook Collection Professional and Applied Computing [Dig. Serial]; Springer-12059 [Dig. Serial]. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1048-6. ???? pp. LCCN ????

Alnaes:2010:ESC

[AM10] Martin Sandve Alnæs and Kent-André Mardal. On the efficiency of symbolic computations combined with code generation for finite element methods. ACM Transactions on Mathematical Software, 37 (1):6:1–6:26, January 2010. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Abi-Mansour:2019:POO

[AM19] Andrew Abi-Mansour. Py-Gran: an object-oriented library for DEM simulation and analysis. SoftwareX, 9 (??):168–174, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301080.

Arrigoni:2021:SPP

[AM21] Marco Arrigoni and Georg K. H. Madsen. Spinney: Post-processing of first-principles calculations of point defects in semiconductors with Python. Computer Physics Communications, 264(??): Article 107946, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521000709.

Augier:2019:FPO

[AMB19] Pierre Augier, Ashwin Vishnu Mohanan, and Cyrille Bonamy. FluidDyn: a Python opensource framework for research and teaching in fluid dynamics by simulations, experiments and data processing. Journal of Open Research Software, 7(1):9–??, April 01, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.237/.

Andreadi:2021:PSS

[AME^{+}21] Nikolai Andreadi, Artem Mitrofanov, Artem Eliseev, Petr Matveev, Stepan Kalmykov, and Vladimir Petrov. PyRad: A software shell for simulating radiolysis with Qball package. Journal of Computational Chemistry, 42(13): 944–950, May 15, 2021. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Aguirre-Mesa:2020:MLC

[AMGM20] Andres M. Aguirre-Mesa, Manuel J. Garcia, and Harry Millwater. MultiZ: a library for computation of high-order derivatives using multicomplex or multidual numbers. ACM Transactions on Mathematical Software, 46(3):23:1– 23:30, September 2020. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3378538.

Aumasson:2014:HFB

[AMPH14]  Jean-Philippe  Aumasson,

Willi Meier, Raphael C.-W. Phan, and Luca Henzen. The Hash Function BLAKE. Information Security and Cryptography. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44756-8 (print), 3-662-44757-6 (e-book). ISSN 1619-7100 (print), 2197-845X (electronic). xviii + 228 + 18 pp. LCCN QA76.9.H36 A96 2014.

Angell:1999:PTE

[Ang99] Kirby W. Angell. Programmer’s toolchest: Examining JPython: a Java test engine puts Python to the test. Dr. Dobb’s Journal of Software Tools, 24(4):78, 81–83, April 1999. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/1999/
1999_04/jpython.txt;   http:
//www.ddj.com/ftp/1999/
1999_04/jpython.zip.

Ang:2000:WBL

[Ang00a] Cheng-Chai Ang. A Webbased lunch ordering system. Linux journal, 79:176– ??, November 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Angell:2000:PSPa

[Ang00b] Kirby W. Angell. Python server pages: Part 1. Dr. Dobb’s Journal of Software Tools, 25(1):44, 46– 47, 50, January 2000. CO-DEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2000/
                                        2000_01/psp1.txt;   http:
                                        //www.ddj.com/ftp/2000/
                                        2000_01/psp100.zip.

Angell:2000:PSPb

[Ang00c] Kirby W. Angell. Python server pages: Part II. Dr. Dobb’s Journal of Software Tools, 25(2):54, 57–61, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:

                                        //www.ddj.com/ftp/2000/
                                        2000_02/psp2.txt.

Anonymous:1995:NIP

[Ano95] Anonymous. NIST investigates Python programming language. Journal of research of the National Institute of Standards and Technology, 100(1):101, January/February 1995. CO-DEN JRITEF. ISSN 1044-677X (print), 2165-7254 (electronic). URL

http://nvl.nist.gov/pub/
nistpubs/jres/jrescont.
pdf.

Anonymous:1997:BRPe

[Ano97a] Anonymous. Book review:

Programming Python: By Mark Lutz. O’Reilly, Sebastopol, CA. (1996). 880 pages. $44.95. Computers and Mathematics with Applications, 33(5):132, March 1997. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122197829525.

Anonymous:1997:BRI

[Ano97b]  Anonymous.  Book re-

view: Internet Programming with Python. Linux journal, 42:??, October 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//www.linuxjournal.com/
2152.html.

Anonymous:1997:PIP

[Ano97c] Anonymous, editor. Proceedings of the 6th International Python Conference: October 14–17, 1997, San Jose, California. Corporation for National Research Initiatives, 1895 Preston White Drive, Suite 100, Reston, VA 20191, USA, 1997. ISBN ???? LCCN ???? Anonymous:1997:PWM

[Ano97d] Anonymous, editor. Proceedings of the Workshop on Management of Semi-Structured Data: Tucson, Arizona, May 16, 1997. ????, ????, 1997. ISBN ???? LCCN ????

Anonymous:1998:P

[Ano98a] Anonymous. Python. Web Review, ??(??):??, April 10, 1998. URL http:

                                        //webreview.com/wr/pub/
                                        freeware/python.html.

Anonymous:1998:PDA

[Ano98b] Anonymous. The Python DB-API. Linux journal, ?? (??):??, April 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

                                        tronic). URL http://www.
                                        ssc.com/lj/.

Anonymous:1998:PEL

[Ano98c] Anonymous. Python extensions: LLNL-distributed extensions. Technical Report UCRL-128569, Lawrence Livermore National Laboratory, Livermore, CA, USA, 1998.

                                        URL http://xfiles.llnl.
                                        gov/python.htm. Includes

Numerical Python, Python/ Pact-PDB interface, Py-Graphics, PyHistory, RNG (random number generator), CXX Objects (Python extensions in C++).

Anonymous:1998:SLG

[Ano98d] Anonymous. Scripting languages go prime time. Software Development magazine, ??(??):??, April 1998. URL

http://www.sdmagazine.com.

Discusses Python, Perl, and Tcl.

Anonymous:1998:WPF

[Ano98e]  Anonymous.  The whole

Python FAQ. Technical report, Corporation for National Research Initiatives, 1895 Preston White Drive, Suite 100, Reston, VA 20191, USA, 1998. URL http:/

/www.python.org/doc/FAQ.

html. For Python Release 1.5.1.

Anonymous:2000:BRPf

[Ano00a]  Anonymous.  Book re-

view: Python programming on Win32: By Mark Hammond and Andy Robinson. O’Reilly, Sebastopol, CA. (2000). 652 pages. $34.95. Computers and Mathematics with Applications, 40(2– 3):418, July/August 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0898122100901974.

Anonymous:2000:BRLb

[Ano00b] Anonymous. Book reviews:

Linux Programmer’s Reference Second Edition by Ibrahim F. Haddad; Python and Tkinter Programming by Phil Hughes; sendmail for Linux by Russell J. T. Dyer. Linux journal, 77:??, September 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue77/3861.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue77/3989.html;   http:
//noframes.linuxjournal.
com/lj-issues/issue77/4184.
html.

Anonymous:2000:POR

[Ano00c] Anonymous. Products: Oracle releases XDK update; Starbase’s code editing system; Arc Second’s palm PC CAD viewer; Minolta’s network document server for Windows 2000; Borland’s Java development tools for Palm OS; Rational’s code management tools; Blaxxun Interactive’s Web communications platform tools; Informix Software’s Linux database engine; ActiveState updates free Python distribution; KDE 2.0 released. Computer, 33(12): 144–146, December 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2000/pdf/rz144.
pdf.

Anonymous:2001:EIA

[Ano01a] Anonymous. Errata: Izzet Agoren’s Kernel Corner, May 2001, Mitch Chapman’s “Create User Interfaces with Glade” (July 2001). Linux journal, 89:6, September 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). See [Ago01, Cha01].

Anonymous:2001:PPS

[Ano01b] Anonymous. Products: ProxySource’s software design and collaboration application; YesSoftware’s code generation application; Persistence Software’s transactional application server; Instantiation’s Java productivity tools; JCanvas visual rapid application IDE; theKompany.com’s Python development environment; Neu-Vis updates E-business visual modeling tools; LegacyJ’s Java-compliant Cobol compiler. Computer, 34 (3):108–109, March 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//dlib.computer.org/co/
books/co2001/pdf/r3108.
pdf.

Anonymous:2002:CPF

[Ano02]  Anonymous.  Correction:

“Python and Finite Elements”. Dr. Dobb’s Journal of Software Tools, 27(4– 4):10, April 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

See [Ple02].

Anonymous:2012:PSR

[Ano12] Anonymous. Python scripts as a replacement for bash utility scripts. Linux journal, 2012(223):1:1–1:??, November 2012. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2013:BRV

[Ano13] Anonymous. Book review: Violent Python, by T. J. O’Connor. Syngress. ISBN 978-1-59749-957-6. Network Security, 2013(6):4, June 2013. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1353485813700679.

Anonymous:2014:BRP

[Ano14] Anonymous. Book review: Python Forensics, Chet Hosmer, Syngress. ISBN 978-0-12-418676-7 (print), 978-0-12-418683-5 (e-book). Network Security, 2014(9):4, September 2014. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S135348581470087X.

Anonymous:2015:BRB

[Ano15a] Anonymous. Book review: Black Hat Python, Justin Seitz. No Starch Press. ISBN 978-1-59327-590-7. Network Security, 2015(4):4, April 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1353485815300258.

Anonymous:2015:BRD

[Ano15b]  Anonymous.  Book re-

view: Doing Math With Python, Amit Saha. No Starch Press. ISBN 978-1-59327-640-9. Network Security, 2015(10):4, October 2015. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S135348581530088X.

Anonymous:2018:PCC

[Ano18]  Anonymous.  ChebTools:

C++11 (and Python) tools for working with Chebyshev expansions. Journal of Open Source Software, 3 (22):569:1–569:3, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00569.

Anonymous:2019:PSS

[Ano19] Anonymous. Profile: Scott Shawcroft: This developer is squeezing Python into microcontrollers. IEEE Spectrum, 56(4):16, April 2019. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2020:SWS

[Ano20] Anonymous. Skills wanted:

Sql, Java, Python, and AWS top employers’ wish lists — [careers]. IEEE Spectrum, 57 (1):59, January 2020. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:2021:RCR

[Ano21] Anonymous. Replication of computational results report for “Doping Tests for Cyber-Physical Systems”. ACM Transactions on Modeling and Computer Simulation, 31(3):17:1–17:2, July 2021. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3459667.

Anonymous:20xx:PP

[Anoxxa] Anonymous. The Python papers. Web site, 20xx. ISSN 1834-3147. URL http:

                                        //pythonpapers.org/tpp.
                                        html.

Anonymous:20xx:PPA

[Anoxxb] Anonymous. Python papers anthology. Web site, 20xx.

                                        URL http://pythonpapers.
                                        org/.

Anonymous:20xx:PPM

[Anoxxc] Anonymous. Python papers monograph. Web site, 20xx.

                                        URL http://pythonpapers.
                                        org/tppm.html.

Anonymous:20xx:PPS

[Anoxxd] Anonymous. Python papers source codes. Web site, 20xx. ISSN 1836-621X.

                                        URL http://pythonpapers.
                                        org/tppsc.html.

Antao:2015:BPC

[Ant15] Tiago Antao. Bioinformatics with Python cookbook. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78217-511-3, 1-78355-865-2 (e-book). 306 (est.) pp. LCCN QA76.73.P98 .A583 2015.

Anderson:2010:UPS

[APS10] Erik W. Anderson, Gilbert A.

Preston, and Claudio T. Silva. Using Python for signal processing and visualization. Computing in Science and Engineering, 12(4):90– 95, July/August 2010. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Arbuckle:2014:LPT

[Arb14] Daniel Arbuckle. Learning Python testing: a straightforward and easy approach to testing your Python projects. Community experience distilled. Packt Publishing, Birmingham, UK, second edition, 2014. ISBN 1-78355-321-9, 1-78355-322-7. v + 180 pp. LCCN QA76.73.P98 A728 2014.

Ahmadzadeh:2020:MDT

[ASAA20] Azim Ahmadzadeh, Kankana Sinha, Berkay Aydin, and Rafal A. Angryk. MVTS-Data Toolkit: a Python package for preprocessing multivariate time series data. SoftwareX, 12(??):Article 100518, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020300157.

USENIX:1996:ATT

[Ass96] USENIX Association, editor. 4th Annual Tcl/Tk Workshop ’96, July 10–13, 1996. Monterey, CA. USENIX, Berkeley, CA, USA, July 10–13, 1996. ISBN 1-880446-78-

2. LCCN QA76.73.T44 T44 1996.

Ahsen:2020:RPS

[AVS20] Mehmet Eren Ahsen, Robert Vogel, and Gustavo A. Stolovitzky. R/PY-SUMMA: an R/Python package for unsupervised ensemble learning for binary classification problems in bioinformatics. Journal of Computational Biology, 27(9):1337– 1340, September 2020. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2019.0348;
https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2019.0348.

Andersson:2006:PSN

[AW06] Mats Andersson and Robert Wedin. Python scripting for network management: PyMIP–TeMIP made simple. Examensarbete, Högskoleingenjörsprogrammet. Luleå tekniska universitet, Skellefteå, Sweden, 2006. ISSN 1404-5494. URL http:/

/epubl.ltu.se/1404-5494/
2006/043/;   http://epubl.
ltu.se/1404-5494/2006/043/
LTU-HIP-EX-06043-SE.pdf.

Aakerblom:2016:MPP

[ÅW16] Beatrice Åkerblom and Tobias Wrigstad. Measuring polymorphism in Python programs. ACM SIGPLAN Notices, 51(2):114–128, February 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aksman:2021:PPI

[AWO^{+}21] Leon M. Aksman, Peter A. Wijeratne, Neil P. Oxtoby, Arman Eshaghi, Cameron Shand, Andre Altmann, Daniel C. Alexander, and Alexandra L. Young. pySuStaIn: a Python implementation of the Subtype and Stage Inference algorithm. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021001096.

Ayars:2014:FPB

[Aya14]  Eric Ayars.  Finally, a

Python-based computational physics text. Computing in Science and Engineering, 16(1):6–7, January/February 2014. CODEN CSENFA. ISSN 1521-9615.

Aoyama:2024:HWP

[AYI^{+}24] Tatsumi Aoyama, Kazuyoshi Yoshimi, Kota Ido, Yuichi Motoyama, Taiki Kawamura, Takahiro Misawa, Takeo Kato, and Akito Kobayashi. H-wave — a Python package for the Hartree–Fock approximation and the random phase approximation. Computer Physics Communications, 298(??):Article 109087, May 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465524000109.

Abdullahi:2024:DPB

[AZH^{+}24] Asli M. Abdullahi, Jaime Hoefken Zink, Matheus Hostert, Daniele Massaro, and Silvia Pascoli. DarkNews: a Python-based event generator for heavy neutral lepton production in neutrinonucleus scattering. Computer Physics Communications, 297(??):Article 109075, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523004204.

Alon:2018:GPB

[AZLY18] Uri Alon, Meital Zilberstein, Omer Levy, and Eran Yahav. A general path-based representation for predicting program properties. ACM SIG-PLAN Notices, 53(4):404– 419, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Browning:2014:PP

[BA14] J. Burton Browning and Marty Alchin. Pro Python. The expert’s voice in Python. Apress, Berkeley, CA, USA, second edition, 2014. ISBN 1-4842-0335-6, 1-4842-0334-8 (e-book). LCCN QA76.73.P98.

Backer:2007:CPE

[Bäc07]  Arnd Bäcker.  Computa-

tional physics education with Python. Computing in Science and Engineering, 9(3): 30–33, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Badia:2020:SDS

[Bad20] Antonio Badia. SQL for Data Science. Data-Centric Systems and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-57591-8, 3-030-57592-6 (ebook), 3-030-57593-4. ISSN 2197-9723 (print), 2197-974X (electronic). xi + 285 + 16 pp. URL http:

//www.springerlink.com/
content/978-3-030-57592-
2.

Bahgat:2015:PGD

[Bah15]  Karim Bahgat.  Python

geospatial development essentials: utilize Python with open source libraries to build a lightweight, portable, and customizable GIS desktop application. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78217-540-7, 1-78217-441-9 (e-book). 192 (est.) pp. LCCN QA76.73.P98. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781782175407.

Bailly:2008:IPA

[Bai08] Yves Bailly. Initiation à la programmation avec Python et C++. Pearson Education France, Paris, France, 2008. xi + 255 pp. LCCN ????

Baldwin:2012:SPG

[Bal12]  Doug Baldwin.  Special

projects grants awarded. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 44(3):6, July 2012. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Balbaert:2018:JP

[Bal18]  Ivo Balbaert.  Julia 1.0

Programming. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78899-909-

6. iv + 184 pp. LCCN QA76.73.J85 2018. URL

                                        http://international.scholarvox.
                                        com/book/88863229.

Barry:2011:PA

[Bar11] Paul Barry. Python for Android. Linux journal, 2011 (203):1:1–1:??, March 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Barba:2021:PJE

[Bar21]  Lorena A. Barba.  The

Python/Jupyter ecosystem: Today’s problem-solving environment for computational science. Computing in Science and Engineering, 23(3): 5–9, May/June 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bartos:2022:NMI

[Bar22]  Erik Bartos.  Numeri-

cal multidimensional integration with PyMikor. Computer Physics Communications, 270(??):Article 108149, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465521002617.

Bassi:2008:PB

[Bas08]  Sebastian Bassi.  Python

for bioinformatics. Chapman and Hall/CRC mathematical and computational biology series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-929-2 (paperback). ???? pp. LCCN ????

Bassi:2010:PB

[Bas10] Sebastian Bassi. Python for bioinformatics. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-58488-929-2 (paperback). xxv + 584 pp. LCCN QH324.2 .B387 2010.

Bauer:1996:IP

[Bau96] Jeff Bauer. An introduction to Python. Linux journal, 21:??, January 1996. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Baxter:2001:BRC

[Bax01] Michael Baxter. Book reviews: Core Python Programming. Linux journal, 85:100–101, May 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //noframes.linuxjournal.
                                        com/lj-issues/issue85/4564.
                                        html.

Bianchi:2013:WPP

[BB13] Riccardo Maria Bianchi and Renaud Brunelière. Watch-Man project — a Python CASE framework for High Energy Physics data analysis in the LHC era. Journal of Computational Science, 4(5):325–333, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750312000336.

Bonetta:2017:FJF

[BB17] Daniele Bonetta and Matthias Brantner. FAD.js: fast JSON data access using JITbased speculative optimizations. Proceedings of the VLDB Endowment, 10(12): 1778–1789, August 2017. CO-DEN ???? ISSN 2150-8097.

Badalyan:2022:AEC

[BB22a] David Badalyan and Oleg Borisenko. Ansible execution control in Python and Golang for cloud orchestration. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022000826.

Bragagnolo:2022:SPL

[BB22b] Andrea Bragagnolo and Carlo Alberto Barbano. Simplify: a Python library for optimizing pruned neural networks. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001576.

Baumgartner:2002:DIP

[BBB02] Florian Baumgartner, Torsten Braun, and Bharat Bhargava. Design and implementation of a python-based active network platform for network management and control. Lecture Notes in Computer Science, 2546:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/
service/series/0558/bibs/
2546/25460177.htm;   http:
//link.springer.de/link/
service/series/0558/papers/
2546/25460177.pdf.

Boguslawski:2024:PIF

[BBC^{+}24] Katharina Boguslawski, Filip Brzek, Rahul Chakraborty, Kacper Cieślak, Seyedehdelaram Jahani, Aleksandra Leszczyk, Artur Nowak, Emil Sujkowski, Julian Świerczyński, Somayeh Ahmadkhani, Dariusz Kedziera, Maximilian H. Kriebel, Piotr Szymon Żuchowski, and Paweł Tecmer. PyBEST: Improved functionality and enhanced performance. Computer Physics Communications, 297(??):Article 109049, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523003946.

Bauman:2015:PTJ

[BBH^{+}15] Spenser Bauman, Carl Friedrich Bolz, Robert Hirschfeld, Vasily Kirilichev, Tobias Pape, Jeremy G. Siek, and Sam Tobin-Hochstadt. Pycket: a tracing JIT for a functional language. ACM SIG-PLAN Notices, 50(9):22–34, September 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bendarag:2022:PPB

[BBHB22] Abdesadik Bendarag, Jamal Bakkas, Mohamed Hanine, and Omar Boutkhoum.

PyOPAsolver: a Python

based tool for ordinal priority approach operations and normalization. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001443.

Boland:2009:IPD

[BC09] Michael G. Boland and Curtis Clifton. Introducing Py-Lighter: dynamic code highlighter. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):489–493, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Bezanson:2018:JDP

[BCC^{+}18] Jeff Bezanson, Jiahao Chen, Benjamin Chung, Stefan Karpinski, Viral B. Shah, Jan Vitek, and Lionel Zoubritzky. Julia: dynamism and performance reconciled by design. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA):

120:1–120:23, October 2018.

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3276490.

Badia:2022:PIT

[BCE^{+}22] Rosa M. Badia, Javier Conejero, Jorge Ejarque, Daniele Lezzi, and Francesc Lordan. PyCOMPSs as an instrument for translational computer science. Computing in Science and Engineering, 24 (2):79–84, March/April 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Broker:2005:UPL

[BCG05] Oliver Bröker, Oscar Chinellato, and Roman Geus. Using Python for large scale linear algebra applications. Future Generation Computer Systems, 21(6):969–979, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Barrachina:2021:PUF

[BCM21] Sergio Barrachina, Adrián Castelló, and Jose I. Mestre. PyDTNN: A user-friendly and extensible framework for distributed deep learning. The Journal of Supercomputing, 77(9):9971– 9987, September 2021. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s11227-021-03673-z.

Balbaert:2015:GSJ

[BCRS15] Ivo Balbaert, Kevin Colaco, Neeshma Ramakrishnan, and Rashmi Sawant, editors. Getting started with Julia programming: enter the exciting world of Julia, a high-performance language for technical computing. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78328-479-X, 1-78328-480-3 (e-book). 214 pp. LCCN QA297 .B353 2015eb. URL http://

public.ebookcentral.proquest.
com/choice/publicfullrecord.
aspx?p=1973847;   http:/
/site.ebrary.com/id/11025933;
 http://www.vlebooks.com/
vleweb/product/openreader?
id=none\%26isbn=9781783284801.

Barkley:2020:HMP

[BDF^{+}20] S. Barkley, T. G. Dimiduk, J. Fung, D. M. Kaz, V. N. Manoharan, R. McGorty, R. W. Perry, and A. Wang. Holographic microscopy with Python and HoloPy. Computing in Science and Engineering, 22(5):72–82, 2020. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brimicombe:2022:TPT

[BDQ^{+}22] Chloe Brimicombe, Claudia Di Napoli, Tiago Quintino, Florian Pappenberger, Rosalind Cornforth, and Hannah L. Cloke. Thermofeel: a Python thermal comfort indices library. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000176.

Bolz:2013:SSC

[BDT13] Carl Friedrich Bolz, Lukas Diekmann, and Laurence Tratt. Storage strategies for collections in dynamically typed languages. ACM SIG-PLAN Notices, 48(10):167– 182, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOP-SLA ’13 conference proceedings.

Boeira:2020:PPP

[BE20] Emerson Boeira and Diego Eckhard. pyvrft: a Python package for the Virtual Reference Feedback Tuning, a direct data-driven control method. SoftwareX, 11(??): Article 100383, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019302894.

Beazley:1996:SEU

[Bea96] D. M. Beazley. SWIG: an easy to use tool for integrating scripting languages with C and C++. In Association [Ass96], pages 129–139. ISBN 1-880446-78-2. LCCN QA76.73.T44 T44 1996.

Beazley:1998:SAC

[Bea98] David Beazley. SWIG and automated C/C++ scripting extensions. Dr. Dobb’s Journal of Software Tools, 23(2): 30, 32, 34–36, 100, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Beazley:2000:PER

[Bea00] David M. Beazley. Python essential reference. New Riders professional library. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0901-7. xviii + 319 pp. LCCN QA76.73.P98 B43 2000.

Beazley:2001:PER

[Bea01] David M. Beazley. Python essential reference. New Riders Publishing, Carmel, IN, USA, second edition, 2001. ISBN 0-7357-1091-0. xviii + 398 pp. LCCN QA76.73.P98 B43 2001.

Beazley:2006:PER

[Bea06] David M. Beazley. Python essential reference. SAMS Publishing, Indianapolis, IN, USA, third edition, 2006. ISBN 0-672-32862-3 (paperback). xiii + 625 pp. LCCN QA76.73.P98 B43 2006.

Beazley:2009:PGB

[Bea09a] David Beazley. Python 3:

The good, the bad, and the ugly. ;login: the USENIX Association newsletter, 34 (2):??, April 2009. CO-DEN LOGNEM. ISSN 1044-

6397. URL https://www.
usenix.org/publications/
login/april-2009-volume-
34-number-2/python-3-good-
bad-and-ugly.

Beazley:2009:PER

[Bea09b] David M. Beazley. Python essential reference. Addison-Wesley, Reading, MA, USA, fourth edition, 2009. ISBN 0-672-32978-6. xxi + 717 pp. LCCN ????

Beazley:2012:TYP

[Bea12]  David Beazley.  Three

years of Python 3. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2012/three-years-
python-3.

Bennett:2008:PDP

[Ben08] James Bennett. Practical Django projects. The expert’s voice in Web development. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-996-9. xvii + 237 pp. LCCN TK5105.888.B4512; TK5105.888.B4512 2008.

Bernard:2013:RSC

[Ber13] Joey Bernard. Running scientific code using IPython and SciPy. Linux journal, 2013(228):3:1–3:??, April 2013. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Brewer:2018:DDN

[BFM18] Brendon J. Brewer and Daniel Foreman-Mackey. DNest4: Diffusive nested sampling in C++ and Python. Journal of Statistical Software, 86 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v086i07;
https://www.jstatsoft.
org/index.php/jss/article/
view/v086i07/v86i07.pdf.

Bentine:2020:PLP

[BFT20] E. Bentine, C. J. Foot, and D. Trypogeorgos. (py)LIon: a package for simulating trapped ion trajectories. Computer Physics Communications, 253(??):Article 107187, August 2020. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520300369.

Batut:2017:PEP

[BG17] Bérénice Batut and Björn Grüning. ENASearch: a Python library for interacting with ENA’s API. Journal of Open Source Software, 2(18):418:1, October 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00418.

Blackmore:2023:DPP

[BGHC23] Jacob A. Blackmore, Philip D. Gregory, Jeremy M. Hutson, and Simon L. Cor-

nish.  Diatomic-py: a

Python module for calculating the rotational and hyperfine structure of ^{1}Σ molecules. Computer Physics Communications, 282(??): Article 108512, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522002314.

Bell:2012:PSA

[BH12] Nathan Bell and Anil N. Hirani. PyDEC: Software and algorithms for discretization of exterior calculus. ACM Transactions on Mathematical Software, 39(1):3:1– 3:41, November 2012. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bourbeau:2018:PPP

[BHA18] James Bourbeau and Zigfried Hampel-Arias. PyUnfold: a Python package for iterative unfolding. Journal of Open Source Software, 3(26):741:1–741:3, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00741.

Beckham:2016:PWC

[BHF16] Christopher Beckham, Mark Hall, and Eibe Frank. WekaPyScript: Classification, regression, and filter schemes for WEKA implemented in Python. Journal of Open Research Software, 4(1):e33–??, August 08, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.108/.

Borowka:2018:PTN

[BHJ^{+}18] S. Borowka, G. Heinrich, S. Jahn, S. P. Jones, M. Kerner, J. Schlenk, and T. Zirke. pySecDec: a toolbox for the numerical evaluation of multi-scale integrals. Computer Physics Communications, 222(??): 313–326, January 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465517303028.

Brown:2018:PPP

[BHS18] Thomas Brown, Jonas Hörsch, and David Schlachtberger. PyPSA: Python for power system analysis. Journal of Open Research Software, 6 (1):4–??, January 16, 2018. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.188/.

Bielak:1997:UPG

[Bie97] Richie Bielak. Using Python to generate HTML pages. Linux Gazette, 19(??):??, July 1997. URL http:/

                                        /www.ssc.com/lg/issue19/
                                        python.html.

Biekotter:2021:MPP

[Bie21] Thomas Biekötter. munuSSM: a Python package for the μfrom-ν Supersymmetric Standard Model. Computer Physics Communications, 264(??):Article 107935, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521000655.

Brunie:2020:TFP

[BII^{+}20] Hugo Brunie, Costin Iancu, Khaled Z. Ibrahim, Philip Brisk, and Brandon Cook. Tuning floating-point precision using dynamic program information and temporal locality. In IEEE [IEE20], pages 1–14. ISBN 1-72819-998-0, 1-72819-999-9 (printondemand). LCCN QA76.88.

Bisca:2001:UPQ

[Bis01] Mihai Bisca. Using Python to query MySQL over the net. Linux journal, 85:104– 106, 108, May 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Bertalan:2014:ONM

[BISC14] Tom S. Bertalan, Akand W.

Islam, Roger B. Sidje, and Eric S. Carlson. OpenMG: a new multigrid implementation in Python. Numerical Linear Algebra with Applications, 21(5):685–700, October 2014. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Beazley:2014:PCR

[BJ14a] David M. Beazley and Brian K. (Brian Kenneth) Jones. Python cookbook: Recipes for mastering Python 3. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2014. ISBN 1-4493-4037-7 (paperback), 1-4493-5736-9 (e-book). xvi + 687 pp. LCCN QA76.73.P98 B43 2013eb.

Beazley:2014:PR

[BJ14b] David M. Beazley and Brian K. (Brian Kenneth) Jones. Python. Receptury. Wydawnictwo Helion, Gliwice, Poland, third edition, 2014. ISBN 1-4920-1351-X (e-book), 1-4920-1350-1, 83-246-8180-9 (print). LCCN QA76.73.P98 B386 2014. URL http://

proquest.safaribooksonline.
com/?fpi=9781492013501.

Translation of [BJ14a] to Polish by Tomasz Walczak.

Bernardi:2023:TEP

[BJM23] Simona Bernardi, Raúl Javierre, and José Merseguer. tegdet: an extensible Python library for anomaly detection using time evolving graphs. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000596.

Brown:2020:PPP

[BJR^{+}20] Daniel D. Brown, Philip Jones, Samuel Rowlinson, Sean Leavey, Anna C. Green, Daniel Töyrä, and Andreas Freise. Pykat: Python package for modelling precision optical interferometers. SoftwareX, 12(??):Article 100613, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711020303265.

Bingol:2019:NPO

[BK19] Onur Rauf Bingol and Adarsh Krishnamurthy. NURBSpython: an open-source object-oriented NURBS modeling framework in Python. SoftwareX, 9(??):85–94, January/June 2019. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711018301778.

Bucur:2014:PSE

[BKC14] Stefan Bucur, Johannes Kinder, and George Candea.

Prototyping symbolic execution engines for interpreted languages. ACM SIGARCH Computer Architecture News, 42(1):239–254, March 2014. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Bezzam:2022:PPL

[BKH^{+}22] Eric Bezzam, Sepand Kashani, Paul Hurley, Martin Vetterli, and Matthieu Simeoni. pyFFS: a Python library for fast Fourier series computation and interpolation with GPU acceleration. SIAM Journal on Scientific Computing, 44(4):??, ???? 2022. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL https://

epubs.siam.org/doi//doi/
10.1137/21M1448641.

Bird:2009:NLP

[BKL09] Steven Bird, Ewan Klein, and Edward Loper. Natural Language Processing with Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-51649-5. 500 (est) pp. LCCN ????

Blank:2003:PPB

[BKMY03] Douglas Blank, Deepak Kumar, Lisa Meeden, and Holly Yanco. Pyro: a Pythonbased versatile programming environment for teaching robotics. ACM Journal on Educational Resources in Computing (JERIC), 3(4):1– 15, December 2003. CODEN ???? ISSN 1531-4278.

Blank:2004:PPB

[BKMY04] Douglas Blank, Deepak Kumar, Lisa Meeden, and Holly Yanco. Pyro: a pythonbased versatile programming environment for teaching robotics. ACM Journal on Educational Resources in Computing (JERIC), 4(3):1– 15, September 2004. CODEN ???? ISSN 1531-4278.

Barba:2021:SCP

[BKRT21] Lorena A. Barba, Andreas Klöckner, Prabhu Ramachandran, and Rollin Thomas. Scientific computing with Python on high-performance heterogeneous systems. Computing in Science and Engineering, 23(4):5–7, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Beazley:1997:EMP

[BL97a] D. M. Beazley and P. S. Lomdahl. Extensible message passing application development and debugging with Python. In IEEE [IEE97a], pages 650–655. ISBN 0-8186-7793-7. LCCN QA76.58 .I56 1997. IEEE catalog number 97TB100107. IEEE Computer Society Press order number PR07792.

Beazley:1997:BFL

[BL97b] David M. Beazley and Peter S. Lomdahl. Building flexible large-scale scientific computing applications with scripting languages. In Heath et al. [HTA^{+}97], page ???? CODEN PSSCFK. ISBN 0-89871-395-1. LCCN ???? CD-ROM for Windows, Macintosh and UNIX; no paper form published.

Blaess:2002:LSS

[Bla02] Christophe Blaess. Langages de scripts sous Linux: Shell Bash, Sed, Awk, Perl, Tcl, Tk, Python, Ruby. Eyrolles, Paris, France, 2002. ISBN 2-212-11028-6. xx + 733 pp. LCCN QA76.7 B4 2002.

Blaess:2004:SSL

[Bla04] Christophe Blaess. Scripts sous Linux: Shell Bash, Sed, Awk, Perl, TCL, Tk, Python, Ruby. Eyrolles, Paris, France, second edition, 2004. ISBN 2-212-11405-2. xxi + 761 pp. LCCN QA76.76O63; QA76.7.

Bleyer:2020:AFR

[Ble20] Jeremy Bleyer. Automating the formulation and resolution of convex variational problems: Applications from image processing to computational mechanics. ACM Transactions on Mathematical Software, 46(3):27:1– 27:33, September 2020. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3393881.

Bizzego:2021:MLP

[BLE21] Andrea Bizzego, Mengyu Lim, and Gianluca Espos-

ito.  mics-library: a

Python package for reproducible studies on the Multiple Indicator Cluster Survey. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021001187.

Boguslawski:2021:PBB

[BLN^{+}21] Katharina Boguslawski, Aleksandra Leszczyk, Artur Nowak, Filip Brzek, Piotr Szymon Żuchowski, Dariusz Kedziera, and Paweł Tecmer. Pythonic black-box electronic structure tool (PyBEST). An open-source Python platform for electronic structure calculations at the interface between chemistry and physics. Computer Physics Communications, 264(??): Article 107933, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521000643.

Bauer:2021:SPL

[BLP^{+}21] Michael Bauer, Wonchan Lee, Manolis Papadakis, Marcin Zalewski, and Michael Garland. Supercomputing in Python with Legate. Computing in Science and Engineering, 23(4):73–79, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Brun:2023:BPJ

[BLS^{+}23] Yuriy Brun, Tian Lin, Jessie Elise Somerville, Elisha M. Myers, and Natalie Ebner. Blindspots in Python and Java APIs result in vulnerable code. ACM Transactions on Software Engineering and Methodology, 32(3):76:1–76:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3571850.

Boschetti:2015:PDS

[BM15] Alberto Boschetti and Luca Massaron. Python data science essentials: become an efficient data science practitioner by thoroughly understanding the key concepts of Python. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-789-3, 1-78528-042-2. 258 pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781785280429.

Brown:2018:PHP

[BM18] Ryan C. Brown and Joshua Moser. HSImage: a Python and C++ library to allow interaction with ENVI-BIL hyperspectral images. Journal of Open Source Software, 3(25):630:1–630:2, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00630.

Blank:2003:PRE

[BMK03] Douglas Blank, Lisa Meeden, and Deepak Kumar. Python robotics: an environment for exploring robotics beyond LEGOs. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):317–321, January 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Broberg:2018:PPT

[BMZ^{+}18] Danny Broberg, Bharat Medasani, Nils E. R. Zimmermann, Guodong Yu, Andrew Canning, Maciej Haranczyk, Mark Asta, and Geoffroy Hautier. PyCDT: a Python toolkit for modeling point defects in semiconductors and insulators. Computer Physics Communications, 226(??): 165–179, May 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465518300079.

Boeing:2017:POP

[Boe17] Geoff Boeing. OSMnx: a Python package to work with graph-theoretic Open-StreetMap street networks. Journal of Open Source Software, 2(12):215:1–215:4, April 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00215.

Boelens:2023:SST

[Boe23] Arnout M. P. Boelens.

Stplanpy: a sustainable transportation planner for Python. SoftwareX, 22(??): ??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000353.

Borcherds:2007:PLC

[Bor07] P. H. Borcherds. Python: a language for computational physics. Computer Physics Communications, 177(1–2): 199–201, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465507000732.

Bowles:2015:MLP

[Bow15] Michael Bowles. Machine learning in Python: essential techniques for predictive analysis. Wiley, New York, NY, USA, 2015. ISBN 1-118-96174-9 (paperback), 1-119-18360-X (e-book), 1-118-96176-5 (e-book), 1-118-96175-7 (e-book). xxix + 326 pp. LCCN Q325.5.

Barrios:2017:DIG

[BP17] David Barrios and Carlos Prieto. D3GB: an interactive genome browser for R, Python, and Word-Press. Journal of Computational Biology, 24(5): 447–449, May 2017. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2016.0213;
https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2016.0213.

Braun:2013:DAN

[Bra13]  Moritz Braun.  Different

approaches to the numerical solution of the 3D Poisson equation implemented in Python. Computing, 95 (1s):49–60, May 2013. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s00607-013-0300-x.

Briegel:2023:PPP

[BRF^{+}23] Karl D. Briegel, Felix Riccius, Jakob Filser, Alexander Bourgund, Robert Spitzenpfeil, Mirco Panighel, Carlo Dri, Barbara A. J. Lechner, and Friedrich Esch. PyfastSPM: a Python package to convert 1D FastSPM data streams into publication quality movies.

SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S235271102200187X.

Brinkmann:2002:GGG

[Bri02] Peter Brinkmann. Gumbie:

a GUI generator for Jython. Dr. Dobb’s Journal of Software Tools, 27(4):46–50, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2002/2002_04/gumbie.txt;
http://www.ddj.com/ftp/
2002/2002_04/gumbie.zip.

Briggs:2006:IER

[Bri06] Keith Briggs. Implementing exact real arithmetic in python, C++ and C. Theoretical Computer Science, 351 (1):74–81, February 14, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Briot:2012:GLAa

[Bri12a] Emmanuel Briot. Gem #105:

Lady Ada kisses Python — part 1. ACM SIGADA Ada Letters, 32(2):45–46, August 2012. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Briot:2012:GLAb

[Bri12b] Emmanuel Briot. Gem #106:

Lady Ada kisses Python — part 2. ACM SIGADA Ada Letters, 32(2):47–49, August 2012. CODEN AALEE5.

ISSN 1094-3641 (print), 1557-9476 (electronic).

Blundell:2010:RTR

[BRM10] Colin Blundell, Arun Raghavan, and Milo M. K. Martin. RETCON: transactional repair without replay. ACM SIGARCH Computer Architecture News, 38(3):258– 269, June 2010. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Brown:2000:PAA

[Bro00] Martin C. Brown. Python annotated archives. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212104-

1. xxii + 722 pp. LCCN QA76.73.P98 B76 2000. Includes CD-ROM.

Browning:2006:DLP

[Bro06] James Burton Browning. Design, logic, and programming with Python: a hands-on approach. iUniverse, New York, NY, USA, 2006. ISBN 0-595-40810-9. xii + 214 pp. LCCN ????

Brune:2009:PUR

[Bru09] Corey Brune. Python: an untapped resource in system administration. ;login: the USENIX Association newsletter, 34(1):??, February 2009. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/
february-2009-volume-34-

number-1/python-untapped-
resource-system-administration.

Balbaert:2019:JPC

[BS19] Ivo Balbaert and Adrian Salceanu. Julia 1.0 programming complete reference guide: discover Julia, a high-performance language for technical computing. Learning path. Packt Publishing, Birmingham, UK, 2019. ISBN 1-83882-467-

7. viii + 451 pp. LCCN QA76.73.J84. URL http://

proquest.safaribooksonline.
com/?fpi=9781838822248.

Bernardi:2021:CPM

[BS21] Austen Bernardi and Jessica M. J. Swanson. CycFlowDec: a Python module for decomposing flow networks using simple cycles. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021000212.

Bauer:2016:PEM

[BSG^{+}16] Martin Bauer, Florian Schornbaum, Christian Godenschwager, Matthias Markl, Daniela Anderl, Harald Köstler, and Ulrich Rüde. A Python extension for the massively parallel multiphysics simulation framework waLBerla. International Journal of Parallel, Emergent and Distributed Systems: IJPEDS, 31(6):529– 542, 2016. CODEN ???? ISSN 1744-5760 (print), 1744-5779 (electronic).

Balbaert:2016:JHP

[BSS16] Ivo Balbaert, Avik Sengupta, and Malcolm Sherrington. Julia: high performance programming: learning path: leverage the power of Julia to design and develop high performing programs. Learning path. Packt Publishing, Birmingham, UK, 2016. ISBN 1-78712-570-X, 1-78712-610-2 (e-book). 697 pp. LCCN QA76.7 .B353 2016.

Boroojeny:2020:PPA

[BSSz^{+}20] Ali Ebrahimpour Boroojeny, Akash Shrestha, Ali Sharifizarchi, Suzanne Renick Gallagher, Süleyman Cenk Sahinalp, and Hamidreza Chitsaz. PyGTED: Python application for computing graph traversal edit distance. Journal of Computational Biology, 27(3):436– 439, March 2020. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

                                        tronic). URL https://www.
                                        liebertpub.com/doi/abs/
                                        10.1089/cmb.2019.0510;
                                        https://www.liebertpub.
                                        com/doi/pdf/10.1089/cmb.
                                        2019.0510.

Binder:2017:QMP

[BST^{+}17] Jan M. Binder, Alexander Stark, Nikolas Tomek, Jochen Scheuer, Florian Frank, Kay D. Jahnke, Christoph Müller, Simon Schmitt, Mathias H. Metsch, Thomas Unden, Tobias Gehring, Alexander Huck, Ulrik L. Andersen, and Lachlan J. Rogers Fedor Jelezko. Qudi: a modular Python suite for experiment control and data processing. SoftwareX, 6 (??):81–84, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300055.

Bruaset:2006:NSP

[BT06] Are Magnus Bruaset and Aslak Tveito, editors. Numerical Solution of Partial Differential Equations on Parallel Computers, volume 51 of Lecture Notes in Computational Science and Engineering. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSA6. ISBN 3-540-29076-1 (print), 3-540-31619-1 (ebook). ISSN 1439-7358. xii + 482 pp. LCCN QA377 .N87 2006. URL http:/

/link.springer.com/book/
10.1007/3-540-31619-1.

Boelens:2021:QMF

[BT21] Arnout M. P. Boelens and Hamdi A. Tchelepi. QuantImPy: Minkowski functionals and functions with Python. SoftwareX, 16(??):??, December 2021. CODEN ????

ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711021001151.

Brekel:2023:VPF

[BTDT23] Josh Brekel, Kelly R. Thorp, Kendall C. DeJonge, and Thomas J. Trout. Version 1.1.0-pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000328.

Belson:2014:AMP

[BTR14] Brandt A. Belson, Jonathan H. Tu, and Clarence W. Rowley. Algorithm 945: modred

— a parallelized model reduction library. ACM Transactions on Mathematical Software, 40(4):30:1–30:23, June 2014. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bucquet:2014:AIP

[Buc14] Samuel Bucquet. Accessing the IO ports of the Beaglebone Black with Python. Linux journal, 2014(246):3:1– 3:??, October 2014. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Buchanan:2015:PWP

[Buc15] Cameron Buchanan. Python web penetration testing cookbook: over 60 indispensable Python recipes to ensure you always have the right code on hand for web application testing. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-990-6, 1-78439-293-6. 224 (est.) pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781784392932.

Bartlett:2021:HPP

[BUS21] John Bartlett, Chris Uchytil, and Duane Storti. Highproductivity parallelism with Python plus packages (but without a cluster). Computing in Science and Engineering, 23(4):38–46, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bienstman:2007:PNR

[BVB^{+}07] Peter Bienstman, Lieven Vanholme, Wim Bogaerts, Pieter Dumon, and Peter Vandersteegen. Python in nanophotonics research. Computing in Science and Engineering, 9(3):46–47, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Bay:2022:PSV

[BVV22] Mélanie M. Bay, Silvia Vignolini, and Kevin Vynck. PyLlama: a stable and versatile Python toolkit for the electromagnetic modelling of multilayered anisotropic media. Computer Physics Communications, 273(??): Article 108256, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521003684.

Bradbury:2014:LPR

[BW14] Alex Bradbury and R. (Russel) Winder. Learning Python with Raspberry Pi. Wiley, New York, NY, USA, 2014. ISBN 1-118-71705-8 (paperback), 1-118-71702-3 (e-book), 1-118-71703-1 (e-book), 1-306-47299-7 (ebook). xvi + 270 pp. LCCN QA76.76.D47 .B73 2014.

Baczkiewicz:2022:CPL

[BW22] Aleksandra Baczkiewicz and Jarosław Watróbski. Crispyn

— a Python library for determining criteria significance with objective weighting methods. SoftwareX, 19(??): ??, July 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022001029.

Bodker:2022:SPB

[BWMS22] Mikkel S. Bødker, Collin J. Wilkinson, John C. Mauro, and Morten M. Smedskjaer. StatMechGlass: Python based software for composition-structure prediction in oxide glasses using statistical mechanics. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001606.

Baczkiewicz:2023:VCP

[BWS23] Aleksandra Baczkiewicz, Jarosław Watróbski, and Wojciech Sałabun. Version [1.1] — [Crispyn — a Python library for determining criteria significance with objective weighting methods]. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002376.

Brandstetter:2021:KPP

[BYL^{+}21] Dominik Brandstetter, Xiaosheng Yang, Daniel Lüftner, F. Stefan Tautz, and Peter Puschnig. kMap.py: a Python program for simulation and data analysis in photoemission tomography. Computer Physics Communications, 263(??): Article 107905, June 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521000461.

Bao:2023:PEN

[BYWW23] Nana Bao, Xingting Yan, Shiwen Wei, and Zihao Wang. Py-EFIT: a new Python package for plasma equilibrium reconstruction on EAST tokamak. Computer Physics Communications, 282(??):

Article 108549, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522002685.

Caliskan:2023:STC

[ÇA23] Murat Çaliskan and Berk Anbaroglu. Space Time Cube analytics in QGIS and Python for hot spot detection. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023001942.

Cannon:2014:PPB

[Can14] Jason Cannon. Python programming for beginners: an introduction to the Python computer language and computer programming. ????, ????, 2014. ISBN 1-5010-0086-1. 152 pp. LCCN QA76.73.P98 C366 2014.

Cass:2017:WTH

[Cas17] Stephen Cass. Wearable tech for Halloween — the Gemma MO’s embedded Python lets you change your code on the fly [resources tools]. IEEE Spectrum, 54(10):15– 16, October 2017. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cass:2020:TPL

[Cas20] S. Cass. The top programming languages: Our latest rankings put Python on topagain — [careers]. IEEE Spectrum, 57(8):22, August 2020. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chowdhury:2022:UTP

[CASA22] Md Atique Reza Chowdhury, Rabe Abdalkareem, Emad Shihab, and Bram Adams. On the untriviality of trivial packages: an empirical study of npm JavaScript packages. IEEE Transactions on Software Engineering, 48 (8):2695–2708, August 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Crespo:1996:WBB

[CB96] Arturo Crespo and Eric A.

Bier. WebWriter: a browserbased editor for constructing Web applications. Computer Networks and ISDN Systems, 28(7–11):1291–1306, May 1, 1996. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
comnet/cas_sub/browse/browse.
cgi?year=1996&volume=28&
issue=7-11&aid=1616.

Cooper:2014:BES

[CBB14] Christopher D. Cooper, Jaydeep P. Bardhan, and L. A.

Barba. A biomolecular electrostatics solver using Python, GPUs and boundary elements that can handle solvent-filled cavities and Stern layers. Computer Physics Communications, 185(3):720–729, March 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465513003731.

Ching:2022:QLO

[CBLI22] Eric J. Ching, Brett Bornhoft, Ali Lasemi, and Matthias Ihme. Quail: a lightweight open-source discontinuous Galerkin code in Python for teaching and prototyping. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S235271102200005X.

Castro:2024:LHP

[CBST24] Oscar Castro, Pierrick Bruneau, Jean-Sébastien Sottet, and Dario Torregrossa. Landscape of high-performance Python to develop data science and machine learning applications. ACM Computing Surveys, 56(3):65:1– 65:??, March 2024. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL https://

dl.acm.org/doi/10.1145/
3617588.

Chard:2021:EAP

[CBW^{+}21] Kyle Chard, Yadu Babuji, Anna Woodard, Ben Clifford, Zhuozhao Li, Mihael Hategan, Ian Foster, Mike Wilde, and Daniel S. Katz. Extended abstract: Productive parallel programming with Parsl. ACM SIGADA Ada Letters, 40(2):73–75, April 2021. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3463478.3463486.

Cooper:2020:SDS

[CC20] Crispin H. V. Cooper and Alain J. F. Chiaradia. sDNA: 3-d spatial network analysis for GIS, CAD, command line and Python. SoftwareX, 12(??):Article 100525, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019303401.

Cooper:2016:PPP

[CCFB16] Christopher D. Cooper, Natalia C. Clementi, Gilbert Forsyth, and Lorena A. Barba. PyGBe: Python, GPUs and boundary elements for biomolecular electrostatics. Journal of Open Source Software, 1(4):43:1, August 2016. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00043.

Cao:2023:TBD

[CCM^{+}23] Yulu Cao, Lin Chen, Wanwangying Ma, Yanhui Li, Yuming Zhou, and Linzhang Wang. Towards better dependency management: a first look at dependency smells in Python projects. IEEE Transactions on Software Engineering, 49(4):1741–1765, April 2023. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Castanos:2012:BPE

[CEI^{+}12] Jose Castanos, David Edelsohn, Kazuaki Ishizaki, Priya Nagpurkar, Toshio Nakatani, Takeshi Ogasawara, and Peng Wu. On the benefits and pitfalls of extending a statically typed language JIT compiler for dynamic scripting languages. ACM SIGPLAN Notices, 47(10):195–212, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cid-Fuentes:2020:EDH

[CFÁA^{+}20] Javier Álvarez Cid-Fuentes, Pol Álvarez, Ramon Amela, Kuninori Ishii, Rafael K. Morizawa, and Rosa M. Badia. Efficient development of high performance data analytics in Python. Future Generation Computer Systems, 111(??):570– 581, October 2020. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X18321393.

Clementi:2017:PPL

[CFCB17] Natalia C. Clementi, Gilbert Forsyth, Christopher D. Cooper, and Lorena A. Barba. PyGBe-LSPR: Python and GPU boundary-integral solver for electrostatics. Journal of Open Source Software, 2(19):306:1–306:2, November 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00306.

Cartis:2019:IFR

[CFMR19] Coralia Cartis, Jan Fiala, Benjamin Marteau, and Lindon Roberts. Improving the flexibility and robustness of model-based derivative-free optimization solvers. ACM Transactions on Mathematical Software, 45(3):32:1– 32:41, August 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3338517.

Castillo-Felisola:2023:CPA

[CFPS23] Oscar Castillo-Felisola, Dominic T. Price, and Mattia Scomparin. Cadabra and Python algorithms in general relativity and cosmology II: Gravitational waves. Computer Physics Communications, 289(??): Article 108748, August 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523000930.

Chang:2014:CPP

[CFSK14] Yuan-Pin Chang, Frank Filsinger, Boris G. Sartakov, and Jochen Küpper. CMIstark: Python package for the stark-effect calculation and symmetry classification of linear, symmetric and asymmetric top wavefunctions in dc electric fields. Computer Physics Communications, 185(1):339–349, January 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465513003019.

Chilenski:2017:EME

[CFW17] M. A. Chilenski, I. C. Faust, and J. R. Walk. eqtools: Modular, extensible, opensource, cross-machine Python tools for working with magnetic equilibria. Computer Physics Communications, 210(??):155–162, January 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046551630282X.

Ching:2017:XOS

[CG17] Daniel J. Ching and Dogă Gürsoy. XDesign: an open-source software package for designing X-ray imaging phantoms and experiments. Journal of Synchrotron Radiation, 24(2):537–544, 2017. ISSN 0909-0495 (print), 1600-5775 (electronic).

Cass:2023:HPS

[CG23] Stephen Cass and Harry Goldstein. How Python swallowed the world: Lessons from compiling top programming languages. IEEE Spectrum, 60(9):2, September 2023. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cabanero-Gomez:2021:EPM

[CGHGRB21] Luis Cabañero-Gomez, Ramon Hervas, Ivan Gonzalez, and Luis Rodriguez-Benitez. eeglib: a Python module for EEG feature extraction. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021000753.

Catanzaro:2011:CCE

[CGK11] Bryan Catanzaro, Michael Garland, and Kurt Keutzer. Copperhead: compiling an embedded data parallel language. ACM SIGPLAN Notices, 46(8):47–56, August 2011. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPoPP ’11 Conference proceedings.

Cokelaer:2017:PSSb

[CH17] Thomas Cokelaer and Juer-

                                        gen Hasch.  'Spectrum':

Spectral analysis in Python. Journal of Open Source Software, 2(18):348:1–348:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00348.

Chapman:2001:CUI

[Cha01] Mitch Chapman. Create user interfaces with Glade. Linux journal, 87:88, 90–92, 94, July 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //noframes.linuxjournal.
                                        com/lj-issues/issue87/4702.

html. See erratum [Ano01a].

Chan:2015:LPO

[Cha15] Jamie Chan. Learn Python in one day and learn it well: Python for beginners with hands-on project: the only book you need to start coding in Python immediately. CreateSpace Independent Publishing, North Charleston, SC, USA, 2015. ISBN 1-5060-9438-4 (paperback). 123 pp. LCCN QA76.73.P98 C453 2015.

Constantine:2016:PPA

[CHG^{+}16] Paul Constantine, Ryan Howard, Andrew Glaws, Zachary Grey, Paul Diaz, and Leslie Fletcher. Python active-subspaces utility library. Journal of Open Source Software, 1(5):79:1, September 2016. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00079.

Cheng:2024:RKB

[CHM24] Wei Cheng, Wei Hu, and Xiaoxing Ma. Revisiting knowledge-based inference of Python runtime environments: a realistic and adaptive approach. IEEE Transactions on Software Engineering, 50(2):258–279, February 2024. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Christopher:2002:PPP

[Chr02] Thomas W. Christopher.

Python programming patterns. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-040956-1. xix + 538 pp. LCCN QA76.73.P98 C47 2002. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130409561.html.

Chun:2001:CPP

[Chu01] Wesley Chun. Core Python programming. Prentice Hall PTR core series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-026036-3. xxix + 771 pp. LCCN QA76.73.P98 C48 2001. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130260363.html.

Chun:2002:KPR

[Chu02a] Wesley J. Chun. Keeping up with Python: the 2.2 release. Linux journal, 99:??, July 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //www.linuxjournal.com/
                                        article/5597.

Chun:2002:PQG

[Chu02b] Wesley J. Chun. Python 2.2 Q&A with Guido van Rossum, creator of Python. Linux journal, 98:??, June 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

                                        //www.linuxjournal.com/
                                        article.php?sid=5948.

Chun:2007:ACPb

[Chu07a] Wesley J. Chun. Au coeur de Python: Notions avancées. (French) [Core Python: Advanced ideas]. CampusPress, Paris, France, 2007. ISBN 2-7440-2195-4. xxii + 337 pp. LCCN ????

Chun:2007:ACPa

[Chu07b] Wesley J. Chun. Au coeur de Python: Notions fondamentales. (French) [Core Python:

Fundamental ideas]. CampusPress, Paris, France, 2007. ISBN 2-7440-2148-2. xxviii + 645 pp. LCCN ????

Chun:2007:CPP

[Chu07c]  Wesley J. Chun.  Core

Python programming. Prentice Hall core series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 2007. ISBN 0-13-226993-7 (paperback). xxxvii + 1077 pp. LCCN QA76.73.P98 C48

2007. URL http://www.loc.
gov/catdir/toc/ecip0615/
2006019559.html.

Chun:2008:PF

[Chu08] Wesley Chun. Python fundamentals. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-714341-9 (paperback). vi + 94 pp. LCCN ????

Cabrera:1996:PFI

[CI96] L.-F. Cabrera and N. Islam, editors. Proceedings of the Fifth International Workshop on Object-Orientation in Operating Systems: October 27– 28, 1996, Seattle, Washington. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7693-0. LCCN QA 76.76 O63 I59 1996. IEEE catalog number 96TB100089.

Chen:2022:HHI

[CJ22] Qiao Chen and Xiangmin Jiao. HIFIR: Hybrid incomplete factorization with iterative refinement for preconditioning ill-conditioned and singular systems. ACM Transactions on Mathematical Software, 48(3):32:1– 32:??, September 2022. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3536165.

Chaudhary:2023:PPP

[CJGA23] Gaurav Chaudhary, Hicham Johra, Laurent Georges, and Bjørn Austbø. pymodconn: a Python package for developing modular sequenceto-sequence control-oriented deep neural networks. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002959.

Chen:2023:PBC

[CJYH23] Jie Chen, Tao Jiang, Dongjin Yu, and Haiyang Hu. Patternbased circular reference detection in Python. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S016764232300014X.

Chapman:2000:EPT

[CK00] Mitch Chapman and Brian Kelley. Examining the PyGtk toolkit. Dr. Dobb’s Journal of Software Tools, 25(4):82, 84, 86, 88, April 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_04/pygtk.
txt.

Czmil:2023:GPI

[CKC23] Anna Czmil, Jacek Kluska, and Sylwester Czmil. GPR: a Python implementation of an extremely simple classifier based on fuzzy logic and gene expression programming. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000584.

Chun:2013:GPP

[CKK^{+}13] Kyungwon Chun, Huioon Kim, Hyounggyu Kim, Kil Su Jung, and Youngjoo Chung. GMES: a Python package for solving Maxwell’s equations using the FDTD method. Computer Physics Communications, 184(4):1272– 1279, April 2013. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465512004079.

Cai:2006:PPS

[CL06] Xing Cai and Hans Petter Langtangen. Parallelizing PDE solvers using the Python programming language. In Bruaset and Tveito [BT06], pages 295–325. CO-DEN LNCSA6. ISBN 3-540-29076-1 (print), 3-540-31619-1 (e-book). ISSN 1439-7358. LCCN QA377 .N87

2006. URL http://link.
springer.com/content/pdf/
10.1007/3-540-31619-1_9.

Clare:2015:RFS

[Cla15] Amanda Clare. Review of “A functional start to computing with Python”, Ted Herman, CRC Press, 2014, ISBN 978-1-4665-0455-4. Journal of Functional Programming, 25:e15, ???? 2015. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:/

/www.cambridge.org/core/
journals/journal-of-functional-
programming/article/review-
of-a-functional-start-to-
computing-with-python-ted-
herman-crc-press-2014-isbn-
9781466504554/CDD2895726478A185EDA78AE98C30BB3.

Cai:2005:PPP

[CLM05] Xing Cai, Hans Petter Langtangen, and Halvard Moe. On the performance of the Python programming language for serial and parallel scientific computations. Scientific Programming, 13(1): 31–56, ???? 2005. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Coleman:2020:MPJ

[CLMM20] Chase Coleman, Spencer Lyon, Lilia Maliar, and Serguei Maliar. Matlab, Python, Julia: What to choose in economics? Computational Economics, page ??, 2020. CODEN CNOMEL. URL http://

link.springer.com/article/
10.1007/s10614-020-09983-
3.

Cornea:2020:ISC

[CLT20] Marius Cornea, Weiqiang Liu, and Arnaud Tisserand, editors. 2020 27th IEEE Symposium on Computer Arithmetic: ARITH 2020: proceedings: Portland, Oregon, USA, 7–10 June 2020. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020. ISBN 1-72817-120-2, 1-72817-121-0. LCCN ???? URL

https://ieeexplore.ieee.
org/servlet/opac?punumber=
9146973.

Cooper:2007:ERH

[CM07] Jonathan Cooper and Steve McKeever. Experience report: a Haskell interpreter for cellML. ACM SIGPLAN Notices, 42(9):247–250, September 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaparro:2020:PPB

[CM20] Gustavo Chaparro and Andrés Mejía. Phasepy: a Python based framework for fluid phase equilibria and interfacial properties computation. Journal of Computational Chemistry, 41(29): 2504–2526, November 2020. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Ceriotti:2014:PPI

[CMM14] Michele Ceriotti, Joshua More, and David E. Manolopoulos. i-PI: a Python interface for ab initio path integral molecular dynamics simulations. Computer Physics Communications, 185(3): 1019–1026, March 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046551300372X.

Caprile:2022:PPP

[CMS22] Fernando Caprile, Luciano A. Masullo, and Fernando D. Stefani. PyFocus — a Python package for vectorial calculations of focused optical fields under realistic conditions. Application to toroidal foci. Computer Physics Communications, 275(??): Article 108315, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522000339.

Coelho:2017:PJS

[Coe17] Luis Pedro Coelho. Jug: Software for parallel reproducible computation in Python. Journal of Open Research Software, 5(1):30–??, October 27, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.161/.

Conway:1995:PGD

[Con95] Matthew J. Conway. Python:

a GUI development tool. Interactions (New York, N.Y.), 2(2):23–28, April 1995. CO-DEN IERAE3. ISSN 1072-5520.

Copeland:2008:ES

[Cop08] Rick Copeland. Essential SQLAlchemy. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 0-596-51614-2. xi + 215 pp. LCCN QA76.9.W43 C67 2008.

Cottom:2003:USB

[Cot03]  Teresa L. Cottom.  Us-

ing SWIG to bind C++ to Python. Computing in Science and Engineering, 5(2): 88–96, c3, March/April 2003. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:/

/csdl.computer.org/comp/
mags/cs/2003/02/c2088abs.
htm;   http://csdl.computer.
org/dl/mags/cs/2003/02/
c2088.htm;   http://csdl.
computer.org/dl/mags/cs/
2003/02/c2088.pdf.

Cox:2014:RPC

[Cox14]  Tim Cox.  Raspberry Pi

cookbook for Python programmers: over 50 easyto-comprehend tailor-made recipes to get the most out of the Raspberry Pi and unleash its huge potential using Python. Packt Publishing, Birmingham, UK, 2014. ISBN 1-84969-662-4. iii + 388 pp. LCCN QA76.8.R19 .C6.

Conti:2012:TMP

[CR12] Juan José Conti and Alejandro Russo. A Taint mode for Python via a library. Lecture Notes in Computer Science, 7127:210–222, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/978-3-642-27937-9_
                                        15/.

Coelho:2015:BML

[CR15] Luis Pedro Coelho and Willi Richert. Building machine learning systems with Python: get more from your data through creating practical machine learning systems with Python. Community experience distilled. Packt Publishing, Birmingham, UK, second edition, 2015. ISBN 1-78439-277-4, 1-78439-288-X (e-book). xi + 301 pp. LCCN QA76.73.P98 C64 2015.

Carini:2022:CPO

[CR22] Matheus Roman Carini and Marcelo Maia Rocha. CESSIPy: a Python open-source module for stochastic system identification in civil engineering. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022000632.

Criado:2018:MPL

[Cri18] Juan C. Criado. MatchingTools: a Python library for symbolic effective field theory calculations. Computer Physics Communications, 227(??):42–50, June 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465518300456.

Champseix:2021:PPH

[CRL21] Robin Champseix, Laurent Ribiere, and Clément Le Couedic. A Python package for heart rate variability analysis and signal preprocessing. Journal of Open Research Software, 9(1):28–??, October 06, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.305/.

Choirat:2009:EP

[CS09] Christine Choirat and Raffello Seri. Econometrics with Python. Journal of Applied Econometrics, 24(4): 698–704, June–July 2009.

CODEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Capocchi:2021:WBS

[CS21] Laurent Capocchi and Jean-Francois Santucci. A webbased simulation of discreteevent system of system with the mobile application DEVSimPy-mob. SoftwareX, 13(??):Article 100625, January 2021. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711020303381.

Chudoba:2013:UPS

[CSRV13] R. Chudoba, V. Sadílek, R. Rypl, and M. Vorechovský. Using Python for scientific computing: Efficient and flexible evaluation of the statistical characteristics of functions with multivariate random inputs. Computer Physics Communications, 184(2):414–427, February 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465512003086.

Caudai:2019:LSC

[CSZ^{+}19] Claudia Caudai, Emanuele Salerno, Monica Zoppè, Ivan Merelli, and Anna Tonazzini. ChromStruct 4: a Python code to estimate the chromatin structure from Hi-C data. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 16(6):1867–1878, November 2019. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1109/
TCBB.2018.2838669.

Cuesta:2013:PDA

[Cue13] Hector Cuesta. Practical data analysis. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78328-099-9 (print), 1-68015-361-7 (ebook), 1-78328-100-6. 360 pp. LCCN QA76.9.S88 D683 2013eb. URL http://

proquest.safaribooksonline.
com/9781783280995.

Chandra:2015:PRE

[CV15] Rakesh Vidya Chandra and Bala Subrahmanyam Varanasi. Python requests essentials: learn how to integrate your applications seamlessly with web services using Python requests. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-541-2, 1-78439-231-6 (e-book). 134 pp. LCCN QA76.73.P98. URL

http://proquest.safaribooksonline.
com/9781784395414.

Cejnek:2022:PPO

[CV22] Matous Cejnek and Jan Vrba.

Padasip: an open-source Python toolbox for adaptive filtering. Journal of Computational Science, 65: ??, November 2022. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S1877750322002460.

Caponi:2023:BPP

[CVV23] Francesco Caponi, David F. Vetsch, and Davide Vanzo. BASEveg: a Python package to model riparian vegetation dynamics coupled with river morphodynamics. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023000572.

Caron:2023:PPM

[CWC^{+}23] Eddy Caron, Jean-François Witz, Christophe Cuvier, Arnaud Beaurain, Vincent Magnier, and Ahmed El Bartali. PYCASO: Python module for calibration of cameras by Soloff’s method. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102300136X.

Couty:2021:GGA

[CWLG^{+}21] Victor Couty, Jean-François Witz, Pauline Lecomte-Grosbras, Julien Berthe, Eric Deletombe, and Mathias Brieu. GPUCorrel: a GPU accelerated Digital Image Correlation software written in Python. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001102.

Couty:2021:CCR

[CWM^{+}21] Victor Couty, Jean-François Witz, Corentin Martel, François Bari, and Antoine Weisrock. CRAPPY: Command and Real-Time Acquisition in Parallelized Python, a Python module for experimental setups. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001278.

Chen:2023:RPO

[CX23] Zhi Chen and Peng Xiong.

RSOME in Python: An open-source package for robust stochastic optimization made easy. INFORMS Journal on Computing, 35(4):717– 724, July/August 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/
doi/full/10.1287/ijoc.2023.
1291.

Chatelain:2023:PAP

[CYKG23] Yohan Chatelain, Nigel Yong Sao Young, Gregory Kiar, and Tristan Glatard. Py-Tracer: Automatically profiling numerical instabilities in Python. IEEE Transactions on Computers, 72(6): 1792–1803, June 2023. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Chen:2022:THP

[CZ22] Zhiwen Chen and Baorong Zhong. TFInterpy: a highperformance spatial interpolation Python package. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022001479.

Camino:2023:CPI

[CZA^{+}23] Bruno Camino, Huanyu Zhou, Eleonora Ascrizzi, Alberto Boccuni, Filippo Bodo, Alessandro Cossard, Davide Mitoli, Anna Maria Ferrari, Alessandro Erba, and Nicholas M. Harri-

                                        son. CRYSTALpytools: a

Python infrastructure for the Crystal code. Computer Physics Communications, 292(??):Article 108853, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465523001984.

Chen:2019:PPL

[CZZ19] Zhangqi Chen, Qiaofu Zhang, and Ji-Cheng Zhao. pydiffusion: a Python library for diffusion simulation and data analysis. Journal of Open Research Software, 7(1):13–??, April 23, 2019. CODEN ????

ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.255/.

Dahlgren:2018:PCP

[Dah18a] Björn Dahlgren. ChemPy: a package useful for chemistry written in Python. Journal of Open Source Software, 3 (24):565:1–565:2, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00565.

Dahlgren:2018:PPSb

[Dah18b] Björn Dahlgren. pyneqsys:

Solve symbolically defined systems of non-linear equations numerically. Journal of Open Source Software, 3(21):531:1–531:2, January 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00531.

Dahlgren:2018:PPSa

[Dah18c] Björn Dahlgren. pyodesys:

Straightforward numerical integration of ODE systems from Python. Journal of Open Source Software, 3(21): 490:1–490:2, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00490.

Desai:2015:PPA

[DAJ^{+}15] Pratik Desai, Saleem Ahmed, James Jones, Jasmine Nadar, and Vikrant Phadke, editors. Python programming for Arduino: develop practical Internet of Things prototypes and applications with Arduino and Python. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78328-593-1 (paperback), 1-78328-594-X (e-book). 400 (est) pp. LCCN QA76.73.P98.

Dalke:2001:MCE

[Dal01] Andrew Dalke. Making C extensions more Pythonic. Dr. Dobb’s Journal of Software Tools, 26(1):68, 70, 72, 74, 76, January 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_01/cpython.txt;   http:
//www.ddj.com/ftp/2001/
2001_01/cpython.zip.

Daly:2007:NGW

[Dal07] Liza Daly. Next-generation web frameworks in Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51371-2. LCCN QA76.73.P98 L59 2007eb; QA76.73.P98. URL http://

www.oreilly.com/catalog/
9780596513719.

Dan:2018:LJE

[Dan18a]  Toomey Dan.  Learning

Jupyter 5: explore interactive computing using Python, Java, JavaScript, R, Julia, and JupyterLab. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78913-740-3, 1-78913-744-6. 282 pp. LCCN Q183.9; QA76.9.I52 .T666 2018.

Daniel:2018:SG

[Dan18b] John Leon Daniel. Security with Go. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78862-791-1. LCCN QA76.585 .L466 2018; QA76.59.

Darcy:2012:EGP

[Dar12] Jeff Darcy. Extending GlusterFS with Python. Linux journal, 2012(223):2:1–2:??, November 2012. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Davies:2021:SPB

[Dav21] Claire L. Davies. SEDBYS:

a Python-based SED Builder for Young Stars. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000327.

Dawson:2003:PPA

[Daw03] Michael Dawson. Python programming for the absolute beginner. Premier Press Inc., Boston, MA, USA, 2003. ISBN 1-59200-073-8 (paperback), 1-59200-073-8, 1-59200-269-2 (e-book). xviii + 452 pp. LCCN QA76.73.P98 D387 2003. URL http://

proquest.safaribooksonline.
com/1592000738.

Dawson:2006:PPA

[Daw06] Mike Dawson. Python programming for the absolute beginner. Thomson Course Technology, Boston, MA, USA, second edition, 2006. ISBN 1-59863-112-8. xxiv + 447 pp. LCCN QA76.73.P98 D39 2006.

Dawson:2010:PPA

[Daw10] Mike Dawson. Python programming for the absolute beginner. For the absolute beginner. Course Technology Cengage Learning, Boston, MA, USA, third edition, 2010. ISBN 1-4354-5500-2. xxiii + 455 pp. LCCN QA76.73.P98 D39 2010.

URL http://catdir.loc.
gov/catdir/enhancements/
fy1105/2009933304-b.html;
 http://catdir.loc.gov/
catdir/enhancements/fy1105/
2009933304-d.html;   http:
//catdir.loc.gov/catdir/
enhancements/fy1105/2009933304-
t.html;   http://www.courseptr.
com/downloads.

Dayley:2007:PTC

[Day07a] Brad Dayley. Python in tasca: codice e commandi in tasca. Pearson Education, Milano, Italy, 2007. ISBN 88-7192-405-3. vi + 277 pp. LCCN ????

Dayley:2007:PPE

[Day07b]  Brad Dayley.  Python

phrasebook: essential code and commands. Developer’s library. SAMS Publishing, Indianapolis, IN, USA, 2007. ISBN 0-672-32910-7 (paperback). v + 275 pp. LCCN QA76.73.P98 D395 2007. URL http:/

/www.loc.gov/catdir/toc/
fy0706/2006922308.html.

Day:2014:PP

[Day14] Charles Day. Python power.

Computing in Science and Engineering, 16(1):88, January/February 2014. CO-DEN CSENFA. ISSN 1521-9615.

Diem:2017:PVP

[DB17] Alexandra K. Diem and Neil W. Bressloff. VaMpy: a Python package to solve 1D blood flow problems. Journal of Open Research Software, 5(1):17–??, June 08, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.159/.

Dirmeier:2022:SHM

[DB22] Simon Dirmeier and Niko Beerenwinkel. Structured hierarchical models for probabilistic inference from perturbation screening data. Annals of Applied Statistics, 16(3):2010–2029, September 2022. CODEN ???? ISSN 1932-6157 (print), 1941-7330 (electronic). URL https://

projecteuclid.org/journals/
annals-of-applied-statistics/
volume-16/issue-3/Structured-
hierarchical-models-for-
probabilistic-inference-
from-perturbation-screening-
data/10.1214/21-AOAS1580.
full.

Dutra:2021:PEP

[DBdFdSR21] José Diogo L. Dutra, Thiago D. Bispo, Sabrina M. de Freitas, and Marcos V. dos S. Rezende. ParamGULP: an efficient Python code for obtaining interatomic potential parameters for General Utility Lattice Program. Computer Physics Communications, 265(??): Article 107996, August 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521001089.

Dzib:2019:EPC

[DCOC^{+}19] Eugenia Dzib, José Luis Cabellos, Filiberto Ortíz-Chi, Sudip Pan, Annia Galano, and Gabriel Merino. Eyringpy: a program for computing rate constants in the gas phase and in solution. International Journal of Quantum Chemistry, 119(2): e25686:1–e25686:??, January 15, 2019. CODEN IJQCB2.

ISSN 0020-7608 (print), 1097-461X (electronic).

Dang:2023:LGP

[DCRF23] Khanh Dang, Jie Chen, Brian Rodgers, and Saryu Fensin. LAVA 1.0: a general-purpose Python toolkit for calculation of material properties with LAMMPS and VASP. Computer Physics Communications, 286(??):Article 108667, May 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523000127.

Dogaru:2015:UPJ

[DD15] I. Dogaru and R. Dogaru. Using Python and Julia for efficient implementation of natural computing and complexity related algorithms. In 2015 20th International Conference on Control Systems and Computer Science, pages 599–604. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2015.

Dsilva:2019:MRD

[DDK19] Joseph Vinish D’silva, Florestan De Moor, and Bettina Kemme. Making an RDBMS data scientist friendly: advanced in-database interactive analytics with visualization support. Proceedings of the VLDB Endowment, 12 (12):1930–1933, August 2019.

CODEN ???? ISSN 2150-8097.

Deitel:2002:PHP

[DDLW02] Harvey M. Deitel, Paul Deitel, Jonathan Liperi, and Ben Wiedermann. Python How To Program. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-092361-3 (paperback), 0-13-092557-8 (CD-ROM). lviii + 1292 + 6 pp. LCCN QA76.73.P98 P98 2002. US$74.00. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130923613.html.

Dolgopolovas:2014:PSC

[DDMS14]  Vladimiras  Dolgopolovas,

Valentina Dagiene, Saulius Minkevicius, and Leonidas Sakalauskas. Python for scientific computing education: Modeling of queueing systems. Scientific Programming, 22(1):37–51, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

deDeusFilho:2022:PII

[ddSNX22] João Carlos Andrade de Deus Filho, Luiz Carlos da Silva Nunes, and José Manuel Cardoso Xavier.

                                        iCorrVision-2D: an inte-

grated python-based opensource Digital Image Correlation software for inplane measurements (Part 1). SoftwareX, 19(??):??, July 2022. CODEN ????

ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022000851.

Palma:2020:APL

[DDT20] Stefano Dalla Palma, Dario Di Nucci, and Damian A.

Tamburri. AnsibleMetrics:

a Python library for measuring Infrastructure-as-Code blueprints in Ansible. SoftwareX, 12(??):Article 100633, July/December 2020. CO-DEN ???? ISSN 2352-

7110. URL http://www.
sciencedirect.com/science/
article/pii/S2352711020303460.

Decaluwe:2004:MPB

[Dec04] Jan Decaluwe. MyHDL: a Python-based hardware description language. Linux journal, 2004(127):??, November 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Dejanovic:2022:PLG

[Dej22] Igor Dejanović. Parglare: a LR/GLR parser for Python. Science of Computer Programming, 214(??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0167642321001271.

DePalma:2019:PPI

[DEMM19] Barbara De Palma, Marco Erba, Luca Mantovani, and Nicola Mosco. A Python program for the implementation of the Γ-method for Monte Carlo simulations. Computer Physics Communications, 234(??):294–301, January 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465518302534.

Dutt:2000:GBG

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder für das Internet richtig vorbereiten; auf der CD: GIMP für Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMP- und GTK-Versionen, über 300 Plug-ins in C, Perl, tcl, Python und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

Dalcin:2021:MSU

[DF21a] Lisandro Dalcin and Yao-Lung L. Fang. mpi4py: Status update after 12 years of development. Computing in Science and Engineering, 23 (4):47–54, July/August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dunbar:2021:BPP

[DF21b] M. Bekker-Nielsen Dunbar and Thomas J. R. Finnie. bayesint: a Python package for calculating Bayesian credible intervals of ratios of beta distributions. Journal of Open Research Software, 9(1):35–??, December 21, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.283/.

Demidov:2018:UPP

[DFC18] Alexander G. Demidov, Michael E.

Fortunato, and Coray M. Colina. Update 0.2 to “pysimm: a Python package for simulation of molecular systems”. SoftwareX, 7(??):63–69, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018300141.

See [FC17].

Doulis:2019:CMP

[DFSW19] Georgios Doulis, Jörg Frauendiener, Chris Stevens, and Ben Whale. COFFEE — an MPI-parallelized Python package for the numerical evolution of differential equations. SoftwareX, 10 (??):Article 100283, July/ December 2019. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711019300950.

Drummond:2009:PPB

[DGMP09] L. Anthony Drummond, Vicente Galiano, Violeta Migallón, and Jose Penadés. Py-ACTS: a Python based interface to ACTS tools and parallel scientific applications. International Journal of Parallel Programming, 37(1):58–77, February 2009. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

                                        //www.springerlink.com/
                                        openurl.asp?genre=article&
                                        issn=0885-7458&volume=37&
                                        issue=1&spage=58.

Dubois:1996:NP

[DHH96] P. F. Dubois, K. Hinsen, and J. Hugunin. Numerical Python. Computers in Physics, 10(3):262– 267, May/June 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

DiPierro:2013:AAP

[Di 13] Massimo Di Pierro. Annotated algorithms in Python: with applications in physics, biology, and finance. Experts4Solutions, Lexington, KY, USA, 2013. ISBN 0-9911604-0-1 (paperback). 388 pp. LCCN QA76.73.P98 D57 2013.

DiPierro:2014:PPP

[Di 14] Massimo Di Pierro. Portable parallel programs with Python and OpenCL. Computing in Science and Engineering, 16 (1):34–40, January/February 2014. CODEN CSENFA. ISSN 1521-9615.

DiDomenico:2023:NPB

[DLC23] Daniel Di Domenico, João V. F. Lima, and Gerson G. H. Cavalheiro. NAS Parallel Benchmarks with Python: a performance and programming effort analysis focusing on GPUs. The Journal of Supercomputing, 79(8): 8890–8911, May 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-022-04932-3.

Decan:2020:SPL

[DM20] Alexandre Decan and Tom Mens. Sismic — a Python library for statechart execution and testing. SoftwareX, 12(??):Article 100590, July/December 2020. CO-DEN ???? ISSN 2352-

7110. URL http://www.
sciencedirect.com/science/
article/pii/S2352711020303034.

Ding:2015:PPF

[DMC^{+}15] Hong Ding, Bharat Medasani, Wei Chen, Kristin A. Persson, Maciej Haranczyk, and Mark Asta. PyDII: a Python framework for computing equilibrium intrinsic point defect concentrations and extrinsic solute site preferences in intermetallic compounds. Computer Physics Communications, 193(??): 118–123, August 2015. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465515001149.

DeFever:2021:MCC

[DMD^{+}21] Ryan S. DeFever, Ray A. Matsumoto, Alexander W. Dowling, Peter T. Cummings, and Edward J. Maginn. MoS-DeF Cassandra: a complete Python interface for the Cassandra Monte Carlo software. Journal of Computational Chemistry, 42(18): 1321–1331, July 5, 2021. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Doglio:2015:MPH

[Dog15] Fernando Doglio. Mastering Python High Performance. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-930-0, 1-78398-931-9 (e-book). 260 pp. LCCN T55.4-60.8.

Donaldson:2009:P

[Don09] Toby Donaldson. Python. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, second edition, 2009.

ISBN 0-321-58544-5 (paperback). vi + 185 pp. LCCN X09.F00872.

Donat:2014:LRP

[Don14] Wolfram Donat. Learn Raspberry Pi programming with Python. TECHNOLOGY IN ACTION series. Apress, Berkeley, CA, USA, 2014. ISBN 1-4302-6424-1 (paperback0). xxi + 231 pp. LCCN QA76.73.P98 .D663 2014.

Doring:2008:ESL

[Dör08]  Holger Döring.  Evaluat-

ing scripting languages: How Python can help political methodologists. The Political Methodologist, 16(1):8–12, 2008.

dosSantosLessa:2001:PDH

[dos01] André dos Santos Lessa.

Python Developer’s Handbook. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31994-2. xxv + 929 pp. LCCN QA76.73.P98 L47 2001. US$44.99.

Dawe:2017:PRI

[DOS17] Edmund Noel Dawe, Piti Ongmongkolkul, and Giordon Stark. root numpy: The interface between ROOT and NumPy. Journal of Open Source Software, 2(16): 307:1–307:2, August 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00307.

Dougherty:1998:PGV

[Dou98] Dale Dougherty. Python’s Guido van Rossum. Web Review, ??(??):??, April 10, 1998. URL http:

                                        //webreview.com/wr/pub/
                                        Dougherty_Dale.

Downey:2009:PSD

[Dow09] Allen Downey. Python for software design: how to think like a computer scientist. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89811-0, 0-521-72596-8. ???? pp. LCCN QA76.73.P98 D693 2009.

Downey:2015:TPH

[Dow15] Allen B. Downey. Think Python: How to Think Like a Computer Scientist. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-3936-2. ???? pp. LCCN ????

Demidov:2021:UPP

[DPF^{+}21] Alexander G. Demidov, B. Lakshitha A. Perera, Michael E. Fortunato, Sibo Lin, and Coray M. Colina. Update 1.1 to “pysimm: a Python package for simulation of molecular systems”, (PII: S2352711016300395). SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000777.

See [FC17].

Dagkakis:2016:MOS

[DPH16] Georgios Dagkakis, Ioannis Papagiannopoulos, and Cathal Heavey. ManPy: an open-source software tool for building discrete event simulation models of manufacturing systems. Software— Practice and Experience, 46 (7):955–981, July 2016. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

dePuiseau:2023:SPF

[dPPD^{+}23] Constantin Waubert de Puiseau, Jannik Peters, Christian Dörpelkus, Hasan Tercan, and Tobias Meisen. schlably: a Python framework for deep reinforcement learning based scheduling experiments. SoftwareX, 22(??): ??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000791.

Dalcin:2005:MP

[DPS05] Lisandro Dalcín, Rodrigo Paz, and Mario Storti. MPI for Python. Journal of Parallel and Distributed Computing, 65(9): 1108–1115, September 2005. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dalcin:2008:MPP

[DPSD08] Lisandro Dalcín, Rodrigo Paz, Mario Storti, and Jorge D’Elía. MPI for Python: Performance improvements and MPI-2 extensions. Journal of Parallel and Distributed Computing, 68(5): 655–662, May 2008. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Dorado-Rojas:2023:MMP

[DRFB^{+}23] Sergio A. Dorado-Rojas, Fernando Fachini, Tetiana Bogodorova, Giuseppe Laera, Marcelo de Castro Fernandes, and Luigi Vanfretti.

ModelicaGridData: Massive

power system simulation data generation and labeling tool using Modelica and Python. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001765.

Daole:2023:OXF

[DSB^{+}23] Mattia Daole, Alessio Schiavo, José Luis Corcuera Bárcena, Pietro Ducange, Francesco Marcelloni, and Alessandro Renda. OpenFL-XAI: Federated learning of explainable artificial intelligence models in Python. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002017.

Daloze:2018:PDL

[DTM^{+}18] Benoit Daloze, Arie Tal, Stefan Marr, Hanspeter Mössenböck, and Erez Petrank. Parallelization of dynamic languages: synchronizing built-in collections. Proceedings of the ACM on Programming Languages (PACMPL), 2(OOPSLA): 108:1–108:30, October 2018.

URL https://dl.acm.org/
doi/abs/10.1145/3276478.

Demo:2018:PPP

[DTR18] Nicola Demo, Marco Tezzele, and Gianluigi Rozza. PyDMD: Python Dynamic Mode Decomposition. Journal of Open Source Software, 3 (22):530:1–530:3, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00530.

Dubois:2005:NP

[Dub05] Paul F. Dubois. A nest of Pythons. Computing in Science and Engineering, 7(6): 81–84, November/December 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Dubois:2007:GEI

[Dub07] Paul F. Dubois. Guest Editor’s introduction: Python: Batteries included. Computing in Science and Engineering, 9(3):7–9, May/June 2007. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
comp/mags/cs/2007/03/c3007.
pdf.

Duffy:2015:LPT

[Duf15] Christopher Duffy. Learning Penetration Testing with Python. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-232-8, 1-78528-955-1 (e-book). 314 pp. LCCN T55.4-60.8.

Durr:2021:CPP

[DV21] Christoph Dürr and Jill-Jênn Vie. Competitive programming in Python: 128 algorithms to develop your coding skills. Cambridge University Press, Cambridge, UK, 2021. ISBN 1-108-71682-2 (paperback), 1-108-59192-2 (e-pub). LCCN QA76.73.P98. URL https:

//assets.cambridge.org/
97811087/16826/toc/9781108716826_

toc.pdf. Translation to English from the French original by Greg Gibbons and Danièle Gibbons.

Dubois:1996:EPO

[DY96a] P. F. Dubois and T.-Y. Yang. Extending Python [objectoriented language]. Computers in Physics, 10(4): 359–365, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.4822457.

Dubois:1996:SPE

[DY96b] Paul F. Dubois and T.-Y.

Yang. Scientific programming: Extending Python. Computers in Physics, 10(4): 359–??, ???? 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Dubois:1999:SPE

[DY99] Paul F. Dubois and T.-Y.

Yang. Scientific programming: Extending Python with Fortran. Computing in Science and Engineering, 1(5):66–73, September/October 1999. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

//dlib.computer.org/cs/
books/cs1999/pdf/c5066.
pdf;   http://www.computer.
org/cse/cs1999/c5066abs.
htm.

DeChavez:2021:OPM

[DyH21] Danjo De Chavez and Jun

ya Hasegawa. OpenMechanochem:

a Python module for mechanochemical simulations. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001436.

Escudero-Arnanz:2023:DPP

[EAMSR^{+}23] Óscar Escudero-Arnanz, Antonio G. Marques, Cristina Soguero-Ruiz, Inmaculada Mora-Jiménez, and Grego-

rio Robles. dtwParallel:

a Python package to efficiently compute dynamic time warping between time series. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000602.

Eckel:2022:PPP

[EBNS22] Stephen Eckel, Daniel S. Barker, Eric B. Norrgard, and Julia Scherschligt. PyLCP: a Python package for computing laser cooling physics. Computer Physics Communications, 270(??):Article 108166, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521002782.

Ethier:2023:PGA

[ECS23] Jeffrey G. Ethier, Andrés Córdoba, and Jay D. Schieber. pyDSM: GPU-accelerated rheology predictions for entangled polymers in Python. Computer Physics Communications, 290(??):Article 108786, September 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523001315.

Ernesti:2008:PUH

[EK08] Johannes Ernesti and Peter Kaiser. Python: das umfassende Handbuch; [aktuell zu Python 2.5; Einführung, Praxis, Referenz; Sprachgrundlagen, Objektorientierung, Modularisierung; Web-Programmierung mit Django, GUIs, Netzwerkkommunikation u.v.m.] . Galileo Press, Bonn, Germany, 2008. ISBN 3-8362-1110-6. ???? pp. LCCN ????

Ernesti:2009:PUH

[EK09] Johannes Ernesti and Peter Kaiser. Python 3: Das umfassende Handbuch. Galileo Computing. Galileo Press GmbH, Bonn, Germany, second edition, 2009. ISBN 3-8362-1412-1. 870 (est.) pp. LCCN ???? EUR 39.90.

Elsner:2023:TPP

[Els23] Jean Elsner. Taming the Panda with Python: a powerful duo for seamless robotics programming and integration. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002285.

Edhlund:2021:POO

[EML21] Ian Edhlund, Matthew Macauley, and Cindy Lee. PBTK optimizer: an open source application for PBTK model parameter optimization in Python. Journal of Open Research Software, 9(1):4–??, April 14, 2021. CODEN ????

ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.285/.

Eschle:2020:ZSP

[ENCS20] Jonas Eschle, Albert Puig Navarro, Rafael Silva Coutinho, and Nicola Serra. zfit: Scalable Pythonic fitting. SoftwareX, 11(??):Article 100508, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711019303851.

Elia:2023:PPL

[EPF^{+}23] Donatello Elia, Cosimo Palazzo, Sandro Fiore, Alessandro D’Anca, Andrea Mariello, and Giovanni Aloisio. PyOphidia: a Python library for high performance data analytics at scale. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023002340.

Enbody:2009:PCP

[EPM09] Richard J. Enbody, William F. Punch, and Mark McCullen. Python CS1 as preparation for C++ CS2. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1): 116–120, March 2009. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Ernst:1999:TMP

[Ern99] T. Ernst. TRAPping Modelica with Python. Lecture Notes in Computer Science, 1575:288–291, 1999. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ege:1998:PTO

[ESM98] R. Ege, M. Singh, and B. Meyer, editors. Proceedings. Technology of Object-Oriented Languages and Systems, TOOLS-23. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8383-X. LCCN ???? IEEE catalog number 97TB100221.

Ehrenberg:2024:PSC

[ESM24] Melanie Ehrenberg, Shahram Sarkani, and Thomas A. Mazzuchi. Python source code vulnerability detection with named entity recognition. Computers & Security, 140(??):??, May 2024. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167404824001032.

Esterbrook:2001:UMI

[Est01]  Chuck Esterbrook.  Us-

ing mix-ins with Python. Linux journal, 84:114, 116, 118, 120–121, April 2001. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue84/4540.
html.

EurOpen:1991:EUD

[Eur91]  EurOpen, editor.  Eu-

rOpen. UNIX Distributed Open Systems in Perspective. Proceedings of the Spring 1991 EurOpen Conference, Tromsø, Norway, May 20–24, 1991. EurOpen, Buntingford, Herts, UK, 1991. ISBN 1-873611-00-5. LCCN ????

Ettienne:2012:IMA

[EVV12] Mikko Berggren Ettienne, Steen Vester, and Jørgen Villadsen. Implementing a multi-agent system in Python with an auction-based agreement approach. Lecture Notes in Computer Science, 7217:185–196, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-31915-0_
11/.

Farrell:2002:MP

[Far02]  Doug Farrell.  Mediator/

Python. Linux journal, 98:??, June 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//www.linuxjournal.com/
article.php?sid=5858.

Fasnacht:2018:PMP

[Fas18] Laurent Fasnacht. mmappickle:

Python 3 module to store memory-mapped numpy array in pickle format. Journal of Open Source Software, 3(26):651:1–651:2, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00651.

Forcier:2009:PWD

[FBC09] Jeff Forcier, Paul Bissex, and Wesley Chun. Python Web development with Django. Developer’s library. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-13-235613-9 (paperback). 377 pp. LCCN TK5105.8885.D54 F68 2009.

Fortunato:2017:PPP

[FC17] Michael E. Fortunato and Coray M. Colina. pysimm: a Python package for simulation of molecular systems. SoftwareX, 6(??):1–6, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711016300395.

See update [DFC18].

Feeley:2016:CML

[Fee16] Marc Feeley. Compiling for multi-language task migration. ACM SIGPLAN Notices, 51(2):63–77, February 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fehily:2002:VQG

[Feh02] Chris Fehily. Visual quickstart guide: Python. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 2002. ISBN ???? xxvi + 410 pp. LCCN A76.73.P98 F44 2002. URL http://

                                        proquest.safaribooksonline.
                                        com/0201748843.

Filipovich:2022:POS

[FH22] Matthew J. Filipovich and Stephen Hughes. PyCharge: an open-source Python package for self-consistent electrodynamics simulations of Lorentz oscillators and moving point charges. Computer Physics Communications, 274(??):Article 108291, May 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522000091.

Furr:2009:PGS

[FhDAF09] Michael Furr, Jong hoon (David) An, and Jeffrey S. Foster. Profile-guided static typing for dynamic scripting languages. ACM SIG-PLAN Notices, 44(10):283– 300, October 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filguiera:2017:DPF

[FKA^{+}17] Rosa Filguiera, Amrey Krause, Malcolm Atkinson, Iraklis Klampanos, and Alexander

Moreno.  dispel4py: a

Python framework for dataintensive scientific computing. The International Journal of High Performance Computing Applications, 31 (4):316–334, July 2017. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Fettig:2006:TNP

[FL06] Abe Fettig and Glyph Lefkowitz.

Twisted network programming essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10032-9. xix + 213 pp. LCCN QA76.73.P98 F48 2005eb; QA76.73.P98. URL

http://www.oreilly.com/
catalog/9780596100322.

Feinberg:2015:COS

[FL15] Jonathan Feinberg and Hans Petter Langtangen. Chaospy: an open source tool for designing methods of uncertainty quantification. Journal of Computational Science, 11:46–57, November 2015. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750315300119.

Flaig:2008:BPP

[Fla08] Ruediger-Marcus Flaig. Bioinformatics programming in Python: a practical course for beginners. Wiley-VCH, Weinheim, Germany, 2008. ISBN 3-527-32094-6 (paperback). ix + 418 pp. LCCN ????

Fraccaroli:2022:SDT

[FLR22] Michele Fraccaroli, Evelina Lamma, and Fabrizio Riguzzi. Symbolic DNN-Tuner: a Python and ProbLog-based system for optimizing Deep Neural Networks hyperparameters. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021001825.

FiguerasiVentura:2020:PRT

[FLS^{+}20] Jordi Figueras i Ventura, Martin Lainer, Zaira Schauwecker, Jacopo Grazioli, and Urs Germann. Pyrad: a realtime weather radar data processing framework based on Py-ART. Journal of Open Research Software, 8 (1):28–??, October 08, 2020. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.330/.

Foreman-Mackey:2016:PCP

[FM16] Daniel Foreman-Mackey. corner.py: Scatterplot matrices in Python.

Journal of Open Source Software, 1(2):24:1–24:2, June 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00024.

Fichera:2017:PFP

[FMPS17] Loris Fichera, Fabrizio Messina, Giuseppe Pappalardo, and Corrado Santoro. A Python framework for programming autonomous robots using a declarative approach. Science of Computer Programming, 139(??):36–55, June 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167642317300242.

Filho:2022:IIPa

[FNX22] João Filho, Luiz Nunes, and

José Xavier. iCorrVision-3D:

an integrated Python-based open-source Digital Image Correlation Software for inplane and out-of-plane measurements (Part 2). SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102200084X.

Forouli:2022:APP

[FPN^{+}22] Aikaterini Forouli, Anastasios Pagonis, Alexandros Nikas, Konstantinos Koasidis, Georgios Xexakis, Themistoklis Koutsellis, Christos Petkidis, and Haris Doukas. AUGMECON-Py: a Python framework for multi-objective linear optimisation under uncertainty. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022001388.

Fredericks:2021:PPL

[FPSZ21] Scott Fredericks, Kevin Parrish, Dean Sayre, and Qiang Zhu. PyXtal: a Python library for crystal structure generation and symmetry analysis. Computer Physics Communications, 261(??): Article 107810, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520304057.

Fiorio:2021:IPP

[FRdN21] Luan Vinícius Fiorio, Chrystian Lenon Remes, and Yales Rômulo de Novaes. impulseest: a Python package for non-parametric impulse response estimation with input–output data. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021000832.

Ferreira-Saraiva:2023:PPP

[FSMCFP23] Bruno D. Ferreira-Saraiva, João P. Matos-Carvalho, Nuno Fachada, and Manuel Pita. ParShift: a Python package to study order and differentiation in group conversations. SoftwareX, 24 (??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002509.

Frison:2020:BAB

[FSZD20] Gianluca Frison, Tommaso Sartor, Andrea Zanelli, and Moritz Diehl. The BLAS API of BLASFEO: Optimizing performance for small matrices. ACM Transactions on Mathematical Software, 46(2):15:1–15:36, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3378671.

Fulton:1996:IPP

[Ful96] Jim Fulton. Introduction to the Python programming language. In USENIX [USE96], page ?? ISBN 1-880446-77-4. LCCN QA76.64 .U85

1996. URL http://www.
usenix.org/publications/
library/proceedings/coots96/

.

Funk:2023:CLC

[Fun23] Dan Funk. Creating a lowcode business process execution platform with Python, BPMN, and DMN. IEEE Software, 40(1):9–17, January/February 2023. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Fleming:2018:PAA

[FV18] David P. Fleming and Jake

                                        VanderPlas. approxposterior:

Approximate posterior distributions in Python. Journal of Open Source Software, 3(29):781:1–781:2, September 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00781.

Gaddis:2009:SP

[Gad09]  Tony Gaddis.  Starting

out with Python. Pearson Addison Wesley, Boston, MA, USA, 2009. ISBN 0-321-53711-4. xv + 482 pp. LCCN QA76.73.P98 G34 2009. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        fy0804/2008001684.html.

Galanakis:2014:PMP

[Gal14] Robert Galanakis. Practical Maya programming with Python: unleash the power of Python in Maya and unlock your creativity. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-84969-472-9 (paperback), 1-84969-473-7 (e-book), 1-322-00849-3 (e-book). viii + 334 pp. LCCN T385 .G34 2014. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781849694728.

Galassi:2022:PPP

[Gal22] Riccardo Malpica Galassi.

PyCSP: a Python package for the analysis and simplification of chemically reacting systems based on Computational Singular Perturbation. Computer Physics Communications, 276(??): Article 108364, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522000832.

Garberson:1998:LEP

[Gar98] John D. Garberson. Letter to the Editor: Programming Python. ;login: the USENIX Association newsletter, 23(4): ??, June 1998. CODEN LOGNEM. ISSN 1044-6397.

Gardner:2009:DGP

[Gar09] James Gardner. The definitive guide to Pylons: [Pylons is a lightweight web framework emphasizing flexibility and rapid development using standard tools from the Python community; includes SQLAlchemy, JavaScript, and WSG!]. The expert’s voice in web development; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2009. ISBN 1-59059-934-9 (paperback). xxv + 536 pp. LCCN ???? US$46,99.

Gutierrez:2016:IDO

[GAS^{+}16] Christopher N. Gutierrez, Mohammed H. Almeshekah, Eugene H. Spafford, Mikhail J. Atallah, and Jeff Avery. Inhibiting and detecting offline password cracking using ErsatzPasswords. ACM Transactions on Privacy and Security (TOPS), 19(3):9:1–9:??, December 2016. CODEN ???? ISSN 2471-2566 (print), 2471-2574 (electronic).

Gonzalez-Cely:2023:RPB

[GCBDDBF23] Aura Ximena Gonzalez-Cely, Cristian Felipe Blanco-Diaz, Camilo A. R. Diaz, and Teodiano Freire Bastos-Filho. Roborueda: Python-based GUI to control a wheelchair and monitor user posture. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002510.

Gins:2018:ACD

[GdGB^{+}18] W. Gins, R. P. de Groote, M. L. Bissell, C. Granados Buitrago, R. Ferrer, K. M. Lynch, G. Neyens, and S. Sels. Analysis of counting data: Development of the SATLAS Python package. Computer Physics Communications, 222(??):286– 294, January 2018. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465517302990.

Grady:2018:PPP

[GDP18] Maxwell Grady, Zhongwei Dai, and Karsten Pohl. PLEASE: The Python Lowenergy Electron Analysis SuitE — enabling rapid analysis of LEEM and LEED data. Journal of Open Research Software, 6(1): 7–??, February 05, 2018. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.191/.

Graichen:2019:SPT

[GEH19] Uwe Graichen, Roland Eichardt, and Jens Haueisen. SpharaPy: a Python toolbox for spatial harmonic analysis of non-uniformly sampled data. SoftwareX, 10(??):Article 100289, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019301670.

Gezerlis:2020:NMPb

[Gez20]  Alex Gezerlis.  Numeri-

cal Methods in Physics with Python. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-48884-6 (hardcover), 1-108-73893-1 (paperback), 1-108-77231-5 (e-pub). ???? pp. LCCN QC20.7.N86 G49 2020.

Gezerlis:2023:NMP

[Gez23]  Alex Gezerlis.  Numeri-

cal Methods in Physics with Python. Cambridge University Press, Cambridge, UK, second edition, 2023. ISBN 1-009-30385-6 (hardcover), 1-009-30386-4 (paperback). 700 pp. LCCN QC20.7.N86 G49 2023.

Goyal:2018:PGP

[GF18] Palash Goyal and Emilio Ferrara. GEM: a Python package for graph embedding methods. Journal of Open Source Software, 3 (29):876:1–876:2, September 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00876.

Gonina:2014:SMC

[GFB^{+}14] Ekaterina Gonina, Gerald Friedland, Eric Battenberg, Penporn Koanantakool, Michael Driscoll, Evangelos Georganas, and Kurt Keutzer. Scalable multimedia content analysis on parallel platforms using Python. ACM Transactions on Multimedia Computing, Communications, and Applications, 10(2):18:1–18:??, February 2014. CODEN ???? ISSN 1551-6857 (print), 1551-6865 (electronic).

Gunaratne:2021:NAB

[GG21] Chathika Gunaratne and Ivan Garibay. NL4Py: Agentbased modeling in Python with parallelizable NetLogo workspaces. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001035.

Gagunashvili:2018:CCG

[GH18] Nikolay D. Gagunashvili and Helgi Halldorsson. CHIWEI: a code of goodness of fit tests for weighted and unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 231(??):245, October 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465518301449.

Gorjao:2022:MEM

[GHKW22] Leonardo Rydin Gorjão, Galib Hassan, Jürgen Kurths, and Dirk Witthaut. MFDFA: Efficient multifractal detrended fluctuation analysis in Python. Computer Physics Communications, 273(??): Article 108254, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521003660.

Gagunashvili:2019:CCC

[GHN19] Nikolay D. Gagunashvili, Helgi Halldorsson, and Helmut Neukirchen. CHICOM:

Code for comparing weighted or unweighted histograms in Fortran-77, C++, R and Python. Computer Physics Communications, 245(??): Article 106872, December 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465519302590.

Goebel:2006:BPT

[GHT06] John A. Goebel, Adil Hasan, and Francesco Safai Tehran. The book of Python: from the tip of the tongue to the end of the tale. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-103-4 (paperback). 1000 pp. LCCN QA76.73.P98. URL http:/

                                        /www.loc.gov/catdir/toc/
                                        ecip064/2005034382.html.

Goebel:2008:BPR

[GHT08] John A. Goebel, Adil Hasan, and Francesco Safai Tehrani. The book of Python: a real-world reference. No Starch Press, San Francisco, CA, USA, 2008. ISBN 1-59327-103-4. ???? pp. LCCN QA76.73.P98 G62

                                        2008. URL http://www.loc.
                                        gov/catdir/toc/ecip064/
                                        2005034382.html.

Gao:2024:CDL

[GHXZ24] Kai Gao, Runzhi He, Bing Xie, and Minghui Zhou. Characterizing deep learning package supply chains in PyPI: Domains, clusters, and disengagement. ACM Transactions on Software Engineering and Methodology, 33(4):97:1–97:??, May 2024. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3640336.

Giroux:2021:TPP

[Gir21] Bernard Giroux. ttcrpy: a Python package for traveltime computation and raytracing. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001217.

Gift:2008:PUL

[GJ08] Noah Gift and Jeremy M.

Jones. Python for Unix and Linux system administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2008. ISBN 0-596-51582-0. xix + 433 pp. LCCN ????

Gharari:2023:EPP

[GKK^{+}23] Shervan Gharari, Kasra Keshavarz, Wouter J. M. Knoben, Gouqiang Tang, and Martyn P. Clark. EASYMORE: a Python package to streamline the remapping of variables for Earth System models. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023002431.

Goldwasser:2007:INP

[GL07a] Michael H. Goldwasser and David Letscher. Introducing network programming into a CS1 course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):19–22, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Goldwasser:2007:TOO

[GL07b] Michael H. Goldwasser and David Letscher. Teaching object-oriented programming in Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):365– 366, September 2007. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Goldwasser:2008:OOP

[GL08a] Michael H. Goldwasser and David Letscher. Objectoriented programming in Python. Pearson Prentice Hall, Upper Saddle River, NJ, USA, 2008. ISBN 0-13-615031-4. xxii + 666 pp. LCCN QA76.73.P98.G65; QA76.73.P98.G65 2008.

Goldwasser:2008:PGP

[GL08b] Michael H. Goldwasser and David Letscher. A Python graphics package for the first day and beyond. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 326, September 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Goldwasser:2008:TOO

[GL08c] Michael H. Goldwasser and David Letscher. Teaching an object-oriented CS1 -: with Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):42–46, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Goldwasser:2009:GPF

[GL09] Michael H. Goldwasser and David Letscher. A graphics package for the first day and beyond. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):206–210, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Gorbovitski:2010:AAO

[GLS^{+}10] Michael Gorbovitski, Yanhong A. Liu, Scott D. Stoller, Tom Rothamel, and Tuncay K. Tekle. Alias analysis for optimization of dynamic languages. ACM SIG-PLAN Notices, 45(12):27–42, December 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gaete-Morales:2021:DPF

[GMKRS21] Carlos Gaete-Morales, Martin Kittel, Alexander Roth, and Wolf-Peter Schill. DIETERpy: a Python framework for the Dispatch and Investment Evaluation Tool with Endogenous Renewables. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021000947.

Garcia:2021:IHG

[GMN21] Andrés Amaya García, David May, and Ed Nutting. Integrated hardware garbage collection. ACM Transactions on Embedded Computing Systems, 20(5):40:1–40:25, July 2021. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450147.

Gao:2020:JLM

[GMP^{+}20] Kaifeng Gao, Gang Mei, Francesco Piccialli, Salvatore Cuomo, Jingzhi Tu, and Zenan Huo. Julia language in machine learning: Algorithms, applications, and open issues. Computer Science Review, 37: 100254, August 2020. ISSN 1574-0137 (print), 1876-7745 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S157401372030071X.

Gavran:2017:AMR

[GMS17] Ivan Gavran, Rupak Majumdar, and Indranil Saha. Antlab: a multi-robot task server. ACM Transactions on Embedded Computing Systems, 16(5s):190:1–190:??, October 2017. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Goona:2023:DOS

[GNPP23] Nithin Kumar Goona, Shraddha M. Naik, Saidi Reddy Parne, and Anand Paul. DssPyLib: an open-source Python FEM software to solve Poisson equation in 2-D using distributed source scheme. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000043.

Gorelick:2014:HPP

[GO14] Micha Gorelick and Ian Ozsvald. High performance Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 1-4493-6159-5 (paperback). xv + 351 pp. LCCN QA76.73.P98 G67 2014.

Govindaraj:2015:TDP

[Gov15] Siddharta Govindaraj. Testdriven Python development. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-792-8, 1-78398-793-6 (e-book). 264 pp. LCCN QA76.73.P98.

Giuffre:2022:NIP

[GP22] Andrea Giuffre’ and Matteo Pini. NiceProp: an interactive Python-based educational tool for non-ideal compressible fluid dynamics. SoftwareX, 17(??):??, January 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021001527.

Gomes:2023:PPT

[GPEM23] Eduardo Gomes, Lucas Pereira, Augusto Esteves, and Hugo Morais. PyECOM: a Python tool for analyzing and simulating Energy Communities. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002765.

Gonzalez-Perez:2020:PPP

[GPKL^{+}20] V. Gonzalez-Perez, P. Keil, Y. Li, A. Zülke, R. Burrel, D. Csala, and H. Hoster. A Python package to preprocess the data produced by novonix high-precision battery-testers. Journal of Open Research Software, 8 (1):3–??, March 04, 2020. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.281/.

Gotze:2021:UFP

[GPP^{+}21] Jan P. Götze, Yuan-Wei Pi, Simon Petry, Fabian Langkabel, Jan Felix Witte, and Oliver Lemke. A userfriendly, Python-based quantum mechanics/Gromacs interface: gmx2qmmm. International Journal of Quantum Chemistry, 121(3):e26486:1– e26486:??, February 5, 2021. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gladstein:2018:SPF

[GQCP^{+}18] Ariella L. Gladstein, Consuelo D. Quinto-Cortés, Julian L. Pistorius, David Christy, Logan Gantner, and Blake L. Joyce. SimPrily: a Python framework to simplify high-throughput genomic simulations. SoftwareX, 7 (??):335–340, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018301213.

Grayson:2000:PTP

[Gra00] John E. Grayson. Python and Tkinter Programming. Manning Publications, Greenwich, CT, USA, 2000. ISBN 1-884777-81-3. xxiii + 658 pp. LCCN ???? US$49.95.

Gray:2018:PQP

[Gra18] Johnnie Gray. quimb: a Python package for quantum information and manybody calculations. Journal of Open Source Software, 3(29):819:1–819:3, September 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00819.

Greenfield:2007:RSP

[Gre07] Perry Greenfield. Reaching for the stars with Python. Computing in Science and Engineering, 9(3): 38–40, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Greenhill:2016:PPP

[Gre16] Simon J. Greenhill. Phylogemetric: a Python library for calculating phylogenetic network metrics. Journal of Open Source Software, 1(2):28:1, June 2016. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00028.

Green:2018:PDP

[Gre18] Gregory M. Green. dustmaps:

a Python interface for maps of interstellar dust. Journal of Open Source Software, 3(26):695:1–695:2, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00695.

Gronchi:2021:NTS

[GRG21] Giorgio Gronchi, Marco Raglianti, and Fabio Giovannelli. Network theory and switching behaviors: a user guide for analyzing electronic records databases. Future Internet, 13(9):228, August 31, 2021. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/
13/9/228.

Gregor:2019:PGP

[GRKN^{+}19] Luke Gregor, Thomas J.

Ryan-Keogh, Sarah-Anne Nicholson, Marcel du Plessis, Isabelle Giddy, and Sebasti-

aan Swart. GliderTools: a

Python toolbox for processing underwater glider data. Frontiers in Marine Science, 6, December 2019. ISSN 2296-7745.

Groce:2020:Pc

[Gro20] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 45(3):4–5, July 2020. URL https:/

/dl.acm.org/doi/10.1145/
3402127.3402129.

Gerum:2019:CPP

[GRW^{+}19] Richard C. Gerum, Sebastian Richter, Alexander Winterl, Christoph Mark, Ben Fabry, Céline Le Bohec, and Daniel P. Zitterbart.

CameraTransform: a Python

package for perspective corrections and image mapping. SoftwareX, 10(??):Article 100333, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019302018.

Gasiorek:2012:OPP

[GS12] Marcin Gaşiorek and Daniel Simson. One-peak posets with positive quadratic Tits form, their mesh translation quivers of roots, and programming in Maple and Python. Linear Algebra and its Applications, 436(7): 2240–2272, April 1, 2012. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0024379511007555.

Grilli:2021:NPS

[GTC21] Nicolò Grilli, Edmund Tarleton, and Alan C. F. Cocks.

Neper2CAE and PyCiGen:

Scripts to generate polycrystals and interface elements in Abaqus. SoftwareX, 13(??):Article 100651, January 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711020303642.

Guelton:2018:PCP

[Gue18] Serge Guelton. Pythran:

Crossing the Python frontier. Computing in Science and Engineering, 20(2): 83–89, ???? 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

https://ieeexplore.ieee.
org/document/8317992/.

Gundecha:2014:LST

[Gun14] Unmesh Gundecha. Learning Selenium testing tools with Python. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-350-7, 1-78398-351-5 (e-book), 1-322-56850-2 (e-book). 216 (est.) pp. LCCN QA76.73.P98.

Gupta:2015:BWA

[Gup15]  Sumit Gupta.  Building

web applications with Python and Neo4j: develop exciting and real-world Pythonbased web applications with Neo4j using frameworks such as Flask, Py2neo, and Django. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-398-1 (print), 1-78398-399-X (ebook). 184 (est.) pp. LCCN QA76.76.A65. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781783983988.

Gutschmidt:2004:GPP

[Gut04] Tom Gutschmidt. Game programming with Python, Lua, and Ruby. Premier Press, Boston, MA, USA, 2004. ISBN 1-59200-077-0, 1-59200-408-3 (e-book). xxvi + 437 pp. LCCN QA76.76.C672 G88 2004b. URL http://

                                        proquest.safaribooksonline.
                                        com/1592000770.

Guttag:2013:ICP

[Gut13] John Guttag. Introduction to Computation and Programming Using Python. MIT Press, Cambridge, MA, USA, 2013. ISBN 0-262-52500-3 (paperback), 0-262-31219-0. xiv + 298 pp. LCCN QA76.73.P98 G88 2013.

Guttag:2016:ICP

[Gut16] John Guttag. Introduction to Computation and Programming Using Python: with Application to Understanding Data. MIT Press, Cambridge, MA, USA, second edition, 2016. ISBN 0-262-52962-9 (paperback), 0-262-33738-X (e-book). xvii + 447 pp. LCCN QA76.73.P98 G88 2016.

Guzdial:2003:MCC

[Guz03] Mark Guzdial. A media computation course for nonmajors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(3):104–108, September 2003. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gvero:2009:BRPb

[Gve09] Igor Gvero. Book review:

Python for Software Design by Allen B. Downey, and published by Cambridge University Press, 2009, 978-0-521-72596-5, 251pp. ACM SIG-SOFT Software Engineering Notes, 34(6):31–32, November 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Godoy:2023:EPP

[GVLD^{+}23] William F. Godoy, Pedro Valero-Lara, T. Elise Dettling, Christian Trefftz, Ian Jorquera, Thomas Sheehy, Ross G. Miller, Marc Gonzalez-Tallada, Jeffrey S. Vetter, and Valentin Churavy. Evaluating performance and portability of high-level programming models: Julia, Python/Numba, and Kokkos on exascale nodes. In IEEE, editor, 2023 IEEE International Parallel and Distributed Processing Symposium Workshops (IPDPSW), pages 373–382. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Gieseke:2018:PPP

[GWM18] Robert Gieseke, Sven N. Willner, and Matthias Mengel. Pymagicc: a Python wrapper for the simple climate model MAGICC. Journal of Open Source Software, 3 (22):516:1–516:3, February 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00516.

Guyer:2009:FPD

[GWW09] Jonathan E. Guyer, Daniel Wheeler, and James A. Warren. FiPy: Partial differential equations with Python. Computing in Science and Engineering, 11(3): 6–15, May/June 2009. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Guo:2023:OPP

[GYWG23] Junjun Guo, Aijun Ye, Xiaowei Wang, and Zhong-

                                        guo Guan. OpenSeesPyView:

Python programming-based visualization and post-processing tool for OpenSeesPy. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022001960.

Goldbaum:2018:PUH

[GZT^{+}18] Nathan J. Goldbaum, John A. ZuHone, Matthew J. Turk, Kacper Kowalik, and Anna L. Rosen. unyt: Handle, manipulate, and convert data with units in Python. Journal of Open Source Software, 3(28):809:1–809:11, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00809.

Herrmann:2020:HRF

[HA20] Julien Herrmann and Guillaume Pallez (Aupy). H-Revolve: a framework for adjoint computation on synchronous hierarchical platforms. ACM Transactions on Mathematical Software, 46(2):12:1–12:25, June 2020. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3378672.

Holmes:2020:URL

[HAB^{+}20] Josie Holmes, Iftekhar Ahmed, Caius Brindescu, Rahul Gopinath, He Zhang, and Alex Groce. Using relative lines of code to guide automated test generation for Python. ACM Transactions on Software Engineering and Methodology, 29 (4):28:1–28:38, October 2020. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3408896.

Hacihabiboglu:2023:SPP

[Hac23a] Hüseyin Hacihabiboglu. sphstat:

a Python package for inferential statistics on vectorial data on the unit sphere.

SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023002108.

Hanasaki:2023:IRT

[HAC^{+}23b] Kota Hanasaki, Zulfikhar A. Ali, Min Choi, Mauro Del Ben, and Bryan M. Wong. Implementation of real-time TDDFT for periodic systems in the open-source PySCF software package. Journal of Computational Chemistry, 44 (9):980–987, April 5, 2023. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hardt:2000:PPZ

[HAE00] Dick Hardt, Gisle Aas, and Paul Everitt. Perl, Python and Zope. In USENIX [USE00], page ?? ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/
                                        aas.html.

Haghighi:2017:PPP

[Hag17] Sepand Haghighi. Pyrgg: Python Random Graph Generator. Journal of Open Source Software, 2(17):331:1– 331:2, September 2017. CO-DEN ???? ISSN 2475-9066.

                                        URL http://joss.theoj.
                                        org/papers/10.21105/joss.
                                        00331.

Huang:2018:ROO

[HAGH18] Wen Huang, P.-A. Absil, Kyle A. Gallivan, and Paul Hand. ROPTLIB: an object-oriented C++ library for optimization on Riemannian manifolds. ACM Transactions on Mathematical Software, 44(4):43:1– 43:21, August 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3218822.

Hajji:2008:PPE

[Haj08] Farid Hajji. Das Python Praxisbuch — eBook: Der große Profi-Leitfaden für Programmierer. Addison Wesley in Pearson Education Deutschland, München, Germany, 2008. ISBN 3-8273-6182-6. 1328 (est.) pp. LCCN ????

Hamilton:1998:PHM

[Ham98]  Michael Hamilton.  The

Python HTMLgen module. Linux journal, 55:22, 24– 26, November 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL ftp://ftp.

ssc.com/pub/lj/listings/
issue55/2986.tgz.

Handley:2018:PFP

[Han18] Will Handley. fgivenx: a Python package for functional posterior plotting. Journal of Open Source Software, 3(28):849:1–849:4, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00849.

Harris:2007:GPL

[Har07] Andy Harris. Game programming: the L Line: the express line to learning. Wiley, New York, NY, USA, 2007. ISBN 0-470-06822-1 (paperback). xxv + 570 pp. LCCN QA76.76.C672.

URL http://www.loc.gov/
catdir/enhancements/fy0741/
2006936755-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006936755-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0741/2006936755-
t.html.

Harwani:2012:IPP

[Har12] B. M. Harwani. Introduction to Python programming and developing GUI applications with PyQT. Cengage Learning, Boston, MA, USA, 2012. ISBN 1-4354-6097-9, 1-4354-6098-7 (e-book). xv + 393 pp. LCCN QA76.73.P98 H37 2012aeb.

Harris:2015:CSP

[Har15] Naftali Harris. Code snippet: LazySorted: a lazily, partially sorted Python list. Journal of Statistical Software, 65(CS-1):??, June 2015. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v65/
c01.

Haslwanter:2016:ISP

[Has16]  Thomas Haslwanter.  An

Introduction to Statistics with Python: With Applications in the Life Sciences. Statistics and Computing. Springer International Publishing, Cham, Switzerland, 2016. ISBN 3-319-28315-4 (hardcover), 3-319-28316-2 (e-book). ISSN 1431-8784. xvii + 278 pp. LCCN QA276.4 .H38 2016.

Holtgrewe:2016:PVP

[HB16] Manuel Holtgrewe and Dieter Beule. VCFPy: a Python 3 library with good support for both reading and writing VCF. Journal of Open Source Software, 1(6):85:1, October 2016. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00085.

Herring:2020:BPL

[HBA^{+}20] Patrick Herring, Chirranjeevi Balaji Gopal, Muratahan Aykol, Joseph H. Montoya, Abraham Anapolsky, Peter M. Attia, William Gent, Jens S. Hummelshøj, Linda Hung, Ha-Kyung Kwon, Patrick Moore, Daniel Schweigert, Kristen A. Severson, Santosh Suram, Zi Yang, Richard D. Braatz, and Brian D. Storey. BEEP: a Python library for battery evaluation and early prediction. SoftwareX, 11 (??):Article 100506, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711020300492.

Hadjidoukas:2020:TST

[HBS^{+}20] P. E. Hadjidoukas, A. Bartezzaghi, F. Scheidegger, R. Istrate, C. Bekas, and A. C. I. Malossi. torcpy: Supporting task parallelism in Python. SoftwareX, 12(??):Article 100517, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711020300091.

Helmus:2016:PAR

[HC16] Jonathan Helmus and Scott Collis. The Python ARM Radar Toolkit (Py-ART), a library for working with weather radar data in the Python programming language. Journal of Open Research Software, 4(1):e25–??, July 18, 2016. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.119/.

Huang:1995:CEM

[HCPF95] C. C. Huang, G. S. Couch, E. F. Pettersen, and T. E. Ferrin. Chimera: an extensible molecular modeling application constructed using standard components. In Hunter and Klein [HK95], pages 724–?? ISBN 981-02-2578-4. LCCN QH323.5.P33

1996. URL http://www.cgl.
ucsf.edu/psb/psb96/.

Henderson:2008:AMC

[Hen08] Peter B. Henderson. Abstraction, model checking and software correctness. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(2):23–24, June 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL ftp:

//ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/
bibliography/Misc/DBLP/
2008.bib.

Herman:2014:FSC

[Her14]  Ted Herman.  A func-

tional start to computing with Python. Chapman and Hall/CRC textbooks in computing. CRC Press, Taylor and Francis, Boca Raton, FL, USA, 2014. ISBN 1-4665-0455-2 (paperback). xiv + 415 pp. LCCN QA76.73.P98 H47 2014.

Hetland:2002:PP

[Het02] Magnus Lie Hetland. Practical Python. Apress, Berkeley, CA, USA, 2002. ISBN 1-59059-006-6. xxi + 619 pp. LCCN QA76.73.P98 H47 2002.

Hetland:2005:BPN

[Het05] Magnus Lie Hetland. Beginning Python: from novice to professional. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-519-X. xxx + 604 pp. LCCN QA76.73.P98 H48 2005.

Hetland:2008:BPN

[Het08] Magnus Lie Hetland. Beginning Python: from novice to professional. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-59059-982-9. xxx + 656 pp. LCCN A76.73.P98 H48 2008eb.

Hetland:2010:PAM

[Het10] Magnus Lie Hetland. Python Algorithms: mastering basic algorithms in the Python Language. Expert’s voice in open source. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-3237-4. xvi + 316 pp. LCCN QA76.73.P98 H485

                                        2010. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1502/2011287235-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1502/2011287235-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/fy11pdf04/
                                        2011287235.html.

Hetland:2014:PAM

[Het14] Magnus Lie Hetland. Python Algorithms: Mastering Basic Algorithms in the Python Language. EBL-Schweitzer. Apress, Berkeley, CA, USA, second edition, 2014. ISBN 1-4842-0055-1. 303 (est.) pp. LCCN QA75.5-76.95.

Harrison:2006:MSP

[HF06] Guy Harrison and Steven Feuerstein. MySQL stored procedure programming: building high-performance web applications with PHP, Perl, Python, Java & .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 0-596-10089-2. xxiii + 609 pp. LCCN QA76.73.S67 H377 2006eb; QA76.73.S67. URL

http://www.oreilly.com/
catalog/9780596100896.

Hahne:2021:APP

[HFB21] Jens Hahne, Stephanie Friedhoff, and Matthias Bolten. Algorithm 1016: PyMGRIT: a Python package for the parallel-in-time method MGRIT. ACM Transactions on Mathematical Software, 47(2):19:1– 19:22, April 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3446979.

Heusser:2017:PQP

[HFF^{+}17] Andrew C. Heusser, Paxton C. Fitzpatrick, Campbell E. Field, Kirsten Ziman, and Jeremy R. Manning. Quail: a Python toolbox for analyzing and plotting free recall data. Journal of Open Source Software, 2(18):424:1–424:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00424.

Hadjimichael:2020:RPL

[HGHR20] Antonia Hadjimichael, David Gold, David Hadka, and Patrick Reed. Rhodium: Python library for manyobjective robust decision making and exploratory modeling. Journal of Open Research Software, 8(1):12–??, June 09, 2020. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.293/.

Hammer:1997:ESI

[HGMC^{+}97] J. Hammer, H. Garcia-Molina, J. Cho, R. Aranha, Crespo, and A. Extracting semistructured information from the Web. In Anonymous [Ano97d], pages 18–25. ISBN ???? LCCN ????

Hoyer:2017:PXL

[HH17] Stephan Hoyer and Joe Hamman. xarray: N-D labeled arrays and datasets in Python. Journal of Open Research Software, 5(1):10–??, April 05, 2017. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.148/.

Haidri:2022:PPP

[HHB^{+}22] Salman Haidri, Yaksh J. Haranwala, Vania Bogorny, Chiara Renso, Vinicius Prado da Fonseca, and Amilcar Soares. PTRAIL — a Python package for parallel trajectory data preprocessing. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001066.

Ho:2021:DSS

[HHJCRB21] Anson T. Y. Ho, Kim P.

Huynh, David T. Jacho-Chávez, and Diego Rojas-Baez. Data science in Stata 16: Frames, lasso, and Python integration. Journal of Statistical Software, 98 (??):??, ???? 2021. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v098s01;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v098s01/v98s01.pdf.

Hambrusch:2009:MAT

[HHK^{+}09] Susanne Hambrusch, Christoph Hoffmann, John T. Korb, Mark Haugan, and Antony L. Hosking. A multidisciplinary approach towards computational thinking for science majors. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(1):183–187, March 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of SIGCSE ’09.

Holmgren:2018:PPP

[HHM18] William F. Holmgren, Clifford W. Hansen, and Mark A.

                                        Mikofski. pvlib python: a

Python package for modeling solar energy systems. Journal of Open Source Software, 3(29):884:1–884:3, September 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00884.

Hynninen:2016:OOP

[HHP^{+}16] T. Hynninen, L. Himanen, V. Parkkinen, T. Musso, J. Corander, and A. S. Foster. An object oriented Python interface for atomistic simulations. Computer Physics Communications, 198(??): 230–237, January 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465515003483.

He:2021:TPP

[HHVB21] Xu He, Nicole Helbig, Matthieu J. Verstraete, and Eric Bousquet. TB2J: a Python package for computing magnetic interaction parameters. Computer Physics Communications, 264(??): Article 107938, July 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521000679.

Hiam:2015:LBP

[Hia15]  Alexander Hiam.  Learn-

ing BeagleBone Python programming: unleash the potential of BeagleBone using Python. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-970-1, 1-78439-080-1 (e-book). ???? pp. LCCN QA76.73.P98.

Hightower:2003:PPJ

[Hig03] Richard Hightower. Python programming with the Java class libraries: a tutorial for building Web and Enterprise applications with Jython. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-61616-5. xii + 620 pp. LCCN QA76.73.P98 H54 2003.

Higgott:2022:PPP

[Hig22] Oscar Higgott. PyMatching: a Python package for decoding quantum codes with minimum-weight perfect matching. ACM Transactions on Quantum Computing (TQC), 3(3):16:1–16:16, September 2022. CODEN ???? ISSN 2643-6809 (print), 2643-6817 (electronic). URL

https://dl.acm.org/doi/
10.1145/3505637.

Hill:2015:LSP

[Hil15a]  Christian Hill.  Learning

scientific programming with Python. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-07541-6 (hardcover), 1-107-42822-X (paperback). vii + 452 pp. LCCN Q183.9 .H58 2015.

Hilpisch:2015:PF

[Hil15b] Yves J. Hilpisch. Python for finance. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-4528-1 (paperback), 1-4919-4539-7 (e-book). xv + 586 pp. LCCN HG176.5 H55 2015.

Hill:2020:LSPe

[Hil20]  Christian Hill.  Learning

scientific programming with Python. Cambridge University Press, Cambridge, UK, second edition, 2020. ISBN 1-108-74591-1. xi + 557 pp. LCCN Q183.9 .H58 2020.

Himpe:2023:EEG

[Him23] Christian Himpe. emgr — EMpirical GRamian framework version 5.99. ACM Transactions on Mathematical Software, 49(3):31:1– 31:??, September 2023. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3609860.

Hinsen:2002:HLS

[Hin02] K. Hinsen. High-level scientific programming with Python. Lecture Notes in Computer Science, 2331:691– ??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2331/23310691.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2331/23310691.
pdf.

Hinsen:2003:HLP

[Hin03] Konrad Hinsen. High-level parallel software development with Python and BSP. Parallel Processing Letters, 13(3): 473–??, September 2003. CO-DEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Hinsen:2007:PSP

[Hin07]  Konrad Hinsen.  Parallel

scripting with Python. Computing in Science and Engineering, 9(6):82–89, November/December 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Haghighi:2018:PPM

[HJHZ18] Sepand Haghighi, Masoomeh Jasemi, Shaahin Hessabi, and Alireza Zolanvari. PyCM: Multiclass confusion matrix library in Python. Journal of Open Source Software, 3(25):729:1–729:2, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00729.

Heinrich:2022:ERP

[HJJ^{+}22] G. Heinrich, S. Jahn, S. P. Jones, M. Kerner, F. Langer, V. Magerya, A. Põldaru, J. Schlenk, and E. Villa. Expansion by regions with pySecDec. Computer Physics Communications, 273(??): Article 108267, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521003799.

Heinrich:2024:NSA

[HJK^{+}24] G. Heinrich, S. P. Jones, M. Kerner, V. Magerya, A. Olsson, and J. Schlenk. Numerical scattering amplitudes with pySecDec. Computer Physics Communications, 295(??):Article 108956, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465523003016.

Hazelton:2017:PPI

[HJPB17] Bryna J. Hazelton, Daniel C. Jacobs, Jonathan C. Pober, and Adam P. Beardsley. pyuvdata: an interface for astronomical interferometeric datasets in Python. Journal of Open Source Software, 2(10):140:1, February 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00140.

Hunter:1995:PSB

[HK95] Lawrence Hunter and Teri E.

Klein, editors. Pacific Symposium on Biocomputing ’96: Hawaii, USA, 3–6 January, 1996. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2578-4. LCCN QH323.5.P33 1996. URL

http://www.cgl.ucsf.edu/
psb/psb96/.

Hosmer:2015:PPN

[HK15] Chet Hosmer and Gary C.

Kessler, editors. Passive Python network mapping: P2NMAP. Syngress Publishing, Inc., Rockland, MA, USA, 2015. ISBN 0-12-802721-5, 0-12-802742-8 (ebook). LCCN TK5105.59 .H67 2015.

Hoffmann:2021:PPC

[HKGvS21] Christoph G. Hoffmann, George N. Kiladis, Maria Gehne, and Christian von Savigny. A Python package to calculate the OLRbased index of the Madden– Julian-Oscillation (OMI) in climate science and weather forecasting. Journal of Open Research Software, 9(1):9–??, May 14, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.331/.

Holovaty:2008:DGD

[HKM08] Adrian Holovaty and Jacob Kaplan-Moss. The Definitive Guide to Django: Web Development Done Right. Springer eBook Collection Professional and Applied Computing [Dig. Serial]; Springer-12059 [Dig. Serial]. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-725-7. xxxii + 447 pp. LCCN TK5105.888.

Hughes:2015:PSS

[HLR15] Adam Hughes, Zhaowen Liu,

                                        and M. Reeves. Scikit-spectra:

Explorative spectroscopy in Python. Journal of Open Research Software, 3(1):e6–??, June 05, 2015. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.bs/.

Hinsen:2006:UBP

[HLSØ06] Konrad Hinsen, Hans Petter Langtangen, Ola Skavhaug, and Åsmund Ødegård. Using BSP and Python to simplify parallel programming. Future Generation Computer Systems, 22(1–2): 123–157, January 2006. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Hwang:2018:CAC

[HM18] John T. Hwang and Joaquim R. R. A. Martins. A computational architecture for coupling heterogeneous numerical models and computing coupled derivatives. ACM Transactions on Mathematical Software, 44(4):37:1– 37:39, August 2018. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
3182393.

Hebert:2022:NPP

[HM22] Raphaël Hebert and Emese Meglécz. NSDPY: a Python package to download DNA sequences from NCBI. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102200036X.

Hussain:2023:PPP

[HMB23] Md. Manjurul Hussain, Ishtiak Mahmud, and Sheikh Hefzul

Bari. pyHomogeneity: a

Python package for homogeneity test of time series data. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.427.

Hylton:1996:KPS

[HMD^{+}96] J. Hylton, K. Manheimer, F. L. Drake, Jr., B. Warsaw, R. Masse, and G. van Rossum. Knowbot programming: system support for mobile agents. In Cabrera and Islam [CI96], pages 8–13. ISBN 0-8186-7693-0. LCCN QA 76.76 O63 I59 1996. IEEE catalog number 96TB100089.

Hernandez-Olivan:2023:MPL

[HOB23] Carlos Hernandez-Olivan and Jose R. Beltran. Musicaiz: a Python library for symbolic music generation, analysis and visualization. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000614.

Horsen:2015:G

[Hor15] C. V. Horsen. GMPY. Web site, 2015. URL https:

                                        //pypi.python.org/pypi/
                                        gmpy2.

Horton:2022:BRF

[Hor22] Nicholas J. Horton. Book review: Foundations of Statistics for Data Scientists: With R and Python Alan Agresti, and Maria Kateri, Boca Raton. FL: CRC Press, 2022, 446 pp., $99.95 (textbook), ISBN 978-0-367-74845-6. Journal of the American Statistical Association, 117(539):1603–1604, 2022. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Horn:2023:PTD

[Hor23] Logan Bishop-Van Horn. pyTDGL: Time-dependent Ginzburg–Landau in Python. Computer Physics Communications, 291(??):Article 108799, October 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523001443.

Hosmer:2012:GSS

[Hos12] Ben Hosmer. Getting started with Salt Stack — the other configuration management system built with Python. Linux journal, 2012(223):3:1– 3:??, November 2012. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hosmer:2014:PFW

[Hos14] Chet Hosmer. Python Forensics: a workbench for inventing and sharing digital forensic technology. Syngress Publishing, Inc., Rockland, MA, USA, 2014. ISBN 0-12-418683-1, 0-12-418676-9. xxviii + 318 pp. LCCN

????  URL http://
proquest.safaribooksonline.
com/9780124186767.

Howes:1998:TPC

[How98] Brad Howes. Template processing classes for Python. Dr. Dobb’s Journal of Software Tools, 23(2):38, 40, 42, 44–46, 48, 100, February 1998. CODEN DDJOEB. ISSN 1044-789X.

Howard:2006:MYS

[How06] Mike Howard. Maybe you should use Python. ;login: the USENIX Association newsletter, 31(5):??, October 2006. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/
                                        october-2006-volume-31-
                                        number-5/maybe-you-should-
                                        use-python.

Hirschfeld:2012:EUC

[HPH12] Robert Hirschfeld, Michael Perscheid, and Michael Haupt. Explicit use-case representation in object-oriented programming languages. ACM SIGPLAN Notices, 47(2):51– 60, February 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hermann:2016:SNU

[HPT^{+}16] Gunter Hermann, Vincent Pohl, Jean Christophe Tremblay, Beate Paulus, Hans-Christian Hege, and Axel Schild. Software news and updates: ORBKIT: a modular Python toolbox for crossplatform postprocessing of quantum chemical wavefunction data. Journal of Computational Chemistry, 37(16): 1511–1520, June 15, 2016. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Hollmer:2020:JVP

[HQF^{+}20] Philipp Höllmer, Liang Qin, Michael F. Faulkner, A. C. Maggs, and Werner Krauth. JeLLyFysh — version 1.0

— a Python application for all-atom event-chain Monte Carlo. Computer Physics Communications, 253(??): Article 107168, August 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520300254.

Hammond:2000:PPW

[HR00] Mark Hammond and Andy Robinson. Python Programming on Win32: Help for Windows Programmers. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, January 2000. ISBN 1-56592-621-8 (paperback), 1-56592-925-X (e-book). xvii + 652 pp. LCCN QA76.73.P98 H36 2000. US$34.95. URL

http://www.oreilly.com/
catalog/pythonwin32.

Hart:2020:MSM

[HR20] Kenneth A. Hart and Julian J. Rimoli. MicroStructPy: a statistical microstructure mesh generator in Python. SoftwareX, 12(??):Article 100595, July/December 2020.

CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711020303083.

Holt:2006:IPF

[HRS06] Alexander Holt, Sarah Rauchas, and Ian Sanders. Introducing Python into the first year curriculum at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3): 335, September 2006. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Homescu:2012:HTJ

[HS12] Andrei Homescu and Alex Suhan. HappyJIT: a tracing JIT compiler for PHP. ACM SIGPLAN Notices, 47(2):25– 36, February 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hu:2013:UPH

[HS13] Helen H. Hu and Tricia D. Shepherd. Using POGIL to help students learn to program. ACM Transactions on Computing Education, 13(3): 13:1–13:??, August 2013. CO-DEN ???? ISSN 1946-6226.

Hagh:2022:RRA

[HS22] Varda F. Hagh and Mahdi Sadjadi. rigidPy: Rigidity analysis in Python. Computer Physics Communications, 275(??):Article 108306, June 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522000248.

Heldens:2022:LPP

[HSD^{+}22] Stijn Heldens, Alessio Sclocco, Henk Dreuning, Ben van Werkhoven, Pieter Hijma, Jason Maassen, and Rob V. van Nieuwpoort. litstudy: a Python package for literature reviews. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102200125X.

Heath:1997:PES

[HTA^{+}97] Michael Heath, Virginia Torczon, Greg Astfalk, Petter E. Bjørstad, Alan H. Karp, Charles H. Koelbel, Vipin Kumar, Robert F. Lucas, Layne T. Watson, and David E. Womble, editors. Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing. Held in Minneapolis, MN, March 14– 17, 1997. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. CODEN PSSCFK. ISBN 0-89871-395-1. LCCN ???? CD-ROM for Windows, Macintosh and UNIX; no paper form published.

Herath:2020:PPL

[HTH^{+}20] Uthpala Herath, Pedram Tavadze, Xu He, Eric Bousquet, Sobhit Singh, Francisco Muñoz, and Aldo H. Romero. PyProcar: a Python library for electronic structure pre/post-processing. Computer Physics Communications, 251(??):Article 107080, June 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465519303935.

Herman:2017:PSO

[HU17] Jon Herman and Will Usher. SALib: an open-source Python library for sensitivity analysis. Journal of Open Source Software, 2 (9):97:1–97:2, January 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00097.

Hucka:2018:PNN

[Huc18] Michael Hucka. Nostril: a nonsense string evaluator written in Python. Journal of Open Source Software, 3(25):596:1–596:2, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00596.

Hugunin:1997:PJB

[Hug97] Jim Hugunin. Python and Java: The best of both worlds. Discusses an implementation of the Python scripting language in Java. The article and the code are available electronically.,

1997. URL http://www.
python.org/jpython/.

Hughes:1999:LP

[Hug99]  Phil Hughes.  Learning

Python. Linux journal, 66: ??, October 1999. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/noframes.linuxjournal.
com/lj-issues/issue66/3541.
html.

Hughes:2001:BRP

[Hug01]  Phil Hughes.  Book re-

views: Python Developer’s Handbook. Linux journal, 82: 180, February 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Hughes:2014:PPI

[Hug14] Adam Hughes. pyparty:

Intuitive particle processing in Python. Journal of Open Research Software, 2 (1):e26–??, September 23, 2014. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.bh/.

Hughes:2018:PMP

[Hug18] Momar G-O Hughes. MCycle:

a Python package for 1D sizing and analysis of thermodynamic power cycles. Journal of Open Source Software, 3(28):710:1–710:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00710.

Harvey:2019:TPS

[HW19] Christine Harvey and R. S.

Weigel. Transplant2Mongo:

Python scripts that insert organ procurement and transplantation network (OPTN) data in MongoDB. Journal of Open Research Software, 7(1):5–??, March 14, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.229/.

Huang:2021:PAC

[HWJ^{+}21] Sitao Huang, Kun Wu, Hyunmin Jeong, Chengyue Wang, Deming Chen, and Wen-Mei Hwu. PyLog: An algorithmcentric Python-based FPGA programming and synthesis flow. IEEE Transactions on Computers, 70(12):2015– 2028, December 2021. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Humer:2015:DSL

[HWW^{+}15] Christian Humer, Christian Wimmer, Christian Wirth, Andreas Wöß, and Thomas Würthinger. A domainspecific language for building self-optimizing AST interpreters. ACM SIGPLAN Notices, 50(3):123–132, March 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hewitt:2016:MWA

[HYG16] Brett Hewitt, Moi Hoon Yap, and Robyn Grant. Manual Whisker Annotator (MWA): A modular open-source tool. Journal of Open Research Software, 4(1):e16–??, April 28, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.93/.

Hu:2023:ESP

[HZ23] Mingzhe Hu and Yu Zhang.

An empirical study of the Python/C API on evolution and bug patterns. Journal of Software: Evolution and Process, 35(2):e2507:1–e2507:??, February 2023. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Ivezic:2014:SDM

[ICVG14] Željko Ivezić, Andrew (Andrew J.) Connolly, Jacob T. VanderPlas, and Alexander (Alexander G.) Gray. Statistics, data mining, and machine learning in astronomy: a practical Python guide for the analysis of survey data. Princeton series in modern observational astronomy. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15168-7. x + 540 pp. LCCN QB51.3.E43 S72

                                        2014. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1410/2013951369-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1410/2013951369-
                                        d.html.

Idris:2014:PDA

[Idr14]  Ivan Idris.  Python data

analysis: learn how to apply powerful data analysis techniques with popular open source Python modules. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78355-335-9, 1-78355-336-7 (e-book). v + 329 pp. LCCN QA76.73.P98 I37 2014.

IEEE:1997:PIP

[IEE97a] IEEE, editor. Proceedings. 11th International Parallel Processing Symposium, April 1–5, 1997, Geneva, Switzerland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7793-7. LCCN QA76.58 .I56 1997. IEEE catalog number 97TB100107. IEEE Computer Society Press order number PR07792.

IEEE:1997:PAP

[IEE97b] IEEE, editor. Proceedings. Asia Pacific Software Engineering Conference and International Computer Science Conference: December 2– 5, 1997, Hong Kong. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8271-X. LCCN QA76.758.A77 1997. IEEE catalog number 97TB100207. IEEE Computer Society order number PR08271.

IEEE:2020:SPI

[IEE20] IEEE, editor. SC’20: Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (Atlanta, Georgia, November 9–19, 2020). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2020. ISBN 1-72819-998-0, 1-72819-999-9 (printondemand). LCCN QA76.88.

Isaacs:2019:PCL

[IG19] K. E. Isaacs and T. Gamblin. Preserving command line workflow for a package management system using ASCII DAG visualization. IEEE Transactions on Visualization and Computer Graphics, 25(9):2804– 2820, September 2019. CO-DEN ITVGEA. ISSN 1077-2626.

Ito:2023:SPF

[IHT23] Yuma Ito, Masanori Hirose, and Makio Tokunaga.

Slitflow: a Python framework for single-molecule dynamics and localization analysis. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023001589.

Ishizaki:2012:ADT

[IOC^{+}12] Kazuaki Ishizaki, Takeshi Ogasawara, Jose Castanos, Priya Nagpurkar, David Edelsohn, and Toshio Nakatani. Adding dynamically-typed language support to a staticallytyped language compiler: performance evaluation, analysis, and tradeoffs. ACM SIGPLAN Notices, 47(7): 169–180, July 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). VEE ’12 conference proceedings.

Ishak:2017:BRSa

[Ish17]  B. Ishak.  Book review:

Statistics, data mining, and machine learning in astronomy: a practical Python guide for the analysis of survey data, by Zeljko Ivezić, Andrew J. Connolly, Jacob T. VanderPlas and Alexander Gray. Contemporary Physics, 58(1):99, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ishak:2019:BRS

[Ish19] B. Ishak. Book review: A student’s guide to Python for physical modeling. Contemporary Physics, 60(1):88–89, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Izzo:2018:VMT

[ISMA18] Richard Izzo, David Steinman, Simone Manini, and Luca Antiga. The vascular modeling toolkit: a Python library for the analysis of tubular structures in medical images. Journal of Open Source Software, 3(25):745:1–745:5, May 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00745.

Jackson:2002:PPI

[Jac02] Keith R. Jackson. pyGlobus:

a Python interface to the Globus Toolkit TM. Concurrency and Computation: Practice and Experience, 14 (13–15):1075–1083, November/December 2002. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jacobs:2014:PPP

[JAGP14] Christian Jacobs, Alexandros Avdis, Gerard Gorman, and Matthew Piggott. PyRDM: a Python-based library for automating the management and online publication of scientific software and data. Journal of Open Research Software, 2(1):e28–??, October 03, 2014. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.bj/.

Jankowski:2010:BRBa

[Jan10] Richard Jankowski. Book review: Data Structures and Algorithms Using Python and C++, by David M. Reed and John Zelle Franklin, Beedle and Associates 2009. ACM SIGACT News, 41(1): 13–15, March 2010. CO-DEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic). See [RZ09].

Jackson:2020:PPD

[JCL^{+}20] Robert Jackson, Scott Collis, Timothy Lang, Corey Potvin, and Todd Munson. PyDDA: a Pythonic direct data assimilation framework for wind retrievals. Journal of Open Research Software, 8(1):20–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.264/.

Joyner:2011:OSC

[JČMG11] David Joyner, Ondřej Čertík, Aaron Meurer, and Brian E. Granger. Open source computer algebra systems: SymPy. ACM Communications in Computer Algebra, 45(3–4):225–234, September 2011. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jeong:2019:SSG

[JCY^{+}19] Eunji Jeong, Sungwoo Cho, Gyeong-In Yu, Joo Seong Jeong, Dong-Jin Shin, Taebum Kim, and Byung-Gon Chun. Speculative symbolic graph execution of imperative deep learning programs. Operating Systems Review, 53 (1):26–33, July 2019. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jones:2002:PX

[JD02] Christopher A. Jones and Fred L. Drake. Python and XML. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00128-2. xvi + 360 pp. LCCN QA76.73.P98 J66 2002.

James:2018:PDP

[JEC18] Ryan G. James, Christopher J. Ellison, and James P. Crutchfield. dit: a Python package for discrete information theory. Journal of Open Source Software, 3 (25):738:1–738:3, May 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00738.

Jimenez:2017:PPB

[JG17] José Jiménez and Josep Ginebra. pyGPGO: Bayesian optimization for Python. Journal of Open Source Software, 2(19):431:1–431:3, November 2017. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00431.

Joswig:2023:PPF

[JKKN23] Fabian Joswig, Simon Kuberski, Justus T. Kuhlmann, and Jan Neuendorf. pyerrors: a Python framework for error analysis of Monte Carlo data. Computer Physics Communications, 288(??): Article 108750, July 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465523000954.

Jeong:2022:WOO

[JKST22] Injun Jeong, Sunghyun Kang, Stefano Scopel, and Gaurav Tomar. WimPyDD: an object-oriented Python code for the calculation of WIMP direct detection signals. Computer Physics Communications, 276(??): Article 108342, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522000601.

Jalowiecki:2023:PPL

[JLP23] Konrad Jałowiecki, Paulina Lewandowska, and Lukasz Pawela. PyQBench: a Python library for benchmarking gate-based quantum computers. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002546.

Joksas:2020:BPT

[JM20] Dovydas Joksas and Adnan Mehonic. badcrossbar: a Python tool for computing and plotting currents and voltages in passive crossbar arrays. SoftwareX, 12(??):Article 100617, July/December 2020. CO-DEN ???? ISSN 2352-

7110. URL http://www.
sciencedirect.com/science/
article/pii/S2352711020303307.

Johansson:2012:QOS

[JNN12] J. R. Johansson, P. D.

Nation, and Franco Nori. QuTiP: an open-source Python framework for the dynamics of open quantum systems. Computer Physics Communications, 183(8): 1760–1772, August 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465512000835.

Johansson:2013:QPF

[JNN13] J. R. Johansson, P. D. Nation, and Franco Nori. QuTiP 2: a Python framework for the dynamics of open quantum systems. Computer Physics Communications, 184(4):1234–1240, April 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465512003955.

Johnson:1997:BRI

[Joh97] Dwight Johnson. Book review: Internet Programming with Python. Linux journal, 42:??, October 1997. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (elec-

                                        tronic). URL http://www.
                                        ssc.com/lj/issue42/2152.

html. See book [WvA96].

Johnson:2008:EPC

[Joh08] Mark Johnson. Essential Python for corpus linguistics. Blackwell, Oxford, UK, 2008. ISBN 1-4051-4563-3 (hardcover), 1-4051-4564-1 (paperback). 208 pp. LCCN ????

Johnson:2012:CIP

[Joh12] Mark J. (Mark James) Johnson. A concise introduction to programming in Python. Chapman and Hall/CRC textbooks in computing. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012.

ISBN 1-4398-9694-1 (paperback). xi + 205 pp. LCCN QA76.73.P98 J64 2012.

Johansson:2015:NPP

[Joh15] Robert Johansson. Numerical Python: a Practical Techniques Approach for Industry. The expert’s voice in PYTHON; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2015. ISBN 1-4842-0554-5 (paperback), 1-4842-0553-7 (e-book). xxii + 487 pp. LCCN ???? URL http:

//www.gbv.de/dms/tib-ub-
hannover/841207984.pdf.

Joshi:2015:OPE

[Jos15] Prateek Joshi. OpenCV with Python by example: build real-world computer vision applications and develop cool demos using OpenCV for Python. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-393-6, 1-78528-987-X. 296 (est.) pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781785283932.

Joshi:2016:JDS

[Jos16]  Anshul Joshi.  Julia for

data science: explore the world of data science from scratch with Julia by your side. Packt Publishing, Birmingham, UK, 2016. ISBN 1-78355-386-3 (e-book), 1-78528-969-1. 339 pp. LCCN QA76.73.J8; T55.4-60.8.

Jin:2023:PEP

[JPJ^{+}23] Gan Jin, Hongsheng Pang, Yuyang Ji, Zujian Dai, and Lixin He. PYATB: an efficient Python package for electronic structure calculations using ab initio tight-binding model. Computer Physics Communications, 291(??): Article 108844, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465523001893.

Jiang:2020:PPF

[JPOB20] S. Jiang, P. Pan, Y. Ou, and C. Batten. PyMTL3: a Python framework for opensource hardware modeling, generation, simulation, and verification. IEEE Micro, 40 (4):58–66, July/August 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jha:2024:GAT

[JS24] Raghav G. Jha and Abhishek Samlodia. GPU-acceleration of tensor renormalization with PyTorch using CUDA. Computer Physics Communications, 294(??):Article 108941, January 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523002862.

Johansson:2014:PMP

[JT14] Fredrik Johansson and The mpmath Development Team. mpmath: a Python library for arbitrary-precision floatingpoint arithmetic. Web site,

2014. URL http://mpmath.
org/.

Jirjies:2016:PPI

[JWHS16] Saman Jirjies, Garrick Wallstrom, Rolf U. Halden, and Matthew Scotch. pyJacqQ: Python implementation of Jacquez’s Q-statistics for space-time clustering of disease exposure in case-control studies. Journal of Statistical Software, 74(??): ??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v074i06;
https://www.jstatsoft.
org/index.php/jss/article/
view/v074i06/v74i06.pdf.

Kak:2008:SOC

[Kak08] Avinash C. Kak. Scripting with objects: a comparative presentation of objectoriented scripting with Perl and Python. Wiley, New York, NY, USA, 2008. ISBN 0-470-17923-6 (paperback). xxxiv + 1279 pp. LCCN QA76.64.K3555; QA76.64.K3555 2008. US$89.95.

URL http://www.loc.gov/

catdir/enhancements/fy0743/
2007035480-d.html;   http:
//www.loc.gov/catdir/enhancements/
fy0808/2007035480-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0835/2007035480-
t.html.

Karkera:2014:BPG

[Kar14] Kiran R. Karkera. Building probabalistic graphical models with Python solve machine learning problems using probabalistic graphical models implemented in Python with real-world applications. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-900-7, 1-78328-901-5 (e-book), 1-306-90287-8 (e-book). LCCN QA279 K37 2014. URL http://

proquest.tech.safaribooksonline.
de/9781783289004.

Karakoc:2023:BPC

[Kar23] Mesut Karakoç. BiFold: a Python code for the calculation of double-folded (bifold) potentials with density-in/ dependent nucleon–nucleon interactions. Computer Physics Communications, 284(??):Article 108613, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522003320.

Kasampalis:2015:MPD

[Kas15]  Sakis Kasampalis.  Mas-

tering Python design patterns. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-932-7, 1-78398-933-5 (e-book). LCCN QA76.73.P98 K37 2015.

Krauss:2007:PMM

[KB07] Ryan W. Krauss and Wayne J.

Book. A Python module for modeling and control design of flexible robots. Computing in Science and Engineering, 9(3):41–45, May/June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Krebber:2018:PMP

[KB18] Manuel Krebber and Henrik Barthels. MatchPy: Pattern matching in Python. Journal of Open Source Software, 3(26):670:1–670:2, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00670.

Kortbeek:2020:BEB

[KBC^{+}20] Vito Kortbeek, Abu Bakar, Stefany Cruz, Kasim Sinan Yildirim, Przemysław Pawełczak, and Josiah Hester. BFree: Enabling battery-free sensor prototyping with Python. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies (IMWUT), 4(4):135:1– 135:39, December 2020. CO-DEN ???? ISSN 2474-9567 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3432191.

Kundu:2018:PPA

[KBLJ18] Sudipta Kundu, Satadeep Bhattacharjee, Seung-Cheol Lee, and Manish Jain. PASTA: Python Algorithms for Searching Transition stAtes. Computer Physics Communications, 233(??): 261–268, December 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465518302583.

Khatri:2011:MBC

[KCS11] Sujata Khatri, R. S. Chhillar, and V. B. Singh. Measuring bug complexity in object oriented software system. ACM SIGSOFT Software Engineering Notes, 36 (6):1–8, November 2011. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Khan:2022:EST

[KCVM22] Faizan Khan, Boqi Chen, Daniel Varro, and Shane McIntosh. An empirical study of type-related defects in Python projects. IEEE Transactions on Software Engineering, 48(8): 3145–3158, August 2022. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Kuchling:1999:PT

[KD99] Andrew Kuchling and Fred Drake, editors. Python tutorial. toExcel, San Jose, CA, 1999. ISBN 1-58348-375-6 (soft cover). ???? pp. LCCN ????

Konda:2018:TPT

[KDC^{+}18] Pradap Konda, Sanjib Das, Paul Suganthan G. C., Philip Martinkus, Adel Ardalan, Jeffrey R. Ballard, Yash Govind, Han Li, Fatemah Panahi, Haojun Zhang, Jeff Naughton, Shishir Prasad, Ganesh Krishnan, Rohit Deep, and Vijay Raghavendra. Technical perspective: Toward building entity matching management systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 47 (1):33–40, March 2018. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Khatami:2023:CCP

[KF23] Siamak Khatami and Christopher Frantz. Copatrec: a correlation pattern recognizer Python package for nonlinear relations. SoftwareX, 23(??): ??, July 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023001528.

Kaiser:2020:JPP

[KFV20] K. E. Kaiser, A. N. Flores, and C. R. Vernon. Janus: a Python package for agentbased modeling of land use and land cover change. Journal of Open Research Software, 8(1):15–??, June 25, 2020. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.306/.

Khuat:2023:HBP

[KG23] Thanh Tung Khuat and Bog-

dan Gabrys. hyperbox-brain:

a Python toolbox for hyperboxbased machine learning algorithms. SoftwareX, 23(??): ??, July 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023001218.

Kaufmann:2023:APP

[KH23] Josef Kaufmann and Karsten Held. ana cont: Python package for analytic continuation. Computer Physics Communications, 282(??): Article 108519, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522002387.

Korosov:2016:PNS

[KHD^{+}16] Anton Korosov, Morten Hansen, Knut-Frode Dagestad, Asuka Yamakawa, Aleksander Vines, and Maik Riechert. Nansat: a scientistorientated Python package for geospatial data processing. Journal of Open Research Software, 4(1):e39–??, October 24, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.120/.

Kuhring:2019:CBO

[KI19] Lucas Kuhring and Zsolt István. I can’t believe it’s not (only) software!: bionic distributed storage for Parquet files. Proceedings of the VLDB Endowment, 12(12): 1838–1841, August 2019. CO-DEN ???? ISSN 2150-8097.

Kinder:2005:EDP

[Kin05] Ken Kinder. Event-driven programming with Twisted and Python. Linux journal, 2005(131):??, March 2005. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kinser:2009:PB

[Kin09] Jason M. Kinser. Python for bioinformatics. Jones and Bartlett series in biomedical informatics. Jones and Bartlett Publishers, Sudbury, MA, USA, 2009. ISBN 0-7637-5186-3. xvii + 417 pp. LCCN QH324.2.K55; QH324.2.K55 2009.

Kirby:2004:AFN

[Kir04]  Robert C. Kirby.  Algo-

rithm 839: FIAT, a new paradigm for computing finite element basis functions.

ACM Transactions on Mathematical Software, 30(4):502– 516, December 2004. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kiusalaas:2010:NME

[Kiu10] Jaan Kiusalaas. Numerical methods in engineering with Python. Cambridge University Press, Cambridge, UK, second edition, 2010. ISBN 0-521-19132-7 (hardcover). x + 422 pp. LCCN TA345 .K584 2010.

Kiusalaas:2013:NME

[Kiu13] Jaan Kiusalaas. Numerical methods in engineering with Python 3. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03385-3. xi + 423 pp. LCCN TA345 .K584 2013.

Korzen:2014:PPP

[KJ14] Marcin Korzeń and Szymon Jaroszewicz. PaCAL: a Python package for arithmetic computations with random variables. Journal of Statistical Software, 57 (10):??, May 2014. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL http://www.
                                        jstatsoft.org/v57/i10.

Kazil:2015:DWU

[KJ15] Jacqueline Kazil and Katharine Jarmul. Data Wrangling Using Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-4881-7. ???? pp. LCCN ????

Kim:1997:DIW

[KL97] Hyeon Jong Kim and Ki Ho Lee. Design and implementation of Web documents creating system with HTMLsupporting library. Journal of KISS(C) (Computing Practices), 3(4):375–383, August 1997. CODEN CKNCFY. ISSN 1226-2293.

Klatchko:1999:DRS

[Kla99]  Ron Klatchko.  Dynami-

cally reconfigurable servers: Python’s extensibility makes it easy. Dr. Dobb’s Journal of Software Tools, 24(1):80, 82– 84, January 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ddj/ftp/1999/1999_01/cprofile.
zip.

Klaver:2011:PC

[Kla11] Adrian Klaver. Python in the cloud. Linux journal, 2011 (210):7:1–7:??, October 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Koulouri:2014:TIP

[KLM14] Theodora Koulouri, Stanislao Lauria, and Robert D. Macredie. Teaching introductory programming: a quantitative evaluation of different approaches. ACM Transactions on Computing Education, 14(4):26:1–26:??, December 2014. CODEN ???? ISSN 1946-6226.

Koulouri:2015:TIP

[KLM15] Theodora Koulouri, Stanislao Lauria, and Robert D. Macredie. Teaching introductory programming: a quantitative evaluation of different approaches. ACM Transactions on Computing Education, 14(4):26:1–26:??, February 2015. CODEN ???? ISSN 1946-6226.

Kinsley:2015:BPG

[KM15] Harrison Kinsley and Will McGugan. Beginning Python Games Development: With Pygame. SpringerLink: Bücher. Apress, Berkeley, CA, USA, second edition, 2015. ISBN 1-4842-0970-2, 1-4842-0971-0. xxii + 308 + 53 pp. LCCN QA75.5-76.95. URL http:

//www.springerlink.com/
content/978-1-4842-0970-
7.

Krishnamoorthy:2021:EPL

[KMK^{+}21] Aravind Krishnamoorthy, Ankit Mishra, Deepak Kamal, Sungwook Hong, Ken ichi Nomura, Subodh Tiwari, Aiichiro Nakano, Rajiv Kalia, Rampi Ramprasad, and Priya Vashishta. EZFF: Python library for multi-objective parameterization and uncertainty quantification of interatomic forcefields for molecular dynamics. SoftwareX, 13(??):Article 100663, January 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S235271102100008X.

Koot:2023:PVP

[KMLP^{+}23] Paul Koot, Miguel Angel Mendoza-Lugo, Dominik Paprotny, Oswaldo Morales-Nápoles, Elisa Ragno, and Daniël T. H. Worm. PyBanshee version (1.0): a Python implementation of the MAT-LAB toolbox BANSHEE for Non-Parametric Bayesian Networks with updated features. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001972.

Kinder:2015:SGP

[KN15] Jesse M. Kinder and Philip Charles Nelson. A student’s guide to Python for physical modeling. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16958-6 (hardback), 0-691-17050-9 (paperback). xiii + 139 pp. LCCN QA76.73.P98 K54 2015.

Knowlton:2008:PCM

[Kno08]  Jim Knowlton.  Python:

create-modify-reuse. Wrox programmer to programmer. Wiley, New York, NY, USA, 2008. ISBN 0-470-25932-9. xxv + 261 pp. LCCN ????

Kormanyos:2011:APC

[Kor11] Christopher Kormanyos. Algorithm 910: a portable C++ multiple-precision system for special-function calculations. ACM Transactions on Mathematical Software, 37(4):45:1– 45:27, February 2011. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kirsanskas:2017:QOS

[KPK^{+}17] Gediminas Kirsanskas, Jonas Nyvold Pedersen, Olov Karlström, Martin Leijnse, and Andreas Wacker. QmeQ 1.0: an open-source Python package for calculations of transport through quantum dot devices. Computer Physics Communications, 221(??): 317–342, December 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465517302515.

Krasanakis:2022:PPP

[KPKS22] Emmanouil Krasanakis, Symeon Papadopoulos, Ioannis Kompatsiaris, and Andreas L. Symeonidis. pygrank: a Python package for graph node ranking. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001455.

Kratzke:2022:CNO

[Kra22] Nane Kratzke. Cloud-native observability: The manyfaceted benefits of structured and unified logging—a multicase study. Future Internet, 14(10):274, September 26, 2022. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/
14/10/274.

Kemmer:2018:NJE

[KRH18a] Thomas Kemmer, Sergej Rjasanow, and Andreas Hildebrandt. NESSie.jl — efficient and intuitive finite element and boundary element methods for nonlocal protein electrostatics in the Julia language. Journal of Computational Science, 28:193–203, September 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S187775031730738X.

Kemmer:2018:PNJ

[KRH18b] Thomas Kemmer, Sergej Rjasanow, and Andreas Hildebrandt. NESSie.jl: Efficient and intuitive finite element and boundary element methods for nonlocal protein electrostatics in the Julia language. Journal of Computational Science, 28:193–203, 2018. ISSN 1877-7503. URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S187775031730738X.

Kim:2023:FPL

[KRSD^{+}23] Taewon D. Kim, M. Richer, Gabriela Sánchez-Díaz, Ramón Alain Miranda-Quintana, Toon Verstraelen, Farnaz Heidar-Zadeh, and Paul W. Ayers. Fanpy: a Python library for prototyping multideterminant methods in ab initio quantum chemistry. Journal of Computational Chemistry, 44(5):697–709, February 15, 2023. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Karam:2021:BPS

[KS21] Mokbel Karam and Tony

                                        Saad. BuckinghamPy: a

Python software for dimensional analysis. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021001291.

Klarner:2012:TSD

[KSB12] Hannes Klarner, Heike Siebert, and Alexander Bockmayr. Time series dependent analysis of unparametrized Thomas networks. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9(5):1338–1351, September 2012. CODEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Koenka:2014:IOS

[KSH14] Israel Joel Koenka, Jorge Sáiz, and Peter C. Hauser. Instrumentino: an open-source modular Python framework for controlling Arduino based experimental instruments. Computer Physics Communications, 185(10):2724– 2729, October 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465514002112.

Karam:2020:PPS

[KSS20] Mokbel Karam, James C.

Sutherland, and Tony Saad. PyModPDE: a Python software for modified equation analysis. SoftwareX, 12(??):Article 100541, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020300224.

Kuchling:1997:PU

[Kuc97] Andrew Kuchling. Python update. Linux journal, 37:??, May 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kuchling:1998:LIG

[Kuc98a] Andrew Kuchling. LJ interviews Guido van Rossum. Linux journal, 55:18, 20– 21, November 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Kuchling:1998:CFP

[Kuc98b] Andrew M. Kuchling. A CGI framework in Python. Web Techniques, 3(2):43– 46, February 1998. CO-DEN WETEFA. ISSN 1086-556X. URL http:

                                        //www.webtechniques.com/
                                        features/1998/02/kuchling/
                                        kuchling.shtml.

Kuchling:1998:PDA

[Kuc98c] Andrew M. Kuchling. The Python DB-API. Linux journal, 49:??, May 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Klymenko:2021:NEP

[KVSC21] M. V. Klymenko, J. A. Vaitkus, J. S. Smith, and J. H. Cole. NanoNET: an extendable Python framework for semi-empirical tightbinding models. Computer Physics Communications, 259(??):Article 107676, February 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465520303283.

Kailasa:2022:PED

[KWBB22] Srinath Kailasa, Tingyu Wang, Lorena A. Barba, and Timo Betcke. PyExaFMM: an exercise in designing highperformance software with Python and Numba. Computing in Science and Engineering, 24(5):77–84, September/October 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Koutsellis:2023:CWB

[KXK^{+}23] Themistoklis Koutsellis, Georgios Xexakis, Konstantinos Koasidis, Natasha Frilingou, Anastasios Karamaneas, Alexandros Nikas, and Haris Doukas. In-cognitive: a web-based Python application for fuzzy cognitive map design, simulation, and uncertainty analysis based on the Monte Carlo method. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002091.

Lutz:1999:LP

[LA99] Mark Lutz and David Ascher.

Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-464-9. xvi + 366 pp. LCCN QA76.73.P98 L8798 1999. US$29.95. URL

http://www.oreilly.com/
catalog/9781565924642;
http://www.oreilly.com/
catalog/lpython/.

Lutz:2004:LP

[LA04] Mark Lutz and David Ascher. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2004. ISBN 0-596-00281-5. xxvi + 591 pp. LCCN QA76.73.P98 L877 2004. URL http://

                                        www.oreilly.com/catalog/
                                        9780596002817.

Lacheiner:2006:EPB

[Lac06] Hermann Lacheiner. Entwicklung einer auf Python basierenden Rich Client Platform für Linux. (German) [Development of a Pythonbased Rich Client Platform for Linux]. Diplome-Arbeit, Universität Linz, Linz, Austria, 2006. v + 86 pp.

Lutz:2007:EPM

[LAG07] Mark Lutz, David Ascher, and Dinu C.. Gherman. Einführung in Python: [moderne OO-Programmierung; behandelt Python 2.5]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 3-89721-488-1. xxvii + 624 pp. LCCN ???? EUR

                                        39.90. URL http://www.

gbv.de/dms/ilmenau/toc/
527924601.PDF.

Lakshminarayanan:2017:BRL

[Lak17] Vasudevan Lakshminarayanan.

Book review: Learning scientific programming with Python, by Christian Hill. Contemporary Physics, 58 (3):282–284, 2017. CO-DEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Lambert:2015:PPT

[Lam15] Kenneth Alfred Lambert.

Python programming for teens. Cengage Learning PTR, Boston, MA, USA, 2015. ISBN 1-305-27195-5 (paperback), 1-305-27196-3 (e-book). xxviii + 243 pp. LCCN QA76.73.P98 L338 2015.

Langtangen:2006:PSC

[Lan06] Hans Petter Langtangen.

Python scripting for computational science, volume 3 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 3-540-29415-5. xxiv + 736 pp. LCCN ????

Langtangen:2007:PSC

[Lan07] Hans Petter Langtangen.

Python scripting for computational science, volume 3 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2007. ISBN 3-540-73915-7. ???? pp. LCCN ????

Langtangen:2008:PSC

[Lan08] Hans Petter Langtangen. Python Scripting for Computational Science, volume 3 of Springer eBook Collection Mathematics and Statistics [Dig. Serial]; Springer-11649 [Dig. Serial]; Texts in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2008. ISBN 3-540-73915-7, 3-540-73916-5. xxiv + 750 pp. LCCN ????

Langtangen:2009:PSP

[Lan09] Hans Petter Langtangen. A primer on scientific programming with Python, volume 6 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-02475-0, 3-642-02474-2. ISSN 1611-0994. xxvii + 693 pp. LCCN QA76.73.P98 L286 2009.

Langtangen:2011:PSP

[Lan11] Hans Petter Langtangen. A primer on scientific programming with Python, volume 6 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2011. ISBN 3-642-18365-4, 3-642-18366-2 (ebook). xxix + 699 pp. LCCN QA76.73.P98 L36 2011.

Langtangen:2012:PSP

[Lan12] Hans Petter Langtangen. A primer on scientific programming with Python, volume 6 of Texts in computational science and engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2012. ISBN 3-642-30292-0, 3-642-30293-9 (e-book). ISSN 1611-0994. LCCN QA76.73.P98 L36 2012.

Lanaro:2013:PHP

[Lan13] Gabriele Lanaro. Python high performance programming: boost the performance of your Python programs using advanced techniques. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78328-845-0 (paperback), 1-306-25426-4 (e-book). 95 pp. LCCN QA76.73 .P98. URL http://

proquest.tech.safaribooksonline.
de/9781783288458.

Lawhead:2015:QPP

[Law15] Joel Lawhead. QGIS Python programming cookbook: over 140 recipes to help you turn QGIS from a desktop GIS tool into a powerful automated geospatial framework. Community Experience Distilled.

Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-498-8, 1-78398-499-6 (e-book). xii + 315 pp. LCCN QA76.73.P98 .L394 2015.

Layton:2015:LDM

[Lay15] Robert Layton. Learning data mining with Python: harness the power of Python to analyze data and create insightful predictive models. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-605-2 (paper), 1-78439-120-4 (e-book). xiv + 317 pp. LCCN QA76.73.P98 L39 2015.

Lazic:2022:BRB

[Laz22]  Stanley E. Lazic.  Book

review: Bayesian Modeling and Computation in Python Learning. Journal of the Royal Statistical Society. Series A (Statistics in Society), 185(S2):S764–S765, December 2022. CODEN JS-SAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL https://

academic.oup.com/jrsssa/
article/185/Supplement_
2/S764/7069537.

Lekinwala:2022:PPL

[LB22] Nirav L. Lekinwala and Mani Bhushan. pyGNMF: a Python library for implementation of generalised non-negative matrix factorisation method. SoftwareX, 20(??):??, December 2022. CODEN ????

ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001753.

List:2011:FUT

[LC11] Michael List and David Car.

A Fortran unit-testing framework utilizing templating and the PyF95++ toolset. ACM Fortran Forum, 30(1):3–15, April 2011. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Luszczek:2007:HPD

[LD07] Piotr Luszczek and Jack Dongarra. High performance development for high end computing with Python Language Wrapper (PLW). The International Journal of High Performance Computing Applications, 21(3):360– 369, August 2007. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/21/
3/360.full.pdf+html.

Lee:2019:CII

[LD19] Edward D. Lee and Bryan C.

Daniels. Convenient Interface to Inverse Ising (ConIII): a Python 3 package for solving Ising-type maximum entropy models. Journal of Open Research Software, 7 (1):3–??, March 04, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.

metajnl.com/articles/10.
5334/jors.217/.

Loraamm:2020:PTV

[LDAL20] Rebecca Loraamm, Joni Downs, James Anderson, and David S. Lamb. PyST-Prism: Tools for voxelbased space-time prisms. SoftwareX, 12(??):Article 100499, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019303309.

LeBrigant:2023:PIG

[LDCM23] Alice Le Brigant, Jules Deschamps, Antoine Collas, and Nina Miolane. Parametric information geometry with the package Geomstats. ACM Transactions on Mathematical Software, 49(4):34:1– 34:??, December 2023. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3627538.

Luo:2021:CPC

[LDW^{+}21] Chenxing Luo, Xin Deng, Wenzhong Wang, Gaurav Shukla, Zhongqing Wu, and Renata M. Wentzcovitch. cij: a Python code for quasiharmonic thermoelasticity. Computer Physics Communications, 267(??): Article 108067, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046552100179X.

LeClercq:2024:APP

[Le 24] Louis-Stéphane Le Clercq.

ABCal: a Python package for author bias computation and scientometric plotting for reviews and meta-analyses. Scientometrics, 129(1):581– 600, January 2024. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11192-023-04880-6.

Lee:2011:PPF

[Lee11] Kent D. Lee. Python programming fundamentals. Undergraduate topics in computer science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-84996-536-6, 1-84996-537-4 (e-book). xii + 241 pp. LCCN QA76.73.P98 L44 2011.

Leeman:2017:BRS

[Lee17] John R. Leeman. Book review: A Student’s Guide to Python for Physical Modeling. Kinder, Jesse M., and Nelson, Philip. 150 pp. Princeton U.P., Princeton, NJ, 2016. Price: $24.95 (paper). ISBN 978-0-691-17050-

3. American Journal of Physics, 85(5):399, May 2017. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Lentin:2015:LRU

[Len15]  Joseph Lentin.  Learning

robotics using Python: design, simulate, program, and prototype an interactive autonomous mobile robot from scratch with the help of Python, ROS, and Open-CV! Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78328-753-5, 1-68015-749-3 (e-book). xiii + 303 pp. LCCN TJ211.495 .L46 2015.

LeVeque:2009:PTR

[LeV09] Randall J. LeVeque. Python tools for reproducible research on hyperbolic problems. Computing in Science and Engineering, 11(1):19– 27, January/February 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lewis:2009:HPP

[Lew09] Andrew Lewis. High performance Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2009. ISBN 0-596-15996-X. ???? pp. LCCN ????

Lambert:2011:PBO

[LFN^{+}11] Emmanuel Lambert, Martin Fiers, Shavkat Nizamov, Martijn Tassaert, Steven G. Johnson, Peter Bienstman, and Wim Bogaerts. Python bindings for the Open Source Electromagnetic Simulator Meep. Computing in Science and Engineering, 13(3): 53–65, May/June 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Lister:2009:FER

[LFT09] Raymond Lister, Colin Fidge, and Donna Teague. Further evidence of a relationship between explaining, tracing and writing skills in introductory programming. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41(3): 161–165, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Lyden:2021:PPM

[LFT21] Andrew Lyden, Graeme Flett, and Paul G. Tuohy. PyLESA: a Python modelling tool for planning-level Local, integrated, and smart Energy Systems Analysis. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000443.

Liu:2023:TUB

[LFYX23] Di Liu, Yang Feng, Yanyan Yan, and Baowen Xu. Towards understanding bugs in Python interpreters. Empirical Software Engineering, 28 (1):??, January 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10239-x.

Liu:2010:LFI

[LGS10] Yanhong A. Liu, Michael Gorbovitski, and Scott D. Stoller. A language and framework for invariantdriven transformations. ACM SIGPLAN Notices, 45(2):55– 64, February 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lunnikivi:2020:TPR

[LH20] Henri Lunnikivi and Kai JylkkäTimo Hämäläinen. Transpiling Python to Rust for optimized performance. In Embedded Computer Systems: Architectures, Modeling, and Simulation: 20th International Conference, SAMOS 2020, Samos, Greece, July 5–9, 2020, Proceedings, pages 127– 138. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

2020. URL http://link.
springer.com/chapter/10.
1007/978-3-030-60939-9_
9.

Libeskind-Hadas:2014:CBP

[LHB14] Ran Libeskind-Hadas and Eliot Christen Bush. Computing for biologists: Python programming and principles. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-04282-8 (hardcover), 1-107-64218-3 (paperback). x + 207 pp. LCCN QH324.2 .L53 2014. URL http:

//www.cambridge.org/us/
academic/subjects/life-
sciences/genomics-bioinformatics-
and-systems-biology/computing-
biologists-python-programming-
and-principles.

Lee:2021:PPB

[LHH^{+}21] In Seong Lee, Jong-Kwon Ha, Daeho Han, Tae In Kim, Sung Wook Moon, and Seung Kyu Min. PyUNIxMD: a Python-based excited state molecular dynamics package. Journal of Computational Chemistry, 42(24):1755–1766, September 15, 2021. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Logaras:2014:PAE

[LHM14] Evangelos Logaras, Orsalia G.

Hazapis, and Elias S. Manolakos. Python to accelerate embedded SoC design: a case study for systems biology. ACM Transactions on Embedded Computing Systems, 13(4): 84:1–84:??, February 2014. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Lindblad:2006:PP

[Lin06a] Erik Lindblad. Programmering i Python. Studentlitteratur, Lund, Sweden, 2006. ISBN 91-44-04520-4. 406 pp. LCCN ????

Lingl:2006:PKC

[Lin06b] Gregor Lingl. Python für Kids: [mit CD ; auf CD: aktuelles Python 2.5, das Grafik-Modul xturtle und alle Programmbeispiele]. bhv, Redline, Heidelberg, Germany, second edition, 2006. ISBN 3-8266-8622-5. 416 pp. LCCN ???? EUR 19.95.

Lipovetsky:2021:BRLb

[Lip21] Stan Lipovetsky. Book review: Linear Models with Python, Faraway Julian J.. Boca Raton, FL, Chapman and Hall/CRC, Taylor & Francis Group, 2021, 308 pp., 85 b/w illustrations, $99.95 (Hardback), ISBN: 978-1-138-48395-8. Technometrics, 63(3):426–427, 2021. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Lipovetsky:2022:BRH

[Lip22] Stan Lipovetsky. Book review: Handbook of Regression Modeling in People Analytics: With Examples in R and Python, by Keith Mc-Nulty. CRC Press, Taylor & Francis Group, Boca Raton, FL, 2021, ISBN 978-1-032-04174-2, xvi + 255 pp., 48 color illustrations, $63.96 (hbk). Technometrics, 64 (1):143–145, 2022. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Liu:2022:BRD

[Liu22] Shuangzhe Liu. Book review: Data Visualization for Social and Policy Research: a Step-by-Step Approach Using R and Python. Jose Manuel Magallanes Reyes. Cambridge University Press, 2022, 292 pages, $105, hardback. ISBN: 978-1-108-49433-5. International Statistical Review = Revue Internationale de Statistique, 90 (3):626–627, December 2022. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Li:2019:AOP

[LJ19] Yangguang Li and Zhen Ming (Jack) Jiang. Assessing and optimizing the performance impact of the just-in-time configuration parameters — a case study on PyPy. Empirical Software Engineering, 24(4): 2323–2363, August 2019. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-019-09691-z.

Li:2023:APA

[LJ23] Yipeng Li and Xiangmin Jiao.

ARPIST: Provably accurate and stable numerical integration over spherical triangles. Journal of Computational and Applied Mathematics, 420(??):??, March 1, 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042722004204.

Lukasczyk:2023:ESA

[LKF23] Stephan Lukasczyk, Florian Kroiß, and Gordon Fraser. An empirical study of automated unit test generation for Python. Empirical Software Engineering, 28 (2):??, March 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10248-w.

Li:2021:ICM

[LKJC21] Yong Li, Shoaib Kamil, Alec Jacobson, and Yotam Cingold. I[HEART]LA: compilable markdown for linear algebra. ACM Transactions on Graphics, 40(6):264:1– 264:14, December 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https://

dl.acm.org/doi/10.1145/
3478513.3480506.

Lee:2023:PPP

[LKK23] Kyunghoon Lee, Jun Hyeong Kim, and Woo Youn Kim. pyMCD: Python package for searching transition states via the multicoordinate driven method. Computer Physics Communications, 291(??): Article 108831, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523001765.

Litvin:2008:MDA

[LL08] Maria Litvin and Gary Litvin. Mathematics for the digital age and programming in Python. Skylight Pub., Andover, MA, USA, 2008. ISBN 0-9727055-8-9. ???? pp. LCCN ???? URL http://

lccn.loc.gov/2007930214.

Luporini:2020:APD

[LLL^{+}20] Fabio Luporini, Mathias Louboutin, Michael Lange, Navjot Kukreja, Philipp Witte, Jan Hückelheim, Charles Yount, Paul H. J. Kelly, Felix J. Herrmann, and Gerard J. Gorman. Architecture and performance of Devito, a system for automated stencil computation. ACM Transactions on Mathematical Software, 46(1):6:1–6:28, April 2020. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3374916.

Langtangen:2003:UDP

[LM03] H. P. Langtangen and K.- A. Mardal. Using Diffpack from Python scripts. In Langtangen and Tveito [LT03], pages 321–360. CO-DEN LNCSA6. ISBN 3-540-01438-1 (print), 3-642-18237-2 (e-book). ISSN 1439-7358. LCCN QA377 .A45

2003. URL http://link.
springer.com/content/pdf/
10.1007/978-3-642-18237-
2_8.

Logg:2012:ASD

[LMW12] Anders Logg, Kent-Andre Mardal, and Garth Wells, editors. Automated Solution of Differential Equations by the Finite Element Method: The FEniCS Book, volume 84 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSA6. ISBN 3-642-23098-9 (print), 3-642-23099-7 (e-book). ISSN 1439-7358. xiii + 723 pp. LCCN ???? URL http://

fenicsproject.org;   http:
//fenicsproject.org/book/
;   http://link.springer.
com/book/10.1007/978-3-
642-23099-8;   http://www.
springerlink.com/content/
978-3-642-23099-8;   https:
//launchpad.net/fenics-

book. The software developed by the FEniCS Project is free for all to use and modify (licensed under the GNU (L)GPL), and so is this book.

Li:2023:MPM

[LN23] Yunguo Li and Huaiwei Ni. MD2D: a Python module for accurate determination of diffusion coefficient from molecular dynamics. Computer Physics Communications, 284(??): Article 108599, March 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522003186.

Lobianco:2019:JQS

[Lob19] Antonello Lobianco. Julia quick syntax reference: a pocket guide for data science programming. Apress, Berkeley, CA, USA, 2019. ISBN 1-4842-5189-X, 1-4842-5190-3 (e-book). xvii + 216 + 66 pp. LCCN QA76.73.J85. URL http:/

/link.springer.com/book/
10.1007/978-1-4842-5190-
4.

Long:2008:SAR

[Lon08] Philip D. Long. Scalable apprenticeships: reconnecting students through technology. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):3–4, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Lott:2014:MOO

[Lot14a] Steven F. Lott. Mastering object-oriented Python: grasp the intricacies of objectoriented programming in Python in order to efficiently build powerful real-world applications. Community expertise distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-097-2, 1-78328-098-0 (e-book). xii + 609 pp. LCCN QA76.73.P98 L688 2014. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781783280971.

Lott:2014:PSA

[Lot14b] Steven F. Lott. Python for secret agents: analyze, encrypt, and uncover intelligence data using Python, the essential tool for all aspiring secret agents. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-042-7 (paperback), 1-78398-043-5 (e-book). v + 197 pp. LCCN TK5105.8883.

Lott:2015:FPP

[Lot15a]  Steven Lott.  Functional

Python programming. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-699-0, 1-78439-761-X (ebook). 361 (est.) pp. LCCN QA76.73.P98 .L688 2015.

Lott:2015:PEM

[Lot15b] Steven F. Lott. Python essentials: modernize existing Python code and plan code migrations to Python using this definitive guide. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-814-4, 1-78439-034-8. xiv + 271 pp. LCCN ????

Lal:2019:CGA

[LP19] Ratan Lal and Pavithra Prabhakar. Counterexample guided abstraction refinement for polyhedral probabilistic hybrid systems. ACM Transactions on Embedded Computing Systems, 18(5s):98:1– 98:??, October 2019. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3358217.

Ledee:2023:EOP

[LPG^{+}23] François Lédée, Pierryves Padey, Kyriaki Goulouti, Sébastien Lasvaux, and Didier Beloin-Saint-Pierre. EcoDynElec: Open Python package to create historical profiles of environmental impacts from regional electricity mixes. SoftwareX, 23(??): ??, July 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023001814.

Liegeois:2023:PPI

[LPH23] Kim Liegeois, Mauro Perego, and Tucker Hartland. PyAlbany: a Python interface to the C++ multiphysics solver Albany. Journal of Computational and Applied Mathematics, 425(??):??, June 2023.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0377042722006355.

Lopez:2014:MPR

[LR14] Félix López and Víctor Romero. Mastering Python regular expressions: leverage regular expressions in Python even for the most complex features. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-315-7 (paperback), 1-78328-316-5 (e-book). iii + 93 pp. LCCN QA76.73.P98 L67 2014. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781783283156.

Lalor:2023:PIS

[LR23] John Patrick Lalor and Pedro Rodriguez. py-irt: a scalable item response theory library for Python. INFORMS Journal on Computing, 35 (1):5–13, January/February 2023. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:/

                                        /pubsonline.informs.org/
                                        doi/full/10.1287/ijoc.2022.
                                        1250.

Li:2023:PTP

[LRL23] Ye Li, Hongxiang Ren, and Haijiang Li. PyVT: a toolkit for preprocessing and analysis of vessel spatiotemporal trajectories. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000122.

Laura:2018:APL

[LRPD18] Jason Laura, Kelvin Rodriguez, Adam C. Paquette, and Evin Dunn. AutoCNet: a Python library for sparse multi-image correspondence identification for planetary data. SoftwareX, 7(??): 34–36, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271101830013X.

Lervik:2017:SNU

[LRvE17] Anders Lervik, Enrico Riccardi, and Titus S. van Erp. Software news and updates: PyRETIS: a well-done, medium-sized Python library for rare events. Journal of Computational Chemistry, 38 (28):2439–2451, October 30, 2017. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Laird:1997:CSL

[LS97] Cameron Laird and Kathryn Soraiz. Choosing a scripting language: Perl, Tcl, and Python: they’re not your father’s scripting languages. SunWorld online, ??(??):??, October 1997. ISSN 1091-8914.

URL http://www.sun.com/
sunworldonline/swol-10-

                                        1997/swol-10-scripting.
                                        html.

Laird:1998:GTW

[LS98] Cameron Laird and Kathryn Soraiz. GUI toolkits: What are your options? an overview of today’s bestbet GUI toolkits. Sun-World online, ??(??):??, March 1998. ISSN 1091-

                                        8914. URL http://www.sun.
                                        com/sunworldonline/swol-
                                        02-1998/swol-02-python.

html. Discusses Python.

Leskovec:2016:SGP

[LS16] Jure Leskovec and Rok Sosic. SNAP: a general-purpose network analysis and graphmining library. ACM Transactions on Intelligent Systems and Technology (TIST), 8(1): 1:1–1:??, October 2016. CO-DEN ???? ISSN 2157-6904 (print), 2157-6912 (electronic).

Lyonnet:2017:PPT

[LS17] F. Lyonnet and I. Schienbein. PyR@TE 2: a Python tool for computing RGEs at two-loop. Computer Physics Communications, 213(??):181–196, April 2017. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046551630368X.

Langtangen:2003:ATC

[LT03] Hans Petter Langtangen and Aslak Tveito, editors. Advanced Topics in Computational Partial Differential Equations: Numerical Methods and Diffpack Programming, volume 33 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSA6. ISBN 3-540-01438-1 (print), 3-642-18237-2 (e-book). ISSN 1439-7358. LCCN QA377 .A45 2003. URL http:/

/link.springer.com/book/
10.1007/978-3-642-18237-
2;   http://www.diffpack.
com/Book;   http://www.
springerlink.com/content/
978-3-642-18237-2.

Lang:2024:EPN

[LTT^{+}24] Logan Lang, Pedram Tavadze, Andres Tellez, Eric Bousquet, He Xu, Francisco Muñoz, Nicolas Vasquez, Uthpala Herath, and Aldo H. Romero. Expanding PyProcar for new features, maintainability, and reliability. Computer Physics Communications, 297(??): Article 109063, April 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523004083.

Lubanovic:2014:IPM

[Lub14] Bill Lubanovic. Introducing Python: modern computing in simple packages. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-5936-1 (paperback), 1-4493-6119-6 (e-book), 1-4493-6118-8 (e-book). xx + 454 pp. LCCN QA76.73.P98 L83 2015.

Lukyanenko:2023:PAS

[Luk23] Dmitry Lukyanenko. Parallel algorithm for solving overdetermined systems of linear equations, taking into account round-off errors. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

                                        URL https://www.mdpi.
                                        com/1999-4893/16/5/242.

Lundh:2001:PSL

[Lun01] Fredrik Lundh. Python Standard Library. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2001. ISBN 0-596-00096-0. xvi + 281 pp. LCCN QA76.73.P98 L86 2001. US$29.95. URL

                                        http://www.oreilly.com/
                                        catalog/pythonsl.

Lutz:1996:PP

[Lut96] Mark Lutz. Programming Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel:

+1 617 354 5800, Fall 1996. ISBN 1-56592-197-6, 0-585-03222-X (e-book). xxii + 880 pp. LCCN QA76.73.P98 L88 1996. US$39.95. URL http:

//shell.rmi.net/~lutz/;
 http://www.oreilly.com/
catalog/python. Also avail-

able in Japanese translation,

see http://www.oreilly.
co.jp/BOOK/python1.htm.

Lutz:1998:POO

[Lut98a] Mark Lutz. Python: an object oriented scripting language. In Peter H. Salus, editor, Handbook of Programming Languages, page 120 (est.). Macmillan Publishing Company, New York, NY, USA, 1998. ISBN 1-57870-008-6 (vol. 1), 1-57870-009-4 (vol. 2), 1-57870-010-8 (vol. 3), 1-57870-011-6 (vol. 4). LCCN QA76.7 .H363

1998. URL http://shell.
rmi.net/~lutz/whatsnew.
html#S13.

Lutz:1998:PPR

[Lut98b] Mark Lutz. Python Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-500-9. iii + 75 pp. LCCN QA76.73.P98 L882 1998. US$6.95. URL http://

www.oreilly.com/catalog/
9781565925007;   http://

                                        www.oreilly.com/catalog/
                                        pythonpr/.

Lutz:1999:UP

[Lut99] Mark Lutz. Using Python. ;login: the USENIX Association newsletter, 24(1s): ??, January 1999. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        1999-1/python.html. Spe-

cial issue on tools.

Lutz:2001:PPO

[Lut01] Mark Lutz. Programming Python: Object-Oriented Scripting. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, March 2001. ISBN 0-596-00085-5. xxxvii + 1255 pp. LCCN QA76.73.P98 L88 2001. US$54.95. URL

                                        http://www.oreilly.com/
                                        catalog/9780596000851;
                                        http://www.oreilly.com/
                                        catalog/python2. Includes

CD-ROM.

Lutz:2002:PPR

[Lut02] Mark Lutz. Python Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00189-4. iv + 124 pp. LCCN QA76.73.P98 L89 2002. US$11.95. URL http:/

/safari.oreilly.com/0596001894;
 http://www.oreilly.com/
catalog/9780596001896;
http://www.oreilly.com/
catalog/pythonpr2.

Lutz:2005:PPR

[Lut05] Mark Lutz. Python pocket reference. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00940-2 (paperback). ix + 148 pp. LCCN QA76.73.P98 L89 2005. URL http://

www.oreilly.com/catalog/
9780596009403.

Lutz:2006:PP

[Lut06] Mark Lutz. Programming Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2006. ISBN 0-596-00925-9. xlii + 1552 pp. LCCN QA76.73.P98 L88 2006eb; QA76.73.P98 L88 2006; QA76.73.P98. URL

http://www.oreilly.com/
catalog/9780596009250.

Lutz:2008:LP

[Lut08] Mark Lutz. Learning Python.

O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2008. ISBN 0-596-51398-4, 0-596-51398-4. xliv + 700 pp. LCCN QA76.73.P98 L877 2008; QA76.73.P98 L877 2008eb; QA76.73.P98 L8798 2008; QA76.73.P98. URL

                                        http://www.oreilly.com/
                                        catalog/9780596513986.

Lutz:2009:LPa

[Lut09a] Mark Lutz. Learning Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2009. ISBN 0-596-15806-8 (paperback). xlix + 1160 pp. LCCN QA76.73.P98 L877 2009.

Lutz:2009:LPb

[Lut09b] Mark Lutz. Learning Python. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2009. ISBN 0-596-80539-X. ???? pp. LCCN QA76.73.P98 L877 2009.

Lutz:2013:LP

[Lut13] Mark Lutz. Learning Python. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fifth edition, 2013. ISBN 1-4493-5573-0. l + 1540 pp. LCCN QA76.73.P98 L877 2013. URL http://

proquestcombo.safaribooksonline.
com/9781449355722.

Lutz:2014:PPR

[Lut14] Mark Lutz. Python pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fifth edition, 2014. ISBN 1-4493-5701-6 (paperback). vii + 254 pp. LCCN QA76.73.P98 L89 2014.

Lauter:2020:FSA

[LV20] Christoph Lauter and Anastasia Volkova. A framework for semi-automatic precision and accuracy analysis for fast and rigorous deep learning. In Cornea et al. [CLT20], pages 103–110. ISBN 1-72817-120-2, 1-72817-121-0. ISSN 2576-2265. LCCN ???? URL

https://ieeexplore.ieee.
org/servlet/opac?punumber=
9146973.

Li:2018:PTP

[LVH^{+}18] Xinya Li, Chris R. Vernon, Mohamad I. Hejazi, Robert P. Link, Zhongwei Huang, Lu Liu, and Leyang Feng. Tethys — a Python package for spatial and temporal downscaling of global water withdrawals. Journal of Open Research Software, 6(1):9–??, February 09, 2018. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.197/.

Logg:2010:DAF

[LW10] Anders Logg and Garth N. Wells. DOLFIN: Automated finite element computing. ACM Transactions on Mathematical Software, 37(2):20:1– 20:28, April 2010. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Lee:2010:JSD

[LWH^{+}10] Byeongcheol Lee, Ben Wiedermann, Martin Hirzel, Robert Grimm, and Kathryn S. McKinley. Jinn: synthesizing dynamic bug detectors for foreign language interfaces. ACM SIGPLAN Notices, 45(6):36–49, June 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Logg:2012:DCP

[LWH12] Anders Logg, Garth N. Wells, and Johan Hake. DOLFIN: a C++/Python finite element library. In Logg et al. [LMW12], pages 173–225. CODEN LNCSA6. ISBN 3-642-23098-9 (print), 3-642-23099-7 (e-book). ISSN 1439-7358. LCCN ???? URL

                                        http://link.springer.com/
                                        content/pdf/10.1007/978-

3-642-23099-8_10. The soft-

ware developed by the FEniCS Project is free for all to use and modify (licensed under the GNU (L)GPL), and so is this book.

Lym:2020:PMT

[LWV20] Jonathan Lym, Gerhard R.

Wittreich, and Dionisios G. Vlachos. A Python Multiscale Thermochemistry Toolbox (pMuTT) for thermochemical and kinetic parameter estimation. Computer Physics Communications, 247(??): Article 106864, February 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465519302516.

Liu:2018:PPB

[LYX^{+}18] Benyuan Liu, Bin Yang, Canhua Xu, Junying Xia, Meng Dai, Zhenyu Ji, Fusheng You, Xiuzhen Dong, Xuetao Shi, and Feng Fu. pyEIT: a Python based framework for Electrical Impedance Tomography. SoftwareX, 7 (??):304–308, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301407.

Liu:2022:EUF

[LYZ^{+}22] Shun Liu, Junjie Yang, Xianxian Zeng, Haiying Song, Jian Cen, and Weichao Xu. An efficient and user-friendly software tool for ordered multi-class receiver operating characteristic analysis based on Python. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001078.

Millman:2011:PSE

[MA11] K. Jarrod Millman and Michael Aivazis. Python for scientists and engineers. Computing in Science and Engineering, 13(2): 9–12, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

MacLachlan:1991:CCL

[Mac91]  Rob MacLachlan.  CMU

Common Lisp user’s manual. Research paper CMU-CS-91-108, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, February 1991. vi + 168 pp. This is a revised version of Technical Report CMU-CS-87-156.

MacLachlan:1992:CCL

[Mac92a]  Rob MacLachlan.  CMU

Common Lisp user’s manual. Research paper CMU-CS-92-161, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, July 1992. v + 142 pp. Supersedes Technical Reports CMU-CS-87-156 and CMU-CS-91-108.

MacLachlan:1992:PCC

[Mac92b] Robert A. MacLachlan. Python compiler for CMU Common Lisp. In ACM [ACM92], pages 235–246. ISBN 0-89791-483-X, 0-89791-481-

3. LCCN QA76.73.L23A26 1992. Also published as LISP Pointers, vol. V, no. 1, January-March, 1992. ACM order no. 552920.

Muller:2014:SAP

[MAC14] Stefan C. Muller, Gustavo Alonso, and Andre Csillaghy. Scaling astroinformatics: Python + automatic parallelization. Computer, 47(9): 41–47, September 2014. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://csdl.
computer.org/csdl/mags/
co/2014/09/mco2014090041-
abs.html.

Macleod:2021:GPP

[MAC^{+}21] Duncan M. Macleod, Joseph S.

Areeda, Scott B. Coughlin, Thomas J. Massinger, and Alexander L. Urban. GWpy: a Python package for gravitational-wave astrophysics. SoftwareX, 13 (??):Article 100657, January 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000029.

Madhavan:2015:MPD

[Mad15] Samir Madhavan. Mastering Python for data science: explore the world of data science through Python and learn how to make sense of data. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-015-1, 1-78439-262-6 (e-book). ???? pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781784390150.

Mandanici:2021:SPG

[MAFM21] A. Mandanici, S. Alessandro Sarà, G. Fiumara, and G. Mandaglio. Studying physics, getting to know Python: RC circuit, simple experiments, coding, and data analysis with Raspberry Pi. Computing in Science and Engineering, 23(1):93– 96, 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mattson:2021:PPM

[MAG21] Timothy G. Mattson, Todd A. Anderson, and Giorgis Georgakoudis. PyOMP: Multithreaded parallel programming in Python. Computing in Science and Engineering, 23(6):77–80, November/December 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Martelli:2002:PC

[Mar02] David Ascher Alex Martelli, editor. Python Cookbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00167-3. xxix + 574 pp. LCCN QA76.73.P98 P983 2002 Stacks. US$39.95. URL

http://safari.oreilly.com/
0596001673;   http://www.
oreilly.com/catalog/pythoncook.

Martelli:2003:PN

[Mar03] Alex Martelli. Python in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00188-6. xv + 636 pp. LCCN QA76.73.P98 M37 2003. URL http://

www.oreilly.com/catalog/
9780596001889.

Martelli:2006:PN

[Mar06] Alex Martelli. Python in a nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2006. ISBN 0-596-10046-9. xiv + 695 pp. LCCN QA76.73.P98 M37 2006eb; QA76.73.P98 M37 2006; QA76.73.P98. URL

http://www.oreilly.com/
catalog/9780596100469.

Martelli:2007:PC

[Mar07] Alex Martelli. Python en concentré. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 2-84177-452-X. xvi + 802 pp. LCCN ???? French translation by Éric Jacoboni and Yann Serra.

Margolis:2017:PSP

[Mar17] Benjamin W. L. Margolis. SimuPy: a Python framework for modeling and simulating dynamical systems. Journal of Open Source Software, 2(17):396:1, September 2017. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00396.

Marowka:2018:PAH

[Mar18] Ami Marowka. Python accelerators for high-performance computing. The Journal of Supercomputing, 74(4):1449– 1460, April 2018. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Martignano:2021:SAA

[Mar21] Maurizio Martignano. Static analysis for Ada, C/C++ and Python: Different languages, different needs. ACM SIGADA Ada Letters, 41(2): 77–80, December 2021. URL

https://dl.acm.org/doi/
10.1145/3530801.3530807.

McCann:2022:COS

[MAS^{+}22] Maile P. McCann, Dylan L. Anderson, Christopher R. Sherwood, Brittany Bruder, A. Spicer Bak, and Katherine L. Brodie.

CoastalImageLib: an open-

source Python package for creating common coastal image products. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001339.

Matloff:2016:BRN

[Mat16] Norman Matloff. Book review: Numerical Python: a Practical Techniques Approach for Industry. Journal of Statistical Software, 70 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v070b04;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v070b04/v70b04.pdf.

Maurer:2002:CPL

[Mau02] W. Douglas Maurer. The comparative programming languages course: a new chain of development. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1): 336–340, March 2002. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Mayer:2017:PNP

[May17] Andreas Mayer. Noisyopt: a Python library for optimizing noisy functions. Journal of Open Source Software, 2(13):258:1, May 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00258.

May:2021:MLG

[May21]  Simon May.  minimal-

lagrangians: Generating and studying dark matter model Lagrangians with just the particle content. Computer Physics Communications, 261(??):Article 107773, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520303878.

Millman:2007:AFM

[MB07] K. Jarrod Millman and Matthew Brett. Analysis of functional magnetic resonance imaging in Python. Computing in Science and Engineering, 9(3): 52–55, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mahar:2017:PHP

[MB17] Sara Mahar and Matthew Bellis. hmis: a Python tool to visualize and analyze HMIS data. Journal of Open Source Software, 2(18):384:1–384:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00384.

Mohanan:2019:FCA

[MBA19] Ashwin Vishnu Mohanan, Cyrille Bonamy, and Pierre Augier. FluidFFT: Common API (C++ and Python) for Fast Fourier Transform HPC libraries. Journal of Open Research Software, 7(1):10–??, April 01, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.238/.

Marcolini:2022:HPL

[MBA^{+}22] Alessia Marcolini, Nicole Bussola, Ernesto Arbitrio, Mohamed Amgad, Giuseppe Jurman, and Cesare Furlanello. histolab: a Python library for reproducible Digital Pathology preprocessing with automated testing. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001558.

Misra:2009:MCT

[MBK09] Ananya Misra, Douglas Blank, and Deepak Kumar.

A music context for teaching introductory computing. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 41 (3):248–252, September 2009. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’09.

Mohanan:2019:FMO

[MBLA19] Ashwin Vishnu Mohanan, Cyrille Bonamy, Miguel Calpe Linares, and Pierre Augier. FluidSim: Modular, objectoriented Python package for high-performance CFD simulations. Journal of Open Research Software, 7(1):14–??, April 26, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.239/.

Morales:2022:PRO

[MBP^{+}22] Félix Morales, Luis Bernal, Gustavo Pereira, Sandra Pérez-Buitrago, Michael Kammer, and D. H. Stalder. PytuTester: RaspberryPi open-source ventilator tester. HardwareX, 12:??, October 2022. CODEN ???? ISSN 2468-0672. URL http:/

/www.sciencedirect.com/
science/article/pii/S2468067222000797.

Moore:2007:PPF

[MBW07] Dana Moore, Raymond Budd, and William Wright. Professional Python frameworks: Web 2.0 programming with Django and TurboGears. Wrox professional guides. Wiley, New York, NY, USA, 2007. ISBN 0-470-13809-2 (paperback). xxvii + 420 pp. LCCN TK5105.888 .M663

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0741/2007032138-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0741/2007032138-
t.html;   http://www.loc.
gov/catdir/enhancements/
fy0804/2007032138-b.html.

McFee:2016:PRE

[McF16]  Brian McFee.  resampy:

efficient sample rate conversion in Python. Journal of Open Source Software, 1(8):125:1, December 2016. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00125.

Mariano:2022:ATI

[MCF^{+}22] Benjamin Mariano, Yanju Chen, Yu Feng, Greg Durrett, and Isil Dillig. Automated transpilation of imperative to functional code using neuralguided program synthesis. Proceedings of the ACM on Programming Languages (PACMPL), 6(OOPSLA1): 71:1–71:27, April 2022. CO-DEN ???? ISSN 2475-1421 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3527315.

McGrath:1998:XPP

[McG98a] S. McGrath. XML programming in Python. Dr. Dobb’s Journal of Software Tools, 23(2):82–??, 84–87, 101–104, February 1998. CODEN DDJOEB. ISSN 1044-789X.

McGrath:1998:IPX

[McG98b]  Sean McGrath.  Internet

programming: XML programming in Python. Dr. Dobb’s Journal of Software Tools, 23(2):82, 84–87, 101– 104, February 1998. CODEN DDJOEB. ISSN 1044-789X.

McGrath:1998:PBT

[McG98c] Sean McGrath. Programmer’s bookshelf: The Tcl/Tk and Python scripting environments. Dr. Dobb’s Journal of Software Tools, 23(10):143, 145, October 1998. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ddj/1998/1998_10/index.
                                        htm.

McGrath:2000:XPP

[McG00] Sean McGrath. XML processing with Python. The Charles F. Goldfarb series on open information management. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2000. ISBN 0-13-021119-2. xxiv + 527 pp. LCCN QA76.76.H94 M3885 2000. US$44.99. URL http:/

                                        /www.phptr.com/ptrbooks/
                                        ptr_0130211192.html. In-

cludes CD-ROM.

McGugan:2007:BGD

[McG07a]  Will McGugan.  Begin-

ning game development with Python and Pygame: from novice to professional. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2007. ISBN 1-59059-872-5. xxiii + 316 pp. LCCN ????

McGuire:2007:GSP

[McG07b]  Paul McGuire.  Getting

started with pyparsing. O’Reilly shortcuts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51423-9. LCCN QA76.76.A65 M33 2007eb; QA76.76.A65. URL http://

www.oreilly.com/catalog/
9780596514235.

Martin:2019:QPM

[MCGK19] R. D. Martin, Q. Cai, T. Garrow, and C. Kapahi. QExpy: a Python-3 module to support undergraduate physics laboratories. SoftwareX, 10(??):Article 100273, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271101930144X.

McKinney:2012:PDA

[McK12] Wes McKinney. Python for Data Analysis. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-1979-3 (paperback). 400 pp. LCCN ???? EUR 32.00; UK £30.99.

McKinley:2016:PWU

[McK16] Kathryn S. McKinley. Programming the world of uncertain things (keynote). ACM SIGPLAN Notices, 51(1):1– 2, January 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marr:2015:TVP

[MD15] Stefan Marr and Stéphane Ducasse. Tracing vs. partial evaluation: comparing meta-compilation approaches for self-optimizing interpreters. ACM SIG-PLAN Notices, 50(10):821– 839, October 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Makowski:2017:PNP

[MD17] Dominique Makowski and

Léo Dutriaux. Neuropsydia.py:

a Python module for creating experiments, tasks and questionnaires. Journal of Open Source Software, 2 (19):259:1–259:2, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00259.

Melchert:2021:PPT

[MD21] Oliver Melchert and Ayhan Demircan. pyGLLE: a Python toolkit for solving the generalized Lugiato–Lefever equation. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S235271102100073X.

Melchert:2022:PFP

[MD22a] O. Melchert and A. Demircan. py-fmas: a Python package for ultrashort optical pulse propagation in terms of forward models for the analytic signal. Computer Physics Communications, 273(??):Article 108257, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521003696.

Melchert:2022:GPG

[MD22b] Oliver Melchert and Ayhan

Demircan. GNLStools.py: a

generalized nonlinear Schrödinger Python module implementing different models of input pulse quantum noise. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001509.

Malard:2022:EDP

[MDR^{+}22] Florian Malard, Laura Danner, Emilie Rouzies, Jesse G. Meyer, Ewen Lescop, and Stéphanie Olivier-Van Stichelen. EpyNN: Educational Python for Neural Networks. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022000905.

Mehrotra:2018:OSR

[MDRN18] Pavan Mehrotra, Sabar Dasgupta, Samantha Robertson, and Paul Nuyujukian. An open-source realtime computational platform (short WIP paper). ACM SIGPLAN Notices, 53(6):109–112, June 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mehta:2015:MPS

[Meh15] Hemant Kumar Mehta. Mastering Python scientific computing: a complete guide for Python programmers to master scientific computing using Python APIs and tools. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78328-883-3, 1-78328-882-

5. ???? pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781783288823.

Mentis:2009:RAD

[Men09] Alexander S. Mentis. A robotics API dialect for type-safe robots: translating Myro to Ada. ACM SIGADA Ada Letters, 29(3): 91–102, December 2009. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Mertz:2003:TPP

[Mer03] David Mertz. Text processing in Python. Addison-Wes-ley, Reading, MA, USA, 2003. ISBN 0-321-11254-7. xix + 520 pp. LCCN QA76.9.T48 M47 2003.

Maulik:2022:PSD

[MFL^{+}22] Romit Maulik, Dimitrios K.

Fytanidis, Bethany Lusch, Venkatram Vishwanath, and Saumil Patel. Python-FOAM: In-situ data analyses with OpenFOAM and Python. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750322001387.

Meng:2021:MPP

[MFQ^{+}21] Siqin Meng, Zhendong Fu, Jianfei Qin, Xiaobai Ma, Yuqing Li, Lijie Hao, Yuntao Liu, Kai Sun, and Dongfeng

Chen.  magcoilcalc: a

Python package for modeling and optimization of axisymmetric magnet coils generating uniform magnetic field for noble gas spin-polarizers. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021001060.

Mardan:2023:PPP

[MGFO23] Amir Mardan, Bernard Giroux, and Gabriel Fabien-Ouellet. PyFWI: a Python package for full-waveform inversion and reservoir monitoring. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000808.

Myers:2007:PUS

[MGS07] Christopher R. Myers, Ryan N. Gutenkunst, and James P. Sethna. Python unleashed on systems biology. Computing in Science and Engineering, 9(3):34–37, May/June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Minichino:2015:LOC

[MH15] Joe Minichino and Joseph Howse. Learning OpenCV 3 computer vision with Python: unleash the power of computer vision with Python using OpenCV. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-977-2, 1-78528-384-7. LCCN TA1634. URL http://

proquest.safaribooksonline.
com/?fpi=9781785283840.

Moore:2018:PSS

[MH18] Jason K. Moore and Mont

Hubbard. skijumpdesign: A

ski jump design tool for specified equivalent fall height. Journal of Open Source Software, 3(28):818:1–818:3, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00818.

Miller:2014:MTP

[Mil14] Thomas Miller. Modeling techniques in predictive analytics with Python and R: a guide to data science. Pearson Education, Upper Saddle River, NJ, 2014. ISBN 0-13-389206-9 (hardcover), 0-13-389212-3. xviii + 418 pp. LCCN QA76.73.P98.

Miller:2015:MDS

[Mil15] Thomas W. Miller. Marketing data science: modeling techniques in predictive analytics with R and Python. Pearson Education, Old Tappan, NJ, USA, 2015. ISBN 0-13-388766-9, 0-13-388762-6. ???? pp. LCCN HF5415. URL http://

proquest.safaribooksonline.
com/?fpi=9780133887662.

Miller:2018:HDA

[Mil18] Curtis Miller. Hands-On Data Analysis with NumPy and Pandas: Implement Python Packages from Data Manipulation to Processing. Packt Publishing Ltd, Birmingham, 2018. ISBN 1-78953-424-0, 1-78953-079-2. 166 pp. LCCN ???? URL http://

public.eblib.com/choice/
PublicFullRecord.aspx?p=
5446042.

Miranda:2018:PPR

[Mir18] Lester James V. Miranda. PySwarms: a research toolkit for particle swarm optimization in Python. Journal of Open Source Software, 3(21):433:1–433:2, January 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00433.

Mishra:2019:PRP

[Mis19] Pradeepta Mishra. PyTorch Recipes: a Problem–solution Approach. Apress, Berkeley, CA, USA, 2019. ISBN 1-4842-4257-2, 1-4842-4258-0 (ebook), 1-4842-4259-9. LCCN QA76.87. URL http://

www.vlebooks.com/vleweb/
product/openreader?id=none\
%26isbn=9781484242582.

Mitchell:2000:DAS

[Mit00] Scott Mitchell. Designing Active Server Pages: Scott Mitchell’s Guide to Writing Reusable Code. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 0-596-00044-8. xii + 348 pp. LCCN TK5105.8885.A26 M58 2000. US$29.95.

URL http://catdir.loc.
gov/catdir/enhancements/
fy0715/00062331-d.html;
 http://catdir.loc.gov/
catdir/enhancements/fy0912/
00062331-b.html;   http:
//catdir.loc.gov/catdir/
enhancements/fy1008/00062331-
t.html;   http://www.oreilly.
com/catalog/9780596000448;
 http://www.oreilly.com/
catalog/designasp.

Mitchell:2015:WSP

[Mit15] Ryan Mitchell. Web scraping with Python: collecting data from the modern web. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-1029-1 (paperback), 1-4919-1028-

3. xiii + 238 pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/9781491910283.

Mulansky:2016:PPP

[MK16] Mario Mulansky and Thomas Kreuz. PySpike — a Python library for analyzing spike train synchrony. SoftwareX, 5(??):178–182, ???? 2016. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711016300255.

Meller:2016:PDM

[ML16a] Yosef Meller and Alex Liberzon. Particle data management software for 3D particle tracking velocimetry and related applications — the Flowtracks package. Journal of Open Research Software, 4(1):e23–??, June 16, 2016. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.101/.

Mortensen:2016:HPP

[ML16b] Mikael Mortensen and Hans Petter Langtangen. High performance Python for direct numerical simulations of turbulent flows. Computer Physics Communications, 203(??):53–65, June 2016. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465516300200.

Mukha:2018:EPP

[ML18] Timofey Mukha and Mattias Liefvendahl. Eddylicious: a Python package for turbulent inflow generation. SoftwareX, 7(??):107–111, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018300487.

Marques:2022:ISF

[MLB22] Henrique Marques, Nuno Laranjeiro, and Jorge Bernardino.

Injecting software faults in Python applications. Empirical Software Engineering, 27 (1):??, January 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-021-10047-9.

Malakhov:2018:CMT

[MLGW18] Anton Malakhov, David Liu, Anton Gorshkov, and Terry Wilmarth. Composable multi-threading and multiprocessing for numeric libraries. In Fatih Akici, David Lippa, Dillon Niederhut, and M Pacer, editors, Proceedings of the 17th Python in Science Conference, Austin, TX, 9– 15 July 2018, pages 15–21. ????, ????, 2018. URL http:

//conference.scipy.org/
proceedings/scipy2018/anton_
malakhov.html.

Maruch:2006:PD

[MM06] Stef Maruch and Aahz Maruch. Python for dummies. Wiley, New York, NY, USA, 2006. ISBN 0-471-77864-8 (paperback). xx + 410 pp. LCCN QA76.73.P98 M374 2006.

URL http://www.loc.gov/
catdir/enhancements/fy0741/
2006924031-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0741/2006924031-d.html;
 http://www.loc.gov/catdir/
toc/fy0713/2006924031.html.

Martelli:2005:PC

[MMA05] Alex Martelli, Anna Martelli Ravenscroft, and David Ascher, editors. Python cookbook. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2005. ISBN 0-596-00797-3. xxxiii + 807 pp. LCCN QA76.73.P98 P983 2005.

Meinke:2008:SVS

[MMEH08] Jan H. Meinke, Sandipan Mohanty, Frank Eisenmenger, and Ulrich H. E. Hansmann. SMMP v. 3.0 — simulating proteins and protein interactions in Python and Fortran. Computer Physics Communications, 178(6):459–470, March 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465507004614.

Micieli:2019:NTP

[MMG19] Davide Micieli, Triestino Minniti, and Giuseppe Gorini. NeuTomPy toolbox, a Python package for tomographic data processing and reconstruction. SoftwareX, 9(??): 260–264, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018302103.

Migallon:2011:PPL

[MMP11] Héctor Migallón, Violeta Migallón, and José Penadés. A Parallel Python library for nonlinear systems. The Journal of Supercomputing, 58 (3):438–448, December 2011. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=58&
issue=3&spage=438.

Mandanici:2022:SPP

[MMP^{+}22] Andrea Mandanici, Giuseppe Mandaglio, Giovanni Pirrotta, Valeria Conti Nibali, and Giacomo Fiumara. Simple physics with Python: a workbook on introductory physics with open-source software. Computing in Science and Engineering, 24(2):74– 78, March/April 2022. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Martins:2009:POO

[MMT09] Joaquim R. R. A. Martins, Christopher Marriage, and Nathan Tedford. pyMDO: an object-oriented framework for multidisciplinary design optimization. ACM Transactions on Mathematical Software, 36 (4):20:1–20:25, August 2009. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Mushtaq:2014:ACG

[MO14] Asif Mushtaq and Kåre Olaussen. Automatic code generator for higher order integrators. Computer Physics Communications, 185(5):1461–1472, May 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465514000253.

Mohit:2015:PPT

[Moh15] Raj Mohit. Python penetration testing essentials. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-858-6, 1-78439-588-9 (e-book). LCCN QA76.73.P98.

Molina:2021:HOS

[MOM21] Tulio Molina, Juan Ortega, and Juan Muñoz. HELMpy, open source package of power flow solvers, including the holomorphic embedding load flow method (HELM), developed on Python 3. Journal of Open Research Software, 9(1):23–??, August 18, 2021. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.310/.

Mooers:2021:PSL

[Moo21] B. H. M. Mooers. A PyMOL snippet library for Jupyter to boost researcher productivity. Computing in Science and Engineering, 23(2):47– 53, March/April 2021. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Morgan:2017:PLP

[Mor17] Benjamin J. Morgan. lattice mc:

a Python lattice-gas Monte Carlo module. Journal of Open Source Software, 2 (13):247:1–247:2, May 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00247.

Moshiri:2020:TMS

[Mos20a] N. Moshiri. TreeSwift: a massively scalable Python tree package. SoftwareX, 11 (??):Article 100436, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711019300767.

Mosteo:2020:RBA

[Mos20b] Alejandro R. Mosteo. RCLAda, or bringing Ada to the Robot Operating System. ACM SIGADA Ada Letters, 39(2): 35–40, April 2020. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https://

dl.acm.org/doi/10.1145/
3394514.3394518.

Malloy:2019:EAT

[MP19a] Brian A. Malloy and James F.

Power. An empirical analysis of the transition from Python 2 to Python 3. Empirical Software Engineering, 24(2): 751–778, April 2019. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s10664-018-9637-2.

Malloy:2019:GEM

[MP19b] Brian A. Malloy and James F. Power. Grammar engineering for multiple front-ends for Python. Software—Practice and Experience, 49(3): 380–400, March 2019. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Melchor-Placencia:2021:OPC

[MPMC21] Carlos Melchor-Placencia and Christian Málaga-Chuquitaype. OpenMoist: a Python code for transient moisture transfer analysis. SoftwareX, 15 (??):??, July 2021. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711021000571.

Miller:2007:CSP

[MR07] Bradley N. Miller and David L. Ranum. Computer science: the Python programming language. Jones and Bartlett Publishers, Sudbury, MA, USA, 2007. ISBN 0-7637-4316-X. 59 pp. LCCN QA76.73.P98 M537 2007.

Miller:2009:PPC

[MR09] Bradley N. Miller and David L. Ranum. Python programming in context. Jones and Bartlett Publishers, Sudbury, MA, USA, 2009. ISBN 0-7637-4602-9 (paperback). xxv + 492 pp. LCCN QA76.73.P98 M544 2009.

Mulryne:2018:PPP

[MR18] David J. Mulryne and John W.

Ronayne. PyTransport: a

Python package for the calculation of inflationary correlation functions. Journal of Open Source Software, 3(23): 494:1–494:2, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00494.

Magalhaes:2022:PPI

[MR22] Tiago E. C. Magalhães and José M. Rebordão. PyWolf: a PyOpenCL implementation for simulating the propagation of partially coherent light. Computer Physics Communications, 276(??): Article 108336, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522000546.

Martelli:2006:PPE

[MRA06] Alex Martelli, Anna Martelli Ravensroft, and David Ascher. Python par l’exemple. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2006. ISBN 2-84177-379-5. xxv + 525 pp. LCCN ???? French translation by Éric Jacoboni.

Meier:2017:PVM

[MRG17] Remigius Meier, Armin Rigo, and Thomas R. Gross. Parallel virtual machines with RPython. ACM SIGPLAN Notices, 52(2):48–59, February 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meier:2018:VMD

[MRG18] Remigius Meier, Armin Rigo, and Thomas R. Gross. Virtual machine design for parallel dynamic programming languages. Proceedings of the ACM on Programming Languages (PACMPL), 2 (OOPSLA):109:1–109:25, October 2018. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3276479.

Meng:2022:PPL

[MRT^{+}22] Fanwang Meng, Michael Richer, Alireza Tehrani, Jonathan La, Taewon David Kim, Paul W. Ayers, and Farnaz Heidar-Zadeh. Procrustes: a Python library to find transformations that maximize the similarity between matrices. Computer Physics Communications, 276(??): Article 108334, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522000522.

Matthew:2000:PLP

[MS^{+}00] Neil Matthew, Richard Stones, et al. Professional Linux programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Myers:2007:PEC

[MS07] Christopher R. Myers and James P. Sethna. Python for education: Computational methods for nonlinear systems. Computing in Science and Engineering, 9(3): 75–79, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Mertz:2013:GIP

[MS13] Andrew Mertz and William Slough. A gentle introduction to PythonTEX. TUGboat, 34(3):302–312, 2013. ISSN 0896-3207.

URL http://www.tug.org/
TUGboat/tb34-3/tb108mertz.
pdf.

Malthe-Sorenssen:2015:EMU

[MS15] Anders Malthe-Sørenssen. Elementary mechanics using Python: a modern course combining analytical and numerical techniques. Undergraduate lecture notes in physics. Springer, Cham, Switzerland, 2015. ISBN 3-319-19595-6 (print), 3-319-19596-4 (e-book). ISSN 2192-4791 (print), 2192-4805 (electronic). xiii + 590 pp. LCCN QA805 .M35 2015.

Mardal:2007:UPS

[MSL^{+}07] Kent-Andre Mardal, Ola Skavhaug, Glenn T. Lines, Gunnar A. Staff, and Åsmund Ødegård. Using Python to solve partial differential equations. Computing in Science and Engineering, 9(3): 48–51, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Meurer:2017:SSC

[MSP^{+}17] Aaron Meurer, Christopher P. Smith, Mateusz Paprocki, Ondřej Čertík, Sergey B. Kirpichev, Matthew Rocklin, Amit Kumar, Sergiu Ivanov, Jason K. Moore, Sartaj Singh, Thilina Rathnayake, Sean Vig, Brian E. Granger, Richard P. Muller, Francesco Bonazzi, Harsh Gupta, Shivam Vats, Fredrik Johansson, Fabian Pedregosa, Matthew J. Curry, Andy R. Terrel, Štěpán Roučka, Ashutosh Saboo, Isuru Fernando, Sumith Kulal, Robert Cimrman, and Anthony Scopatz. SymPy: symbolic computing in Python. PeerJ Computer Science, 3:e103:1– e103:27, January 2017. ISSN 2376-5992.

Mamidi:2022:PAG

[MSP^{+}22] Nischay Ram Mamidi, Dhruv Saxena, Kumar Prasun, Anil Nemili, Bharatkumar Sharma, and S. M. Deshpande. Performance analysis of GPU accelerated meshfree q-LSKUM solvers in Fortran, C, Python, and Julia. In IEEE, editor, 2022 IEEE 29th International Conference on High Performance Computing, Data, and Analytics (HiPC), pages 156– 165. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022.

Miller:2003:OCP

[MSR03] W. W. Miller, C. Sontag, and J. F. Rose. OPUS: a CORBA pipeline for Java, Python, and Perl applications. Astronomical Society of the Pacific Conference Series, 295: 261–264, 2003. CODEN ???? ISSN 1050-3390.

Moura:2019:UJP

[MSS^{+}19] R. A. R. Moura, M. A. O.

Schroeder, S. J. S. Silva, E. G. Nepomuceno, P. H. N. Vieira, and A. C. S. Lima. The usage of Julia programming in grounding grids simulations : An alternative to MATLAB and Python. In 2019 International Symposium on Lightning Protection (XV SIPDA), pages 1– 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2019.

Menvouta:2023:DPP

[MSV23] Emmanuel Jordy Menvouta, Sven Serneels, and Tim Verdonck. direpack: a Python 3 package for state-of-the-art statistical dimensionality reduction methods. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102200200X.

Mount:2008:PRF

[MSW08] Sarah Mount, James Shuttleworth, and Russel Winder. Python for Rookies: a first course in programming. Thomson Learning (EMEA), London, UK, 2008. ISBN 1-84480-701-0. xxi + 454 pp. LCCN QA76.73.P98 M68 2008.

McNicholas:2018:DSJ

[MT18] Paul D. McNicholas and Peter A. Tait. Data Science with Julia. Taylor and Francis, CRC Press, Boca Raton, FL, USA, 2018. ISBN 1-138-49998-6 (paperback), 1-351-01364-5 (e-book: Mobipocket), 1-351-01365-3 (e-book), 1-351-01366-1 (ebook: PDF), 1-351-01367-X (e-book). 241 pp. LCCN QA76.73.J85 M37 2018.

McNicholas:2019:DSJ

[MT19] Paul D. McNicholas and Peter A. Tait. Data science with Julia. Chapman and Hall/ CRC, Boca Raton, FL, USA, 2019. ISBN 1-138-49999-4, 1-351-01364-5, 1-351-01365-3, 1-351-01366-1, 1-351-01367-X. xix + 220 pp. LCCN QA76.73.J85 M37 2019eb.

Meng:2018:MPP

[MTPHH18] Siqin Meng, Rasmus Toft-Petersen, Lijie Hao, and Klaus Habicht. multiflexxlib: a Python package for data reduction and visualization for the cold-neutron multi energy wide angle analyzer MultiFLEXX. SoftwareX, 7 (??):309–312, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301055.

McCubbine:2018:GPC

[MTS^{+}18] Jack McCubbine, Fabio Caratori Tontini, Vaughan Stagpoole, Euan Smith, and Grant O’Brien. Gsolve, a Python computer program with a graphical user interface to transform relative gravity survey measurements to absolute gravity values and gravity anomalies. SoftwareX, 7(??):122– 128, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018300566.

Mueller:2014:BPP

[Mue14] John Mueller. Beginning programming with Python for dummies. Wiley, New York, NY, USA, 2014. ISBN 1-118-89145-7 (paperback), 1-118-89147-3 (e-book), 1-118-89149-X (ePDF). xiv + 382 pp. LCCN QA76.73.P98 M839 2014. URL http:/

                                        /www.dummies.com/how-to/
                                        content/beginning-python-
                                        for-dummies-cheat-sheet.
                                        html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1513/2014935516-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1513/2014935516-
                                        t.html.

Mullner:2013:FFH

[Mül13] Daniel Müllner. fastcluster: Fast hierarchical, agglomerative clustering routines for R and Python. Journal of Statistical Software, 53(9):??, May 2013. CODEN JSSOBK. ISSN 1548-7660. URL http:

                                        //www.jstatsoft.org/v53/
                                        i09.

Muller:2023:PSC

[Mül23] Matthias Müller. PyblioNet — software for the creation, visualization and analysis of bibliometric networks. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023002613.

Murray:2018:PPP

[Mur18] Steven G. Murray. powerbox:

a Python package for creating structured fields with isotropic power spectra. Journal of Open Source Software, 3(28):850:1–850:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00850.

Ma:2020:PPP

[MWK^{+}20] He Ma, Wennie Wang, Siyoung Kim, Man-Hin Cheng, Marco Govoni, and Giulia Galli. PyCDFT: a Python package for constrained density functional theory. Journal of Computational Chemistry, 41(20):1859–1867, July 30, 2020. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Mauro:2020:KTC

[MWM20] Yihong Z. Mauro, Collin J.

Wilkinson, and John C. Mauro. KineticPy: a tool to calculate long-time kinetics in energy landscapes with broken ergodicity. SoftwareX, 11 (??):Article 100393, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711019303516.

Mirzaev:2018:PEP

[MWS18] Inom Mirzaev, Drew F. K.

Williamson, and Jacob G. Scott. egtplot: a Python package for three-strategy evolutionary games. Journal of Open Source Software, 3(26):735:1–735:4, June 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00735.

Myridis:2015:IPA

[Myr15] Nikolaos E. Myridis. Image processing and acquisition using Python, by Ravishankar Chityala and Sridevi Pudipeddi, Scope: textbook. Level: general readership, undergraduate, teacher. Contemporary Physics, 56(2):243, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ma:2023:PPT

[MZL^{+}23] Jinfeng Ma, Hua Zheng, Ruonan Li, Kaifeng Rao, Yanzheng Yang, and Weifeng

Li. PyVecContour: a Python

toolkit for vectorized isosurface mapping. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000134.

Nagel:2006:EPY

[Nag06]  William Nagel.  Embed-

ding Python in your C programs. Linux journal, 2006 (142):8, February 2006. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Nagar:2017:BJP

[Nag17]  Sandeep Nagar.  Begin-

ning Julia Programming: For Engineers and Scientists. Apress, Berkeley, CA, USA, 2017. ISBN 1-4842-3170-8, 1-4842-3171-6. xxi + 351 + 20 + 18 pp. LCCN QA76.7-76.73; QA76.76.C65. URL

http://www.springerlink.
com/content/978-1-4842-
3171-5.

Nair:2014:GSB

[Nai14] Vineeth G. Nair. Getting Started with Beautiful Soup: build your own web scraper and learn all about web scraping with Beautiful Soup. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-956-2, 1-78328-955-4, 1-4619-5720-6 (e-book), 1-306-40146-1 (e-book). LCCN QA76.9.D343. URL http://

proquest.tech.safaribooksonline.
de/9781783289554.

Niemeyer:2017:PAJ

[NCS17] Kyle E. Niemeyer, Nicholas J.

Curtis, and Chih-Jen Sung. pyJac: Analytical Jacobian generator for chemical kinetics. Computer Physics Communications, 215(??): 188–203, June 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465517300462.

Nunez-Elizalde:2018:PCS

[NEGZG18] Anwar O. Nunez-Elizalde, James S. Gao, Tianjiao Zhang, and Jack L. Gal-

                                        lant. cottoncandy: sci-

entific Python package for easy cloud storage. Journal of Open Source Software, 3(28):890:1–890:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00890.

Nelli:2015:PDA

[Nel15] Fabio Nelli. Python data analytics: data analysis and science using Pandas, matplotlib, and the Python programming language. The expert’s voice in Python. Apress, Berkeley, CA, USA, 2015. ISBN 1-4842-0959-1 (paperback), 1-4842-0958-3 (e-book). xxi + 337 pp. LCCN QA76.73.P98.

Nguyen:2008:IPCa

[Ngu08a] Binh Q. Nguyen. The implementation of a Python class for structuring network data collected in a test bed. Technical report D-arl-tr-4423, AD-a479 698, United States Army Research Lab, Adelphi, MD, USA, 2008. 30 pp.

Nguyen:2008:IPCb

[Ngu08b] Binh Q. Nguyen. An introduction to Python (a onehour tour). Technical report AD-arl-tn-0328, ADa484 316, United States Army Research Lab, Adelphi, MD, USA, 2008. 20 pp.

Nguyen:2008:PPE

[Ngu08c] Binh Q. Nguyen. pyGFC — a Python extension to the C++ Geodesy Foundation Classes. Technical report AD-arl-tr-4623, AD-a488 020, United States Army Research Lab, Adelphi, MD, USA, 2008.

Nguyen:2008:CSL

[Ngu08d] Quan Nguyen. CAD scripting language: a collection of Perl, Ruby, Python, TCL and Skill scripts. Ramacad, San Jose, CA, USA, 2008. ISBN 0-9777812-2-4. ???? pp. LCCN ????

Nguyen:2008:SLC

[Ngu08e] Quan Nguyen, editor. Scripting languages: a collection of Perl, Ruby, Python, TCL and Unix. Ramacad, San Jose, CA, USA, 2008. ISBN 0-9777812-3-2. ???? pp. LCCN ????

Nilsen:2007:PSC

[Nil07a] Jon K. Nilsen. Python in scientific computing: Applications to Bose–Einstein condensates. Computer Physics Communications, 177(1–2): 45, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465507001312.

Nilsen:2007:MIQ

[Nil07b] Jon Kristian Nilsen. MontePython: Implementing quantum Monte Carlo using Python. Computer Physics Communications, 177(10): 799–814, November 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465507003141.

Nixon:2015:GSP

[Nix15] Dan Nixon. Getting started with Python and Raspberry Pi. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78355-159-3. ???? pp. LCCN ????

Niskanen:2023:CPB

[NL23] Matti Niskanen and Timo Lähivaara. COMPOSTI: a Python-based program for seismic trans-dimensional inversion. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022002163.

Nguyen:2022:EPQ

[NM22] Thien Nguyen and Alexander J. McCaskey. Extending Python for quantumclassical computing via quantum just-in-time compilation. ACM Transactions on Quantum Computing (TQC), 3(4):24:1–24:25, December 2022. CODEN ???? ISSN 2643-6809 (print), 2643-6817 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3544496.

Naecker:2017:PPP

[NMGB17] Benjamin Naecker, Niru Maheswaranathan, Surya Ganguli, and Stephen Baccus. Pyret: a Python package for analysis of neurophysiology data. Journal of Open Source Software, 2(9):137:1, January 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00137.

Nolp:2023:SNC

[NO23] Klaus Nölp and Lena Oden.

Simplifying non-contiguous data transfer with MPI for Python. The Journal of Supercomputing, 79(17):20019– 20040, November 2023. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s11227-023-05398-7.

Nordh:2017:PPF

[Nor17] Jerker Nordh. pyParticleEst:

a Python framework for particle-based estimation methods. Journal of Statistical Software, 78(??):??, ???? 2017. CODEN JSSOBK. ISSN 1548-

7660. URL https://www.
jstatsoft.org/index.php/
jss/article/view/v078i03;
 https://www.jstatsoft.
org/index.php/jss/article/
view/v078i03/v78i03.pdf.

Nino-Ruiz:2023:ADA

[NRC23] Elías D. Nino-Ruiz and Randy Consuegra. AMLCS-DA: a data assimilation package in Python for Atmospheric General Circulation Models. SoftwareX, 22(??): ??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023000705.

Nabhani:2022:TOO

[NS22] Abbas Nabhani and Hanne K. Sjølie. TreeSim: an objectoriented individual tree simulator and 3D visualization tool in Python. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S235271102200139X.

Nazarov:2024:HLG

[NT24] N. A. Nazarov and V. V. Terekhov. High level GPUaccelerated 2D PIV framework in Python. Computer Physics Communications, 295(??):Article 109009, February 2024. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465523003545.

Nourisa:2022:COS

[NZPWR22] Jalil Nourisa, Berit Zeller-Plumhoff, and Regine Willumeit-Römer. CppyABM: an opensource agent-based modeling library to integrate C++ and Python. Software—Practice and Experience, 52(6): 1337–1351, June 2022. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Omar:2017:PSF

[OA17] Cyrus Omar and Jonathan Aldrich. Programmable semantic fragments: the design and implementation of typy. ACM SIGPLAN Notices, 52(3):81–92, March 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Overberg:2023:EOS

[OBWM23] Florian A. Overberg, Philipp C.

Böttcher, Dirk Witthaut, and Simon Morgenthaler. Emipy: an open-source Python-based tool to analyze industrial emissions in Europe. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023001541.

OConnor:2013:VPC

[O’C13] T. J. O’Connor. Violent Python: a cookbook for hackers, forensic analysts, penetration testers and security engineers. Syngress Publishing, Inc., Rockland, MA, USA, 2013. ISBN 1-59749-957-9 (paperback). xxv + 262 pp. LCCN QA76.73.P98 O26 2013.

Ortner:2020:MFP

[OC20] Michael Ortner and Lucas Gabriel Coliado Bandeira. Magpylib: a free Python package for magnetic field computation. SoftwareX, 11 (??):Article 100466, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711020300170.

Ochem:2009:MLP

[Och09]  Quentin Ochem.  Multi-

language programming with Ada. ACM SIGADA Ada Letters, 29(3):19–20, December 2009. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Ortin:2021:CPA

[OE21] Francisco Ortin and Javier Escalada. Cnerator: a Python application for the controlled stochastic generation of standard C source code. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S235271102100056X.

Olivier:2020:UGP

[OGA^{+}20a] Audrey Olivier, Dimitrios Giovanis, B. S. Aakash, Mohit Chauhan, Lohit Vandanapu, and Michael D. Shields. UQpy: a general purpose Python package and development environment for uncertainty quantification. Journal of Computational Science, 47, 2020. ISSN 1877-7503 (print), 1877-7511 (electronic).

Olivier:2020:PUG

[OGA^{+}20b] Audrey Olivier, Dimitris G.

Giovanis, B. S. Aakash, Mohit Chauhan, Lohit Vandanapu, and Michael D. Shields. UQpy: a general purpose Python package and development environment for uncertainty quantification. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750320305056.

Otis:2017:PPC

[OL17] Richard Otis and Zi-Kui Liu. pycalphad: CALPHADbased computational thermodynamics in Python. Journal of Open Research Software, 5(1):1–??, January 09, 2017. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.140/.

Oliphant:2007:PSC

[Oli07] Travis E. Oliphant. Python for scientific computing. Computing in Science and Engineering, 9(3):10–20, May/ June 2007. CODEN CSENFA.

ISSN 1521-9615 (print), 1558-

366X (electronic).

Orella:2021:HTA

[OLRLB21] Michael Julian Orella, McLain Evan Leonard, Yuriy Román-Leshkov, and Fikile Richard Brushett. High-throughput analysis of contact angle goniometry data using DropPy. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021000108.

Ortin:2014:SDL

[OMGDG14] Francisco Ortin, Sheila Mendez, Vicente García-Díaz, and Miguel Garcia. On the suitability of dynamic languages for hot-reprogramming a robotics framework: a Python case study. Software—Practice and Experience, 44(1): 77–104, January 2014. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Obonna:2023:DMM

[OOM^{+}23] Ugochukwu Onyekachi Obonna, Felix Kelechi Opara, Christian Chidiebere Mbaocha, Jude-Kennedy Chibuzo Obichere, Isdore Onyema Akwukwaegbu, Miriam Mmesoma Amaefule, , and Cosmas Ifeanyi Nwakanma. Detection of man-in-the-middle (MitM) cyber-attacks in oil and gas process control networks using machine learning algorithms. Future Internet, 15 (8):280, August 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/8/
280.

Oxvig:2014:PMP

[OPA^{+}14] Christian Oxvig, Patrick Pedersen, Thomas Arildsen, Jan Østergaard, and Torben Larsen. Magni: a Python package for compressive sampling and reconstruction of atomic force microscopy images. Journal of Open Research Software, 2(1):e29–??, October 07, 2014. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.bk/.

Orban:2018:BRI

[Orb18] Dominique Orban. Book review: Introduction to Computation and Programming Using Python. Second Edition, with Application to Understanding Data, by John V. Guttag. SIAM Review, 60(2): 483–485, ???? 2018. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Orlowski:1997:NSC

[Orl97] A. Orlowski. And now for something completely different? [Python language]. .EXE: the software developers’ magazine, 12(2):34–35, 37, 39, 41, July 1997. CO-DEN EXEEE5. ISSN 0268-6872.

Orr:2005:RDP

[Orr05] Mike Orr. Review: Dive into Python. Linux journal, 2005 (130):??, February 2005. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Osais:2017:CSF

[Osa17] Yahya E. Osais. Computer Simulation: a Foundational Approach Using Python, volume 101 of Chapman and Hall/CRC computer and information science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2017. ISBN 1-315-12029-1 (e-book), 1-351-63708-8 (ebook: Mobi), 1-4987-2682-8 (hardcover), 1-4987-2683-6 (e-book PDF). LCCN QA76.9.C65 O83 2017.

Ottoni:2018:HJP

[Ott18] Guilherme Ottoni. HHVM JIT: a profile-guided, regionbased compiler for PHP and hack. ACM SIGPLAN Notices, 53(4):151–165, April 2018. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ostrouchov:2018:PPA

[OZW18] Christopher Ostrouchov, Yanwen Zhang, and William J. Weber. pysrim: Automation, analysis, and plotting of SRIM calculations. Journal of Open Source Software, 3(28):829:1–829:3, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00829.

Pierce:1997:AEU

[PAB^{+}97] J. S. Pierce, S. Audia, T. Burnette, K. Christiansen, D. Cosgrove, M. Conway, K. Hinckley, K. Monkaitis, J. Patten, J. Shothet, D. Staack, B. Stearns, Sturgill, C., G. Williams, and R. Pausch. Alice: easy to use interactive 3D graphics. In ACM [ACM97], pages 77–78. ISBN 0-89791-881-9. LCCN ????

Palach:2014:PPP

[Pal14] Jan Palach. Parallel programming with Python develop efficient parallel systems using the robust Python environment. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-839-6, 1-78328-840-X (e-book). LCCN QA76.642. URL http://

proquest.tech.safaribooksonline.
de/9781783288397.

Parkin:2011:DEL

[Par11] Tom Parkin. Debugging embedded Linux platforms with DGB and Python. Linux journal, 2011(206):2:1–2:??, June 2011. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Park:2022:PPC

[Par22] Jin Seok Park. pyBaram: Parallel compressible flow solver in high-performance Python for teaching and research. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S235271102200190X.

Pasca:2022:PPP

[PARS22] Dag Pasquale Pasca, Angelo Aloisio, Marco Martino Rosso, and Stefanos Sotiropoulos. PyOMA and PyOMA GUI: a Python module and software for Operational Modal Analysis. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022001340.

Pataky:2016:RSO

[Pat16] Todd C. Pataky. rft1d: Smooth one-dimensional random field upcrossing probabilities in Python. Journal of Statistical Software, 71 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

                                        7660. URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v071i07;
                                         https://www.jstatsoft.
                                        org/index.php/jss/article/
                                        view/v071i07/v71i07.pdf.

Payne:2011:BPU

[Pay11]  James Payne.  Beginning

Python: using Python 2.6 and Python 3.1. Wrox programmer to programmer. Wiley, New York, NY, USA, 2011. ISBN 0-470-41463-4. xxxiv + 588 pp. LCCN ????

Psarras:2022:LAM

[PBB22] Christos Psarras, Henrik Barthels, and Paolo Bientinesi. The linear algebra mapping problem. Current state of linear algebra languages and libraries. ACM Transactions on Mathematical Software, 48(3):26:1– 26:??, September 2022. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3549935.

Pflanzner:2022:LAB

[PBK22] Tamas Pflanzner, Hamza Baniata, and Attila Kertesz. Latency analysis of blockchainbased SSI applications. Future Internet, 14(10):282, September 29, 2022. CO-DEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/10/282.

Pradal:2009:PPB

[PBN^{+}09] C. Pradal, F. Boudon, C. Nouguier, J. Chopard, and C. Godin. PlantGL: a Python-based geometric library for 3D plant modelling at different scales. Graphical models, 71(1):1–21, January 2009. CODEN GRMOFM. ISSN 1524-0703 (print), 1524-0711 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1524070308000143.

Pintor:2022:SSE

[PDS^{+}22] Maura Pintor, Luca Demetrio, Angelo Sotgiu, Marco Melis, Ambra Demontis, and Battista Biggio. secml: Secure and explainable machine learning in Python. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000656.

Pelletier:2001:Z

[Pel01]  Michel Pelletier.  Zope,

                                        2001.  URL http://db.
                                        usenix.org/publications/
                                        library/proceedings/lisa2001/

tech/. Unpublished invited talk, LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA.

Percival:2014:TDD

[Per14a] Harry Percival. Test-driven development with Python. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2014. ISBN 1-4493-6482-9 (paperback). xxviii + 449 pp. LCCN QA76.73.P98 P46 2014.

Perkins:2014:PTP

[Per14b] Jacob Perkins. Python 3 text processing with NLTK 3 cookbook: over 80 practical recipes on natural language processing techniques using Python’s NLKT 3.0. Packt Publishing, Birmingham, UK, second edition, 2014. ISBN 1-78216-785-4, 1-78216-786-2 (e-book). iii + 288 pp. LCCN QA76.73.P98 P43 2014.

Petrone:2002:DPP

[Pet02] Jason Petrone. 3-D programming with Python. Linux journal, 94:89–94, February 2002. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Petukhova:2022:TPP

[PF22] Alina Petukhova and Nuno Fachada. TextCL: a Python package for NLP preprocessing tasks. SoftwareX, 19(??): ??, July 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022000802.

Pape:2016:LIS

[PFH^{+}16] Tobias Pape, Tim Felgentreff, Robert Hirschfeld, Anton Gulenko, and Carl Friedrich Bolz. Language-independent storage strategies for tracing JIT-based virtual machines. ACM SIGPLAN Notices, 51(2):104–113, February 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Piasini:2021:EPP

[PFLG21] Eugenio Piasini, Alexandre L. S. Filipowicz, Jonathan Levine, and Joshua I. Gold. Embo: a Python package for empirical data analysis using the information bottleneck. Journal of Open Research Software, 9(1):10–??, May 31, 2021. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.322/.

Perez:2007:ISI

[PG07] Fernando Pérez and Brian E. Granger. IPython: a system for interactive scientific computing. Computing in Science and Engineering, 9(3): 21–29, May/June 2007. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Perez:2011:PES

[PGH11] Fernando Perez, Brian E. Granger, and John D. Hunter. Python: an ecosystem for scientific computing. Computing in Science and Engineering, 13(2):13– 21, March/April 2011. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Puzyrev:2019:PCF

[PGM19] Vladimir Puzyrev, Mehdi Ghommem, and Shiv Meka. pyROM: a computational framework for reduced order modeling. Journal of Computational Science, 30:157– 173, January 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1877750318307518.

Padmavathi:2023:WCE

[PH23] P. Padmavathi and J. Harikiran. Wireless capsule endoscopy infected images detection and classification using MobileNetV2-BiLSTM model. International Journal of Image and Graphics (IJIG), 23(05):??, September 2023. ISSN 0219-

4678. URL https://www.
worldscientific.com/doi/
10.1142/S0219467823500419.

Patil:2010:PBS

[PHF10] Anand Patil, David Huard, and Christopher J. Fonnesbeck. PyMC: Bayesian stochastic modelling in Python. Journal of Statistical Software, 35(4):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http:/

/www.jstatsoft.org/v35/
i04.

Pool:2012:SNU

[PHH^{+}12] René Pool, Jaap Heringa, Martin Hoefling, Roland Schulz, Jeremy C. Smith, and K. Anton Feenstra. Software news and updates: Enabling grand-canonical Monte Carlo: Extending the flexibility of GROMACS through the GromPy Python interface module. Journal of Computational Chemistry, 33(12): 1207–1214, May 5, 2012. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Phillips:2010:POO

[Phi10] Dusty Phillips. Python 3 object oriented programming. Packt Publishing, Birmingham, UK, 2010. ISBN 1-84951-126-8, 1-84951-127-6 (e-book). v + 388 pp. LCCN QA76.73.P98 P45 2010eb.

Phillips:2015:POO

[Phi15] Dusty Phillips. Python 3 object-oriented programming: unleash the power of Python 3 objects. Packt Publishing, Birmingham, UK, second edition, 2015. ISBN 1-78439-878-0, 1-78439-878-0. xii + 431 pp. LCCN QA76.73.P98.

Pilgrim:2004:DP

[Pil04] Mark Pilgrim. Dive into Python. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-356-1 (paperback). xviii + 413 pp. LCCN QA76.73.P98 P55 2004.

Pilgrim:2009:DP

[Pil09] Mark Pilgrim. Dive into Python 3. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2009.

ISBN 1-4302-2415-0. xlix + 360 pp. LCCN QA76.73.P98

P57 2009. URL http://www.
loc.gov/catdir/enhancements/
fy1502/2011377607-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1502/2011377607-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1502/2011377607-t.html.

Pinon:2024:HRM

[Piñ24] Adelson Piñón. Handbook of regression modeling in people analytics, with examples in R and Python. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):544–545, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

academic.oup.com/jrsssa/
article/187/2/544/7256154.

Pippi:2015:PGA

[Pip15] Massimiliano Pippi. Python for Google App Engine: master the full range of development features provided by Google App Engine to build and run scalable web applications in Python. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-819-5 (paperback), 1-78439-237-5 (e-book). 198 (est) pp. LCCN TK5105.8885.G643 .P577 2015.

Pitkin:2018:PPP

[Pit18] Matthew Pitkin. psrqpy: a Python interface for querying the ATNF pulsar catalogue. Journal of Open Source Software, 3(22):538:1–538:2, February 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00538.

Pelletier:2021:GJP

[PKDM21] Michel Pelletier, Will Kimmerer, Timothy A. Davis, and Timothy G. Mattson. The GraphBLAS in Julia and Python: the PageRank and triangle centralities. In IEEE, editor, 2021 IEEE High Performance Extreme Computing Conference (HPEC), pages 1–7. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2021.

Pluta:2020:EOP

[PL20] Adam Pluta and Ontje

                                        Lünsdorf. esy-osmfilter —

a Python library to efficiently extract OpenStreetMap data. Journal of Open Research Software, 8(1):19–??, September 01, 2020. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.317/.

Pletzer:2002:PFE

[Ple02] Alexander Pletzer. Python and finite elements. Dr. Dobb’s Journal of Software Tools, 27(3):36, 38–40, March 2002. CODEN DDJOEB.

ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2002/
2002_03/ellipt2d.txt;
http://www.ddj.com/ftp/
2002/2002_03/ellipt2d.zip.

See correction [Ano02].

Plosch:1997:DCP

[Plo97] R. Plosch. Design by contract for Python. In IEEE [IEE97b], pages 213–219. ISBN 0-8186-8271-X. LCCN QA76.758.A77 1997. IEEE catalog number 97TB100207. IEEE Computer Society order number PR08271.

Pokuri:2022:APS

[PLRG22] Balaji Sesha Sarath Pokuri, Alec Lofquist, Chad Risko, and Baskar Ganapathysubramanian. Algorithm 1025: PARyOpt: a software for parallel asynchronous remote Bayesian optimization. ACM Transactions on Mathematical Software, 48(2):24:1– 24:15, June 2022. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3529517.

Paramitha:2023:TLA

[PM23] Ranindya Paramitha and Fabio Massacci. Technical leverage analysis in the Python ecosystem. Empirical Software Engineering, 28 (6):??, November 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.

                                        springer.com/article/10.
                                        1007/s10664-023-10355-2.

Pimentel:2017:NTC

[PMBF17] João Felipe Pimentel, Leonardo Murta, Vanessa Braganholo, and Juliana Freire. noWorkflow: a tool for collecting, analyzing, and managing provenance from Python scripts. Proceedings of the VLDB Endowment, 10(12):1841–1844, August 2017. CODEN ???? ISSN 2150-8097.

Politz:2013:PFM

[PMM^{+}13] Joe Gibbs Politz, Alejandro Martinez, Matthew Milano, Sumner Warren, Daniel Patterson, Junsong Li, Anand Chitipothu, and Shriram Krishnamurthi. Python: the full monty. ACM SIGPLAN Notices, 48(10):217–232, October 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). OOPSLA ’13 conference proceedings.

Poore:2016:API

[Poo16] Geoffrey Poore. Advances in PythonTEX with an introduction to fvextra. TUGboat, 37(2):187–192, 2016. ISSN 0896-3207.

Pop:2010:ERH

[Pop10] Iustin Pop. Experience report: Haskell as a reagent: results and observations on the use of Haskell in a Python project. ACM SIGPLAN Notices, 45(9):369–374, September 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Porter:2003:MDX

[Por03] Brett Porter. Managing devices with XML-RPC. Dr. Dobb’s Journal of Software Tools, 28(4):66, 68–70, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/documents/
s=7827/ddj0304h/.

Pandey:2023:PNI

[PP23] Vivek Pandey and Sudhir K.

Pandey. PY-Nodes: an abinitio Python code for searching nodes in a material using Nelder–Mead’s simplex approach. Computer Physics Communications, 283(??): Article 108570, February 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522002892.

Pedroni:2002:JE

[PR02] Samuele Pedroni and Noel Rappin. Jython Essentials. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00247-5. xx + 277 pp. LCCN QA76.73.J38 P43 2002 Stacks. US$24.95.

                                        URL http://www.oreilly.
                                        com/catalog/jythoness.

Prechelt:2000:ECS

[Pre00] Lutz Prechelt. An empirical comparison of seven programming languages. Computer, 33(10):23–29, October 2000. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

                                        //dlib.computer.org/co/
                                        books/co2000/pdf/rx023.
                                        pdf;   http://www.computer.
                                        org/computer/co2000/rx023abs.
                                        htm.

Prechelt:2003:SLG

[Pre03] L. Prechelt. Are scripting languages any good? A validation of Perl, Python, Rexx, and Tcl against C, C++, and Java. Advances in Computers, 57:207–271, 2003. CO-DEN ???? ISSN 0065-2458.

Polimis:2017:CIR

[PRH17] Kivan Polimis, Ariel Rokem, and Bryna Hazelton. Confidence intervals for random forests in Python. Journal of Open Source Software, 2 (19):124:1–124:4, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00124.

Prasanna:2021:DOS

[PSGL21] Krishnamohan G. Prasanna, Rahul Sunil, Kapil Gupta, and Seung-Cheol Lee. DJ-Mol: an open-source modeling platform for computational chemistry and materials science with a Python interpreter. Journal of Computational Chemistry, 42 (29):2116–2129, November 5, 2021. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Pulleyn:2000:EPM

[Pul00] Ivan Pulleyn. Embedding Python in multi-threaded C/ C++ applications. Linux journal, 73:??, May 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Price-Whelan:2017:PGP

[PW17] Adrian M. Price-Whelan.

Gala: a Python package for galactic dynamics. Journal of Open Source Software, 2(18):388:1–388:2, October 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00388.

Price-Whelan:2017:PSU

[PWFM17] Adrian M. Price-Whelan and Daniel Foreman-Mackey. schwimmbad: a uniform interface to parallel processing pools in Python. Journal of Open Source Software, 2(17):357:1–357:2, September 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00357.

Qin:2019:QPP

[QZWU19] Tian Qin, Qi Zhang, Renata M. Wentzcovitch, and Koichiro Umemoto. qha: a Python package for quasiharmonic free energy calculation for multi-configuration systems. Computer Physics Communications, 237(??): 199–207, April 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465518303953.

Radenski:2006:PFL

[Rad06] Atanas Radenski. “Python first”: a lab-based digital introduction to computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(3):197–201, September 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Radenski:2008:DCS

[Rad08] Atanas Radenski. Digital CS1 study pack based on Moodle and Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3): 325, September 2008. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Richardson:2001:LEO

[RAH^{+}01] Bruce Richardson, Anonymous, Nathan Hokanson, Ken O. Burtch, Jim V., Jerel Crosland, Paul Taylor, Sheldon Dubrowin, Paul Dale Roberts, Dean Provins, Kathy Lynn, and Andre Lessa. Letters to the editor: Offended; A real bastard; common misconception; Ada boy!; wacky names; penultimate Linux box?; SuSe too loosa; LJ interactive; sold on Soldier; groff is great; what’s up with Ogg?; changes to the Python Developer’s Handook. Linux journal, 83: 6, 141–142, March 2001. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Ramalho:2015:FPC

[Ram15]  Luciano Ramalho.  Flu-

ent Python: clear, concise, and effective programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-4600-8 (paperback), 1-4919-4623-7, 1-4919-4625-3 (e-book), 1-4919-4626-1 (ebook). xxiv + 743 pp. LCCN AA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781491946237.

Ramachandran:2018:APB

[Ram18a] Prabhu Ramachandran. automan: a Python-based automation framework for numerical computing. Computing in Science and Engineering, 20(5):81–97, September/October 2018. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/
                                        cs/2018/05/mcs2018050081-
                                        abs.html.

Ramalho:2018:PVG

[Ram18b] Luciano Ramalho. Python vs. Go. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2018. ISBN 1-4920-3317-0, 1-4920-3318-9. ???? pp. LCCN QA76.73.P98.

Raschka:2018:PMP

[Ras18] Sebastian Raschka. MLxtend: Providing machine learning and data science utilities and extensions to Python’s scientific computing stack. Journal of Open Source Software, 3 (24):638:1–638:2, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

                                        /joss.theoj.org/papers/
                                        10.21105/joss.00638.

Rasmussen:2020:PLR

[Ras20] Leandro Lima Rasmussen. UnBlocks^{gen}: a Python library for 3D rock mass generation and analysis. SoftwareX, 12(??):Article 100577, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020302909.

Raymond:1998:SPO

[Ray98] Eric Raymond. Stop the presses: Open source summit. Linux journal, 50:??, 1998. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:/

/www.ssc.com/lj/issue50/

2918.html. Discusses Perl, Python, and Tcl.

Ramachandran:2021:PPB

[RBP^{+}21] Prabhu Ramachandran, Aditya Bhosale, Kunal Puri, Pawan Negi, Abhinav Muta, A. Dinesh, Dileep Menon, Rahul Govind, Suraj Sanka, Amal S. Sebastian, Ananyo Sen, Rohan Kaushik, Anshuman Kumar, Vikas Kurapati, Mrinalgouda Patil, Deep Tavker, Pankaj Pandey, Chandrashekhar Kaushik, Arkopal Dutt, and Arpit Agarwal. PySPH: a Pythonbased framework for smoothed particle hydrodynamics. ACM Transactions on Mathematical Software, 47(4):34:1– 34:38, December 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

dl.acm.org/doi/10.1145/
3460773.

Raychev:2016:PMC

[RBV16] Veselin Raychev, Pavol Bielik, and Martin Vechev. Probabilistic model for code with decision trees. ACM SIG-PLAN Notices, 51(10):731– 747, October 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roubeyrie:2018:PWP

[RC18] Lionel Roubeyrie and Sébastien Celles. Windrose: a Python

                                        Matplotlib, Numpy library to

manage wind and pollution data, draw windrose. Journal of Open Source Software, 3(29):268:1–268:5, September 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00268.

Ramon-Cortes:2020:AAP

[RCAE^{+}20] Cristian Ramon-Cortes, Ramon Amela, Jorge Ejarque, Philippe Clauss, and Rosa M.

                                        Badia. AutoParallel: Auto-

matic parallelisation and distributed execution of affine loop nests in Python. The International Journal of High Performance Computing Applications, 34(6):659– 675, November 1, 2020. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/
                                        doi/full/10.1177/1094342020937050.

Rickett:2006:RPF

[RCRS06] Christopher D. Rickett, Sung-Eun Choi, Craig E. Rasmussen, and Matthew J. Sottile. Rapid prototyping frameworks for developing scientific applications: a case study. The Journal of Supercomputing, 36(2): 123–134, May 2006. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//www.springerlink.com/
openurl.asp?genre=article&
issn=0920-8542&volume=36&
issue=2&spage=123.

Ramos-Carreno:2023:DDC

[RCT23] Carlos Ramos-Carreño and José L. Torrecilla. dcor: Distance correlation and energy statistics in Python. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000225.

Ravasio:2021:OOO

[RDB21] Claudio S. Ravasio, Lyndon Da Cruz, and Christos Bergeles. oflibnumpy

and oflibpytorch: Opti-

cal flow handling and manipulation in Python. Journal of Open Research Software, 9(1):31–??, November 26, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.380/.

Ramalhinho:2023:FSP

[RDBC23] João Ramalhinho, Thomas Dowrick, Ester Bonmati, and Matthew J. Clarkson. Fan-Slicer: a Pycuda package for fast reslicing of ultrasound shaped planes. Journal of Open Research Software, 11(1):??, ???? 2023. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.422.

Ramm:2007:RWA

[RDS07] Mark Ramm, Kevin Dangoor, and Gigi Sayfan. Rapid Web applications with TurboGears: using Python to create Ajaxpowered sites. Prentice Hall open source software development series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2007. ISBN 0-13-243388-5 (paperback). xxvii + 472 pp. LCCN TK5105.888 .R355 2007.

Reinartz:2022:PPP

[RE22] Christopher Reinartz and Thomas T. Enevoldsen. pyTEP: a Python package for interactive simulations of the Tennessee Eastman process. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000449.

Reed:2004:RAD

[Ree04]  David Reed.  Rapid ap-

plication development with Python and Glade. Linux journal, 2004(123):??, July 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Rempt:2001:PPT

[Rem01a] Boudewijn Rempt. Python’s PyQt toolkit. Dr. Dobb’s Journal of Software Tools, 26 (1):88, 90, 92, 94, January 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_01/pyqt.txt.

Rempt:2001:SJP

[Rem01b] Boudewijn Rempt. Scripting with Java and Python: Building a Python console window in a Java application. Dr. Dobb’s Journal of Software Tools, 26(10): 56, 60–61, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/ftp/2001/
2001_10/pyconsol.txt;
http://www.ddj.com/ftp/
2001/2001_10/pyconsol.zip.

Ren:2023:PSA

[Ren23] Xiquan Ren. Python shared atomic data types. Special focus section: Benchmarking, Experimentation Tools, and Reproducible Practices for Data-Intensive Systems from Edge to Cloud: Software: Practice and Experience, 53 (12):2393–2407, December 2023. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ren:2016:JTS

[RF16] Brianna M. Ren and Jeffrey S. Foster. Just-in-time static type checking for dynamic languages. ACM SIG-PLAN Notices, 51(6):462– 476, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossant:2013:PPL

[RFG13] Cyrille Rossant, Bertrand Fontaine, and Dan F. M. Goodman. Playdoh: a lightweight Python library for distributed computing and optimisation. Journal of Computational Science, 4(5):352–359, September 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1877750311000561.

Ramu:2000:CCP

[RG00] Chenna Ramu and Christina Gemuend. cgimodel: CGI programming made easy with Python. Linux journal, 75:??, July 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Rhodes:2010:FPN

[RG10] Brandon Rhodes and John Goerzen. Foundations of Python network programming: the comprehensive guide to building network applications with Python. The expert’s voice in open source. Apress, Berkeley, CA, USA, second edition, 2010. ISBN 1-4302-3003-7 (paperback), 1-4302-3004-5 (e-book). xx + 345 pp. LCCN QA76.73.P98 G64 2010eb. URL http://

proquest.safaribooksonline.
com/9781430230038.

Rhodes:2014:FPN

[RG14] Brandon Rhodes and John Goerzen. Foundations of Python network programming. Apress, Berkeley, CA, USA, third edition, 2014. ISBN 1-4302-5854-3 (paperback), 1-4302-5855-1 (e-book). xxi + 388 pp. LCCN QA76.73.P98 R48

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1603/2015458068-b.html;
http://www.loc.gov/catdir/
enhancements/fy1603/2015458068-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1603/2015458068-t.html.

Ramasubramani:2018:PRP

[RG18] Vyas Ramasubramani and Sharon C. Glotzer. rowan: a Python package for working with quaternions. Journal of Open Source Software, 3(27):787:1–787:3, July 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00787.

Reizenstein:2020:AIL

[RG20] Jeremy F. Reizenstein and Benjamin Graham. Algorithm 1004: The Iisignature library: Efficient calculation of iterated-integral signatures and log signatures. ACM Transactions on Mathematical Software, 46(1):8:1–8:21, March 2020. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3371237.

Robertson:2022:RPP

[RGP^{+}22] Hayden Robertson, Isaac J. Gresham, Stuart W. Prescott, Grant B. Webber, Erica J. Wanless, and Andrew Nelson. refellips: a Python package for the analysis of variable angle spectroscopic ellipsometry data. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001431.

Richardson:2021:TLP

[RGS^{+}21] A. S. Richardson, D. F. Gordon, S. B. Swanekamp, I. M. Rittersdorf, P. E. Adamson, O. S. Grannis, G. T. Morgan, A. Ostenfeld, K. L. Phlips, C. G. Sun, G. Tang, and D. J. Watkins. TurboPy: a lightweight Python framework for computational physics. Computer Physics Communications, 258(??): Article 107607, January 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465520302897.

RodasdePaz:2015:PGP

[RH15] Alejandro Rodas de Paz and Joseph Howse. Python game programming by example: a pragmatic guide for developing your own games with Python. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-153-4, 1-78528-391-X (e-book). 230 (est.) pp. LCCN QA76.76.C672. URL http:/

/proquest.safaribooksonline.
com/?fpi=9781785281532.

Ramirez:2022:PPP

[RHLTG^{+}22] Erick Ramírez, Sergio Hernández-López, Enelio Torres-Garcia, Karla Reyes-Morales, and Jorge Balmaseda. pICNIK: a Python package with isoconversional computations for non-isothermal kinetics. Computer Physics Communications, 278(??):Article 108416, September 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522001357.

Rathgeber:2017:FAF

[RHM^{+}17] Florian Rathgeber, David A.

Ham, Lawrence Mitchell, Michael Lange, Fabio Luporini, Andrew T. T. Mcrae, Gheorghe-Teodor Bercea, Graham R. Markall, and Paul H. J. Kelly. Firedrake: Automating the finite element method by composing abstractions. ACM Transactions on Mathematical Software, 43(3):24:1– 24:??, January 2017. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL https://dl.
acm.org/citation.cfm?id=
2998441.

Richardson:2013:GSB

[Ric13]  Matt Richardson.  Get-

ting started with BeagleBone. Maker Media, Sebastopol, CA, 2013. ISBN 1-4493-4537-9 (paperback), 1-4493-4536-0, 1-4493-4535-2 (e-book), 1-4493-4533-6 (e-book). xiii + 126 pp. LCCN TK7895.E42 R43 2013.

Richardson:2014:BEL

[Ric14] Matt Richardson. Beagle-Bone für Einsteiger: [Linuxbasierte Elektronik-Projekte mit Python und JavaScript]. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2014. ISBN 3-95561-409-3. xii + 134 pp. LCCN ????

Richardson:2015:AP

[Ric15] Craig Richardson. Adventures in Python. Wiley, New York, NY, USA, 2015. ISBN 1-118-95185-9, 1-118-95179-4. 282 pp. LCCN QA76.73.P98.

Riehl:2009:LEO

[Rie09] Jonathan Riehl. Language embedding and optimization in Mython. ACM SIG-PLAN Notices, 44(12):39–48, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ravitch:2009:AGL

[RJAL09] Tristan Ravitch, Steve Jackson, Eric Aderhold, and Ben Liblit. Automatic generation of library bindings using static analysis. ACM SIG-PLAN Notices, 44(6):352– 362, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rose:2019:PSB

[RK19] Michael E. Rose and John R.

Kitchin.  pybliometrics:

Scriptable bibliometrics using a Python interface to Scopus. SoftwareX, 10(??):Article 100263, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019300573.

Ryan-Keogh:2021:PPU

[RKR21] Thomas J. Ryan-Keogh and Charlotte M. Robinson. Phytoplankton photophysiology utilities: a Python toolbox for the standardization of processing active chlorophyll-a fluorescence data. Frontiers in Marine Science, 8, July 2021. ISSN 2296-7745.

Ragan-Kelley:2014:OPP

[RKVL14] Benjamin Ragan-Kelley, John P. Verboncoeur, and Ming-Chieh Lin. Optimizing physical parameters in 1-D particle-in-cell simulations with Python. Computer Physics Communications, 185(10):2487–2494, October 2014. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465514001994.

Rao:2019:NLP

[RM19] Delip Rao and Brian McMahan. Natural Language Processing with PyTorch: Build Intelligent Language Applications Using Deep Learning. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2019. ISBN 1-4919-7818-X, 1-4919-7820-1 (e-book), 1-4919-7822-8, 1-4919-7823-6 (paperback). xiii + 238 pp. LCCN QA76.9.N38 R36 2019.

Raj:2023:FPT

[RMM23] Surbhi Raj, Jimson Mathew, and Arijit Mondal. FDT: a Python toolkit for fake image and video detection. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000912.

Ranum:2006:SAT

[RMZG06] David Ranum, Bradley Miller, John Zelle, and Mark Guzdial. Successful approaches to teaching introductory computer science courses with Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):396–397, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Ranathunga:2017:MPP

[RNR17] D. Ranathunga, H. Nguyen, and M. Roughan. MGtoolkit: a Python package for implementing metagraphs. SoftwareX, 6(??):85–90, ???? 2017. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300080.

Raschka:2015:PML

[RO15a] Sebastian Raschka and Randal S. Olson. Python machine learning: unlock deeper insights into machine learning with this vital guide to cutting-edge predictive analytics. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78355-514-9, 1-78355-513-0. ???? pp. LCCN QA76.73.P98. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781783555130.

Redondo:2015:CEC

[RO15b] Jose Manuel Redondo and Francisco Ortin. A comprehensive evaluation of common Python implementations. IEEE Software, 32(4):76–84, July/August 2015. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/
                                        so/2015/04/mso2015040076-
                                        abs.html.

Roald:2023:MLC

[Roa23] Marie Roald. MatCoupLy: Learning coupled matrix factorizations with Python. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022002102.

Rohit:2016:JC

[Roh16]  Jalem Raj Rohit.  Julia

Cookbook. Packt Publishing, Birmingham, UK, 2016. ISBN 1-78588-201-5, 1-78588-363-1 (e-book). v + 157 pp. LCCN QA76.73.J8; T55.4-60.8. URL http:/

                                        /sbiproxy.uqac.ca/login?
                                        url=https://international.
                                        scholarvox.com/book/88843406.

Romero:2014:MPR

[Rom14] Victor Romero. Mastering Python regular expressions. Shroff Publishers, ????, 2014.

ISBN 93-5110-550-4. LCCN ????

Rook:1997:LCC

[Roo97] David Rook. A language collector comments on: Java, Perl & Python, October 1997.

URL http://www.chips.
navy.mil/chips/archives/
97_oct/file12.htm.

Rooker:1998:BRP

[Roo98] Terry Rooker. Book review: Programming Python. ;login: the USENIX Association newsletter, 23(2): ??, April 1998. CODEN LOGNEM. ISSN 1044-6397.

URL http://www.usenix.
org/publications/login/
1998-4/python.html.

Rosenberg:2008:DCT

[Ros08] Scott Rosenberg. Dreaming in code: Two dozen Programmers, three years, 4,732 bugs, and one quest for transcendent software. Three Rivers Press, New York, NY, USA, 2008. ISBN ???? 403 pp. LCCN ????

Rossant:2013:LII

[Ros13] Cyrille Rossant. Learning IPython for interactive computing and data visualization: Learn IPython for interactive Python programming, highperformance numerical computing, and data visualization. Open source: community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78216-993-8 (paperback), 1-78216-994-6 (e-book), 1-299-54508-4 (e-book). iv + 123 pp. LCCN QA76.73.P98 .R677 2013.

Rossant:2014:IIC

[Ros14] Cyrille Rossant. IPython interactive computing and visualization cookbook: over 100 hands-on recipes to sharpen your skills in highperformance numerical computing and data science with Python. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-481-1, 1-78328-482-X (e-book), 1-322-16622-6 (e-book). v + 494 pp. LCCN QA76.73.P98 R677 2014.

Rose:2018:PCP

[Ros18] Brian E. J. Rose. CLIMLAB: a Python toolkit for interactive, process-oriented climate modeling. Journal of Open Source Software, 3 (24):659:1–659:2, April 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00659.

Rodriguez:2023:SPN

[RPL^{+}23] Jose S. Rodriguez, Robert B. Parker, Carl D. Laird, Bethany L. Nicholson, John D. Siirola, and Michael L. Bynum. Scalable parallel nonlinear optimization with PyNumero and Parapint. INFORMS Journal on Computing, 35(2):509– 517, March/April 2023. CO-DEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https://

pubsonline.informs.org/
doi/full/10.1287/ijoc.2023.
1272.

Ruiz-Rosero:2019:SSS

[RRRGVD19] Juan Ruiz-Rosero, Gustavo Ramirez-Gonzalez, and Jesus Viveros-Delgado. Software survey: ScientoPy, a scientometric tool for topics trend analysis in scientific publications. Scientometrics, 121(2):1165–1188, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s11192-019-03213-w.

Raamana:2017:PPC

[RS17] Pradeep Reddy Raamana and Stephen C. Strother. Python class defining a machine learning dataset ensuring key-based correspondence and maintaining integrity. Journal of Open Source Software, 2(17):382:1, September 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00382.

Robertson:2021:AEP

[RSPJ21] E. J. Robertson, N. Sibalić, R. M. Potvliege, and M. P. A. Jones. ARC 3.0: an expanded Python toolbox for atomic physics calculations. Computer Physics Communications, 261(??): Article 107814, April 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520304136.

Rogers:2019:SLB

[RSRT19] Samuel Rogers, Joshua Slycord, Ronak Raheja, and Hamed Tabkhi. Scalable LLVM-based accelerator modeling in gem5. IEEE Computer Architecture Letters, 18(1):18–21, January/ June 2019. CODEN ???? ISSN 1556-6056 (print), 1556-6064 (electronic).

Russell:2014:MSW

[Rus14] Matthew A. Russell. Mining the social web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2014. ISBN 1-4493-6761-5 (paperback), 1-4493-7045-4 (e-book), 1-4493-6821-2 (e-book), 1-4493-6822-0 (ebook). xxiv + 421 pp. LCCN QA76.9.D343 R87 2013.

Ravasi:2020:PLO

[RV20] Matteo Ravasi and Ivan Vasconcelos. PyLops — a linearoperator Python library for scalable algebra and optimization. SoftwareX, 11 (??):Article 100361, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711019301086.

Reed:2009:DSA

[RZ09] David M. Reed and John M.

Zelle. Data structures and algorithms using Python and C++. Franklin, Beedle and Associates, Inc., Wilsonville, OR, USA, 2009. ISBN 1-59028-233-7. ???? pp. LCCN QA76.73.P98 R44 2009.

Saadah:2023:BRA

[Sa’23] Aminatus Sa’adah. Book review: Artificial Intelligence with Python, Teik Toe Teoh and Zheng Rong, Springer, 2022, 336 pp., EUR 16.99, ISBN 978-981-16-8615-3. Technometrics, 65(3):451–452, 2023. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Silva:2018:HFP

[SAA18] D. J. Silva, J. S. Amaral, and V. S. Amaral. Heatrapy: a flexible Python framework for computing dynamic heat transfer processes involving caloric effects in 1.5D systems. SoftwareX, 7 (??):373–382, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301298.

Salkovic:2020:POB

[SAB^{+}20] Edin Salkovic, Mostafa M. Abbas, Samir Brahim Belhaouari, Khaoula Errafii, and Halima Bensmail. OutPyR: Bayesian inference for RNA-Seq outlier detection. Journal of Computational Science, 47:??, November 2020. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1877750320305433.

Sleem:2020:PPB

[SABEh20] Ahmed Sleem, Mohamed Abdel-Baset, and Ibrahim Elhenawy. PyIVNS: a Python based tool for intervalvalued neutrosophic operations and normalization. SoftwareX, 12(??):Article 100632, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711020303459.

Schoofs:2022:ATA

[SAD22] Ebert Schoofs, Mehrdad Abdi, and Serge Demeyer. AmPyfier: Test amplification in Python. Journal of Software: Evolution and Process, 34(11):e2490:1– e2490:??, November 2022. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Saha:2012:PPC

[Sah12] Amit Saha. Parallel programming in C and Python. Linux journal, 2012(217):4:1–4:??, May 2012. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Saha:2015:DMP

[Sah15] Amit Saha. Doing math with Python: use programming to explore algebra, statistics, calculus, and more! No Starch Press, San Francisco, CA, USA, 2015. ISBN 1-59327-640-0 (paperback). xvii + 244 pp. LCCN QA20.C65 S24 2015.

Sale:2014:TPA

[Sal14] David Sale. Testing Python:

applying unit testing, TDD, BDD, and acceptance testing. Wiley, New York, NY, USA, 2014. ISBN 1-118-90122-3 (paperback), 1-118-90125-8 (e-book), 1-118-90124-X (ePDF). xv + 222 pp. LCCN QA76.73.P98 .S254 2014eb.

Salceanu:2018:JPP

[Sal18] Adrian Salceanu. Julia programming projects: learn Julia 1.x by building apps for data analysis, visualization, machine learning, and the Web. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78829-725-3. ix + 482 pp. LCCN QA76.73.J85. URL http://

proquest.safaribooksonline.
com/?fpi=9781788292740.

Sander:2013:ESP

[San13] Leonard M. (Leonard Michael) Sander. Equilibrium statistical physics: with computer simulations in Python. CreateSpace Independent Publishing, North Charleston, SC, USA, 2013. ISBN 1-4910-6651-2. xii + 321 pp. LCCN QC174.8 .S36 2013.

Sanderson:2015:PGA

[San15] Dan Sanderson. Programming Google App Engine with Python. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0025-3, 1-4919-0024-5, 1-4919-0368-6 (e-book), 1-4919-0367-8 (ebook). xxi + 439 pp. LCCN TK5105.88813.

Sarker:2014:PNP

[Sar14] M. O. Faruque Sarker. Python network programming cookbook over 70 detailed recipes to develop pratical solutions for a wide range of real-world network programming tasks. Quick answers to common problems. Packt Publishing, Birmingham, UK, 2014. ISBN 1-84951-346-

5. LCCN ???? URL http://

proquest.tech.safaribooksonline.
de/9781849513463.

Sauter:2023:BRF

[Sau23] Roger Sauter. Book review:

Foundations of Statistics for Data Scientists: With R and Python by Alan Agresti and Maria Kateri, Boca Raton, FL: Chapman and Hall/CRC, 2022, xv + 453 pp., $136.46 (hcb). Technometrics, 65 (1):132–133, 2023. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Savage:2023:NRP

[Sav23] Neil Savage. News: Revamping Python for an AI world. Communications of the ACM, 66(12):13–14, December 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https://

dl.acm.org/doi/10.1145/
3624987.

Stevens:2015:PPB

[SB15] Tim Stevens and Wayne Boucher. Python programming for biology, bioinformatics, and beyond. Cambridge University Press, Cambridge, UK, 2015. ISBN 0-521-89583-9 (hardcover), 0-521-72009-5 (paperback). viii + 702 + 8 pp. LCCN QH324.2 .S727 2014.

Smigaj:2015:SBI

[ŚBA^{+}15] Wojciech Śmigaj, Timo Betcke, Simon Arridge, Joel Phillips, and Martin Schweiger. Solving boundary integral problems with BEM++. ACM Transactions on Mathematical Software, 41(2):6:1– 6:40, January 2015. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Serrano:2017:MIP

[SBC^{+}17] E. Serrano, J. G. Blas, J. Carretero, M. Abella, and M. Desco. Medical imaging processing on a big data platform using Python: Experiences with heterogeneous and homogeneous architectures. In 2017 17th IEEE/ACM International Symposium on Cluster, Cloud and Grid Computing (CCGRID), pages 830– 837. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017.

Strickland:2014:PPM

[SBMD14] Christopher Strickland, Robert Burdett, Kerrie Mengersen, and Robert Denham. PySSM: a Python module for Bayesian inference of linear Gaussian state space models. Journal of Statistical Software, 57(6):??, April 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v57/i06.

Sandha:2019:DDM

[SCAK^{+}19] Sandeep Singh Sandha, Wellington Cabrera, Mohammed Al-Kateb, Sanjay Nair, and Mani Srivastava. In-database distributed machine learning: demonstration using Teradata SQL engine. Proceedings of the VLDB Endowment, 12 (12):1854–1857, August 2019. CODEN ???? ISSN 2150-8097.

Schillaci:2017:PP

[Sch17] Michael Jay Schillaci. Perfectly Python. Computing in Science and Engineering, 19(6):51–53, November/December 2017. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL https://www.
computer.org/csdl/mags/
cs/2017/06/mcs2017060051.
html.

Schick:2021:UPT

[Sch21] Daniel Schick. udkm1Dsim — a Python toolbox for simulating 1D ultrafast dynamics in condensed matter. Computer Physics Communications, 266(??):Article 108031, September 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521001430.

Scott:2009:PLP

[Sco09] Michael L. Scott. Programming Language Pragmatics. Morgan Kaufmann Publishers, San Francisco, CA, USA, third edition, 2009. ISBN 0-12-374514-4. xxx + 910 pp. LCCN ???? Many sections of the book are relegated to the accompanying CD-ROM.

Scott:2017:PSI

[Sco17] Camille Scott. shmlast: an improved implementation of conditional reciprocal best hits with LAST and Python. Journal of Open Source Software, 2(9):142:1–142:4, January 2017. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00142.

Semper:2020:ERF

[SDP^{+}20] S. Semper, M. Döbereiner, S. Pawar, M. Landmann, and G. Del Galdo. eadf: Representation of far-field antenna responses in Python. SoftwareX, 12(??):Article 100583, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S235271102030296X.

Scherer:2000:SPV

[SDS00] David Scherer, Paul Dubois, and Bruce Sherwood. Scientific programming: VPython: 3D interactive scientific graphics for students. Computing in Science and Engineering, 2(5):56–62, September/October 2000. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http:

                                        //dlib.computer.org/cs/
                                        books/cs2000/pdf/c5056.
                                        pdf.

Siqueira:2016:PPP

[SdSS16] Abel Soares Siqueira, Raniere Costa da Silva, and Luiz-Rafael

Santos.  Perprof-py: a

Python package for performance profile of mathematical optimization software. Journal of Open Research Software, 4(1):e12–??, April 22, 2016. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.81/.

Segaran:2007:PCI

[Seg07] Toby Segaran. Programming collective intelligence: building Smart Web 2.0 applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-52932-5, 0-596-55068-5. xxi + 334 pp. LCCN T58.5 .S43 2007. URL http://

www.oreilly.com/catalog/
9780596529321.

Seitz:2009:GPP

[Sei09]  Justin Seitz.  Gray Hat

Python: Python programming for hackers and reverse engineers. No Starch Press, San Francisco, CA, USA, 2009. ISBN 1-59327-192-1, 1-59327-224-3 (e-book). ???? pp. LCCN QA76.9.A25 S457 2009. URL http://

                                        proquest.safaribooksonline.
                                        com/9781593271923.

Seitz:2014:BPP

[Sei14]  Justin Seitz.  Black hat

Python: Python programming for hackers and pentesters. No Starch Press, San Francisco, CA, USA, 2014. ISBN 1-59327-590-0. xviii + 170 pp. LCCN QA76.73.P98 S45 2015.

Severance:2013:PIE

[Sev13] Charles Severance. Python for informatics: exploring information. CreateSpace Independent Publishing, North Charleston, SC, USA, 2013. ISBN 1-4923-3924-5. xii + 224 + 12 pp. LCCN QA76.73.P98 S48 2013b.

Severance:2015:GVRa

[Sev15a] Charles Severance. Guido van Rossum: The early years of Python. Computer, 48(2): 7–9, February 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        co/2015/02/mco2015020007.
                                        html.

Severance:2015:GVRb

[Sev15b] Charles Severance. Guido van Rossum: The modern era of Python. Computer, 48 (3):8–10, March 2015. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/

co/2015/03/mco2015030008.
html.

Souza:2023:EPB

[SFC23] Paulo S. Souza, Tiago Ferreto, and Rodrigo N. Calheiros. EdgeSimPy: Pythonbased modeling and simulation of edge computing resource management policies. Future Generation Computer Systems, 148(??): 446–459, November 2023. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167739X23002340.

Sanders:2007:SMM

[SG07] Ian Douglas Sanders and Vashti C. Galpin. Students’ mental models of recursion at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):317, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Smith:2018:POP

[SG18] Daniel G. A. Smith and Johnnie Gray. opt einsum

— a Python package for optimizing contraction order for einsum-like expressions. Journal of Open Source Software, 3(26):753:1–753:3, June 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00753.

Shakir:2017:PGG

[SGPHD^{+}17] Dzhoshkun Ismail Shakir, Luis Carlos García-Peraza-Herrera, Pankaj Daga, Tom Doel, Matthew J. Clarkson, Sébastien Ourselin, and Tom Vercauteren. GIFT-Grab: Real-time C++ and Python multi-channel video capture, processing and encoding API. Journal of Open Research Software, 5(1):27–??, October 09, 2017. CODEN ???? ISSN 2049-9647. URL https:

                                        //openresearchsoftware.
                                        metajnl.com/articles/10.
                                        5334/jors.169/.

Sun:2023:UUG

[SGZ23] Yiming Sun, Daniel German, and Stefano Zacchiroli. Using the uniqueness of global identifiers to determine the provenance of Python software source code. Empirical Software Engineering, 28 (5):??, September 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s10664-023-10317-8.

Suess:2017:PMM

[SH17] Daniel Suess and Milan Holzäpfel. mpnum: a matrix product representation library for Python. Journal of Open Source Software, 2(20):465:1–465:2, December 2017. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00465.

Steppa:2019:HPH

[SH19a] Constantin Steppa and Tim L.

Holch. HexagDLy — processing hexagonally sampled data with CNNs in Py-Torch. SoftwareX, 9(??): 193–198, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018302723.

Svatunek:2019:APT

[SH19b] Dennis Svatunek and Kendall N.

Houk. autoDIAS: a Python tool for an automated distortion/interaction activation strain analysis. Journal of Computational Chemistry, 40 (28):2509–2515, October 30, 2019. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shannon:2003:ABF

[Sha03] Christine Shannon. Another breadth-first approach to CS I using Python. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(1):248– 251, January 2003. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Shalabh:2024:AUB

[Sha24] Shalabh. Applied univariate, bivariate, and multivariate statistics using Python: a Beginner’s guide to advanced data analysis. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187(2):548, April 2024. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://

                                        academic.oup.com/jrsssa/
                                        article/187/2/548/7282112.

Shell:1997:PDS

[She97] Jeffrey P. Shell. Python does scripts and objects — Python is a platform-independent OOP language with capabilities ranging from simple scripting to sophisticated object libraries. Byte Magazine, 22(2):63–64, February 1997. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Shein:2015:NPB

[She15a] Esther Shein. News: Python for beginners. Communications of the ACM, 58(3): 19–21, March 2015. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://cacm.
                                        acm.org/magazines/2015/
                                        3/183588/fulltext.

Sherrington:2015:MJD

[She15b] Malcolm Sherrington. Mastering Julia: develop your analytical and programming skills further in Julia to solve complex data processing problems. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78355-331-6 (paperback), 1-78355-332-4 (e-book). xiv + 385 pp. LCCN QA76.7 .S547 2015; QA76.73.J8 S54 2015.

Sega:2018:PPP

[SHFJ18] Marcello Sega, György Hantal, Balázs Fábián, and Pál Jedlovszky. Pytim: a Python package for the interfacial analysis of molecular simulations. Journal of Computational Chemistry, 39(25): 2118–2125, September 30, 2018. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Shi:2007:PIG

[Shi07] Xuan Shi. Python for Internet GIS applications. Computing in Science and Engineering, 9(3):56–59, May/ June 2007. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Sipper:2023:EKE

[SHTE23] Moshe Sipper, Tomer Halperin, Itai Tzruia, and Achiya Elyasaf. EC-KitY: Evolutionary computation tool kit in Python with seamless machine learning integration. SoftwareX, 22(??): ??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023000778.

Siek:2017:CPT

[Sie17]  Jeremy Siek.  Challenges

and progress toward efficient gradual typing (invited talk). ACM SIGPLAN Notices, 52(11):2, November 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sileika:2014:PPS

[Sil14] Rytis Sileika. Pro Python system administration. The expert’s voice in Python. Apress, Berkeley, CA, USA, second edition, 2014. ISBN 1-4842-0218-X (paperback), 1-4842-0217-1 (e-book). xxvi + 399 pp. LCCN QA76.73.P98 S55 2014.

Sitaker:2002:PPW

[Sit02]  Kragen Sitaker.  Python

or Perl: Which is better? ;login: the USENIX Association newsletter, 27(3): ??, June 2002. CODEN LOGNEM. ISSN 1044-6397.

                                        URL http://www.usenix.
                                        org/publications/login/
                                        2002-06/pdfs/sitaker.pdf.

Sharma:2021:DDS

[SJK^{+}21] Moolchand Sharma, Bhanu Jain, Chetan Kargeti, Vinayak Gupta, and Deepak Gupta. Detection and diagnosis of skin diseases using residual neural networks (RESNET).

International Journal of Image and Graphics (IJIG), 21 (05):??, December 2021. ISSN 0219-4678. URL https://

www.worldscientific.com/
doi/10.1142/S0219467821400027.

Stoliaroff:2018:PEU

[SJL18] Adrien Stoliaroff, Stéphane Jobic, and Camille Latouche. PyDEF 2.0: an easy to use post-treatment software for publishable charts featuring a graphical user interface. Journal of Computational Chemistry, 39(26): 2251–2261, October 5, 2018. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Spiegelberg:2019:TRE

[SK19] Leonhard F. Spiegelberg and Tim Kraska. Tuplex: robust, efficient analytics when Python rules. Proceedings of the VLDB Endowment, 12 (12):1958–1961, August 2019. CODEN ???? ISSN 2150-8097.

Stubblebine:2008:RAK

[SKS08] Toni Stubblebine, Peter Klicman, and Lars Schulten. Reguläre Ausdrücke — kurz and gut [für Perl, Ruby, PHP, C#, Python, Java and .NET]. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel:

+1 617 354 5800, second edition, 2008. ISBN 3-89721-535-7. 131 pp. LCCN ???? EUR 9.90.

Shekhovtsov:2023:VPU

[SKS23] Andrii Shekhovtsov, Bartłomiej Kizielewicz, and Wojciech Sałabun. Version [1.1] — [pymcdm — the universal library for solving multicriteria decision-making problems]. SoftwareX, 24(??): ??, December 2023. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711023002157.

Sanders:2008:SPP

[SL08] Ian D. Sanders and Sasha Langford. Students’ perceptions of Python as a first programming language at Wits. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):365, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Stansbury:2020:PAF

[SL20] Conrad Stansbury and Alessandra Lanzara. PyARPES: an analysis framework for multimodal angle-resolved photoemission spectroscopies. SoftwareX, 11(??):Article 100472, January/June 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019301633.

Schwarz:2021:ICC

[SL21] Diana Estévez Schwarz and René Lamour. InitDAE: Computation of consistent values, index determination and diagnosis of singularities of DAEs using automatic differentiation in Python. Journal of Computational and Applied Mathematics, 387(??): Article 112486, ???? 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0377042719304893.

Slatkin:2015:EPS

[Sla15]  Brett Slatkin.  Effective

Python: 59 specific ways to write better Python. Effective software development series. Addison-Wesley, Reading, MA, USA, 2015. ISBN 0-13-403441-4. ???? pp. LCCN QA76.73.P98.

Szustak:2023:POP

[SLA^{+}23] Lukasz Szustak, Marcin Lawenda, Sebastian Arming, Gregor Bankhamer, Christoph Schweimer, and Robert Elsässer. Profiling and optimization of Pythonbased social sciences applications on HPC systems by means of task and data parallelism. Future Generation Computer Systems, 148(??): 623–635, November 2023.

CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167739X23002571.

Shah:2023:QPB

[SLB^{+}23] S. A. Shah, Hao Li, Eric R. Bittner, Carlos Silva, and Andrei Piryatinski. QuDPy: a Python-based tool for computing ultrafast non-linear optical responses. Computer Physics Communications, 292(??):Article 108891, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465523002369.

Singh:2021:MPL

[SLDF^{+}21] Sobhit Singh, Logan Lang, Viviana Dovale-Farelo, Uthpala Herath, Pedram Tavadze, François-Xavier Coudert, and Aldo H. Romero. MechElastic: a Python library for analysis of mechanical and elastic properties of bulk and 2D materials. Computer Physics Communications, 267(??): Article 108068, October 2021. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521001806.

Sinz:2014:PNP

[SLGB14] Fabian H. Sinz, Jörn-Philipp Lies, Sebastian Gerwinn, and Matthias Bethge. Natter: a Python natural image statistics toolbox. Journal of Statistical Software, 61(5): ??, November 2014. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v61/i05.

Safa:2022:TPP

[SLP^{+}22] Ibrahim Safa, Jeffrey Lazar, Alex Pizzuto, Oswaldo Vasquez, Carlos A. Argüelles, and Justin Vandenbroucke. TauRunner: a public Python program to propagate neutral and charged leptons. Computer Physics Communications, 278(??):Article 108422, September 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522001412.

Stubblebine:2004:SHD

[SM04] Tony Stubblebine and Junko Mishima. Seiki hyogen desukutoppu rifarensu: regular expressions for Perl, C, PHP, Python, Java, and .NET. Orairi Japan, Tokyo, Japan, 2004. ISBN 4-87311-170-6. vi + 96 pp. LCCN ????

Small:2012:SPB

[Sma12] Alex Small. Scientific Python for both expert and novice programmers. Computing in Science and Engineering, 14 (2):6–7, March/April 2012.

CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Shen:2023:VAO

[SMD23] Yang Shen, Robert H. Moore, and Ankit Deo. Visualizing Abaqus output database in ParaView: a universal converter in Python and C++. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711023000274.

Schoedl:2023:PMA

[SMF^{+}23] Nathan W. Schoedl, Emma J. MacKie, Michael J. Field, Eric A. Stubbs, Allan Zhang, Matthew Hibbs, and Mathieu Gravey. A Python multiprocessing approach for fast geostatistical simulations of subglacial topography. Computing in Science and Engineering, 25(3):42–49, March 2023. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Smith:2015:C

[Smi15] Kurt W. Smith. Cython. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2015. ISBN 1-4919-0155-1. 240 pp. LCCN ???? URL http://

proquest.safaribooksonline.
com/9781491901731.

Smith:2017:PUP

[Smi17] Andrew P. Smith. UKCensusAPI:

Python and R interfaces to the nomisweb UK census data API. Journal of Open Source Software, 2 (19):408:1, November 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00408.

Smith:2018:PHC

[Smi18] Andrew P. Smith. humanleague:

a C++ microsynthesis package with R and Python interfaces. Journal of Open Source Software, 3(25):629:1, May 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00629.

Schroeder:2006:VTO

[SML06] Will Schroeder, Ken Martin, and Bill Lorensen. The visualization toolkit: an objectoriented approach to 3D graphics [visualize data in 3D — medical, engineering or scientific; build your own applications with C++, Tcl, Java or Python; includes source code for VTK (supports UNIX, Windows and Mac)]. Kitware, Clifton Park, NY, fourth edition, 2006. ISBN 1-930934-19-X. xvi + 512 pp. LCCN ????

Smith:2022:ICC

[SMM^{+}22] Matthew J. Smith, Mohammad A. Mansournia, Camille Maringe, Paul N. Zivich, Stephen R. Cole, Clémence Leyrat, Aurélien Belot, Bernard Rachet, and Miguel A. Luque-Fernandez. Introduction to computational causal inference using reproducible Stata, R, and Python code: a tutorial. Statistics in Medicine, 41(2): 407–432, January 30, 2022. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sulzer:2021:PBM

[SMT^{+}21] Valentin Sulzer, Scott G. Marquis, Robert Timms, Martin Robinson, and S. Jon Chapman. Python battery mathematical modelling (Py-BaMM). Journal of Open Research Software, 9(1):14–??, June 08, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.309/.

Serang:2012:FMS

[SN12] Oliver Serang and William Stratford Noble. Faster mass spectrometry-based protein inference: Junction trees are more efficient than sampling and marginalization by enumeration. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 9 (3):809–817, May 2012. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Shajii:2019:SHP

[SNB^{+}19] Ariya Shajii, Ibrahim Numanagić, Riyadh Baghdadi, Bonnie Berger, and Saman Amarasinghe. Seq: a high-performance language for bioinformatics. Proceedings of the ACM on Programming Languages (PACMPL), 3(OOPSLA): 125:1–125:29, October 2019.

URL https://dl.acm.org/
doi/abs/10.1145/3360551.

Sinaie:2017:PMP

[SNNB17] Sina Sinaie, Viet Ha Nguyen, Chi Thanh Nguyen, and Stéphane Bordas. Programming the material point method in Julia. Advances in Engineering Software, 105 (??):17–29, March 2017. CO-DEN AESODT. ISSN 0965-9978 (print), 0141-1195 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0965997816302769.

Sihi:2023:TPC

[SP23] Antik Sihi and Sudhir K.

Pandey. TRACK: a Python code for calculating the transport properties of correlated electron systems using Kubo formalism. Computer Physics Communications, 285(??): Article 108640, April 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522003599.

Speck:2019:APP

[Spe19] Robert Speck. Algorithm 997: pySDC-prototyping spectral deferred corrections. ACM Transactions on Mathematical Software, 45(3):35:1– 35:23, August 2019. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

                                        tronic). URL https://dl.
                                        acm.org/citation.cfm?id=
                                        3310410.

Spielman:2018:PPP

[Spi18] Stephanie J. Spielman. phyphy: Python package for facilitating the execution and parsing of HyPhy standard analyses. Journal of Open Source Software, 3(21):514:1, January 2018. CODEN ???? ISSN 2475-9066. URL http:

                                        //joss.theoj.org/papers/
                                        10.21105/joss.00514.

Spotz:2012:PRA

[Spo12] William F. Spotz. PyTrilinos: Recent advances in the Python interface to Trilinos. Scientific Programming, 20 (3):311–325, ???? 2012. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Russell:2011:MSW

[sR11] Matthew A. (Computer scientist) Russell. Mining the social web. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2011. ISBN 1-4493-8834-5 (paperback), 1-4493-0416-8 (e-book), 1-4493-0393-5 (e-book), 1-4493-9475-

2. xx + 332 pp. LCCN QA76.9.D343 R87 2011. URL

http://twitter.com/\#!/
SocialWebMining.

Saad:2019:PPS

[SR19] Tony Saad and Giovanna Ruai. PyMaxEnt: a Python software for maximum entropy moment reconstruction. SoftwareX, 10(??):Article 100353, July/December 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019302456.

Shajii:2023:CCH

[SRS^{+}23] Ariya Shajii, Gabriel Ramirez, Haris Smajlović, Jessica Ray, Bonnie Berger, Saman Amarasinghe, and Ibrahim Numanagić. Codon: A compiler for high-performance Pythonic applications and DSLs. In Proceedings of the 32nd ACM SIG-PLAN International Conference on Compiler Construction. ACM Press, New York, NY 10036, USA, February 2023. URL https:

//docs.exaloop.io/codon/

                                        ;   https://github.com/
                                        exaloop/codon; https://
                                        github.com/exaloop/codon.
                                        git.

Stefik:2013:EIP

[SS13] Andreas Stefik and Susanna Siebert. An empirical investigation into programming language syntax. ACM Transactions on Computing Education, 13(4):19:1–19:??, November 2013. CODEN ???? ISSN 1946-6226.

Sartore:2021:P

[SS21] Lohan Sartore and Ingo Schienbein. PyR@TE 3. Computer Physics Communications, 261(??):Article 107819, April 2021. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520304124.

Silvestri:2022:SFP

[SSD^{+}22] Luciano G. Silvestri, Lucas J. Stanek, Gautham Dharuman, Yongjun Choi, and Michael S. Murillo. Sarkas: a fast pure-Python molecular dynamics suite for plasma physics. Computer Physics Communications, 272(??): Article 108245, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046552100357X.

Sala:2008:PHP

[SSH08] Marzio Sala, W. F. Spotz, and M. A. Heroux. PyTrilinos: High-performance distributedmemory solvers for Python. ACM Transactions on Mathematical Software, 34(2):7:1– 7:33, March 2008. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Simon:2016:PIA

[SSH16] Cory M. Simon, Berend Smit, and Maciej Haranczyk. pyI-AST: Ideal adsorbed solution theory (IAST) Python package. Computer Physics Communications, 200(??): 364–380, March 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465515004403.

Sharma:2022:UOS

[SSS22] Pankajeshwara Nand Sharma, Bastin Tony Roy Savarimuthu, and Nigel Stanger. Unearthing open source decisionmaking processes: a case study of Python enhancement proposals. Software—Practice and Experience, 52(10): 2312–2346, October 2022. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Singh:2023:BEC

[SST23] Raman Singh, Sean Sturley, and Hitesh Tewari.

Blockchain-enabled Chebyshev polynomial-based group authentication for secure communication in an Internet of Things network. Future Internet, 15(3):96, February 28, 2023. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/
15/3/96.

Scheible:2007:MPR

[ST^{+}07] Jürgen Scheible, Ville Tuulos, et al. Mobile Python: rapid prototyping of applications on the mobile platform. Wiley, New York, NY, USA, 2007. ISBN 0-470-51505-8 (paperback). xx + 327 pp. LCCN QA76.73.P98 S34

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0739/2007029113-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0833/2007029113-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0833/2007029113-t.html.

Stadler:2017:CCP

[Sta17]  Konstantin Stadler.  The

country converter coco — a Python package for converting country names between different classification schemes. Journal of Open Source Software, 2(16): 332:1–332:2, August 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00332.

Staub:2019:XHA

[Sta19] Florian Staub. xSLHA: an Les Houches Accord reader for Python and Mathematica. Computer Physics Communications, 241(??):132– 138, August 2019. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465519300918.

Stadler:2021:PPB

[Sta21] Konstantin Stadler. Pymrio — a Python based multiregional input-output analysis toolbox. Journal of Open Research Software, 9(1):8–??, May 11, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.251/.

Stelter:2006:BHA

[Ste06] Fred Stelter. Building a home automation and security system with Python. Linux journal, 2006(142):4, February 2006. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Stewart:2014:PS

[Ste14] John Stewart. Python for Scientists. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06139-3 (hardcover), 1-107-68642-3. xii + 220 pp. LCCN Q183.9 .S865

2014. URL http://assets.

cambridge.org/97811070/
61392/cover/9781107061392.
jpg.

Stone:2013:BRT

[Sto13] James V. Stone. Bayes’ rule: a tutorial introduction to Bayesian analysis. Sebtel Press, Lexington, KY, USA, 2013. ISBN 0-9563728-4-8 (paperback). 170 pp. LCCN QA279.5 .S766 2013.

Sharma:2018:RWF

[STS^{+}18] Abhishek Sharma, Yuan Tian, Agus Sulistya, Dinusha Wijedasa, and David Lo. Recommending who to follow in the software engineering Twitter space. ACM Transactions on Software Engineering and Methodology, 27(4):16:1–16:??, November 2018. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Stubblebine:2003:REP

[Stu03] Tony Stubblebine. Regular expression pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2003. ISBN 0-596-00415-X. vi + 93 pp. LCCN QA76.9.T48 S78 2003. URL

http://www.oreilly.com/
catalog/9780596004156.

Stubblebine:2007:REP

[Stu07] Tony Stubblebine. Regular expression pocket reference: Regular expressions for Perl, Ruby, PHP, Python, C, Java, and .NET. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-51427-1 (paperback). vii + 117 pp. LCCN QA76.9.T48 S78 2007. URL http://

proquest.safaribooksonline.
com/9780596514273;   http:
//www.loc.gov/catdir/toc/
fy0802/2007281074.html.

Suchojad:2013:ZAE

[Suc13] Dariusz Suchojad. Zato — agile ESB, SOA, REST and cloud integrations in Python. Linux journal, 2013(235):2:1– 2:??, November 2013. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Summerfield:2008:RGP

[Sum08] Mark Summerfield. Rapid GUI programming with Python and Qt: the definitive guide to PyQt programming. Prentice Hall open source software development series. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2008. ISBN 0-13-235418-7 (hardcover). xiv + 625 pp. LCCN QA76.9.U83 S89 2008.

Summerfield:2009:PPC

[Sum09]  Mark Summerfield.  Pro-

gramming in Python 3: a complete introduction to the Python language. Developer’s library. Addison-Wes-ley, Reading, MA, USA, 2009. ISBN 0-13-712929-7 (paperback). xiv + 525 pp. LCCN QA76.73.P98 S86 2009.

Summerfield:2010:PPC

[Sum10] Mark Summerfield. Programming in Python 3: a complete introduction to the Python language. Developer’s library. Addison-Wesley, Reading, MA, USA, second edition, 2010. ISBN 0-321-68056-1 (paperback). xvi + 630 pp. LCCN QA76.73.P98 S86 2010.

Schwarz:2021:PPF

[SUM21] Sebastian Schwarz, Sebastian Alexander Uerlich, and

                                        Antonello Monti. pycity scheduling

— a Python framework for the development and assessment of optimisationbased power scheduling algorithms for multi-energy systems in city districts. SoftwareX, 16(??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711021001230.

Suzuki:1997:P

[Suz97]  Junichi Suzuki.  Python.

Japanese Dr Dobbs, ??(??):

??, ???? 1997. A series of seven articles on Python.

Sandner:2014:CMC

[SV14] Raimar Sandner and András Vukics. C++QEDv2 Milestone 10: a C++/Python application-programming framework for simulating open quantum dynamics. Computer Physics Communications, 185(9):2380–2382, September 2014. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465514001349.

Shacham:2009:CAS

[SVY09] Ohad Shacham, Martin Vechev, and Eran Yahav. Chameleon: adaptive selection of collections. ACM SIG-PLAN Notices, 44(6):408– 418, June 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sabia:2014:PTV

[SW14] Martino Sabia and Cathy Wang. Python tools for Visual Studio leverage the power of the Visual Studio IDE to develop better and more efficient Python projects. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-868-X, 1-78328-869-8. LCCN QA76.73.P98. URL http://

                                        proquest.tech.safaribooksonline.
                                        de/9781783288687.

Sarker:2015:LPN

[SW15] M. O. Faruque Sarker and Sam Washington. Learning Python network programming: utilize Python 3 to get network applications up and running quickly and easily. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-600-1, 1-78439-115-8 (e-book). 320 (est.) pp. LCCN QA76.73.P98. URL http://

                                        proquest.safaribooksonline.
                                        com/9781784396008.

Sedgewick:2015:IPP

[SWD15] Robert Sedgewick, Kevin Daniel Wayne, and Robert Dondero. Introduction to programming in Python: an interdisciplinary approach. Addison-Wesley, Reading, MA, USA, 2015. ISBN 0-13-407653-2, 0-13-407643-5. ???? pp. LCCN QA76.73.P98. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9780134076539.

Sweigart:2009:PPL

[Swe09] Albert Sweigart. Playing with Python: learn to program by making games. No Starch Press, San Francisco, CA, USA, 2009. ISBN 1-59327-198-0 (paperback). ???? pp. LCCN ????

Sweigart:2013:HSC

[Swe13] Al Sweigart. Hacking Secret Ciphers with Python: [a beginner’s guide to cryptography and computer programming with Python]. CreateSpace Independent Publishing, North Charleston, SC, USA, 2013. ISBN 1-4826-1437-5. 436 pp. LCCN ????

URL http://swbplus.bsz-
bw.de/bsz40205301Xinh.htm.

Sweigart:2014:ABS

[Swe14] Al Sweigart. Automate the boring stuff with python: practical programming for total beginners. No Starch Press, San Francisco, CA, USA, 2014. ISBN 1-59327-599-4. xxi + 479 pp. LCCN QA76.73.P98. URL http://

proquestcombo.safaribooksonline.
com/9781457189906.

Swinnen:2009:APA

[Swi09]  Gérard Swinnen.  Ap-

prendre á programmer avec Python: objet, multithreading, événements, bases de données, programmation web, programmation réseau, Unicode. Eyrolles, Paris, France, 2009. xviii + 341 pp. LCCN ????

Shi:2021:CMS

[SYGY21] Yanjun Shi, Hao Yu, Yijia Guo, and Zhiheng Yuan. A collaborative merging strategy with lane changing in multilane freeway on-ramp area with V2X network. Future Internet, 13(5):123, May 10, 2021. CODEN ???? ISSN 1999-5903. URL https:/

                                        /www.mdpi.com/1999-5903/
                                        13/5/123.

Song:2023:PPC

[SZNW23] Shanshan Song, Mingyu Zhu, Hongcheng Ni, and Jian

                                        Wu. PyStructureFactor: a

Python code for the molecular structure factor in tunneling ionization rates. Computer Physics Communications, 292(??):Article 108882, November 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0010465523002278.

Singh:2022:DMT

[SZW^{+}22] Navjot Singh, Zecheng Zhang, Xiaoxiao Wu, Naijing Zhang, Siyuan Zhang, and Edgar Solomonik. Distributedmemory tensor completion for generalized loss functions in Python using new sparse tensor kernels. Journal of Parallel and Distributed Computing, 169(??): 269–285, November 2022. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0743731522001708.

Tabba:2010:ACP

[Tab10] Fuad Tabba. Adding concurrency in Python using a commercial processor’s hardware transactional memory support. ACM SIGARCH Computer Architecture News, 38(5):12–19, December 2010. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Tadj:2023:EID

[TAD23] Timothy Tadj, Reza Arablouei, and Volkan Dedeoglu. On evaluating IoT data trust via machine learning. Future Internet, 15(9):309, September 12, 2023. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/
15/9/309.

Talbot:2000:WGT

[Tal00] Hugues Talbot. wxPython, a GUI toolkit. Linux journal, 74:??, June 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Tan:2001:JWP

[Tan01] C. K. Tan. A JDBC wrapper — in Python! Dr. Dobb’s Journal of Software Tools, 26(8):50, 52, 54, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL

http://www.ddj.com/ftp/
2001/2001_08/jdbcpy.txt;
http://www.ddj.com/ftp/
2001/2001_08/jdbcpy.zip.

Tanimoto:2006:IPA

[Tan06] Steven L. Tanimoto. Introduction to Python for Artificial Intelligence. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN ???? ???? pp. LCCN ???? US$19.00. URL http://

www.computer.org/portal/
pages/ieeecs/ReadyNotes/
tanimotoabstract.html.

Tauber:2016:PPP

[Tau16] J. K. Tauber. pyuca: a Python implementation of the Unicode Collation Algorithm. Journal of Open Source Software, 1(1):21:1, May 2016. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00021.

Taylor:2018:PPP

[Tay18] Shawn David Taylor. pyPhenology: a Python framework for plant phenology modelling. Journal of Open Source Software, 3(28):827:1–827:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00827.

Tejedor:2017:PPC

[TBA^{+}17] Enric Tejedor, Yolanda Becerra, Guillem Alomar, Anna Queralt, Rosa M. Badia, Jordi Torres, Toni Cortes, and Jesús Labarta. Py-COMPSs: Parallel computational workflows in Python. The International Journal of High Performance Computing Applications, 31(1):66– 82, January 2017. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Thorp:2023:VPF

[TBD23] Kelly R. Thorp, Josh Brekel, and Kendall C. DeJonge. Version 1.2.0 — pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002145.

Turcani:2018:SPT

[TBJ18] Lukas Turcani, Enrico Berardo, and Kim E. Jelfs. stk: a Python toolkit for supramolecular assembly. Journal of Computational Chemistry, 39(23):1931–1942, September 5, 2018. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Terven:2021:KAK

[TCE21] Juan R. Terven and Diana M.

Córdova-Esparza. KinZ: an Azure Kinect toolkit for Python and Matlab. Science of Computer Programming, 211(??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0167642321000952.

Tan:2018:EPB

[TEG18] Jianchao Tan, Jose Echevarria, and Yotam Gingold. Efficient palette-based decomposition and recoloring of images via RGBXY-space geometry. ACM Transactions on Graphics, 37(6):262:1– 262:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Telles:2006:PPC

[Tel06] Matthew A. Telles. Python power!: the comprehensive guide. Thomson Course Technology PTR, Boston, MA, USA, 2006. ISBN 1-59863-158-6. xx + 508 pp. LCCN QA76.73.P98 T45 2006.

Thrussell:2023:EPL

[TF23] Jasper Thrussell and Jim Michael Ferguson. ExactPack: a Python library of exact analytic solutions. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102300256X.

Tan:2021:ETD

[TFAL21] Jie Tan, Daniel Feitosa, Paris Avgeriou, and Mircea Lungu. Evolution of technical debt remediation in Python: a case study on the Apache Software Ecosystem. Journal of Software: Evolution and Process, 33(4):e2319:1–e2319:??, April 2021. CODEN ???? ISSN 2047-7473 (print), 2047-7481 (electronic).

Tohline:2009:CPM

[TGEA09] Joel E. Tohline, Jinghya Ge, Wesley Even, and Erik Anderson. A customized Python module for CFD flow analysis within VisTrails. Computing in Science and Engineering, 11(3):68–73, May/June 2009. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Theis:2009:EPC

[The09]  Thomas Theis.  Einstieg

in Python 3: [auf CD: Python 3 und alle Code-Beispiele des Buchs ; für Programmanfänger und Umsteiger ; mit vielen Beispielen und èUbungsaufgaben ; inkl. objektorientierter Programmierung, Datenbanken, Internet u.v.m.]. Galileo computing. Galileo Press, Bonn, Germany, second edition, 2009. ISBN 3-8362-1406-7. 399 pp. LCCN ???? EUR 24.90.

Tesse:2023:GMP

[THG^{+}23] Robin Tesse, Cédric Hernalsteens, Eustache Gnacadja, Nicolas Pauly, Eliott Ramoisiaux, and Marion Vanwelde. Georges: a modular Python library for seamless beam dynamics simulations and optimization. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023002753.

Thorp:2022:PFE

[Tho22] Kelly R. Thorp. pyfao56: FAO-56 evapotranspiration in Python. SoftwareX, 19(??): ??, July 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022001261.

Thurman:2022:YTA

[Thu22] Marnanel Thurman. yex: a TEX-alike typesetter in Python. TUGboat, 43(2): 134–135, 2022. CODEN ???? ISSN 0896-3207. URL

                                        https://tug.org/TUGboat/
                                        tb43-2/tb134thurman-yex.
                                        pdf.

Tidwell:2007:XMX

[Tid07] Doug Tidwell. XSLT: mastering XML transformations. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2007. ISBN 0-596-52721-7 (paperback). xviii + 965 pp. LCCN QA76.73.X58 T53 2008. US$31.99.

Tien:2016:PPM

[Tie16] Vivienne Tien. Python and physical modeling. Computing in Science and Engineering, 18(3):8–10, May/June 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Turner:2021:OLE

[TLR21] Ross J. Turner, Rebecca B.

Latto, and Anya M. Reading. An ObsPy library for event detection and seismic attribute calculation: Preparing waveforms for automated analysis. Journal of Open Research Software, 9(1):29–??, October 19, 2021. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.365/.

Tritt:2023:SGO

[TMH^{+}23] Alex Tritt, Joshua Morris, Joel Hochstetter, R. P. Anderson, James Saunderson, and L. D. Turner. Spinsim: a GPU optimized Python package for simulating spinhalf and spin-one quantum systems. Computer Physics Communications, 287(??): Article 108701, June 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523000462.

Toms:2015:AAG

[Tom15]  Silas Toms.  ArcPy and

ArcGIS-geospatial analysis with Python: use the ArcPy module to automate the analysis and mapping of geospatial data in ArcGIS. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78398-866-5, 1-78398-867-3 (e-book). 224 (est.) pp. LCCN G70.212. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781783988662.

Thieu:2023:MOS

[TOPC23] Nguyen Van Thieu, Diego Oliva, and Marco Pérez-

                                        Cisneros.  MetaCluster:

an open-source Python library for metaheuristic-based clustering problems. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023002935.

Tosi:2009:MPD

[Tos09] Sandro Tosi. Matplotlib for Python developers: build remarkable publication quality plots the easy way. From technologies to solutions. Packt Publishing, Birmingham, UK, 2009. ISBN 1-84719-790-6. iv + 293 pp. LCCN QA76.73.P48 T67 2009.

Toriyama:2023:VPT

[TQGE23] Michael Y. Toriyama, Jiaxing Qu, Lídia C. Gomes, and Elif Ertekin. VTAnDeM: a Python toolkit for simultaneously visualizing phase stability, defect energetics, and carrier concentrations of materials. Computer Physics Communications, 287(??): Article 108691, June 2023.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046552300036X.

Tawfik:2022:PPP

[TR22] Sherif Abdulkader Tawfik and Salvy P. Russo.

PyPhotonics: a Python

package for the evaluation of luminescence properties of defects. Computer Physics Communications, 273(??): Article 108222, April 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465521003349.

Troan:1996:FSSb

[Tro96a] Erik Troan. Free software solutions: Basic X programming in Python. The X Journal: Computing Technology with the X Window System, 5(6):84–??, June 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSc

[Tro96b] Erik Troan. Free software solutions: Meta widgets in Python. The X Journal: Computing Technology with the X Window System, 5(7): ??, July/August 1996. CO-DEN XJOUEA. ISSN 1056-7003.

Troan:1996:FSSa

[Tro96c] Erik Troan. Free software solutions: The Python language. The X Journal: Computing Technology with the X Window System, 5(5):96– ??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Troan:1997:FSS

[Tro97] Eric Troan. Free software solutions: From Python to Java. UNIX Developer, 1 (1):77–78, January/February 1997. ISSN 1090-2279.

Trevino:2014:ANP

[TS14] Jeffrey Treviño and Craig Sapp. Automated notation of piano recordings for historic performance practice study. Journal on Computing and Cultural Heritage (JOCCH), 7(3):17:1–17:??, August 2014. CODEN ???? ISSN 1556-4673 (print), 1556-4711 (electronic).

Tartarini:2020:PPP

[TS20] Federico Tartarini and Ste-

                                        fano Schiavon. pythermalcomfort:

a Python package for thermal comfort research. SoftwareX, 12(??):Article 100578, July/December 2020. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711020302910.

Thiruvathukal:2002:WPT

[TSC02] George K. (George Kuriakose) Thiruvathukal, John P. Shafaee, and Thomas W. Christopher. Web programming: techniques for integrating Python, Linux, Apache, and MySQL. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2002. ISBN 0-13-041065-9. xviii + 745 pp. LCCN QA76.625 .T48

2002. URL http://www.
phptr.com/ptrbooks/ptr_
0130410659.html.

Takikawa:2012:GTF

[TSD^{+}12] Asumu Takikawa, T. Stephen Strickland, Christos Dimoulas, Sam Tobin-Hochstadt, and Matthias Felleisen. Gradual typing for first-class classes. ACM SIGPLAN Notices, 47(10):793–810, October 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tsapetis:2023:UVU

[TSG^{+}23] Dimitrios Tsapetis, Michael D.

Shields, Dimitris G. Giovanis, Audrey Olivier, Lukas Novak, Promit Chakroborty, Himanshu Sharma, Mohit Chauhan, Katiana Kontolati, Lohit Vandanapu, Dimitrios Loukrezis, and Michael Gardner. UQpy v4.1: Uncertainty quantification with Python. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002571.

Theisen:2021:FTM

[TT21] Lambert Theisen and Manuel Torrilhon. fenicsR13: a tensorial mixed finite element solver for the linear R13 equations using the FEniCS computing platform. ACM Transactions on Mathematical Software, 47(2):17:1– 17:29, April 2021. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3442378.

Toth:2023:VBP

[TTHZ23] Róbert Tóth, Bálint Tóth, Miklós Hoffmann, and Mari-

                                        anna Zichar. viskillz-blender

— a Python package to generate assets of Mental Cutting Test exercises using Blender. SoftwareX, 22(??): ??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711023000249.

Tatsubori:2010:EJT

[TTS^{+}10] Michiaki Tatsubori, Akihiko Tozawa, Toyotaro Suzumura, Scott Trent, and Tamiya Onodera. Evaluation of a justin-time compiler retrofitted for PHP. ACM SIGPLAN Notices, 45(7):121–132, July 2010. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Toby:2013:GIG

[TV13] Brian H. Toby and Robert B. Von Dreele. GSAS-II: the genesis of a modern opensource all purpose crystallography software package. Journal of Applied Crystallography, 46(2):544–549, 2013. CODEN JACGAR. ISSN 0021-8898 (print), 1600-5767 (electronic).

Ucoluk:2012:IPC

[ÜK12] Göktürk Üçoluk and Sinan Kalkan. Introduction to programming concepts with case studies in Python. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-7091-1342-3, 3-7091-1343-1 (e-book). LCCN QA76.6 .U26 2012.

Ulloa:2015:KIA

[Ull15] Roberto Ulloa. Kivy — interactive applications and games in Python: create responsive cross-platform UI/UX applications and games in Python and using the open source Kivy library. Community Experience Distilled. Packt Publishing, Birmingham, UK, second edition, 2015. ISBN 1-78528-692-7, 1-78528-438-X. 206 (est.) pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781785286926.

USENIX:1996:PSUb

[USE96] USENIX, editor. Proceedings of the Second USENIX Conference on Object-Oriented Technologies and Systems (COOTS), June 17–21, 1996, Toronto, Canada. USENIX, Berkeley, CA, USA, 1996. ISBN 1-880446-77-4. LCCN QA76.64 .U85 1996. URL

                                        http://www.usenix.org/publications/
                                        library/proceedings/coots96/

.

USENIX:2000:PAL

[USE00]  USENIX, editor.  Pro-

ceedings of the 4th Annual Linux Showcase and Conference, Atlanta, October 10–14, 2000, Atlanta, Georgia, USA. USENIX, Berkeley, CA, USA, 2000. ISBN 1-880446-17-0. LCCN

                                        ????  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/als2000/

.

UUIG:1995:VBA

[UVa95] UVa User Interface Group. VR blackboard: Alice: Rapid prototyping for virtual reality. IEEE Computer Graphics and Applications, 15(3): 8–11, May 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Vaingast:2009:BPV

[Vai09] Shai Vaingast. Beginning Python visualization: crafting visual transformation scripts. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1843-6 (paperback). xx + 363 pp. LCCN QA76.73.P98 V35 2009.

Vaingast:2014:BPV

[Vai14] Shai Vaingast. Beginning Python visualization crafting visual transformation scripts. Expert’s voice in Python. Apress, Berkeley, CA, USA, second edition, 2014. ISBN 1-4842-0053-5. LCCN ???? URL http://

proquest.tech.safaribooksonline.
de/9781484200520.

vanRossum:1995:EEP

[van95a] Guido van Rossum. Extending and embedding the Python interpreter. Report CS-R9527, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, April 1995. i + 22 pp.

vanRossum:1995:PLR

[van95b] Guido van Rossum. Python library reference. Report CS-R9524, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, April 1995. iv + 186 pp.

URL http://www.python.
org/doc/lib/lib.html.

vanRossum:1995:PRM

[van95c] Guido van Rossum. Python reference manual. Report CS-R9525, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, April 1995. ii + 54 pp.

URL http://www.python.
org/doc/ref/ref-1.html.

vanRossum:1995:PT

[van95d] Guido van Rossum. Python tutorial. Report CS-R9526, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, April 1995. iii + 65 pp. URL http:/

                                        /www.python.org/doc/tut/
                                        tut.html.

vanRossum:1997:P

[van97a] Guido van Rossum. Python. World Wide Web Journal, 2 (2):??, Spring 1997. CO-DEN WWWFFI. ISSN 1085-

                                        2301. URL http://www.ora.
                                        com/catalog/wj6/.

Vanaken:1997:WCS

[Van97b] Michel Vanaken. Writing CGI scripts in Python. Linux journal, 34:??, February 1997. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

vanRossum:1998:TPL

[van98a] G. van Rossum. A tour of the Python language. In Ege et al. [ESM98], pages 370–?? ISBN 0-8186-8383-X. LCCN ???? IEEE catalog number 97TB100221.

vanRossum:1998:EEP

[van98b] Guido van Rossum. Extending and embedding the Python interpreter. Technical report, Corporation for National Research Initiatives, 1895 Preston White Drive, Suite 100, Reston, VA 20191, USA, April 14, 1998.

URL http://www.python.
org/doc/ext/ext.html. For

Python Release 1.5.1.

vanRossum:1998:PCA

[van98c] Guido van Rossum. Python/C API reference manual. Technical report, Corporation for National Research Initiatives, 1895 Preston White Drive, Suite 100, Reston, VA 20191, USA, 1998. URL http:/

/www.python.org/doc/api/

api.html. For Python Release 1.5.1.

Vanderplas:2016:PDS

[Van16] Jake Vanderplas. Python Data Science Handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2016. ISBN 1-4919-1205-7. ???? pp. LCCN ????

Verstraelen:2021:IPL

[VAP^{+}21] Toon Verstraelen, William Adams, Leila Pujal, Alireza Tehrani, Braden D. Kelly, Luis Macaya, Fanwang Meng, Michael Richer, Raymundo Hernández-Esparza, Xiaotian Derrick Yang, Matthew Chan, Taewon David Kim, Maarten Cools-Ceuppens, Valerii Chuiko, Esteban Vöhringer-Martinez, Paul W. Ayers, and Farnaz Heidar-Zadeh. IOData: a Python library for reading, writing, and converting computational chemistry file formats and generating input files. Journal of Computational Chemistry, 42(6):458– 464, March 5, 2021. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Varley:2016:EPP

[Var16]  Ryan Varley.  ExoData:

a Python package to handle large exoplanet catalogue data. Computer Physics Communications, 207(??): 298–309, October 2016. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465516301254.

Vallisneri:2008:PXA

[VB08] Michele Vallisneri and Stanislav Babak. Python and XML for agile scientific computing. Computing in Science and Engineering, 10(1):80– 87, January/February 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Villaverde:2018:PTI

[VBB18] Alejandro F. Villaverde, Kolja Becker, and Julio R. Banga. PREMER: a tool to infer biological networks. IEEE/ACM Transactions on Computational Biology and Bioinformatics, 15(4):1193– 1202, July 2018. CO-DEN ITCBCY. ISSN 1545-5963 (print), 1557-9964 (electronic).

Velichko:2023:NNE

[VBI^{+}23] Andrei Velichko, Maksim Belyaev, Yuriy Izotov, Murugappan Murugappan, and Hanif Heidari. Neural network entropy (NNetEn): Entropy-based EEG signal and chaotic time series classification, Python package for NNetEn calculation. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.
com/1999-4893/16/5/255.

Vrbancic:2018:PNP

[VBM^{+}18] Grega Vrbančič, Lucija Brezočnik, Uroš Mlakar, Dušan Fister, and Iztok Fister, Jr. NiaPy: Python microframework for building natureinspired algorithms. Journal of Open Source Software, 3(23):613:1–613:3, March 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00613.

Vidmar:2018:QPP

[VC18] R. Vidmar and N. Creati.

QCOBJ: a Python package to handle quantity-aware configuration files. SoftwareX, 7 (??):347–351, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018302383.

Valiev:2022:CPP

[VCF22] Marat Valiev, Gennady N. Chuev, and Marina V. Fedotova. CDFTPY: a Python package for performing classical density functional theory calculations for molecular liquids. Computer Physics Communications, 276(??): Article 108338, July 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046552200056X.

Vishart:2021:PPB

[VCLS21] Jonas Lynge Vishart, Jaime Castillo-León, and Winnie E. Svendsen. pyEIA: a Pythonbased framework for data analysis of electrochemical methods for immunoassays. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711021000637.

Vivien:2017:GDP

[VCR17] Vladimir Vivien, Mario Contreras, and Mat Ryer. Go: Design Patterns for Real-World Projects. Packt Publishing, Birmingham, UK, 2017. ISBN 1-78839-055-5, 1-78839-287-6. 1091 pp. LCCN ????

vanRossum:1991:LSG

[vd91] G. van Rossum and J. de Boer. Linking a stub generator (AIL) to a prototyping language (Python). In EurOpen [Eur91], pages 229–

247. ISBN 1-873611-00-5. LCCN ????

vanderHam:2018:PSD

[vdH18] Ruud van der Ham. salabim:

discrete event simulation and animation in Python. Journal of Open Source Software, 3(27):767:1–767:2, July 2018. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00767.

vandenOord:2020:PID

[vdOJP^{+}20] Gijs van den Oord, Fredrik Jansson, Inti Pelupessy, Maria Chertova, Johanna H. Grönqvist, Pier Siebesma, and Daan Crommelin. A Python interface to the Dutch Atmospheric Large-Eddy Simulation. SoftwareX, 12(??):Article 100608, July/December 2020. CO-DEN ???? ISSN 2352-

7110. URL http://www.
sciencedirect.com/science/
article/pii/S2352711020303216.

Velasquez:2023:TAB

[Vel23] Juan D. Velasquez. TechMiner:

Analysis of bibliographic datasets using Python. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S235271102300153X.

Verma:2022:PBT

[Ver22a] Ashok Kumar Verma. A Python-based tool for constructing observables from the DSN’s closed-loop archival tracking data files. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022001145.

Verschelde:2022:EAS

[Ver22b] Jan Verschelde. Exporting Ada software to Python and Julia. ACM SIGADA Ada Letters, 42(1):76–78, June 2022. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL

https://dl.acm.org/doi/
10.1145/3577949.3577961.

Varner:2019:WPP

[VEV^{+}19] James F. Varner, Noor Eldabagh, Derek Volta, Reem Eldabagh, and Jonathan J. Foley IV. WPTherml: a Python package for the design of materials for harnessing heat. Journal of Open Research Software, 7(1):28–??, August 19, 2019. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.
metajnl.com/articles/10.
5334/jors.271/.

vonLowis:1997:PB

[vF97] Martin von Löwis and Nils Fischbeck. Das Python-Buch. Addison-Wesley Longman, Reading, MA, USA, 1997. ISBN 3-8273-1110-1. 495 pp. DM 69,90; ATS 510; CHR 63,00. Includes CD-ROM.

VanHensbergen:2008:HAR

[VFMM08] Eric Van Hensbergen, Charles Forsyth, Jim McKie, and Ron Minnich. Holistic aggregate resource environment. Operating Systems Review, 42(1): 85–91, January 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Virmani:2016:CFU

[Vir16]  Vineet Virmani.  Com-

putational finance using QuantLib-Python. Computing in Science and Engineering, 18(2):78–88, March/ April 2016. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vitousek:2015:DEG

[VKSB15] Michael M. Vitousek, Andrew M. Kent, Jeremy G. Siek, and Jim Baker. Design and evaluation of gradual typing for Python. ACM SIGPLAN Notices, 50(2):45– 56, February 2015. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Velikajne:2022:PRP

[VM22] Nina Velikajne and Miha

Moškon. RhythmCount: a

Python package to analyse the rhythmicity in count data. Journal of Computational Science, 63:??, September 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S1877750322001429.

Vieira:2017:WCC

[VMFG17] Camilo Vieira, Alejandra J. Magana, Michael L. Falk, and R. Edwin Garcia. Writing incode comments to self-explain in computational science and engineering education. ACM Transactions on Computing Education, 17(4):17:1–17:??, September 2017. CODEN ???? ISSN 1946-6226.

Vergara:2023:SSP

[VMRFC23] J. M. Vergara, M. E. Mora-Ramos, E. Flórez, and J. D. Correa. SPIN: P[S]imple [P]ython [I]pywidgets [N]otebook interface to obtain the optoelectronic properties of materials employing DFT. Computer Physics Communications, 284(??):Article 108614, March 2023. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465522003332.

Verma:2019:CAH

[VPO19] Pradeepika Verma, Sukomal Pal, and Hari Om. A comparative analysis on Hindi and English extractive text summarization. ACM Transactions on Asian and Low-Resource Language Information Processing (TALLIP), 18(3):30:1–30:??, July 2019. CODEN ???? ISSN 2375-4699 (print), 2375-4702 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3308754.

vanRossum:2003:IPR

[vRD03] Guido van Rossum and Fred L. Drake, Jr., editors. An introduction to Python: release 2.2.2. Network Theory Ltd., Bristol, UK, 2003. ISBN 0-9541617-6-9. ii + 115 pp. LCCN ????

vanRossum:2005:PLR

[vRD05a] Guido van Rossum and Fred L. Drake, Jr. Python Language Reference Manual. Network Theory Ltd., Bristol, UK, 2005. ISBN 0-9541617-8-5. ii + 112 pp. LCCN QA76.73.P98.

URL http://www.network-
theory.co.uk/python/language/

.

vanRossum:2005:PTI

[vRD05b] Guido van Rossum and Fred L. Drake, Jr. The Python Tutorial — An Introduction to Python. Network Theory Ltd., Bristol, UK, 2005. ISBN 0-9541617-6-

9. 124 (est.) pp. LCCN ????

vanRossum:1999:PT

[vRDK99] Guido van Rossum, Fred L. Drake, and Andrew Kuchling. Python tutorial. Open Docs Library, Lincoln, NE, USA, 1999. ISBN 1-58348-375-6 (soft cover). LCCN QA76.73.P98 V36 2000b.

                                        URL ftp://uiarchive.cso.
                                        uiuc.edu/pub/etext/gutenberg/

.

Vitousek:2017:BTL

[VSS17] Michael M. Vitousek, Cameron Swords, and Jeremy G. Siek. Big types in little runtime: open-world soundness and collaborative blame for gradual type systems. ACM SIG-PLAN Notices, 52(1):762– 774, January 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Verschelde:2015:PHC

[VY15] Jan Verschelde and Xiangcheng Yu. Polynomial homotopy continuation on GPUs. ACM Communications in Computer Algebra, 49 (4):130–133, December 2015. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Wai:2021:XA

[Wai21] Richard Wai. XERIS/APEX. ACM SIGADA Ada Letters, 40(2):65–69, April 2021.

CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3463478.3463484.

Welker:2022:CPF

[WAK22] Simon Welker, Muhamed Amin, and Jochen Küpper. CMInject: Python framework for the numerical simulation of nanoparticle injection pipelines. Computer Physics Communications, 270(??):Article 108138, January 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0010465521002502.

WalnutCreek:1997:PA

[Wal97] Walnut Creek. The Python archive. CD ROM., June 1997. US$39.95. URL http:

//www.cdrom.com/titles/
prog/python.htm.  In-

cludes precompiled versions of Python for: AIX, BeOS, BSDI, Digital Unix (DEC OSF/1), DGUX, FreeBSD, HPUX, IRIX, Linux, Mac OS, OpenVMS (alpha and VAX), OS/2 (emx), RS6000, SCO, Sequent PTS, SGI IRIX, Solaris x86, Sparc/Solaris, Sparc/SunOS, Ultrix, VMS/VAX, and Windows 95/NT.

Walerowski:2007:PSV

[Wal07] Peter Walerowski, editor.

Python: 5 Stunden Video-Training: PC, Mac und TV, volume 6080 of Video2Brain; Addison-Wesley; Open source library. Addison-Wesley, Reading, MA, USA, 2007. ISBN 3-8273-6080-3. LCCN ???? EUR 49.95. One DVD.

Walker:2022:POP

[WAN^{+}22] S. D. Walker, A. Abramov, L. J. Nevay, W. Shields, and S. T. Boogert. Pyg4ometry: a Python library for the creation of Monte Carlo radiation transport physical geometries. Computer Physics Communications, 272(??): Article 108228, March 2022. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521003404.

Warner:2018:IPP

[War18] Mellissa S. C. Warner. Introduction to PySPLIT: a Python toolkit for NOAA ARL’s HYSPLIT model. Computing in Science and Engineering, 20(5):47–62, September/October 2018. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/cs/2018/05/mcs2018050047-
                                        abs.html.

Watters:1995:TAN

[Wat95] Aaron R. Watters. Tutorial article no. 005: The what, why, who, and where of Python. UnixWorld Online, ??(??):??, ???? 1995.

URL http://www.wcmh.com/
uworld/archives/95/tutorial/
005.html.

Wanzenbock:2024:CFS

[WBK^{+}24] Ralf Wanzenböck, Florian Buchner, Péter Kovács, Georg K. H. Madsen, and Jesús Carrete. Clinamen2: Functional-style evolutionary optimization in Python for atomistic structure searches. Computer Physics Communications, 297(??):Article 109065, April 2024. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S0010465523004101.

Wagner:2022:CAT

[WBR^{+}22] Felix Wagner, Daniel Bartolot, Damir Rizvanovic, Florian Reindl, Jochen Schieck, and Wolfgang Waltenberger. Cait: Analysis toolkit for cryogenic particle detectors in Python. Computing and Software for Big Science, 6 (1):??, December 2022. CO-DEN ???? ISSN 2510-2036 (print), 2510-2044 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s41781-022-00092-4.

Wacha:2021:PPE

[WBS21] András Wacha and Tamás Beke-Somfai. PmlBeta: a Py-MOL extension for building β-amino acid insertions and β-peptide sequences. SoftwareX, 13(??):Article 100654, January 2021. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711020303678.

Watrobski:2022:PMR

[WBS22] Jarosław Watróbski, Aleksandra Baczkiewicz, and Wojciech Sałabun. pyrepo-mcda

— reference objects based MCDA software package. SoftwareX, 19(??):??, July 2022. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022000711.

Weigend:2006:OPP

[Wei06a] Michael Weigend. Objektorientierte Programmierung mit Python: [Klassen, Objekte, Vererbung und Polymorphie praktisch angewendet; XML, GUI-Programmierung, Threads und CGI-Scripting; Übungen mit Musterlösungen zu jedem Kapitel]. Programmierung. mitp, Bonn, Germany, third edition, 2006. ISBN 3-8266-1660-X. 700 pp. LCCN ???? EUR

                                        36.95. URL http://www.
                                        gbv.de/dms/ilmenau/toc/
                                        510475221.PDF.

Weigend:2006:PGP

[Wei06b] Michael Weigend. Python Ge-Packt: [schneller Zugriff auf Module, Klassen und Funktionen; XML, Tkinter, Datenbanken, Internet-Programmierung; objektorientierte Programmierung und New-Style-Klassen]. Ge-Packte Referenz. mitp, Heidelberg, Germany, third edition, 2006. ISBN 3-8266-1659-6. 618 pp. LCCN ???? EUR 17.95. URL http:/

/www.gbv.de/dms/bsz/toc/
bsz254799760inh.pdf.

Weigend:2009:OPP

[Wei09] Michael Weigend. Objektorientierte Programmierung mit Python 3.0. mitp bei Redline. REDLINE, Heidelberg, Neckar, Germany, fourth edition, 2009. ISBN 3-8266-1750-9. 752 pp. LCCN ???? EUR 39.95.

Weiming:2015:MPF

[Wei15] James Ma Weiming. Mastering Python for finance: understand, design, and implement state-of-the-art mathematical and statistical applications used in finance with Python. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78439-451-3. xv + 312 pp. LCCN ????

Weppner:2015:DPS

[Wep15] Stephen Weppner. A different perspective on scientific programming [review of “Annotated algorithms in Python; with applications in physics, biology, and finance’ (Di Pierro, M.; 2013)]. Computing in Science and Engineering, 17(1):6–7, January/February 2015. CO-DEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

                                        tronic). URL http://csdl.
                                        computer.org/csdl/mags/
                                        cs/2015/01/mcs2015010006.
                                        pdf.

Westra:2015:PGA

[Wes15] Erik Westra. Python geospatial analysis essentials: process, analyze, and display geospatial data using Python libraries and related tools. Community experience distilled. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78355-389-8, 1-78217-451-6. ume pp. LCCN QA76.73.P98. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781782174516.

Wette:2020:SPO

[Wet20]  Karl Wette.  SWIGLAL:

Python and Octave interfaces to the LALSuite gravitational-wave data analysis libraries. SoftwareX, 12(??):Article 100634, July/ December 2020. CO-DEN ???? ISSN 2352-

                                        7110. URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/S2352711020303472.

Whalley:2018:PEE

[Wha18] Lucy D. Whalley. effmass: An effective mass package. Journal of Open Source Software, 3(28):797:1–797:2, August 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00797.

Willner:2017:PPP

[WHG17] Sven N. Willner, Corinne Hartin, and Robert Gieseke. pyhector: a Python interface for the simple climate model Hector. Journal of Open Source Software, 2 (12):248:1–248:2, April 2017. CODEN ???? ISSN 2475-9066. URL http:/

/joss.theoj.org/papers/
10.21105/joss.00248.

Wiecha:2018:PPP

[Wie18] Peter R. Wiecha. pyGDM — a Python toolkit for fullfield electro-dynamical simulations and evolutionary optimization of nanostructures. Computer Physics Communications, 233(??): 167–192, December 2018. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046551830225X.

Willison:1997:BEP

[Wil97a]  Frank Willison.  Bleed-

ing edge: Python: It’s not just for laughs. Web Review, ??(??):??, January 3, 1997. URL http:

//webreview.com/wr/pub/
Willison_Frank.

Wilson:1997:PBP

[Wil97b] Gregory V. Wilson. Programmer’s bookshelf: Perusing the bookshelf. Dr. Dobb’s Journal of Software Tools, 22(11): 125, 127, November 1997. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBP

[Wil00] Gregory V. Wilson. Programmer’s bookshelf: Python, C++, and other religions. Dr. Dobb’s Journal of Software Tools, 25(8):145–147, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2001:PBT

[Wil01] Gregory V. Wilson. Programmer’s bookshelf: Time warps. Dr. Dobb’s Journal of Software Tools, 26(2):159– 160, February 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Wilson:2002:PBS

[Wil02] Gregory V. Wilson. Programmer’s bookshelf: Sometimes you get what you want. Dr. Dobb’s Journal of Software Tools, 27(2):107–170, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:

//www.ddj.com/.

Wilson:2005:DCS

[Wil05] Greg Wilson. Data crunching: solve everyday problems using Java, Python and more. The pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-7-1. viii + 193 pp. LCCN QA63 .W55 2005.

Wimmer:2012:AEN

[Wim12] M. Wimmer. Algorithm 923:

Efficient numerical computation of the Pfaffian for dense and banded skew-symmetric matrices. ACM Transactions on Mathematical Software, 38 (4):30:1–30:17, August 2012. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Wittek:2015:ANS

[Wit15] Peter Wittek. Algorithm 950: Ncpol2sdpa — sparse semidefinite programming relaxations for polynomial optimization problems of noncommuting variables. ACM Transactions on Mathematical Software, 41(3):21:1– 21:12, June 2015. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Witherden:2021:PPP

[Wit21] Freddie D. Witherden. Python at petascale with PyFR or: How I learned to stop worrying and love the snake. Computing in Science and Engineering, 23(4):29–37, July/ August 2021. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Wieckowski:2022:PPL

[WKS22] Jakub Wieckowski, Bartłomiej Kizielewicz, and Wojciech Sałabun. pyFDM: a Python library for uncertainty decision analysis methods. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001893.

Wieckowski:2023:HDM

[WKS23] Jakub Wieckowski, Bartłomiej Kizielewicz, and Wojciech Sałabun. Handling decisionmaking in Intuitionistic Fuzzy environment: PyIFDM package. SoftwareX, 22(??):??, May 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023000407.

Walter:2013:ADP

[WL13] Sebastian F. Walter and Lutz Lehmann. Algorithmic differentiation in Python with AlgoPy. Journal of Computational Science, 4(5):334– 344, September 2013. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S1877750311001013.

Williams:2020:PSP

[WL20] Brendan Williams and Michael Lindner. pyfMRIqc: a software package for raw fMRI data quality assurance. Journal of Open Research Software, 8(1):23–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https:

//openresearchsoftware.

metajnl.com/articles/10.
5334/jors.280/.

Wilkinson:2021:EPM

[WM21] Collin J. Wilkinson and John C. Mauro. Explorer.py: Mapping the energy landscapes of complex materials. SoftwareX, 14(??):??, June 2021. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711021000285.

Wilbers:2012:IJT

[WMA12] Ilmar M. Wilbers, Kent-Andre Mardal, and Martin S. Alnæs. Instant: justin-time compilation of C/ C++ in Python. In Logg et al. [LMW12], pages 257–

272. CODEN LNCSA6. ISBN 3-642-23098-9 (print), 3-642-23099-7 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
content/pdf/10.1007/978-
3-642-23099-8_14. The soft-

ware developed by the FEniCS Project is free for all to use and modify (licensed under the GNU (L)GPL), and so is this book.

Wiecha:2022:PNF

[WMA^{+}22] Peter R. Wiecha, Clément Majorel, Arnaud Arbouet, Adelin Patoux, Yoann Brûlé, Gérard Colas des Francs, and Christian Girard. “pyGDM”

— new functionalities and major improvements to the Python toolkit for nanooptics full-field simulations.

Computer Physics Communications, 270(??):Article 108142, January 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S001046552100254X.

Wilkinson:2018:RPC

[WMM18] Collin J. Wilkinson, Yihong Z. Mauro, and John C. Mauro. RelaxPy: Python code for modeling of glass relaxation behavior. SoftwareX, 7(??):245–254, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711018301146.

Williamson:2014:PPP

[WO14] Todd Williamson and Ronald A. Olsson. PySy: a Python package for enhanced concurrent programming. Concurrency and Computation: Practice and Experience, 26 (2):309–335, February 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Widyasari:2022:RWP

[WPH^{+}22] Ratnadira Widyasari, Gede Artha Azriadi Prana, Stefanus Agus Haryono, Shaowei Wang, and David Lo. Real world projects, real faults: evaluating spectrum based fault localization techniques on Python projects. Empirical Software Engineering, 27 (6):??, November 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10189-4.

Weise:2020:PSU

[WPM^{+}20] Konstantin Weise, Lucas Poßner, Erik Müller, Richard Gast, and Thomas R. Knösche. Pygpc: a sensitivity and uncertainty analysis toolbox for Python. SoftwareX, 11 (??):Article 100450, January/ June 2020. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711020300078.

Walter:2021:MML

[WRBT21] Vivien Walter, Céline Ruscher, Olivier Benzerara, and Fabrice Thalmann. MLLPA: a machine learning-assisted Python module to study phase-specific events in lipid membranes. Journal of Computational Chemistry, 42(13): 930–943, May 15, 2021. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Wagner:2022:FEL

[WSK22] Christoph W. Wagner, Sebastian Semper, and Jan Kirchhof. fastmat: Efficient linear transforms in Python. SoftwareX, 18(??): ??, June 2022. CODEN ???? ISSN 2352-7110. URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S2352711022000218.

Wu:2013:HSC

[Wu13] Youfeng Wu. HW/SW codesigned acceleration of dynamic languages. ACM SIGPLAN Notices, 48(5): 1–2, May 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wittreich:2022:PGA

[WV22] Gerhard R. Wittreich and Dionisios G. Vlachos. Python Group Additivity (pGrAdd) software for estimating species thermochemical properties. Computer Physics Communications, 273(??):Article 108277, April 2022. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465521003891.

Watters:1996:IPP

[WvA96] Aaron Watters, Guido van Rossum, and James C. Ahlstrom. Internet programming with Python. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55851-484-8. xviii + 477 pp. LCCN QA76.73.P98 W38 1996. US$34.95. URL http:

                                        //www.fsbassociates.com/
                                        books/python.htm.

Wainer:2018:CEP

[WX18] Jacques Wainer and Eduardo C. Xavier. A controlled experiment on Python vs C for an introductory programming course: Students’ outcomes. ACM Transactions on Computing Education, 18(3): 12:1–12:??, September 2018. CODEN ???? ISSN 1946-6226.

Willcox:2018:PPI

[WZ18] Donald E. Willcox and Michael Zingale. pynucastro: an interface to nuclear reaction rates and code generator for reaction network equations. Journal of Open Source Software, 3(23):588:1–588:3, March 2018. CODEN ???? ISSN 2475-9066. URL http:

//joss.theoj.org/papers/
10.21105/joss.00588.

Wang:2023:PPP

[WZZ^{+}23] Hongjin Wang, Jingyi Zhuang, Zhen Zhang, Qi Zhang, and Renata M. Wentzcovitch. pgm: a Python package for free energy calculations within the phonon gas model. Computer Physics Communications, 291(??): Article 108845, October 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/S001046552300190X.

Xie:2021:RMC

[XDR21] Yihui Xie, Christophe Dervieux, and Emily Riederer. R Markdown Cookbook. The R series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2021. ISBN 0-367-56382-7 (hardcover), 0-367-56383-5 (paperback), 1-00-029080-8 (e-book), 1-00-029084-0 (Mobipocket e-book), 1-00-029088-3 (e-book), 1-00-309747-2 (e-book). xxix + 329 pp. LCCN QA276.45.R3 X54 2021.

Xu:2023:PPT

[XGW23] Jiachen Xu and Moritz Grosse-Wentrup. PyTES: a Python toolbox for closedloop transcranial electrical stimulation. SoftwareX, 23(??):??, July 2023. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711023000997.

Xu:2018:PPP

[XL18] Yang Xu and Xiao-Chun Luo. PyPathway: Python package for biological network analysis and visualization. Journal of Computational Biology, 25 (5):499–504, May 2018. CO-DEN JCOBEM. ISSN 1066-5277 (print), 1557-8666 (elec-

tronic). URL https://www.
liebertpub.com/doi/abs/
10.1089/cmb.2017.0199;
https://www.liebertpub.
com/doi/pdf/10.1089/cmb.
2017.0199.

Xia:2010:CSP

[XMW10] Xiao-Qin Xia, Michael Mc-Clelland, and Yipeng Wang. Code snippet: PypeR, a Python package for using R in Python. Journal of Statistical Software, 35(CS-2):??, July 2010. CO-DEN JSSOBK. ISSN 1548-

7660. URL http://www.
jstatsoft.org/v35/c02.

Yaacov:2023:BBP

[Yaa23] Tom Yaacov. BPpy: Behavioral programming in Python. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002522.

Yan:2014:PFB

[Yan14] Yuxing Yan. Python for finance: build real-life Python applications for quantitative finance and financial engineering. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78328-438-2, 1-78328-437-4. ix + 386 pp. LCCN QA76.73.P98 Y36 2014.

Yesylevskyy:2015:SNU

[Yes15] Semen O. Yesylevskyy. Software news and updates: Pteros 2.0: Evolution of the fast parallel molecular analysis library for C++ and Python. Journal of Computational Chemistry, 36(19): 1480–1488, July 15, 2015. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Yang:1998:SOO

[YFD98] T.-Y. Brian Yang, Geoffrey Furnish, and Paul F. Dubois. Steering object-oriented scientific computations. In Ege et al. [ESM98], pages 112– 119. ISBN 0-8186-8383-X. LCCN ???? IEEE catalog number 97TB100221.

Yuan:2021:PPP

[YH21] Zhenfei Yuan and Taizhong Hu. pyvine: The Python package for regular vine copula modeling, sampling and testing. Communications in Mathematics and Statistics, 9 (1):53–86, March 2021. CO-DEN ???? ISSN 2194-6701 (print), 2194-671X (elec-

                                        tronic). URL http://link.
                                        springer.com/article/10.
                                        1007/s40304-019-00195-2.

Yang:2016:PDI

[YHA^{+}16] Jean Yang, Travis Hance, Thomas H. Austin, Armando Solar-Lezama, Cormac Flanagan, and Stephen Chong. Precise, dynamic information flow for databasebacked applications. ACM SIGPLAN Notices, 51(6): 631–647, June 2016. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yang:2022:MPF

[YHX22] Yilin Yang, Tianxing He, and Baowen Xu. Mining Python fix patterns via analyzing fine-grained source code changes. Empirical Software Engineering, 27 (2):??, March 2022. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-021-10087-1.

Yadav:2019:PPB

[YKKD19] Vinay Yadav, Subhankar Karmakar, Pradip P. Kalbar, and A. K. Dikshit. Py-TOPS: a Python based tool for TOPSIS. SoftwareX, 9 (??):217–222, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018302279.

Yu:2023:REP

[YMA23] Jiaxin Yu, Tapan Mukerji, and Per Avseth. rockphypy: an extensive Python library for rock physics modeling. SoftwareX, 24(??):??, December 2023. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711023002637.

Yoon:2023:PFO

[YMCF23] Tae Jun Yoon, Katie A.

Maerzke, Robert P. Currier, and Alp T. Findikoglu. PyOECP: a flexible open-source software library for estimating and modeling the complex permittivity based on the open-ended coaxial probe (OECP) technique. Computer Physics Communications, 282(??): Article 108517, January 2023. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465522002363.

Younker:2008:FAP

[You08] Jeff Younker. Foundations of agile Python development: [Python, agile project methods, and a comprehensive open source tool chain!]. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-981-

0. xxi + 393 pp. LCCN

QA76.73.P98.

Yang:2016:VAV

[YPB16] Yuting Yang, Sam Prestwood, and Connelly Barnes. VizGen: accelerating visual computing prototypes in dynamic languages. ACM Transactions on Graphics, 35 (6):206:1–206:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zaccone:2015:PPP

[Zac15] Giancarlo Zaccone. Python parallel programming cookbook: master efficient parallel programming to build powerful applications using Python. Quick answers to common problems. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-958-6, 1-78528-672-2. 286 (est.) pp. LCCN QA76.73.P98. URL http://

proquest.safaribooksonline.
com/?fpi=9781785289583.

Zdybal:2020:PPS

[ZAPS20] Kamila Zdybał, Elizabeth Armstrong, Alessandro Parente, and James C. Sutherland. PCAfold: Python software to generate, analyze and improve PCA-derived low-dimensional manifolds. SoftwareX, 12(??):Article 100630, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711020303435.

Zhang:2020:GPP

[ZC20] Yifei Zhang and Jia Cao.

GSimPy: a Python package for measuring group similarity. SoftwareX, 12(??):Article 100526, July/December 2020. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711019303590.

Zhao:2020:ITC

[ZGL20] Yafei Zhao, Paolo Vincenzo Genovese, and Zhixing Li. Intelligent thermal comfort controlling system for buildings based on IoT and AI. Future Internet, 12(2):30, February 10, 2020. CODEN ???? ISSN 1999-5903. URL https:/

/www.mdpi.com/1999-5903/
12/2/30.

Zhao:2023:PFE

[ZGV^{+}23] Yanli Zhao, Andrew Gu, Rohan Varma, Liang Luo, Chien-Chin Huang, Min Xu, Less Wright, Hamid Shojanazeri, Myle Ott, Sam Shleifer, Alban Desmaison, Can Balioglu, Pritam Damania, Bernard Nguyen, Geeta Chauhan, Yuchen Hao, Ajit Mathews, and Shen Li. Py-Torch FSDP: Experiences on scaling fully sharded data parallel. Proceedings of the VLDB Endowment, 16 (12):3848–3860, August 2023. CODEN ???? ISSN 2150-8097. URL https://

                                        dl.acm.org/doi/10.14778/
                                        3611540.3611569.

Zhou:1997:CCJ

[Zho97] Ling Zhou. A comparison of C++, Java and Python. Thesis (m.s.), Department of Computer Science, Texas A&M University, College Station, TX, USA, 1997. vi + 44 pp.

Ziade:2006:PP

[Zia06] Tarek Ziadé. Programmation Python. Eyrolles, Paris, France, 2006. ISBN 2-212-11677-2. xxxviii + 537 pp. LCCN ????

Ziegenhagen:2019:CLP

[Zie19a] Uwe Ziegenhagen. Combining L^{A}TEX with Python. TUGboat, 40(2):126–128, ???? 2019. CODEN ???? ISSN 0896-3207. URL https:/

/tug.org/TUGboat/tb40-2/
tb125ziegenhagen-python.
pdf.

Zienert:2019:CTP

[Zie19b] Tilo Zienert. cp-tools: a Python library for predicting heat capacity of crystalline substances. SoftwareX, 9 (??):244–247, January/June 2019. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711018301791.

Ziade:2008:EPP

[ZKB^{+}08] Tarek Ziadé, Paul Kennedy, Shannon Behrens, Wendy Langer, and Siddharth Mangarole. Expert Python programming: learn best practices to designing, coding, and distributing your Python software. Packt Publishing, Birmingham, UK, 2008. ISBN 1-84719-494-X, 1-84719-495-8 (ebook). LCCN A76.73.P98 Z53 2008. URL http://

proquest.safaribooksonline.
com/9781847194947.

Zia:2020:PPH

[ZL20] Haseeb Zia and Brice Lecampion. PyFrac: a planar 3D hydraulic fracture simulator. Computer Physics Communications, 255(??): Article 107368, October 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0010465520301582.

Zhao:2023:CMP

[ZL23] Yang Zhao and Qing Liu. Causal ML: Python package for causal inference machine learning. SoftwareX, 21(??):??, February 2023. CODEN ???? ISSN 2352-7110. URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S2352711022002126.

Zhang:2014:AIO

[ZLBF14] Wei Zhang, Per Larsen, Stefan Brunthaler, and Michael Franz. Accelerating iterators in optimizing AST interpreters. ACM SIGPLAN Notices, 49(10):727–743, October 2014. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zerouali:2021:UJP

[ZMD21] Ahmed Zerouali, Tom Mens, and Coen De Roover. On the usage of javascript, Python and Ruby packages in Docker Hub images. Science of Computer Programming, 207(??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167642321000460.

Zhu:2018:OPL

[ZMS18] Minjie Zhu, Frank McKenna, and Michael H. Scott. OpenSeesPy:

Python library for the OpenSees finite element framework. SoftwareX, 7 (??):1–5, January/June 2018. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711017300584.

Zolotov:2021:PFO

[ZRK21] Oleg Zolotov, Yulia Romanovskaya, and Maria Knyazeva. pyFIRI — a free and open source Python software package of the nonauroral Earth’s lower ionosphere. SoftwareX, 16 (??):??, December 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021001461.

Zolotov:2022:VPF

[ZRK22] Oleg Zolotov, Yulia Romanovskaya, and Maria Knyazeva. Version 2.0pyFIRI — a free and open source Python software package of the non-auroral Earth’s lower ionosphere. SoftwareX, 20(??):??, December 2022. CODEN ???? ISSN 2352-7110. URL http:

//www.sciencedirect.com/
science/article/pii/S2352711022001819.

Zampieri:2021:ARA

[ZTC^{+}21] Matteo Zampieri, Andrea Toreti, Andrej Ceglar, Pierluca De Palma, Thomas Chatzopoulos, and Melania Michetti. Analysing the resilience of agricultural production systems with ResiPy, the Python production resilience estimation package. SoftwareX, 15(??):??, July 2021. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711021000716.

Zaman:2022:PPL

[ZTT22] Mashiyat Zaman, Kotaro Tanahashi, and Shu Tanaka. PyQUBO: Python library for mapping combinatorial optimization problems to QUBO form. IEEE Transactions on Computers, 71(4): 838–850, April 2022. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Zukowski:1997:ISU

[Zuk97] Monty Zukowski. Implementing a selective undo framework in Python. In Anonymous [Ano97c], pages 69–75. ISBN ???? LCCN ????

URL http://www.python.
org/workshops/1997-10/proceedings/
zukowski.html.

Zhang:2019:SSE

[ZV19] Zhiping Zhang and Jeffrey D. Varner. SEML: a simplified English modeling language for constructing biological models in Julia. IFAC-PapersOnLine, 52 (26):121–128, 2019. ISSN 2405-8963. URL https:

//www.sciencedirect.com/
science/article/pii/S2405896319321299.

8th Conference on Foundations of Systems Biology in Engineering FOSBE 2019.

Zhang:2022:PIP

[ZWCQ22] Yinsheng Zhang, Haiyan Wang, Yongbo Cheng, and Xiaolin Qin. pyCLAMs: an integrated Python toolkit for classifiability analysis. SoftwareX, 18(??):??, June 2022. CODEN ???? ISSN 2352-7110. URL http:/

/www.sciencedirect.com/
science/article/pii/S2352711022000188.

Zhang:2023:RRB

[ZXX23] Qiang Zhang, Lei Xu, and Baowen Xu. RegCPython: a register-based Python interpreter for better performance. ACM Transactions on Architecture and Code Optimization, 20(1):14:1–14:??, March 2023. CODEN ???? ISSN 1544-3566 (print), 1544-3973 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3568973.

Zhang:2024:PMJ

[ZXX24] Qiang Zhang, Lei Xu, and Baowen Xu. Python meets JIT compilers: a simple implementation and a comparative evaluation. Software— Practice and Experience, 54 (2):225–256, February 2024. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Zhang:2022:QIO

[ZXZX22] Qiang Zhang, Lei Xu, Xiangyu Zhang, and Baowen Xu. Quantifying the interpretation overhead of Python. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0167642321001520.